A Complete Bibliography of Publications in Ambix: Journal of the Society for the History of Alchemy and

Chemistry

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.15

Title word cross-reference

$115 [Mor06]. $119 [Ano07b]. $15 [Ano06b]. 1800 [Lyk18]. $186 [Sha05]. $232 [Num07]. $24 [Ano06b]. $24.00 [Ano97c]. $24.95 [Ano97c]. $29.95 [Hal11]. $30 [Ano06e, Vau04]. $34.50 [Las96]. $37.50 [Ano05b]. $48 [Ano07e]. $49.95 [Ano97c]. $69.95 [Vic94]. $80 [Rat92a].

-forskning [Nie16]. -talets [Lyk18].

0 [Ano97c, Ano05c, Ano06c, Ano06d, Ano07a, Bro97c, Dob82a, Don88, Kni04, Kra16a, Las96, Rat92a, Tra05, Vau04, Vic94].

0-19-511229-6 [Kni04]. 0-19-938334-0 [Kra16a]. 0-226-75018-3 [Las96]. 0-262-13427-6 [Tra05]. 0-262-51134-7 [Tra05]. 0-333-97001-2 [Ano05c]. 0-333-97005-5 [Ano05c]. 0-374-11679-2 [Vau04]. 0-465-02775-X [Ano06d]. 0-521-23143-4 [Dob82a]. 0-521-36144-3 [Vic94]. 0-521-49636-5 [Bro97c]. 0-521-57199-5 [Ano06c]. 0-521-66295-8 [Ano06e]. 0-7185-0020-2 [Ano97c]. 0-7546-3352-7 [Ano06e]. 0-8047-3972-2 [Ano06a]. 0-8135-2234-X [Ano97c]. 0-8135-2235-8 [Ano97c]. 0-8229-4156-2 [Tra05]. 0-8412-1776-9 [Ano97c]. 0-85404-464-7 [Ano07a]. 0-941901-30-0 [Ano06e]. 0220 [Hom19]. 0220-5521 [Hom19]. 0940 [Bro20a]. 0940-3426 [Bro20a].

1 [Ano68a, Ano05b, Ano07b, Bar19a, Bur02, McE78a, Mor06, Per03, Roo23, Sha05]. 1-4020-1039-7 [Ano07b]. 1-4039-2033-8 [Mor06]. 1-85196-109-7 [Bur02]. 1-85196-123-9 [Bur02]. 1-85196-522-X [Bur02]. 1-86094-151-6 [Ano07e]. 1-86094-228-8 [Ano07e]. 1-86094-389-6 [Ano06b]. 1-86094-390-X [Ano06b]. 1-86094-506-6 [Ano07e]. 1-86094-576-7 [Ano07e]. 1-873132-72-7 [Ano05b]. 100Jährigen [Bro20a]. 11 [Cra10]. 1190/$1950 [Bur02]. 1386-4328 [Bro02]. 14-volume [Bur02]. 1433 [Bro02]. 1433-5158 [Bro02]. 15th [Wel82]. 1627-91 [Pri16b]. 1639 [Iac22]. 1650s [Pri13]. 16th [Rap18, Sin59, SvOM93]. 16th-Century [SvOM93]. 1720 [Rag22]. 1750-1850 [Hij21]. 1770s [Ant22]. 1780s [NW18]. 1790s [Ant22]. 17th [Ell22, Str79]. 17th-Century [Ell22, Str79]. 1800 [Kno22]. 1820s [Hor18]. 1834 [Bro86a]. 1840 [Gol18]. 1860s [MT92]. 1867 [Bro16c]. 18th [MTKT^{+}15]. 1900 [Low18]. 1908 [Fer09]. 1920s [Cec04]. 1930s [Car16a, SMGL18, Hor18]. 1950 [Ano97c]. 1953-2020 [Kra21]. 1960s [Cec04, NW18, Qui20]. 1970s [Car16a]. 1973 [Ano75c]. 1975 [CS19]. 1989 [vM90, Rat92a]. 1990s [Qui20]. 19th [Lyk18, Rap18].

2 [Cob22, FFS74a, HS23, Jam19b, Kin82, KCU81, Per04, Pul19, Tay23, Vau04]. 2004 [Die04a, Mor06]. 2006 [LP05]. 2008 [Ano07f, Ano07g]. 2012 [Hom17b]. 2012/Building [Hom17b]. 2017 [Ano16a, Ano16b, Ano16c, Ano16d]. 2018 [Ano17b]. 2019 [Ano19a]. 2020 [Ano19b, Ano19c, Ano20b]. 2021 [Ano20a, Ano21a, Ano21b]. 2022 [Ano22c]. 2023 [Ano22a, Ano22b, Ano23a, Ano23b].

20th [Orl16, Ram18, Sim16a]. 21st [Bro20d].

3 [Ano04a, Ano06e, Don88, FFS74b, Fer22, Kno22, Las96, Lau19, Run20, Vic94]. 3-406-50962-2 [Mor06]. 3-406-51067-1 [Ano05a]. 3-406-51484-7 [Mor06]. 3-88208-060-0 [Don88]. 3-907770-89-7 [Ano06e]. 3-929134-34-9 [Ano04a]. 3426 [Bro20a]. 3r [Fer09].

4 [Dob82a, FFS75, Her18, Iac22]. 495/$740 [Bur02].

5 [Ano05c, Ano06c, Bro97c, Cha23, CG22, Cra10, FFS76, Hal11, Tho22]. 5-93808-082-7 [Ano06d]. 5158 [Bro02]. 5521 [Hom19].

6 [FFS77, Kni04, New23, Rag22, Tra05, Num07]. 6-volume [Cha23, Duf23, HS23, New23, Roo23, Tay23]. 6.Das [Bro20a]. 60 [Nor15]. 69 [Hom19]. 6r [Fer09].

7

[Ano05b, Ano06e, Ano07a, Ano07b, Hed22]. 74.00 [Bar19b]. 75th [Mor12].

8 [Mor06, Vag19, Van22]. 84-472-1046-4 [Num07]. 84-7491-792-1 [Num07].

9 [Ano97c, Ano04a, Ano08a, Bur02, Duf23, Kim20, Ree22]. 90 [Rat92a, Sha05, Num07]. 90-04-09287-0 [Rat92a]. 90-04-11516-1 [Sha05]. 90-04-14128-6 [Num07]. 965-555-149-0 [Mor06]. 978 [Ano08a, Bar19a, Cha23, Cob22, CG22, Duf23, Fer22, Hal11, Hed22, HS23, Her18, Iac22, Kim20, Kno22, Lau19, New23, Pul19, Rag22, Ree22, Roo23, Run20, Tay23, Tho22, Vag19, Van22]. 978-0-19-066853-2 [Pul19]. 978-0-19-537613-5 [Hal11]. 978-0-22-668086-6 [Rag22]. 978-0-22-672705-9 [Ree22].

978-0-26-253869-5 [Mos22]. 978-0-7546-5181-9 [Ano08a]. 978-0-8014-4546-0 [Hal11]. 978-0-8142-1338-4 [Her18]. 978-1-42-143995-2 [Cob22]. 978-1-4696-3714-3 [CS19]. 978-1-4696-3715-0 [CS19]. 978-1-4696-3716-7 [CS19]. 978-1-4742-9453-9 [Duf23]. 978-1-4742-9454-6 [New23]. 978-1-4742-9459-1 [Roo23]. 978-1-4742-9465-2 [Tay23]. 978-1-4742-9480-5 [Cha23]. 978-1-4742-9481-2 [HS23]. 978-1-48-750177-8 [Van22]. 978-1-84-384593-5 [Tho22]. 978-2-36781-323-3 [Run20]. 978-3-402-16005-3 [Lau19]. 978-3-447-11265-9 [Kim20]. 978-8-82-226566-1 [Bar19a]. 978-84-9888-976-5 [CG22]. 978-9-08-704968-3 [Kno22]. 978-9-81-123353-1 [BS22]. 978-9-81-123353-8 [BS22]. 978-90-04-31522-8 [Vag19]. 978-90-04-36032-7 [Hed22]. 978-90-04-46511-4 [Iac22]. 978-90-04-50337-3 [Fer22]. 981-256-466-7 [Ano07e]. 981-256-600-7 [Ano07e].

= [Ano04a].

A. [Ben70, Jac00, Ser76]. A.D. [Füc51]. AA [Bro02]. år [Nie16]. Abbot [Bra79]. Abditis [Num07]. Aberdeen [Rea61]. Abiogenesis [Abr64]. Abraham [Pri16a]. Abstractions [Mor21a]. Academia [RS04, Tya04, dV22]. Academic [Ano07b, Bro02, Roc94, Sme87c]. Academie [Str79, Byc21, Kim96]. Academies [BS14]. Academy [DLS86, And10, Kon13, Nie16, Sme61]. Accademiche [Car17b]. Acceptance [Kim91]. Accomplishment [Kni86].

According [Bol76, Füc51, Ryd94]. Accounts [Bro72, RW88]. Acetic [BN80]. Achievement [Cro09]. Achievers [Ano06e, BCS03]. Acid [BN80, Le 73, NG20, War76a]. Acids [Hun99, Kau73]. Across [Riv22]. Active [Lev21, McG68]. Activities [App77]. Activity} [Tre79]. Acts [Sme68]. Adam [PW69]. Adaptation [Pri16a]. Added [Sme87c]. Addenda [Rob66, GC80]. Adding [Hom17a]. Additions [Tay14a, Tay14b]. Adept [PZ18, Sch20, Zub21]. Administration [Kon13]. Adolph [Kel14]. Adolphe [Car93, Roc94]. Adoption [Hed21]. Adsorption [Pal10]. Adulation [Cob19]. Adulteration [Cob19, Cob20, SW14]. Advantage [Pal10]. Advocate [Fis96]. Aerial [McE78a, McE78b, McE78c, Mce79]. Aesthetic [Rag22]. Aesthetics [Ful19]. Aether [Hei73, Kra89]. Aethereus [Pri86]. Affair [Van22]. Affinities [Leh10]. Affinity [Dun70, Lev70, Sme63, Tay08]. After [All92, Gay12, Fau11, Van04, War97]. Aftermath [Joh11]. AG [Die04a, Mor06, Die04a]. Again [FPS16]. Age [Cob20, Deb81, Hem15, Kra22a, Mon63, Roq16, Bro21, CB23, HS23, Roo23]. Agents [Qui16a]. Ages [KF70, Wil75, New23]. Agricola [Roc85]. Agricultural [Fin91, Oga08, SB92, Jam15]. Agriculture [Mun91, Jam15]. Agrippa [MJ75, New82]. Agustí [CG22]. Aha [Gro21]. Ahmad [KF73]. Air [BB04, Fic71, Hom17b, Koh75, Le 73, Sco70, Sme68, Tra12a, Wal80]. Airs [Tom09]. Akademi [Nie16]. Al [SLT49, Abr84, Hey38a, Ryd94]. Al-Jawbari [Abr84]. Al-Khwarizmi [Ryd94]. Al-Razi [Hey38a]. Al-Waraqi [SLT49]. Albert [Gru06]. Albertan [Gru06]. Albertus [Gru09, New14, New15, Par37a]. alchemiae [Mou09]. Alchemical [App77, Bay05, Bri66, Clu05b, Dav87, Duf15, Eam21, Fer09, Fig73, For05, Gru06, Hay77, Hey38b, Jos49b, Jos63, Kah10, Kah11, Kar88, Kar90b, Kib60, KLPG06, Lan22, Mar09, Mor85, Nor20, Num01, Obr09, Pap90, Par37b, PVPP17, PHW22, PLB23, Pla82, Pri92, Pri13, Pri16d, PZ18, Ram08, Ram10, Rob66, Roo17a, Sal17, She60, SAHL62, Suc16, Taa14, Tay37a, Tay38, Tay49a, Tay49b, Tay56, The99, Tho56, Tho63a, TCPYN59, Wel82, Wil63a, Wil66, Wuj84, Zub21, dS22, Fer17]. Alchemie [Lau19, Vin09]. Alchemist [Bar19b, Bil13, Bur92, Clu17b, Dam77, DN37, Fig65, Fri59, Kar90a, Kau83, Nor20, Pat88, Pri21b, Run18, Sch73, Spa20, The91, Wil63b]. Alchemists [Abr84, Cha07, For13, HN59, Lin77, New14, New15, New20b, Wel51, Wil62]. Alchemy [All03, Ano84c, Ano16d, Ano17c, Ano17b, Ano19a, Ano19c, Ano21b, Ano22c, Ano23b, Bal93, Bil13, Bra85, Bre76, Bro11, Bro48, Bur92, Cam01, CGT18, Cha07, Clu17a, Clu17b, Cri73, Del20, Dob73, Dob74, Dub61, Duf20, Dym08, Fer22, Fer16, Fra17, Fri21a, Füc51, Gan20a, Geo57a, Geo57b, Gia18, Gre38, Gru09, Ham82, Has61, Hed13a, Hed13b, Hed21, Hei16, Her87, Hey37b, Hey38a, Jan20, Jon18, Jon21, Kar03, Kel12, Lan66, Lau19, Lei16, Lin84, Mac67, Mah76, Mah77, Mar17, Mar21, Mar75, Mar04, MT11, Maw16, Maw18, Mer20, Mer90, Mon37, Mon63, Mor22a, Mor22b, Mou09, MJ75, New90a, New20c, Num07, Num13, Par37a, Par38b, Pat82, Pay89, Pet15, Pow98, Pri87, Ram16, Rat66, Rat92a, Rea38, Ric09, Rin17, Rya90, Ryd94]. Alchemy [Sad75, Sad77, Sán16a, Sec79, She57, She58, She62, She70b, She85a, Sim87, Sta53, Szu86, Szy93, Tar18, Tay37b, Tho38b, Tra05, Vin09, Vrt23, Wes61, Wes84, Wey76, Zim84, Zin84, Zub18, vM90, Hed17, LP05, MTKT^{+}15, Pio18, She59, vM90, Clu07, Rat92a]. Alchimie [Pio18, Fra17]. alchimique [Fer17]. Alchimy [Rei57]. Alchymie [Fri59]. Alchymists [Geo62]. Aldershot [Ano06e, Ano08a]. Aleksandr [Rae89]. Alembics [SvOM93]. Alexander [Rag22, Kik09, Lar67, Pri86, Roc22, AGJ08]. Alexandria [Pap90, Tay37a, Tay38]. Alfonso [KM90]. Alfred [Kau72, Van22, Kau73]. Algorithm [Lan22]. ’Ali [KF73]. Alison [Ano06e]. Alkali [Ree12, Ree15]. Alkaloid [Uss03]. Alkaloids [Hom17a]. Allen [Tra05, Clu07, Pri09]. Allin [Bil13]. Almighty [Bax13]. Alongside [Roc21]. Aloys [Kea92]. Alpha [Tre79]. Alquimia [LP05, Num07]. already [Las03]. Alter [Bro20a]. Alternative [Ric04]. Aluminium [Le 15]. am [Bro16b, Dav87, SL22]. Amadeo [Car17b]. Amalgamated [RGG23]. Amalgamation [Vrt23]. Amalgams [Mar23]. Amanuenses [Fri22]. Amateur [SG23]. Amazonian [RGG23]. ambiental [CG22]. Ambix [Ano68a, Ano90b, Clu07, Mor12, Ram13, Sme87a]. Ambrose [Zub21]. Amedeo [Cia01, Sán17]. America [Mor06, GW17, Han76, Ros96, Tra04, Wil63a, Wil66, Whi18]. American [Ano97c, Mor06, Ano97c, Con96, Far20, Fis18, Kop76, Mei95, PT17, Tra04]. among [New14, New15]. Amsterdam [Sne79]. Amy [Ano06e]. an-Nadim [Füc51]. Anaesthetic [Rae16]. Analogy [Rus19, Tat19]. Analyses [SMGL18, Uss03]. Analysis [DFB22, Edd01, Jam83a, KR14, Lan22, Pol15, Pow14, Rae97, Sne79, Sou92, SW14, Tra14]. Analytical [Bro19a, Hij22, Hom99, Kle05, SW14]. Anatomical [Kno06]. Ancient [Ano75c, Ebe07, GL17, Hal11, Mar17, Mar21, Hor60, Kel71, Lev57, Lev58, MB21]. Andean [RGG23]. Andreas [Bor15, Don88, dV22]. Andrews [Sha02]. Anew [Tom16b]. Anfängen [Bro20a]. Angelo [Gel94]. Anilin [Kea92]. Aniline [RT97]. anima [Mou09]. Animal [Col71, Col96, Lef18a]. Anna [Num01]. Annales [Cou72]. Anne [Ant22].

Anniversary [Mor12]. annotated [Bro16c]. Annotations [Jos64]. Announcement [Ano89a, Ano89b, Ano94a, Ano97a, Ano00b, Ano00a, Ano01a, Ano19a, Hud01, Lew87, Sut83]. Announcements [Ano86a, Ano86b, Ano87a, Ano88a, Ano90a, Ano90b, Ano92, Ano93, Ano94b, Ano96, Ano97b, Ano98a, Ano00c, Ano01b, Hud98]. Anomaly [Wol72]. Anselme [Mer03]. Anselmo [Lei16]. Answers [Abr64]. Anthocyanin [Sie08]. Anthony [Mor06]. Anthropocene [Sán16c]. Antibiotics [Pui04]. Antiphlogistic [Jac00]. antiphlogistique [Per84]. Antiquity [For38, Hen17, MTKT^{+}15, Mer20, Sta53, Duf20, Rin17, Duf23, Kir18]. Antonio [Bur02]. Antwerp [Jos64]. Apothecarie [Sim14a]. Apothecaries [Sim06]. Apparatus [HB84, Joh18]. Applications [BG00, CT88]. Appreciation [Bro97a, Bro19b, Pri09, Rus01]. Apprehensions [Del20]. Apprentice [For08]. Approach [VMT22]. Approaches [Ano12a, Cec04]. Appropriation [Duc08]. April [Rat92a, vM90]. Aqueous [Nor07, TCPYN59]. Aquinas [Tar18]. Arab [Has61]. Arabic [Hal11, AGJ08, Bur92, Del21, Fer09, Füc51, Hal09, Ham82, KF73, Mar17, MT16, Ple72, Van09]. Arbiter [Sme68]. Archaeological [MT12, Mos15, Per19]. Archéologie [Hom19]. Archetype [Ano05b]. Archibald [Ano07a, Mac88]. Ardent [GDNN72]. Area [Sán16a]. Areas [Fig73]. Argon [Hir81]. Argument [BV16a]. Aristotelian [Man61]. Aristotelianism [Kim20]. Aristotle [Bol76, KF70, Szu86, Via15]. Armstrong [Ric04]. Ars [Pow98, Vin09]. Art [Ano05b, Bal02, Cla01, Fer16, Mar21, MTKT^{+}15, Vau04, Wee07, Hed22]. arte [Mou09, Taa14, Pow98]. Artemii [Fig65]. Artephius [Pol20]. Arthur [Col02, Rus01, Tra05, Abr94, App77, App79, DeK80, Fig65, Sch95]. Articulating [Van04]. Artifact [KCU81]. Artificial [Col84, Far20, Mer04, Tra95]. Artisan [GB91]. Artisanal [Fri22]. Artistry [New20b]. Aschendorff [Lau19]. Asclepius [Vic94, Cop92]. Ascription [Dun65, Tim06]. Ashgate [Ano06e, Ano08a]. Ashmolean [MT12, VMT22]. Asian [BN80]. Aso [CS19]. Aspects [Bro48, HJ02, Mal08, Pri16b, Sch95]. Assay [Bor15, Bur92]. Assistants [Gay00]. Assmann [Hal11]. Association [Bro86a, BS90, Ric04]. Associations [App79]. Assyria [Tho38a]. Astrologer [Bur92]. Astrology [Mon63]. Atlantic [Riv22, Sim22, WW07]. Atmospheric [DM97, Far71, Sán16c]. Atom [Cia01, Roq17, Sin87, Tho22]. Atomic [Col78, Gro21, Pra10, She70a, Sin85, Sin87, War97]. Atomism [Car93, Hor60, Kel14]. Atoms [Cle21, HJ02, Sim14b]. Attendance [Ano22b]. Attested [Ric09]. Attitude [Geo57b, MJ75]. Attitudes [Bra79]. Attractionibus [Sme59]. attribuée [Fer17]. Attributed [Ram10, Sal17, SAHL62, Suc16, Fer17]. Aubervilliers [Kra21]. Aufbruch [Hag16]. August [Bro16b, Mor22a, RW88]. Augustan [Gol83]. Auguste [Lev21]. Augustus [Kin84]. Aurae [Kel14]. Aurea [Hey37a]. Aurifaction [Mul75]. Aurifiction [Mul75]. Aurum [Kel14]. Australia [Rae95, Rae05]. Australian [Mac89]. Austria [SvOM93]. Authentic [Ple69]. Author [Ano38a, GB91]. Authority [Kel12, Mor21a, Mur22, Sim02]. Authorship [Nye14, Sim02]. Autobiography [Kra18]. Autonomous [GL17]. Avicenna [Mou09, New14, New15, SAHL62]. Avicennae [Big17]. Avogadro [Car17b, Cia01, Mor80, Sán17]. Award [Ano87b, Ano05d, Ano07c, Ano07d, Ano16d, Ano17a, Ano17b, Ano19a, Ano19c, Ano21a, Ano21b, Ano22c, Ano22b, Ano23b, Ano20a].

Awards [Ano83]. Awkward [Bee94]. Azure [Gru06].

B [Gor83, KF73, Per84, Sme57, Sal17]. B.C. [Tho38a]. Babel [BV16b]. Back [Arm12, FPS16]. Background [Web66, Web67a, Wil98]. Backhouse [Jos49c]. Bacon [Sad75, Bre76, Gre38, Mat13, New14, New15, Ree75a, Ree75b, Ree77a, Ree77b, Ree79, Wee07, Wes61]. Bacteria [Bro17b, Štr16]. Bad [New20a]. Baden [Jam95]. Baekeland [Mos22]. Bakelite [Mos22]. Baking [Jon93]. Balancing [Tei82]. Balazs [Ano07e]. Balcarres [PVPP17]. Balduin [Kel14]. Ball [Vau04]. Balloon [Tom17b]. Baltimore [Cob22]. Balzac [Sou92]. Band [Bro20a]. Banfi [TJ53, TJ56, App77]. Banfihunyadi [Gom77]. Banned [Pér16]. Baptist [Hed21, Pag62, Roo17a]. Baptista [Hei95, Jen04]. Barbara [Tra05]. Barcelona [CG22]. Barchusen [Han67]. Basel [Ano06e]. Bases [Kau73]. Basic [Ano06d, GW17]. Basil [Pri87]. Basilisks [New20a]. Basingstoke [Ano05c, Mor06]. Basis [Bra85, Deb80]. Batteries [Dol98]. Battery [Owe01]. Batthyány [Pri21b]. Battle [Mur22]. Baumann [Roc21]. Became [Sim16a]. Beck [Ano05a, Mor06]. Becoming [Ant22]. Becquerel [Owe01]. Beddoes [GS61, Lev81, ML08]. Bedfellows [Bal93]. Bedfords [WK16]. Bedi [Tra05]. Beer [Ano04a]. Before [App79, Fis73a, Fis73b, Gay12, War97, Jen81, Per03, Per04]. Beginning [Has61]. Beginnings [Bur16, Con22, Her87, Pra10]. Beguin [Kah21]. Below [Zub21]. Bench [Ano07b, HRR03]. Benito [Lei16]. Bento [CA15]. Bentos [Bro17d]. Benzene [Cam91, Gol80, Wil12]. Beretta [Duf23]. Bergman [Jam22, Sme59]. Bergner [Kar90a]. Berichte [Bro20a]. Berlin [Ano04a, BR00, Fer09]. Bernard [Kir61]. Bernhard [Mor06, Gel78, Sho03]. Béroalde [Zin84]. beroepsgroep [Kno22]. Berthelot [Ant19]. Berthollet [GI97, Le 75, Sme77, Sme78]. Best [Dob82a]. Beth [Ano06e]. Better [O’N21]. Between [Kar93, Kle95, Rag12, Sad77, Sel07, Sme68, Taa14, Tar18, Van22, Bro05, ML08, Ros02, GL21]. Beyond [And10, Ant17, Mos22]. Bibl [Tho63a]. Biblical [Hed13b]. Bibliographical [GC80, Sme57, Sme77, Wil73]. Bibliography [Ber92, Hey37b, GC80]. Biblioteca [LP05]. Bibliotheca [Ano84a]. Bibliothek [Mor22a]. biddes [Gru06]. Big [Pet15]. Bildungsbürgertum [Mun90]. Bill [Col02]. Biochemical [Nic12, Sco70]. Biochemistry [CA15, Van04]. Biographical [Jam19a, Kra22b, Poo09]. biographies [Sim14b]. Biography [Bro20d, Dob82a, Hud99, Mor09, Sal13, Sim87, Söd07, Wes80, Ano08a]. Biological [Gay12, Sha20, Tei82]. Biologics [Qui16a]. Biology [Cec04, Cre17]. Biomedical [Nor15, Sla04]. Biomedicine [Gab16]. Biomolecular [Mas91]. Bionomica [Don88]. Bionomica-Verlag [Don88]. Bioproperty [Gab16]. Biotechnology [GS16]. Birth [Ram00]. bis [Bro16c, Bro20a, Hom17b, LE03, Mor06]. Bischoöfl. [Tho63a]. Bitters [AF16]. Bitumen [For38]. Blachford [War84]. Black [And06, Don78, Per82]. Bladel [Hal11]. Blomfild [Sch73]. Blood [Ver22]. Bloomsbury [Cha23, Duf23, HS23, New23, Roo23, Tay23]. Blowpipe [Sal19]. Blowpipes [Dol98]. Bloxam [DLS86]. Board [Jam15]. Bodies [Her14, Mar23, Mat13, RGG23]. Bodleian [Jos49b]. Body [Cis08, Dan06]. Boë [Rag12]. Boerhaave [And06, Chr94, Gib58, Han67, Hen18, Kno01, Kno06, Lov72, Pow14]. Boerhaavian [Sum07]. Böhme [Ano05a]. Bohr [Roq17, Sim14b]. Boldizsár [Pri21b]. Bologna [Pri16c, Sel07, Tho56]. Bomb [War97]. Bombay [Tya04]. Bonaventura [Suc16]. Bonjour [Sme59]. Bonn [Ano04a, BR00]. Bonus [Cri73]. Book [Ano90b, Ano04a, Ano05b, Ano05a, Ano05c, Ano06c, Ano06a, Ano06b, Ano07a, Ano07b, Ano08a, Bág16, Bar19a, Bar19b, Bay05, BS22, Bro17b, Bro19a, Bro20a, Cha23, CS19, Cob22, CG22, Dob74, Duf23, Fer22, Gol18, Gor19, GL18, Hed22, Hen19, HS23, Her18, Hom19, Hor18, Iac22, Jew19, Kah19, Kim20, Kir18, Kni04, Kni16a, Kno22, Kra16a, Kra18, Lau19, Low18, Maw18, Min18, Mor18a, Mos22, New23, Num07, Pri79, Pul19, Rag22, Ram18, Rap18, Ree22, Roo18, Roo23, Roq16, Run18, Run20, Sch19, Tak17, Tay19, Tay23, Tho22, Tom18, Vag19, Van22, Vau04, Whi18, AGJ08]. Books [Ano70e, Ano70f, Ano71a, Ano72a, Ano72d, Ano73a, Ano73b, Ano73e, Ano73f, Ano74a, Ano74c, Ano75a, Ano75b, Ano75g, Ano77c, Ano05b, Ano06d, Cla01, Dym08, Gia18, Sme87c, LP05, Num07, Pio18]. Bordoni [Vag19]. Borodin [Rae89]. Børsen [BS22]. Bostock [Nev60]. Bostok [Har00]. Boston [Ano07b, Num07, Sha05]. Boundaries [Bud20]. Bowden [Ano06e]. Bowels [Ree79]. Box [Bro02]. Boyle [Bou22, BGH78, Bur02, Cle21, Eam80, HD00, HCP01, Kim91, MP02, Pil38, Pri92, Pri16b, Wal80, Web66, Wes61, Zub21, Bur02]. Brahm [Dam77]. Branco [Lei16]. Brass [Pag14]. Brazil [dMC15]. Brazilian [Fil92, LK12]. Breeders [Mor72]. Breeding [Bro03a]. Brendels [Pri21a]. Brevis [HS78]. Brewer [Tho22]. Brewing [Sum07]. Bridging [Maw16]. Briefe [Bro16c]. Bright [Bal02, Vau04]. Brill [Fer22, Hed22, Iac22, Num07, Rat92a, Sha05, Vag19]. Bristol [Oga08]. Britain [Bro00a, Bro22, Di 22, Dol98, Duc08, Gay12, PG21, Jam21, Kee13, Mal08, NG16, RK06, MR20, SW14]. British [Hor18, Bro86a, BS90, Bro13, Bro17a, Bro20e, Bur16, Hor11, Qui16b, Rea61]. Britton [Cam77]. Broad [Tra95]. Brodie [Iri13]. Brown [Lar67]. Browne [Deb62b]. Bruce [Roo23]. Brunschwig [Taa14]. Brunswick [Ano97c]. Bucquet [McD66, Per89]. Building [Hom17b]. Bullock [Bro20b]. Bungay [Sad75]. Bunsen [Bro13, Bro16b, Jam95, Naw14]. Burn [Tom09]. Burnett [New23]. Business [GB91, Mos22]. Business-Leader [GB91]. Butlerov [Lar71]. Bypass [TR99].

C [Ano05a, Mor06, Har00, QR20]. C. [For08, Le 75]. C.-L. [Le 75]. ca [Col84, Mer22, Ram10, Rat92a]. ca. [Jon18, Sal17]. Cabala [For13, Wal76, For13]. Cabalistica [For13]. Cabinets [Kee13]. Cabral [CA15]. Cal [Ser76]. Cal-Tech [Ser76]. Calamines [Iri18]. Calcination [McK38, Per69]. Calcinatus [Per69]. Calco [Mor06, Tra04]. Calculus [Iri13]. California [Ano06a]. Call [Ano05d, Ano07c, Ano07d, Ano17a, Ano20a]. Called [Gol80]. Calvinist [Kno01]. Calx [Koh75]. Cambridge [Ano75c, Ano06c, Ano06e, Ano07a, Bro97c, Dob82a, Mos22, Tra05, Vic94, Con22, Nye02, Ple59]. Campaign [Ant22]. Candid [HH00, HH05, Ano07e]. Candles [Kee13]. Canon [Ram08]. Canonis [Big17]. Caprolactam [vR07]. Capture [Mår19]. Capuchins [Riv18]. Carbohydrates [Gel78]. Carbon [Nic12]. Carcinogenesis [Arm12]. Cardiac [Hei16]. Career [Clu05b, BJ21, Num01, MR20, Sim20, Wat95]. Careers [Bud20, VG04]. Carl [Ano06b, Tay19, Jam22]. Carneades [MP02]. cárnica [Bro17d]. Caro [Tra91]. Carolina [CS19, McV75]. Carolyn [Ree22]. Carr [Shm94]. Carrick [Her18]. Carson [Tra12b]. Cartesianism [Rag12]. Cartoon [Sin59]. Case [Car16b, Fri21b, Fri22, GB91, Mor96, Olt05, Sie08, Tre79, VP18]. Cassandra [Tho22]. Castelo [Lei16]. catalog [LP05]. catálogo [LP05]. Catalogue [Ram10]. Cataloguing [Fig73].

Catalysis [Col75, Col76b]. Catalyst [Per82]. Catalytic [ZH05]. Catena [Hey37a]. causes [FH05, Num07]. Causis [Num07]. Caveat [Geo62]. Cavendish [Clu11, Gol17]. celebrada [LP05]. Celebrating [Wil12]. Celebrity [Hem15, Roq16]. Censorship [Tar18]. Centennial [KP78]. Center [Mor06]. Central [Szy93, Szy96, Ree15]. Centre [Kra21]. Centres [Hom05]. Centuries [BV16a, DM97, Kar80, WK16, Old74, Rap18]. Centuries-Long [BV16a]. Century [Ano97c, BSGBBV02, BG00, Bro20d, Deb81, Dol98, Fis18, GB14, Goo69, Hei73, Hol89, Hom05, Joh11, Kra89, Las96, Lef18a, Lyk18, Mar93, Pag14, Pri21a, Rae16, Ram18, MR20, Sal13, Sch89, Sim16a, Sin59, SvOM93, Sud78, Szu88, Tho38a, Tho22, Tra14, Wel82, Whi18, GL21, Leh09, Orl16, Pio18, RK06, Sch06, Abr79, And10, Ave86, Bal93, Ber92, BS14, Big17, Bor15, Bro22, Bur16, Cam01, Car17a, CC80, Chr94, Clu93, DeK09, Di 22, Ell22, Fil92, SG23, Gor77, GW17, Gue15, HVD20, Hed13b, Hen18, Hey38b, Hij22, Jam95, Kon13, Lac17, Lan66, Le 82, Le 15, Leh10, Lei16, Lin84, MTKT^{+}15, Mei83, Nor07, Nye14, O’N21, Olt05, Per13, Ple69, Poo09, Pow98, Pow14, Pri16c, Qui16a, Rag12, Ree77a, Rei15, Ric09]. Century [Rob16, Roo01, Sad77, Sal19, Sco70, Ser13, Ser22, Sha94, Sie08, Sim06, Sim14a, Str79, Suc16, Szu86, Szy96, Tei82, Tha68, Tho38b, VMT22, Wer13, Wey76, WW07, Wil68, Wil70a, PG21, Cha23, Tay23, Vag19, Ram18]. Ceramic [Cha81]. Ceramics [Pol15]. Châtelier [Las94]. Chadarevian [Ano06a]. Chair [Lev81]. Chairman [Ano94c]. chaleur [Ban07]. Challenge [Sla01]. Chamberlain [PZ18]. Change [GI97, Jam21, Joh11, Kin84, Sán16c]. Changed [Bro21, Ree22]. Changes [BGH78]. Changing [Ano06d, Ber11, Dmi04, Pri21a, Rus87]. Chaos [Bax13]. Chapel [CS19]. Chapter [Deb63]. chapters [Tho63b]. Charitable [Riv18]. Charles [FFS77, New23, DLS86, Jam81, Kik09, War84]. Charleton [Gel71]. Charlottenburg [Ano04a]. Charnock [Pri79]. Chatto [Bur02]. Chaucer [Tim06, Dun65, Run18]. Cheating [Cob20]. Chelovek [Dmi04, Ano06d]. Chemia [For13]. Chemiatry [Mor85]. Chemical [AGCF^{+}15, All03, Amb22, Ano97c, Ano06e, Arm12, Ave86, BSGBBV02, Bol76, Bor15, Bro67, Bro86a, Bro16b, Bro20d, Cam77, Cec04, Cha81, Cle21, Clu93, Cob19, Col78, Cra10, Cre17, Cro09, De 85, DeK09, Deb61, Deb68, Deb85, Don78, Dun81, FFS75, Fic71, Fra11, Gay97, Gay12, Gib58, Gor98, Gor77, Gor83, GW17, Gou83, GI97, Gro21, Gue76, Gue15, GL18, Han76, HS18, HS20, Hij22, Hom99, Hor18, Hud99, Iri13, Iri18, Jam83a, Jam19b, Jam21, BJ21, Joh18, Kea92, Kik11, Kim91, Kim96, Kin81, Kin82, Kin84, Kle95, Kle13, Lar71, Lla85, Lyk22, Mac88, Mau70, McE78a, McE78b, McE78c, Mce79, Mer03, Mor06, Moy89, New20c, Nye14, O’N21, Pag14, Par37b, Per10b, PHW22, Pol15, Pom62, Qui20, Rae95, Ram18, Rap18, Ree20, Ree22, Roo01]. Chemical [RA02, Ros96, Rus98, RR05, SC90, Sha20, She48, Sim06, Sim20, Sim98, Sim02, Sin87, Sla01, Sme63, Sme66, Sou92, SMGL18, Sud80, Sum07, Szy96, Tay19, Tra04, Tra05, Tya04, VP18, War76b, Wei18, Wer13, Wil98, vL48, Lyk18, Nie16, Tho63b, Ben70, Bro20b, Bro21, Lac17, MT92, Nie16, Pri87, Pri06, Sme89c, Ano07a, Ree23]. Chemically [Sch22]. Chemicals [Jas14]. Chemicus [Abr94, All19b, Wes75]. Chemie [Ano04a, FM07, LE03, Mor06]. Chemiker [Ano04a, BR00, Joh18, Joh18]. Chemische [Joh18]. Chemischen [LE03, Mor06, Bro16b]. Chemist [BR00, DU79, Fil92, Fin91, GB91, HB84, KM90, Kno01, Mor72, Ree20, Roc22, Sha03, Sho03, Sla01, Sme71, Tos01, Wat95, Kno22]. Chemistry [And06, And10, And11, And15, And17, Ano82, Ano84c, Ano85a, Ano16d, Ano17b, Ano19a, Ano19c, Ano21b, Ano22c, Ano23b, Ant17, Arm12, Bee94, BBS10, Ben72, Ber92, Ber11, BS22, BG00, BJJK81, Bro00a, Bro02, Bro17a, Bro17c, Bro22, Bro87, Bro97c, Bro97d, Bro05, Büt00, CT88, Cha17, Cha21, Cha23, Chr94, Clu07, Cob22, Col88, Col96, dMC15, DLS86, DU79, Deb63, Deb67, Deb71, DM97, Dol98, Duf23, Edd08, Ede93, Ede96, Edm19, Ell22, Fau11, Fig65, Fis96, Fis18, Fra11, FM07, Fre10, Fur21, GB14, Gay00, Gol83, Gol16, Gou82, Gre38, PG21, Hen18, HS15, HS18, HS23, Her18, Hij22, Hom99, Hom05, Hom95, Hor18, Iri17, Jac00, Jam22, Joh11, Kaj13, Kar92, Kea92, Kee13, Kle05, Kni67, Kni74, Kni86, Kni16c]. Chemistry [Kno06, Kon13, Kop76, Kra89, Kra22b, Lan81, Las03, Le 75, Le 82, Leh09, Leh10, LR16, Lev70, Lev58, Löw80, LK12, Lyk20, Maw16, McV75, Mei83, Mil68, Mor69, Mor11, Mor16a, Mos15, Mun90, Mun91, Naw14, New20a, New23, Olt05, Pag69, Par38a, Per03, Per04, Per10a, Per10b, Per13, Per82, Pri06, Rae89, Rae05, Ram00, Ram17, Rea61, QR20, RS04, Rei15, RK06, Rob06, Roc87, Roc21, Roo23, Ros02, Rus37, Rus87, Rus96, Sal19, Sán16b, Sán16c, SB92, Sch89, Sch94, Sel07, Ser13, SL22, Ser22, Ser76, Sme61, SW14, Sum07, Taa23, Tay14a, Tay14b, Tay23, Tha68, Tho38a, Tom16a, Tom17a, Via15, Wei18, Wer13, Wey76, Wil68, Wil70a, Zac56, ZH05, Ano17c, Ant22, Bro02, Bro11, Cra10, Fra17, Lev81, Bro02, Bro19a, Old73]. Chemistry [Tak17]. Chemists [Ave86, Bro69, Bro86a, Bro03a, Bro20e, Bro21, Bro68, Col71, Fau11, HH00, Hor11, Kik09, Kim05, Mac89, Mer22, Min18, Roc87, Sch06, Sut76, TR99, Ano07e, Joh18]. Chemotherapy [Sla04]. Chenevix [DU79]. Cheney [Clu93]. Chernyaev [GLCK76]. Chesne [Kah13]. Chicago [Las96, Rag22, Ree22]. Chile [Tra14]. Chimie [BG00, Hom19, Per84, Fra17, Sme89b, Cou72]. Chimique [Sme77, Sme78, Lef18b]. China [Cra10, Gor77, Kel71, Pol15]. Chinese [HN59, Pay89, She85a, TCPYN59, Ano75c]. Chivalrous [Edm19]. Chlorinated [Tra12b]. Cholera [FFS76]. Christian [Kel14, Num13, Roo17a]. Christoph [Sha05, Kar90a]. Chrysocolla [Mar23]. Chung [Ano07e]. Chymeutike [Zac56]. Chymia [Roc85]. Chymiatria [Mor21b]. Chymica [For13]. Chymical [Cle21, Kle14, MT12, Mor21b, Pri16c, Rag12, Riv18, VMT22, Web67a, WW07]. Chymicall [Clu93]. Chymicarum [GS89]. Chymie [Bai00, Nev62]. Chymist [Poo09, Rue22, Web66]. Chymistry [Bro11, Byc21, Clu11, Duc08, Fri21b, Hir21, Jam15, Kah21, Pri21a, Rag12, Roo07]. Cibinensis [KLPG06]. ciencia [LP05, Num07]. Cilac [Hom19]. Cinchona [Riv22]. Ciphers [Pri92]. Circle [Mor85, Pri13, Szu88, Bay05, Clu93, LR16]. Circulation [Min16, Tei82]. Circumstances [Lyk22]. Citizens [All03, Tra05]. City [Ser13]. Civilisation [Cra10]. Civility [Las96, Sha94]. Claims [Rae97]. Clarified [App79]. Clarke [Ano05b, Dol98]. Classical [Roc87]. Classification [Edd08, Fis73a, Fis73b, Fis74, Iri13, Kap69, Kle05, NW18]. Classifications [BSGBBV02]. Classrooms [BS14]. Clavis [Dob82b]. Clericuzio [Bur02]. Clifford [Bay05]. Climate [BS22, Sán16c]. Climbing [Mun90]. Clinical [Büt00]. Clo [Fra11]. Clock [BV16a, Sha20]. cloth [Her18]. Club [Lac17, Lac17]. Co [Bro22, Nye14, The91]. Co-Authorship [Nye14]. Co-Creator [The91]. Co-operation [Bro22]. Co. [Tra91]. Coal [Mus17]. Cobbold [Ree22]. code [Car17a]. Codes [Pri92]. codice [Car17a]. Codicological [Maw18]. Coercion [Di 22]. Coins [Kar88]. Cold [Min16, Bud20]. Coley [Bro20c]. Colin [Ano07a, Bro97c]. Collaboration [McD66, ML08, Nye14, Per89]. Collaborator [Rei10]. Colleague [Pra91]. Collecting [Edd08]. Collection [Mac67, McV75]. Collections [Con22]. College [Ano06b, Ano07e, Bro17c, Dun65, Gay00, Kik09, Mac89, Tho59b, Tim06, Ben72, Gay00, BBS10]. Colloid [Ede96, TR99]. Colloid/Macromolecule [TR99]. Colloidal [Ede93]. Colloids [Ede96]. Colonial [New20c, CE17, Del20, Wil63a, Wil66]. Colony [Nor20]. Color [Bal02, Kir23, Vau04]. coloso [Bro17d]. Colossus [Bro17d]. Colour [BGH78, Deb62b, Eam80, NG15, Wer13]. Colours [Ano05b, Cla01, HVD20, Tho59a, Tho60b]. Columbus [Her18]. Combination [Bol76, Iri18]. Combinations [Jam19b]. Combrune [Sum07]. Combustion [Lei75, Lys37, McK38]. Coming [GDNN72]. Commemorative [Res23]. Commentaria [Big17]. Commentary [Mar04, Bur99]. Comments [Bro00a, Deb72]. Commerce [Di 22]. Commercial [Maw18, Sum07]. Committee [Bro86a]. Communication [Gol83, Mal08, Mor85, Zot93]. Communications [Ano17c, Car16b, Sme87c]. Community [Kik11, VMT22, Wei18]. Como [Abr79]. Companies [QR20]. Companion [HBB^{+}03, Kni04]. Company [Mor06, Ree15, Tra04]. Comparison [BN80]. Compendia [Mar17]. Competing [Wer13]. Compilation [Hal09]. Complex [Nor15]. Complutense [Num07]. Composition [Del21]. Compostelle [Suc16]. Compound [Clu71, MB21, Gol18]. Compounds [Kau72, Far75]. Compromise [Deb60, Mor21b]. Concealments [Pri92]. Concept [Cia01, Dan06, Hen20, Lov72, Rin17]. Conception [Cis08, Cri73, GI97].

Concerning [Mou09, Sin59, War76b, Wil73]. Conference [Rat92a, Roq17, vM90]. Conflict [Hir81, SW14]. Congrès [Hed22]. Congress [HS15]. Conjecture [Per89]. connection [Sie89]. Conrad [Fri59, Han67]. Consequences [Hom99, GL18]. Conservation [Pom62, Tei82]. Consideration [Lar71]. Considerations [Mou09]. Conspiracy [Bro97c, Rus96]. Constant [Owe01]. Constellation [Mon63]. Constituents [Str79]. construcción [CG22]. Constructing [Jam19a, Kim96]. Construction [Per10a]. Consultancy [MR20]. Consultant [Lyk20, Qui20, Ree20]. Consultants [QR20]. Consulting [Sim20, Wat95]. consumer [Orl16]. Contagion [FFS76]. Contaminated [Tra02]. Contamination [CB23]. Contemporaries [Böh63]. Contemporary [Deb72, Han67]. Content [Uss03]. Context [GI97, Le 82]. Contexts [Kim91, Kle05]. Contextualising [For17]. Continental [Ber11]. Continuous [Sim20]. Contrasting [Bra79]. Contribution [Lla85]. Contributions [Bro20e, Mos15, Sán16b]. Contributors [Ano04b, Ano05e, Ano05f, Ano05g]. Control [Gay12, GL17]. Controversies [Tit38]. Controversy [Bro97c, Cam01, Car93, Kau72, Owe01, Pal10, Rat64, Roo17b, Rus96, Sha20]. Convenience [HS20]. Conversations [Ano07e, HH00, HH05]. Conversion [Le 75]. Converted [Sme68]. Cook [Arm12]. Copenhaver [Vic94]. Copies [McK58, Mck61b]. Copley [Mck61a]. Coptic [Ric09]. Copy [Dob79]. Cordus [Pri86]. Cores [Abr79]. Cornell [Hal11]. Cornwall [Jam19b]. Corporate [vR07, LE03, Mor06]. Corporation [Die04a, Die04b, Mor06, Ser76]. Corporeal [Kle14]. Corpus [Duf15, PW69, Ram10, Vic94, Cop92, KF73]. Corpuscles [Gor19]. Corpuscular [LMN01, Sha05]. Corpuscularism [Big17]. Correction [Ano68a, Win98]. Corrective [O’N21]. Correspondence [Bur02, Clu93, HCP01, Pri21b, Van22]. Correspondents [Sha02]. Corridor [All03, Tra05]. Corrigenda [Rob66]. Corundum [Iri17]. Cosmic [Mas91]. Cosmogonic [Bur99]. Cosmology [Pap90, Ree75a, Ree75b, Ree77a]. Cossa [KM90]. Costa [Kra16a]. Counterculture [vR07]. Counterfeits [And16]. Countries [SC90, Vag19]. Country [Jew19]. Course [Kin84, Sme61]. Courses [BBS10, Kah21, Leh09, Leh10, Per10a, Per10b]. Court [mKC21, Mor21a, Lin77]. Courtney [Jam19b]. Courtrooms [Sán21]. Courts [BS14]. Cradle [Gol16]. Craft [Taa14]. Craig [Qui20]. Cramer [Bor15]. Crater [Car16a]. Creating [Per03, Per04]. Creativity [Joh11]. Creator [The91]. Creiling [Fri59]. Crisis [Joh11]. Critical [Hen20]. Criticism [dV22]. Critique [Clu11, Lei16]. Critiques [Ras18]. Cronstedt [Sme66]. Crookes [Jam81, Sin85]. Cross [Gan20a, Mon63, Pag69]. Cross-Roads [Pag69]. Crosthwaite [Pra91]. Crow [Ano06e]. Crucible [Gan20a, Mon63]. Crucibles [SvOM93]. Crum [Lar67]. Crust [Deb61]. Crutzen [Sán16c]. Crystallographic [Iri17]. Crystallography [Goo69, Sal13]. Crystals [Iri18, Bro16b]. Cucurbits [SvOM93]. Cullen [Chr94, Tay14a, Tay14b]. Culpeper [Clu93]. Cultic [Wil98]. Cultural [Ano97c, BB04, Cha23, Duf23, HS23, Jam95, Mei95, New23, RS04, Roo23, Sim16b, Taa23, Tay23, Zim84, dS22]. Culture [Clu17a, Hem15, Hen17, Nye14, Roq16, VMT22]. Cultures [PLB23]. Cupellation [MT16]. Curare [Car16b, Car16b]. Cure [Pat92, SM17]. Curie [Roq16, Hem15]. Curious [Bro20b]. Currents [Sch89]. Cusa [Lau19]. Cyanamid [Mor06, Tra04]. Czech [Kar90b].

D [Ano06d, Ano07e, Bro02, Tho22, Ram10, Sal17]. D-76128 [Bro02]. D. [Alm98, Dmi04]. D.2.8 [Dun65, Tim06]. D.I. [Dmi04]. Dagognet [HS18]. d’Alchimie [Run20]. Dale [Tra91]. Dalton [Col78, FFS74b, Gro21, Pra91, Pra92, Pra10]. Daltonian [Mau70, Sme77]. Daltoniana [Nev60]. Daniel [Dol98, Kim20, Tay23, Hir21, Kle15]. Daniell [Owe01]. dans [Run20]. Dante [Bar19b]. Dark [Edd08, Far20]. Darkness [DP03, Ano06b]. Darzens [Las94]. Dastin [Jos49a, Man61, The91, The99, The86, Thi08]. Data [Per19]. Date [Bai00]. Dating [Dob82b, Sta96]. Daughters [Wal16]. David [Ano06b, Bri66, Bro19b, Dav87, PVPP17]. Davidson [Rea61]. Davis [Bur02, Ree20, Ree23]. Davisson [Kah21]. Davy [Col75, Dol98, Ful19, Ful75, Gol19, Jam15, Jam19a, Jam19b, JR19, Kni16b, Kni19, Lac19, Rus19, Tat19]. Day [Ano07e, Hay77]. Dayton [Her18]. DC [Ano97c]. dccxvii [Bur02]. Death [Ram00, War84]. Debate [BG00, Hir81, Mei83, RA02, TR99]. Debates [Hir21]. Deborah [Ano07e]. Debus [Clu07, Pri09]. Debye [Rei10]. Decade [Far89]. Decay [Sin87]. Deception [And16]. Deciphering [PLB23]. Decknamen [Lan22]. Declaratio [HS78]. Decline [Lan81]. Dedicated [Pri79]. Dee [Bay05, Abr94, App77, App79, Clu17a, Clu17b, Clu71, Clu05b, Fig65, For05, For17, Har00, Jos64, Luh86, Mer17, Nor05, Van17, Wal76]. Defense [Con96]. Definition [She85a]. del [Car17a, GL21]. Delayed [Wei18]. Dele [Rag12]. Deliberative [Gab16]. della [Car17a]. dell’alchimia [Hed17]. d’Émile [Chi17]. Democritus [Her87, Mar09]. Demonstrating [Fre10]. Denmark [Jac00]. Departing [Mor12].

Department [Ben70, Bro17c, Tya04]. Deployment [GL18]. Deposition [DM97]. Deregulation [RA02]. Descartes [Hei16]. Descendants [Wil63a, Wil66]. Desertion [Olt05]. Designer [HB84]. Despotism [Cro95]. Detecting [DeK09, Tra12b]. Detection [Jam81]. Detectives [Whi18]. Déterminer [Mor80]. Deutsche [Joh18]. deutschen [Don88, SK87]. Deutscher [Joh18]. Devastation [Tar03, Tar05, Tra05]. Developing [DeK09]. Development [Ben70, Bro87, Eli11, Gel71, Hom99, Jon93, Kin81, Kin82, Pul20, Sch94, Zot93, vR07]. Developments [MT11, Ram13]. Devilish [Sin59]. Devitalising [Sou00]. Dewar [Gol80]. Dexter [Ano83, Ano85b, Ano87b]. d’Hermès [Hed22]. Diagnôsis} [Via15]. Diagnosis [BP04, Ano06e]. Diagrammatic [Clu17b]. Diagrams [HS15]. Dialectical [Lar71]. Dialogic [Ful19]. Dictionary [Fer09, Mor09, Tak17]. Didier [Fer22]. Dieffenbach [Bro00a]. Dienel [Mor06]. Diet [dMC15]. Different [Kle95]. Digby [Dob71, Dob73, Dob74, Pri13]. Digital [Lan22, Mor22a]. Dii [Fig65]. Dijon [Sme61]. Dimension [Ano06a, dCH04]. Dining [Lac17]. Dionysian [Wil98]. Disadvantage [Pal10]. Disciplinary [Ram00, Sim98]. Discipline [Kni16c]. Disciplines [Nor15]. Discourse [Gab16]. Discover [Nic12]. Discovered [LRT14, Pri13]. Discovery [Eam80, Hir81, Jam85, Kar80, KP78, Pom62, Sce22, Wil12, Bro16b]. Discussion [Par37b]. Disease [Bro21]. diseases [FH05, Num07]. Disinfectants [Rae06]. Display [RC16]. Disposal [Tar02]. Disputation [Cha07]. Disputations [Mor21b]. Dispute [Ful75]. Disputes [All03, Tra05]. Disquisitio [Sme59]. Disrupted [Lef18a]. Dissemination [BJJK81]. Dissent [Mor96]. Dissociation [Bro69, Ric04]. Dissymmetry [Mas91]. distillandi [Taa14]. Distillation [BN80, Kra16b]. Distilling [Taa14]. Distinct [Cha17]. Distinction [Wee07]. Distinguishing [Tom09]. Diversity [And10]. Divine [Mar09]. Djerassi [Ano06b]. DM [Don88]. Dmitriev [Ano06d]. Dmitrii [Ano06d, Gor04, Sta98]. Döbereiner [Col76b]. Docimastic [SvOM93]. Doctoral [Lar67]. Doctrine [Mcg67, Edd01, Sco70]. Document [KCU81, Lef18b]. Documentary [Ful75]. Does [HS15]. Dogma [Kni86]. domination [Joh18]. Dominican [Tar18]. Donne [Hay77, Sad77]. Don’t [Gue15, Mer22]. Door [For08, Hed17]. Dordrecht [Ano07b, Bro02]. Doubt [Sme77]. Doubts [Mau70]. Down [Rae05]. Dr [Ano04a]. Dr. [Gib60a, Lev81, Tay49a]. Dragon [Dav87]. Dramatic [DP03, Ano06b]. Draw [Gor18]. Drawing [Bri66]. Dream [RW88]. Dreams [Ano97c, Lev21]. Dripping [Dav87]. Dritten [Joh18]. Drugs [Mur22, PT17, Sim22, Sim16b]. Drycleaning [Cam91]. DSM [vR07]. Dublin [Dun65, Tim06]. ducal [Bar19a]. Duke [Zub18]. Dumas [Kap69, Run20]. d’une [Mor80]. During [Ree15, Tho38b, KF70, Löw80, Sel07, Sla04]. Dutch [Rob06]. Duveen [GC80]. DVY [Jam19b]. DVY/2 [Jam19b]. Dye [Kir18, MT92]. Dyes [Cob19, Ree22, RT97, Tra95, Tra04]. Dyestuffs [Tra02]. Dynamical [Kni67].

E-mail [Ano17c]. E. [Kle95, Sme71]. Earl [Shi49]. Earliest [DN37]. Early [AF16, And16, Arm12, Bro11, Bro16a, Bur92, CT88, Cha07, Clu07, Dym08, Edd08, Ede93, Far71, Fer22, Fis18, For05, For13, Fri21a, Gia18, Gol16, Goo69, Gou82, Iri20, Jam19b, Kim20, Lac17, Lev70, LMN01, Mah76, McK38, Mer20, Mil70, Mos97, New20b, Nor07, Pri21b, Pul20, Ras18, Roo23, Sch19, Sch89, Sha05, TCPYN59, Zub18, All19a, And10, Ave86, BG00, Cam01, CGT18, Dol98, Dri16, Duf20, Dup23, Hed13a, HN59, Kah10, Kah11, Kee13, KR14, Maw16, New20c, PT17, Sie08, Sim22, Wal16, Zha23, Ano97c]. Early-Modern [Zub18]. Earth [BB04, Tra12a, Vau04, Bal02, Bro21, CB23, Deb61, Ree79, Tra02, Vrt23]. Earths [Kra21]. East [BN80]. Ebeling [Hal11]. Eben [Jon93]. ebook [BS22, CS19, Kim20]. École [Lan81]. Ecological [CS19]. Ecology [Sán16a]. Economic [Ree20]. Ed [Ano75c]. Eddy [Tay23]. Edelstein [Mor06, Ano05d, Ano07c, Ano07d]. Edinburgh [Gol16, Mor69, Per82]. Edited [Ano97c, Ano04a, Ano05c, Ano06c, Ano06a, Ano06e, Ano07a, Ano07b, Ano07e, Ano08a, BS22, Bro02, Bro20a, Bur02, Cha23, Clu07, CG22, Duf23, Fer22, HS23, Kim20, Kni04, New23, Num07, Rat92a, Roo23, Sha05, Tay23, Tra05, Van22, Pul19]. Editing [Ree23]. Edition [Jam19b, Num07, Fin91]. Editiones [Ano06e]. Editor [Ano65a, Ano67a, Ano68b, Ano94c, Cou77, Dic72, Fig73, Gib66, dVH58, Mor12, War76a]. Editorial [Ano62a, Bro03b, Clu05a, MT02, Mor04, Sut83]. Editors [Ano09a]. Eduard [Bro20a]. Education [Han76, Kop76, Sán16b, Nie16]. Educator [Gor83]. Edward [Bur02, Bro86a, Dol98, Rus96, Bro97c]. Effects [RS04]. Egede [Nor20]. Egg [She58]. ego [Ano06d]. Egypt [Wil98]. Egyptian [Mar17]. Eighteenth [DM97, Hei73, Hol89, Lef18a, Leh09, Pag14, RK06, Sud78, Tay23, Ave86, Ber92, Bor15, Chr94, Deb81, Di 22, Fil92, Gue15, Hen18, Hey38b, Kon13, Le 82, Leh10, Lei16, Mei83, Per13, Pow98, Pow14, Sal19, Sch89, Sco70, Ser13, Ser22, Sim06, Tei82, Tha68, Wer13]. Eighteenth-Century [Hei73, Lef18a, Pag14, Sud78, RK06, Gue15, Hen18, Kon13, Lei16, Mei83, Pow98, Sal19, Sco70, Ser13, Ser22, Tha68]. Eightness [PW69]. Eirenaeus [New90a, Wil64, Wil72]. Eleazar [Pri16a]. Electivis [Sme59]. Electricity [Fis18, Sud78]. Electro [Sud80].

Electro-Chemical [Sud80]. Electrolytes [Wol72]. electrons [Sim14b]. Elémens [BG00]. Élémentaire [Bai00]. Élémentaires [Mor80]. Elemental [Uss03]. Elemente [BB04, Ano05a]. Elements [BB04, Dun70, FCO15, Hij22, Kar80, Kle14, Kra16a, PW74, Per73, Pow14, Sce22, Sou00]. Elixir [Gia18, Mah76, Mat13]. Elixirs [Mah77]. Elizabeth [Pri79]. Elizabethan [Con22, Deb60, Sch73]. Ellen [Ano06e]. Embrace [Wei18]. Emergence [Cle21, GS16, Roc87, Sta98]. Emerging [Hom05, Tra14]. Emil [Moy89, Sal13]. Émile [Chi17]. Emma [Shm94]. Emmens [Kau83]. Emperor [PZ18]. Empire [CE17, Mor18a, Mus17, Tom17b, Jan20, Ros02]. Empiricism [Gor19]. Empiricus [Del20]. Enabling [Lyk22]. enactment [For21]. Encounter [Zub21]. Encyclopedia [Suc16]. Encyclopedic [Tak17]. End [Pow98, Bil13]. enero [LP05]. Engel [Ano04a]. Engineer [Ree20]. Engineering [GS16, BS22]. Engineers [Lyk22]. England [All19a, Deb60, Del20, Ell22, Gol83, Hed13a, Hen18, LR16, Poo09, Rat64, Roo01, Rus96, Sha94, Wil62, Bro97c, Las96]. English [Cop92, Vic94, Ave86, Bai00, BV16b, Con22, Gru06, Jon18, Jon21, Rob66, SM17, War76b, Web67a]. Englishing [Ram16]. Engraver [Fri22]. Enlightenment [Deb81, SL22, Hen17, LK12]. Enlightenments [Tom17b]. Enquiry [Ful19]. Entdeckung [Bro16b, Sal13]. Entre [GL21]. Entrepreneurial [Mor16a]. Entwicklung [Hom17b]. Environment [MB03, VP18, Tra05]. Environmental [DeK09, Jas14, RGG23, Tar03, Tar05, Tra05, Vrt23]. Episode [Fig65, Wel82]. Episteme [Fri22]. Epistemological [HJ02]. Epistemology [Vag19]. Epistolary [Kle15]. Epoch [Ano06d, Dmi04]. epokh [Ano06d]. Epokhi [Dmi04]. Equipment [Ede93, HN59]. Equivalence [Mau70]. Era [SL22, Maw16]. Erasmus [Rue22]. Erbe [Lau19]. Erde [Ano05a, BB04]. Erdmann [Bro16c, Bro16c]. Eric [Bro02, Pul19, McK60]. Erika [Van22]. Erinnerungen [SK87, Don88]. Erker [Mor06]. Ernest [McK63]. Ernst [Bro00a]. Erratum [Mor02, New15]. Error [Sta96]. Errors [Kar80]. ersten [Bro16b]. Eschatology [Lin84]. Esoteric [Mer90]. Esotericism [Kah19]. Essai [Mor80, Per84, Sme77, Sme78]. Essay [Ano75c, Ano97c, Ano06d, Ano06e, Ano07e, Bro02, Bro97c, Bur02, Cra93, Dob82a, Don88, Hal11, Hom95, Jam19b, Kle04, Las96, Mor06, Pat92, Rat92a, Rat96, Rus98, Sha05, Sme89a, Tra05, Vic94, Deb80]. Essays [McK58]. Establishing [Ram08]. Establishment [Jam83a, O’N21, SB92]. etc [Dob82a, Rat92a]. Ethanol [BN80]. Ether [Gor98, Pri86, Sin87]. Etherification [Pri86]. Ethics [BS22]. Etymology [HS20]. EUR [Ano04a, Ano05a, Ano07b, Bro20a, Lau19, Mor06, Num07, Sha05]. Europe [Ano05c, Ber11, CGT18, Dri16, SG23, Kah10, Kah11, KR14, Num13, Ost02a, Ost02b, PT17, Tom17b, Wil09, Zha23, Sch19]. European [Ano82, Pay89]. Evan [Iac22]. Everyday [All19a, Pri21b, Sch22, Mor18a]. Evidence [Dob82b, Ful75, New86]. Evolution [And06, Sin85]. Evolving [Amb22]. Examination [Far75, Old73]. Example [MT16, SD87, Tra14]. Examples [Per03, Per04, Pri87]. Excavated [SvOM93]. executioner [GL21]. Exemplar [Tya04]. Exhalation [Nor06]. exhibition [LP05]. Exhumed [Mau70]. Existence [Hir81]. Exotica [Pri16c]. Experience [Dup23]. Experiences [Ban07]. Experiment [FPS16, Jos63, Mur22, New20b, Pas20, Sne79]. Experimental [BGH78, BN80, Dob74, Dol98, Jon21, Kni16b, Web67b]. Experimentalisierung [Car16b]. Experimentalization [Car16b]. Experimentation [MT16, Wil98, Run20]. Experiments [Bur99, Jam19b, Kin81, Kin82, Leh10, Shi49, Tre74, VMT22]. Expert [Wat95]. Expertise [BV17]. Experts [All03, Kle13, Tra05]. Explanation [Sco70]. Explanations [Fra11]. Explorations [Ano75c]. Exploring [Bro11, Fur21, Mor22a, Zub18]. exposición [LP05]. Expressed [Cri73]. Exquisite [WW07]. Extraordinary [Gol17]. Eymerich [Tar18].

F [Ano97c, Kle95, Sme71]. F. [Sme59]. F.-J. [Sme59]. F.R.S. [Mac88]. Face [Ano06e, BCS03]. Facem [Abr79]. Facticity [Fre10]. Factor [Ree77b]. Factory [And17, FFS77]. Facts [Fre10]. Fahrenheit [Gol57]. Fairy [NG16]. Fake [Gan20b, Mor18b]. Fakes [And16]. Fall [BV16b, Car16b, Mor11]. False [Abr84]. Fama [Luh86]. Familiar [Kee13]. Famine [Bro21]. Famous [Ano07e, HH00, HH05, Zot93]. Faraday [Mos15, Wil12]. farmacia [Bar19a]. Farrar [Vau04]. Fasciculus [Abr94]. Fate [Lev81, Mau02, Ree77a]. Father [Cam01]. Fatherland [Sch89]. Faveur [Nev62]. Fear [Cob22]. Features [Las03]. Feijoo [Lei16]. Felice [Wel82]. Feliciano [Wel82]. Fellow [WW07]. Female [Mer22, Ser13]. Feminist [Tos01]. Fen [Big17]. Fergus [BJ21]. Fernel [Num07, FH05]. Ferrara [Cri73]. Ferrous [Cra10]. Fertilisers [Eli11]. Festschrift [Bro20a]. Feuer [BB04, Ano05a]. ffor [Gru06]. Fictional [Gor19]. Field [Cha21]. Fields [Fri21b]. Fifteenth [Szu86]. Fifty [Sme87a, Sme87b]. Fight [Cob22]. Figuiers [Hun79]. Figures [Tay49b]. Find [Gue15]. Finney [Bro08]. Fire [Amb22, Jen81, Jon21, Kno01, Lov72, Pow14, Rap18, War76b, BB04]. Firenze [Bar19a]. Fireworks [SG23]. First [Bro69, Bro11, Gib60a, Kah21, Qui20, Rea61, SC90, Sme87a, Sme87b, Tho38b, Vin09, Bro16b, VG04]. fiscal [GL21]. Fischer [Moy89]. Fitness [Gol19]. Fits [Bee94].

Five [Bro11, McK58]. Fixation [Far71]. fixe [Fra17]. Fixed [Le 73, Fra17]. fl [App87, Poo09]. Flame [Kni76]. Flamel [Pri16a]. Flames [DeK80]. Flanders [Mal08]. Flexibility [Nor15]. Florence [Bar19a, Hom19, Bar19a, Tho56]. Florian [Hal11]. Flos [Lon07]. Flour [PG21]. Fludd [Bur99, Deb67, Fri22, HS78, Jos49b, Jos63]. Fluid [Bud20]. Fluorescence [DeK09]. Flürscheim [Sho03]. fols [Fer09]. Fontani [Kra16a]. Food [Cob20, Orl16, Cob22, Ree22]. Foods [dMC15]. Fool [Roo04]. Force [Bro68, McG68, Sou00]. Forced [War76b]. Fordyce [Tay14a, Tay14b]. Forensic [Bur16]. Forest [Sán16a]. foreword [Hal11]. Forgeries [KH20]. Forgery [Fer22]. Forgotten [Deb63, Hor18, HB84, Sha03]. Formation [Deb61, Pri86]. Forms [FM07, FH05, Num07]. Formula [Gol80, HS15, Pla82]. Formulae [Uss03]. Forrester [Num07]. forskning [Nie16]. Fortune [Zub18]. Fought [Bro21]. Foundation [SC90, GW17, Pri06]. Foundations [Bro02]. Founding [Pat82]. Founts [Mar23]. Fourcroy [CS79, Gue76]. Fragments [Mar17]. France [BBS10, BSGBBV02, BG00, Deb81, PG21, Le 15, Leh09, Leh10, Lla85, Mal08, Per03, Per04, Pow98, Riv18, Roc94, Sán21]. Francis [Gre38, Mer79, Ree75a, Ree75b, Ree79, Wee07, Wes61]. Franciscus [Rag12]. Francois [Leh09, HS18, Leh09, Olt05]. Frank [Hol56, Sim87]. Frankland [Bro97c, Rus96]. Franziska [Bro20a]. Frater [Jon18]. Fraternitatis [Luh86]. Fraud [Zub21]. Fray [Bro17d]. Frederic [Ano07b]. free [Kim20]. French [Vag19, BS14, Chi17, Cro95, Fau11, Fer17, Fic71, Fra17, Kra16b, Pio18, Sme89c, SMGL18]. French-speaking [Vag19]. Friar [Sad75, Suc16, Sad75]. Friars [Jon18]. Friedrich [Sou00, Bro16b]. Friend [Har00, Pra91]. Friendship [Bro22, Jam22].

Frontispiece [For17]. Fryer [Gor77]. Fuit [Del20]. Functions [Dun70]. Furnace [WW07]. Furnaces [HVD20]. Further [App77, Ful75, Wil72]. Future [Ram17].

G [Clu07, Kin84, Pri09]. G. [Old73]. Gabriel [DeK80, Deb61, Geo62, Leh09]. Gabriel-Francois [Leh09]. Galambos [Ano06e]. Galan [CG22]. Galen [Tho63b]. Galenic [MB21, Mor21b]. Galenism [Rag12]. Galenist [Rat64]. Galileo [She70a]. Galvanism [Fis18]. Gandharan [BN80]. Gas [BS22, FFS75, Pag62, Pal10, Rae16]. Gaseindustrie [Hom17b]. Gases [Col78, Hom17b]. Gatekeeper [Rus98]. Gay [Cro81]. Gay-Lussac [Cro81]. Geber [Ple69]. Gegenwart [Bro20a]. Geiger [Tho88]. Gems [Iri20]. Gender [SL22, Ser22]. Gendered [Sim16b]. Genealogy [Cha21]. General [App87, dV22]. Genese [Orl16]. Genesis [Nor07, Orl16]. Genetic [GS16]. Geneva [HS15]. Geneviève [Run20]. Genth [Bro16b, Bro16b]. Gentleman [Clu93]. Geoffroy [Kle95, Sme71]. Geomancy [KF73]. Geometrical [Wal76]. Georg [Mor96]. George [Bro20c, Bra79, DeK80, Ram08, Ram10, Ree20, Ree23, Tay14a, Tay14b, Wil63b, Wil73]. Gerard [Ano72b]. German [BR00, BB04, Bro16b, Bro16c, Bro19c, Car16b, Die04a, Fig73, FM07, Hag16, Han76, Hom99, Joh18, Kle14, Lau19, Löw80, Orl16, Poo09, Sch89, SK87, Wei18, Zub21]. German-Speaking [Fig73]. Germany [Bro02, Ben72, Bro22, Car16a, Fre10, Wal16]. Gernot [Ano05a]. Gerrylynn [Ano07a]. Geschichte [Ano04a, Die04a, LE03, Mor06]. geschiedenis [Kno22]. Gesellschaft [Bro20a, Joh18]. Gestation [Pay89]. gewonnen [Hom17b]. Ghost [Wuj84]. Giessen [Bro20a, Bro03a, Roc03]. Gift [Orl16]. Giovanni [Bar19a]. Girl [Hor18].

Giroux [Vau04]. gjennom [Nie16]. Glances [Sad77]. Glands [Kno06]. Glass [HVD20, MC06]. Gld [Rat92a]. Glitters [Kar07]. Global [Hom05, Iri20, Joh11]. globalen [Hom17b]. Globalising [Tra17]. Globalization [NG20]. Glosses [Pol20]. Glutenophilia} [PG21]. Gnomon [Sta57]. Gnostic [PW69, Pag60]. Gnosticism [She57]. Gnostics [Wel51]. God [Kno01]. Godfrey [Zub21]. Gold [Bil23, mKC21, Gia18, Kar07, Kra22a, Mac88, Mer90, Oak23, Roo04, RGG23, Vrt23]. Gold-Making [mKC21, Mer90]. Golden [Jos49b]. Good [Gue15, O’N21]. Gordin [Ano06d]. Gordon [Mor16b]. Gorman [Tho22]. Got [Bud18]. Göttingen [BR00, Ano04a]. Göttling [Sme66]. Gottlob [Kir03]. Governance [Gab16, Gol18, LE03, Mor06]. Government [Jam21, Moy89]. Grading [Clu71]. Graduate [Kop76]. Graeco [Mar17]. Graham [Kik09]. grand [Bar19a]. grand-ducal [Bar19a]. granducale [Bar19a]. Grants [Ano22b]. Grapevine [Las10]. Graph [HS18]. Gravius [Poo09]. Great [Gol17, Ree75b, PG21, RK06]. Greece [Hor60]. Greek [Cop92, Vic94, Bro48, Duf15, Duf20, Her87, Mah76, Mah77, Mer20, Pla82, Tay37b, Wel51]. Green [Wer13]. Greene [Sad75]. Greenland [Nor20]. Grégoire [Hed22]. Gregory [BS90]. Greville [Jam81]. Grew [Roo07]. Grids [Gor18]. Griffin [GB91]. Grocers [Sel07]. Groningen [Rat92a, vM90]. Ground [Low18, RA02]. Grounds [Gor21]. Group [Maw18]. Grove [Owe01]. Growing [Bro05, Rae05]. Growth [Kni16c, Mår19]. Grüneburg [Ano12a]. Guide [Ano06e, RST03]. Guillaume [Leh09]. Guillem [CG22]. Guillem-Llobat [CG22]. Guillermo [Pul19]. Guinea [Sán16a]. Gunpowder [Sin59]. Guns [Sin59]. Günther [Ano04a]. Gurvitsch [SD87]. Guy [Col88]. Guyton [Sme57, Sme61, Sme63, Sme66, Sme68, Sme87c].

H [Ano05a, Kau83, KCU81, Mor06]. Habsburg [Kon13, Ros02]. Hachez [Hom19]. Hachez-Leroy [Hom19]. Hair [Sou92]. Hales [Bax13]. Half [Tho38b, Vag19]. Hall [Sim14a]. Handedness [Mas91]. Hands [Bro17b, Štr16]. Hans [Ano07b, Mor06, PZ18, Nor20]. Hans-Jörg [Ano07b]. Hans-Liudger [Mor06]. Hanson [Jam81]. Harcourt [Kin84]. hardback [Ano97c]. hardcover [BS22, CS19, Cob22]. Hare [Fis18]. Hargittai [Ano07e]. Hariot [Sea56]. Harrassowitz [Kim20]. Harris [Bow75]. Harrow [Jos49b]. Harteck [Far89]. Hartlib [Clu93, Wil68, Wil70a]. Hartmut [Ano05a]. Harvey [Cob22]. Hathor [Mar21]. Hauch [Jac00]. Haüy [Mau70]. Haven [Car17a]. Hawliczek [Dic72]. Hayyan [Ple69]. hbk [Ano06b, Ano06e, Ano07e, Hal11, Mor06]. Healers [Wal16]. Healing [Sim22, SM17]. Health [Di 22, Gol19]. Heat [Don78, Fis18, Hei16, Jam19b]. Hehrenrettung [Fri59]. Heidelberg [Naw14]. Heilbron [Kni04]. Heinrich [Bro22, Tra91]. held [LP05]. Hellenic [Zac56]. Hellenistic [BN80, She59, Wil98]. Helmholtz [Jam95]. Helmont [Duc08, Hed21, Hei95, Jen04, Mer79, Pag62, Roo17a, Wal80]. Helmontian [Clu11, Deb72, Deb80, Rat64]. Helmstedt [Fri21b]. Help [MT16]. Henk [Kno22]. Henriques [Fil92]. Henry [Jew19, Num07, FFS73, FFS74a, FFS74b, FFS75, FFS76, FFS77, Geo57a, Gol17, Gui80, McK63, Ric04, Web67b]. Henrys [FFS73, FFS74a, FFS74b, FFS75, FFS76, FFS77]. Heretic [Pat88]. Heritage [Ano06e, Pri06]. Herman [Ano97c, Kno01, Lov72, Pow14]. Hermann [Chr94]. Hermeneutics [Mer90]. Hermes [Hal11, Ebe07, SLT49, Van09, vL38].

Hermetic [For17, Kel14, Mar21, Mor85, She60]. Hermetica [Cop92, Ano84a, vL38]. Hermeticae [PLB23]. Hermeticism [AGJ08, Ebe07, Hal11]. Hermeticum [Vic94, Cop92, KF73]. Hermitica. [Vic94]. Herrschaftsapparat [Joh18]. Herzog [Mor22a]. Hess [Van22]. Hessen [Mor85]. Hessen-Kassel [Mor85]. Heterogeneous [Col75, Col76b]. Hexagon [Mor06]. Heydon [Kel12]. Heym [Ano72b]. Hickrill [GW17]. hidden [FH05, Num07]. Hieroglyphic [For17, Nor05]. Hieroglyphica [Clu17b, Clu05b, For05, Jos64, Luh86, Mer17, Van17, Wal76]. Hieronymus [Taa14]. High [New14, New15]. Higher [PW74]. Hill [CS19]. Hills [Ful19]. Hilversum [Kno22]. Him [Bax13]. Hirai [Fer22]. Hiro [Fer22]. Histoires [Tom18]. Historians [FPS16]. Historic [Jan20]. Histórica [LP05]. Historical [Ano06e, Bro02, DN37, GL15, HJ02, KCU81, RST03, Sch06, Tar02, Car17a, LP05, Sim16b]. historically [GF23]. históricamente [GF23]. Histories [Gol18, Kra22b, Oak23, RGG23, Vag19, Wat17, Wey76]. Historiography [Ano97c, Ber92, Ber11, Chr94, Cre17, MT11, Vag19]. History [And11, And17, Ano82, Ano84c, Ano06c, Ano06e, Ano07a, Ano07b, Ano08a, Ano09a, Ano12a, Ano16d, Ano17c, Ano17b, Ano19a, Ano19c, Ano21b, Ano22c, Ano23b, BV16a, BV17, BB04, Bre76, Bro17c, Bro19a, BGH78, Bur16, BP04, Cam91, Cha17, Cha23, CS19, Col96, CB23, Deb71, Deb85, Die04a, Die04b, Duf23, Ebe07, Far20, Fig65, Fur21, Hal11, Ham17, HBB^{+}03, Hen20, HS23, Hij21, HRR03, Jas14, Kah13, Kar07, Kir23, Kni04, Kni17, Las96, Lla85, Mar75, McV75, Mei95, Mer04, Mor06, Mor11, Mor18a, Mor18b, New23, NG15, Nye02, Obr09, Pag14, Pas20, Pér16, Ple72, Pol15, Qui16a, Ram17, Rap18, Ras18, Rat66, Rat92a, Roc94, Roo23, Rus37, RR05, Sán16b, Sce22, Sha94, Söd07, Taa23, Tak17, Tar03, Tar05, Tay23, Tom16a]. History [Tom17a, Tra05, Ver22, Whi18, Zub21, vM90, Ano97c, Pri06]. Hitherto [Mah76, McK58]. Hoaxes [And16]. Hofmann [Ben70, Ben72]. Holding [Štr16, Bro17b]. Holmes [Ano07b]. Holmyard [McK60]. Holy [Jan20]. Homagius [Mor96]. Homberg [Byc21, Str79]. Homeri [Hey37a]. Homunculus [Mur20]. Honourable [Sim20]. Hooke [Lys37]. Hoosier [Min18]. Hope [Nor20]. Hopkins [Cob22, Han76]. Hopwood [Ano06a]. Horsford [Jon93]. Horst [Ano04a]. Horus [dS22]. Hospital [Col88]. House [MT12]. Household [All19a]. Houtou [Olt05]. HPO [vR07]. Hüls [LE03, Mor06]. Human [Ano06e, BCS03, Ver22]. Humanism [Edd08, Wel82]. Humanities [For21, PHW22]. Humboldt [Hag16]. Humboldts [Hag16]. Humphry [Col75, Dol98, Gol19, Jam15, Jam19a, Jam19b, JR19, Kni16b, Rus19, Tat19]. Hundred [GL18, Roq17]. Hunger [Bur02]. Hunter [Bur02, BJ21]. Hunyades [App77, TJ53, TJ56]. Hutton [Don78]. Hyacinth [Mal08]. Hyde [Kni16a, Uss15]. Hydrocarbon [Tra12b]. Hydrocarbons [FFS75]. Hydrolysis [Kau73]. Hyle [Bro02]. Hyle. [Bro02]. Hypothesis [Bro69, She70a].

Iatrochemist [App87, Gel94]. Iberia [Lei16]. Ibn [SLT49, Ple69, SLT49]. Icaria [CG22]. ICHC13 [Ano22b]. Iconography [Hil75]. Idea [vF65]. Ideal [Rae16, Van17]. Ideas [Cis08, Pag69, PW69]. Identification [Hij22]. Identifications [Sud78]. identités [Chi17]. identities [Chi17]. Identity [KLPG06, Lon07, Ram00, Wil64, Wil72]. If [Gue15]. Ignorance [Pal10]. ignorancia [CG22]. Igor [Ano06d]. Iguanas [Dav87]. II [Bri66, Dob73, Fis73b, Hor11, Kah11, McE78b, PZ18, Sin88b, Sla04, Sme87b, Tay14b, Wil70a]. III [Dob74, McE78c]. Il’ich [GLCK76]. Ille [Del20]. Illumination [Abr79]. Illuminators [Cla01, Ano05b]. illus [Ano97c, Ano04a, Ano05a, Ano05c, Ano06c, Ano06a, Ano06d, Ano06e, Ano07b, Ano07e, Ano08a, Bar19a, Bar19b, Bro20a, Bro97c, Cha23, CS19, Cob22, CG22, Duf23, Fer22, Hal11, HS23, Her18, Hom19, Iac22, Kim20, Kni04, Kno22, Lau19, Mor06, Mos22, New23, Num07, Pul19, Rag22, Ree22, Roo23, Run20, Tay23, Tra05, Van22, Vau04]. Illustrated [Kni17]. illustrations [Dob82a]. Il’ya [GLCK76]. im [Bro20a, Hag16, Joh18, Lau19]. Image [Jam19a, Wey76]. Imagery [Hay77]. Images [Szu86, Hal09]. Imagination [Run18]. Imagined [Tom17b]. imaginibus [Mur20]. Imbalanced [Sch22]. Imitations [Kar07]. Immutability [Mcg67]. Impact [Car16a, Kir23, Ram18, RS04]. Imperial [Ano06b, Ano07e, Bro17c, dMC15, Cro95, Qui20]. Importance [Wes61]. Impurities [Pri87]. Inauguration [Tra17]. Including [App79, Sco70]. Incomparable [Bou22]. Independent [Hor18]. Index [Ano38a, Ano38d, Ano77a, Ano04a, Ano05a, Ano05c, Ano06c, Ano06a, Ano06e, Ano07a, Ano07e, Ano08a, Bar19a, Bar19b, BS22, Bur02, Cha23, CS19, Cob22, CG22, Duf23, Fer22, Hal11, Hed22, HS23, Her18, Iac22, Kni04, Kno22, Lau19, Mor06, Mos22, New23, Num07, Pul19, Rag22, Ree22, Roo23, Run20, Sha05, Tay23, Tho22, Tra05, Vag19, Van22, Vau04, All19b, Wes75]. India [Gor83, Hor60, Wat17]. Indian [BN80]. Indications [Geo57b]. Indicators [Deb62b, Eam80]. industri [Lyk18, Nie16]. industria [Bro17d]. Industrial [Hom17b, Tar02, Tra14, War76a, vR07, CB23]. Industriel [Hom19]. Industries [Qui20]. Industry [Ano06e, Bro17d, Bro05, Cec04, FFS75, Ham17, Hom17b, Jon93, Le 15, Mår19, Mer22, Mer04, MT92, Mor18a, Moy89, Pui04, Ree15, RS04, RST03, Sta98, Tra02, Tra17, Tya04, VP18, Lyk18, Nie16]. Inexactitude [Nor15]. infancy [Jam15]. Inflammable [Tom09]. Influence [Büt00, Gel78, Hed21, Hen18, Mer79, Szy96, Tay08]. Influences [Old74, War76b]. Information [App77, Gom77, Hem15, Las10, Roq16]. Infra [CT88]. Infra-Red [CT88]. Ingenious [Kau83]. Ingenuity [Zha23]. Ingold [Sch94]. Inhale [ML08]. Initiatives [Eli11]. Innovation [SG23, Leh10, Mer20, Pui04]. Inorganic [Naw14, Sin85]. Inquiry [Bra79]. Inquisition [Tar18]. Insecticide [Cec04]. Insights [Ras18]. Inspectorate [Ree12]. Inspiring [Ano07e, Chu14, Chu13]. Instauration [Ree75b]. Institut [Bro16b]. Institute [Bro02, Lyk20, Lyk22, Bro16b, CA15]. Institutet [Nor15]. Institution [Jam19b, CC80, Jam15]. Institutional [Sel07]. Instruction [Mil70]. Instructor [GB91]. Instrumentation [DeK09]. Instruments [RC16]. Integrated [Gay12]. Integrating [Per19]. Intellectual [Gel71, Hem15, Roq16, Sim87]. Intelligence [Kni16a, Uss15]. Interaction [Fri22]. Interchange [Tya04]. Interdisciplinary [Bro02]. interest [Tho63b]. Interests [Clu93]. International [Bro02, Car16b, Fau11, Hom17b, Kik11, Mår19, Qui16b, Rat92a, vM90]. internationale [Car16b]. Interpretation [Hed13b, Luh86, Mar75, Nor05, Sie82]. Interpreter [Fis96, New82]. Intertwined [Ham82]. Intervention [RS04]. Interwar [Tra17]. Introduction [Bur99, Cop92, Deb63, Fri21a, Hey37b, Jac00, JR19, Jos64, KH20, KR14, Mon38, QR20, RT97, Tom16a, Don88, Vic94]. Introductory [Kni86]. Inventing [Jon21, MB03, Tra05]. Invention [Bal02, FM07, Vau04, Mos22]. Invisible [Dan06, McG68]. Invisibles [CG22]. Iodine [Ful75]. Ira [Han76, War08]. Iron [Wil09, Pag14]. Isaac [Dob82a, Tay56, All19b, Wes75, Wes80, Roo18]. ISBN [Ano97c, Ano04a, Ano05b, Ano05a, Ano05c, Ano06c, Ano06a, Ano06b, Ano06d, Ano06e, Ano07a, Ano07b, Ano07e, Ano08a, Bar19a, BS22, Bro97c, Bur02, Cha23, CS19, Cob22, CG22, Dob82a, Don88, Duf23, Fer22, Hal11, Hed22, HS23, Her18, Iac22, Kim20, Kni04, Kno22, Kra16a, Las96, Lau19, Mor06, Mos22, New23, Num07, Pul19, Rag22, Rat92a, Ree22, Roo23, Run20, Sha05, Tay23, Tho22, Tra05, Vag19, Van22, Vau04, Vic94]. Iseo [Suc16]. Isis [dS22]. Islamic [Ham82, Ryd94]. Isolated [Ban07, Far75]. Isolation [Jam81]. Isomerism [HJ02]. Isomorphismus [Sal13]. ISSN [Bro02, Bro20a, Hom19]. Issue [Hom19]. István [Ano07e]. Italian [Bar19a, Car17a, Hed17, KM90]. Italy [Deb63, Gue15, Kah21, Wel82]. Ithaca [Hal11]. Itinerant [Gib60b]. IV [Rea38, Mce79, Mar04]. Ivanovich [Fig65, Sta98]. Ivory [Tay49b]. ix [Ano06e, Vau04].

J [Kni04, Rat92a, Rob06, Sán16c, Sin87, Sin88a, Sin88b]. J. [Per84, Sin87, Sin88a, Sin88b, Sme59]. Jabir [Ple69]. Jabirian [Wil88]. Jacques [BBS10, Sho03]. Jahre [Bro20a]. Jahrhunderts [Orl16]. James [Mil70, Arm12, Don78, HS78, Iri18, Kni74, ML08, Rea38, Tom16b]. Jan [Hal11, Hed21, Roo17a]. Jane [Gol19, Kni86]. Janos [Gom77]. January [LP05]. Japan [Fur21, Roc22]. Japanese [Kik11]. Jawbari [Abr84]. Jean [FH05, Kah21, McK58, Num07]. Jeffrey [Ano97c]. Jena [Pri21a]. Jersey [Ano97c]. Jerusalem [Mor06]. Jesuits [Cam01]. Jesus [Bal93]. Jewess [Pat82]. Jiten [Tak17]. Jo [Ano06c]. Joachim [BS22, Bro02, KN14, Pag69]. Joan [Jen04]. Joaquim [Fil92]. Joaquín [Num07]. Joel [Tra05]. Jofuku [DN37]. Johann [Bor15, Col76b, Fri59, Han67, Hei95, Sou00]. Johannes [Cam01, Don88, App77, TJ53, TJ56]. John [Edd01, Jos49a, Kel12, McK60, Num07, Sha05, Bay05, Bil13, Böh63, Bow75, Bro20b, Clu17a, Clu17b, Clu71, Clu05b, Deb62a, FFS74b, For05, For17, GB91, Gor77, Gro21, GS89, Har00, Hay77, Jos49a, Jos64, Lac19, Mal08, Man61, Mer17, Nev60, Nor05, Pag62, Pra91, Pra92, Sou00, The91, The99, The86, Thi08, Tra95, Wal76, Wil63a, Wil66]. Johns [Cob22, Han76]. Johnson [Gib60a]. Joint [Ano09a]. Jonas [DeC15, Bág16]. Jonathan [Cob22]. Jonson [Lin77]. Jordan [Qui20]. Jörg [Ano07b]. Joris [Mos22]. Josef [Dic72]. Joseph [GB91, Olt05, Con96, Don78, Kah13, McE78a, McE78b, McE78c, Mce79, Mck61a, Per82, Sch67]. Joshi [DN37]. Joshua [Amb22]. Journal [Bro02, Hey38b, CS79, Ree23]. Journals [Ano89c, Ano09a]. Journey [Ano06e, BP04]. Joyce [Tra05]. Jr. [Wil63a, Wil66]. Jubiläum [Bro20a]. Julie [Ano06e]. Julius [Ant17]. Jungian [vF65]. Jungius [Pag69]. Jürgen [Ano07b]. Just [Ano97c, HS15]. Justification [Tre74]. Justus [Bro16c, Bro20a, Bro16c, Büt00, Fin91, Jon93, Mun90, Tho88, Zot93].

Kagakushi [Tak17]. Kahn [Fer22].

Kaleidoscopic [Sla04]. Karl [Kir03, Kea92]. Karlsruhe [Bro02]. Karolinska [Nor15]. Kassel [Mor85]. Kasteleyn [Rob06]. Kastner [Kir03]. Kazan [Bro95]. Keill [GS89]. Kekule [WR83, Fis73a, Fis73b, Fis74, RW88, WR83]. kemisk [Lyk18]. Kenelm [Dob71, Dob73, Dob74, Pri13]. Kevin [Hal11]. Key [Jos63, Pol20]. Khalid [Sal17]. Khimizdat [Ano06d]. Khwarizmi [Ryd94]. Killed [Bro21]. Kind [Ric09, Ros96]. Kinetics [Kin81, Kin82]. King [Jew19]. Kirchhoff [Jam95, Rae97]. Kirwan [Mau02, Sme68, Tay08]. Kjemiutdanning [Nie16]. Klagenfurt [Tho63a]. Klein [Tay23]. Kleinert [Don88]. Klickstein [GC80]. Kluwer [Ano07b, Bro02]. Knight [Bro19b, HB84]. Knowledge [AGCF^{+}15, All19a, BV17, Dup23, Iri20, Jam19b, KF70, Kno06, Min16, Roo17a, Ser76, Sim22, Sim16a, Van17, dS22, Kim20, Hen19]. Known [Pag14]. Kokshan [VP18]. Kolbe [Roc87]. Köln [Sha05]. kommentierte [Bro16c]. Kommunikationen [Car16b]. Kopp [Roc21]. Koreas [Kim05]. Kristalle [Bro16b]. Kues [Lau19]. Kulturgeschichte [Ano05a, BB04]. Kunskap [Lyk18].

L [Ano07e, Kni04, Tra05]. L. [Le 75, Sme57]. Lab [Nor15]. Labillardière [Olt05]. Laboratories [Bro17a, CC80, MTKT^{+}15, Sán21, SC90, Sim14a, Sme66, Wer13]. Laboratorium [Bro20a, Bro20a]. Laboratory [And17, Ano06e, Bil13, Bro72, BP04, Ede93, FPS16, Gue15, Hil75, HN59, Mil70, New86, Nye14, Pil38, Sal19, Sim06, SvOM93, VMT22, MT12, Naw14, Ree15, Tra95]. Lacaze [Hed22]. Lady [Bay05, Bou22, SL22]. Language [BV16b, Dun81, Kar90b, Lan22]. Lapidary [KF70]. Lapidibus [Tho60a]. Laplace [Ban07]. Lapworth [Sch95]. Large [Ano97c, Mar93]. Last [Kar90a, Wil62]. Late [DeK09, Fil92, Kra89, LMN01, Mer20, Sch89, Tra14, Ser13, CGT18, Duf20, Gel94, Gui80, Rin17, MR20, Sha05]. Latin [Cop92, Kib60, Vic94, BV16b, Bur92, Del21, KF70, Sta70, Tho63b, Vin09]. Laureate [BR00]. Laurent [Lev21]. Lavoisier [Cro95, Ant22, Bai00, Ban07, Cro09, Dun70, Fra17, GC80, Gou83, Jen81, Kle05, Koh75, Le 75, McD66, Mck61b, Per69, Per73, Per82, Per89, Sie82, Sie89, Sme89b, Sme89c]. Law [Mor16b, Pom62]. Lawrence [Ano07b, Bur02]. Lay [BV17]. Layers [Low18]. Lead [DeK09, Sán21]. Leader [GB91]. Learned [Kle14, Mer03]. Learning [Mor21b, Taa14]. Lebenserinnerungen [Ano04a, BR00]. lebt [Bro20a]. Lecturers [Gib60b]. Lectures [Bro97d, Fre10, Mil68]. Leeds [WK16]. Legacy [Jew19, Lau19, RGG23, Tra02, Zub18]. Legend [Kir61]. Legends [Pri16a]. Leibniz [Mer79, Roo17b]. Leibniz-Stahl [Roo17b]. Leicester [Ano97c]. Leiden [Fer22, Hed22, Iac22, Num07, Rat92a, Sha05, Vag19]. LEMCO [Bro17d]. Lemery [Pow98]. Leo [Bar19a, Mos22]. Leonhard [Mor22b]. Leroy [Hom19]. Lesley [Ano06e]. Let [Tom09]. Lethal [Mor18b]. Letter [Bro86a, Dic72, Fig73, Pra10, The86, Tre79, War76a, Jos49a, dS22]. Lettere [Sán17]. Letters [Ano65a, Ano67a, Ano68b, App79, Cou77, DFB22, Gib66, Gol19, dVH58, Jam81, LRT14, Nev60, Pol15, PZ18, Thi08, Bro16c]. Levene [Hun99]. Lewis [Pag14]. Lexicography [Con22]. Liaison [Fis96]. Liaison-Officer [Fis96]. Libavius [dV22]. Liber [Mur20, Suc16, Taa14]. Liberation [Sou00]. Librarianship [Ros96]. Library [Bil13, Cam77, Car17a, Cro81, FPS16, Jam19b, LP05, Sme87c, Wes84, Wil63a, Wil66, Jos49b, Pri06]. Libri [Big17]. Libro [Abr79]. libros [LP05, Num07]. Licence [Geo57a]. Liebig [Bro16c, Bro20a, Bro72, BS90, Bro16c, Büt00, Fin91, Jon93, Kir03, Mor72, Mun90, Mun91, Oga08, SB92, Son77, Tho88, Uss03, Zot93]. Liebig-Gesellschaft [Bro20a]. Liebig-Laboratorium [Bro20a]. Liebigs [Bro20a]. Liebigs-Museum [Bro20a]. Life [Ano07e, Ant17, Bou22, Bro17b, Bud18, Chu14, Chu13, DeC15, Eam21, Fil92, Gab16, Gom77, Hed13b, Her16, Jew19, Kin84, Kni16a, Kra21, Mor16b, Pri16b, Pri21b, Roc21, Shm94, Sim20, Štr16, Uss15, Bág16]. Ligand [BJJK81]. Light [Eam80, Far20, Gan20b, Hom95, Lac19, LR16, Lin77, Pri16d, vF65, Jam19b].

Lightfoot [Tra95]. Lights [Gru09]. Like [Mer22]. Lima [New20c]. Limits [Kim91, Kle05]. Linda [Her18]. Linde [Die04a, Mor06, Die04a, Die04b, Mor06]. Lindsay [PVPP17, PVPP17]. Line [Jam85]. Linear [War08]. Linguistic [Maw18]. Link [BS22, She85a]. Links [Bro05]. Linné [Bro16c]. Liquid [Wil09]. Lisbon [CA15, LK12]. List [Dun70]. Listening [AGJ08]. Lister [Roo04]. Literary [War97]. Literature [Bar19b, Füc51, MT16, Ple72, RR05, Sch06, Ano07a]. Literaturpolitik [Kea92]. Little [Pag14]. Liudger [Mor06]. Lived [Dup23]. Liverpool [Mor16a]. Liversidge [Mac88]. Lives [Bro20e, Jam22, Min18, Roo18]. Living [BV16a, Bil23, Mar23, Qui16a]. Livre [Duv48]. livres [Pio18]. Llobat [CG22]. Lloyd [Bro20b]. Local [Car16a]. Lockyer [Bro69]. locus [Zin84]. Lomonosov [Lei75, Pom62]. London [Ano75c, Ano97c, Ano05b, Ano06b, Ano07b, Ano07e, Bur02, Cha23, Dob82a, Duf23, Hal11, HS23, New23, Rag22, Roo23, Tay23, Tra05, Bro67, Bro17c, Ham17, Kik09, Mac89, MT92]. Long [BV16a, Di 22]. Look [Kim91]. Looking [BSGBBV02, PHW22]. López [Num07]. Lord [PVPP17]. Lorentz [Mor06]. Lorenz [Tho88]. Lost [FCO15, Kra16a]. Lothar [Roc19]. Loudon [DLS86]. Louis [Ano06e, BBS10, Riv18]. Louisiana [Tra05, All03]. Louvain [Ano85a, Hed21]. Louvre [Riv18]. Lucies [Hay77]. Lucifers [Far20]. Luft [Ano05a, BB04, Hom17b]. Lullian [Szu88]. Luso [Fil92]. Lussac [Cro81]. Lutheran [Mon63]. Lüthy [Sha05]. Lux [Kah19]. lxxxiii [Vic94].

M [Bur02, Mor06, Num07, Rat92a]. M.I.T [Ano75c]. M.P. [Har00]. MA [Cob22, SLT49]. Macbridean [Sco70]. Machine [Lan22]. Macmillan [Mor06]. Macneven [Mil70]. Macro [vF65]. Macro- [vF65]. Macrobiotics [Mul75].

Macromolecule [TR99]. Mad [Sch06]. Madame [Sme89c, Sme89b]. madder [Far75]. Made [Bud18, Kar88, Mor21a, Qui16a, Mor06, Tra04]. Madrid [Num07, Ser13]. Magdolna [Ano07e]. Magellan [Mal08]. Magi [Shi49]. Magic [Hen17, Kar93, Rat66, Rya90, Hed17]. Magica [Hed17]. Magical [Gui80]. Magisterio [Del21]. Magnesia [FFS77]. Magnus [Gru09, New14, New15, Par37a]. Maier [LRT14]. mail [Ano17c]. Major [Rue22]. Make [Pri16c, Gru06]. Makes [BV16a]. Making [AF16, mKC21, CS19, Fri21b, Gor98, HVD20, Hem15, Kni16b, Kno01, Kno06, Mer90, Mor21a, Oak23, Pas20, PZ18, Pui04, Ree15, Roc22, Sim16a, Tom16b, Zha23, Iac22, Roq16]. Malaria [Sla04]. Malcolm [Ano05c]. Man [Ano06d, Dmi04, For08, Kni16b, LK12]. Management [Gay12]. Managing [Sán21]. Manchester [FFS73, FFS74a, FFS74b, FFS75, FFS76, FFS77, Pra92]. Manhattan [Her18]. Manière [Mor80]. Manipulation [Ram18]. Mannheim [Don88]. Manoel [Fil92]. Mansfield [War84]. Manufactory [Kle13]. Manufactured [Cob20]. Manufacturer [Mer03]. Manufacturing [Mer22]. Manuscript [Abr79, Jam19b, Kar90b, Maw18, Mor80, Tho63a]. Manuscripts [Fig73, Kah13, PVPP17, Pri13, Ric09, Tho56, Zub18, Pio18]. manuscrits [Pio18]. many [Dob82a]. Marburg [Bro16b, Mor21b]. Marburger [Bro16b]. Marcet [Kni86, LR16]. March [Kir23]. Marco [Duf23, Kra16a]. Marcus [Bro19b]. Margaret [Bay05, Clu11]. Margarita [Cri73]. Marginalia [Big17]. Maria [Num01, Pat82]. Mariagrazia [Kra16a]. Marie [Roq16, Ant22, Hem15, Tos01]. Marie-Anne [Ant22]. Marjory [Bro17b, Štr16]. Mark [Ano97c, Ano05b]. Markets [Her14]. Marqués [LP05]. Marrano [Pat88]. Martels [Rat92a].

Martin [Roo04]. Mary [Ano06c, Ano06e, Kra16a]. Mason [Bro08]. Mass [Ano75c, KLPG06, Tra05, Fer17]. Massachusetts [Mos22]. Masses [Mor80]. Master [App87]. Mastery [Sch22]. Material [Nye14, Sal13, Szu88, VMT22]. Materialist [Clu11]. Materials [Gol18, HVD20, Iri20, Kle05, Pas20]. Mateu [BS14, BG00, GL21]. Mathematical [Ano06c, Nye02]. Mathematics [Clu71, Deb68]. matière [Pio18]. Matte [Nev62]. Matter [AGJ08, And17, Bow75, CE17, Cis08, Dan06, Gor19, Jen81, Lev21, LMN01, Pag61, PW74, Pol20, Ree77b, Sha05, She62, Tei82, Zha23, Pio18, Pom62, Tha68]. Matthew [Tay23]. May [Ano22b]. Mayerne [App79]. Mayow [Böh63]. Meandering [Kra21]. Meaning [Gol57]. Measure [Bro19a]. Meat [Bro17d]. Mechanica [GS89]. Mechanical [Fra11, Kim91, OO74]. Mechanicism [Cle21]. Mechanics [Kno06]. Mechanism [Deb80]. Mechanistic [Las03]. Medal [Mck61a]. Medals [Kar88]. Media [Cob19]. Mediaeval [TCPYN59, HN59]. Mediator [Moy89]. Medical [BG00, Bra79, Car17a, Col88, Hed13b, Mor22b, Nor15, Web67a]. Médicale [BG00]. Medication [Sch22]. Medicine [AF16, All19a, Ano06e, Ano09a, dMC15, Duc08, Fer22, Fig65, Kle15, Maw18, ML08, Mor21b, Sec79, Sim16b, VG04, FH05, Num07]. Medicines [Di 22, New20c, Qui16b, Sim06]. Medics [Hol89]. Medicus [Jon18]. Medieval [Ano05b, Cla01, LMN01, Pol20, Run18, Sha05, Bra85, CGT18, Gru06, Kah10, Kah11, KR14, New14, New15, Obr09, Ram16, Sin59, Tho59a, Wil09]. Médiévale [Run20]. Mediterranean [MB21]. Méditerranée [Run20]. Medulla} [PLB23]. Meeting [Hud99, Oga08, Ano84c]. Meetings [Sme87b]. Meikle [Ano97c]. Melancholy [Bra85]. Melchior [KLPG06, Fer17]. Mellon [Mac67]. Melting [dS22]. Members [Ano17c, Sme87b]. Mémoire [Hom19, Ban07]. Mémoires [Sme89b]. Memoirs [BR00, Gou83, Roc21]. Memories [SK87]. Men [Mer22]. Mendeleev [Ano06d, Alm98, Bro98, Dmi04, Gor98, Gor04, Gor18, Ric98, Sta98, Pul19]. Mendeleeve [Ano06d]. Mendeleyeve [Dmi04]. Mendicant [CGT18]. Mercellis [Mos22]. Mercenarius [Del20]. Merchants [Di 22]. Merck [Ano06e, VG04]. Mercurius [Per69]. Mercury [Bil23, Mar23, Mer79, New14, Oak23, Rap18, Vrt23, Koh75, New15, Lin77]. Mesopotamia [Lev57]. Mesopotamian [Lev58]. messe [Fer17]. Metal [Kar88]. Metallogenesis [Dym08, Nor06]. Metallurgists [Lan66]. Metallurgy [Cra10, Kar92]. Metals [Hij22, Kar03, Mar23, Per19]. Metamorphosis [Mun91]. Metaphysics [CE17, McE78a, McE78b, McE78c, Mce79]. Metempsychosis [Sec79]. Meteorologica [Mar04]. Method [Jam83a, Sim02]. Méthode [Lef18b]. Methodology [Bro87, McE78a, McE78b, McE78c, Mce79]. Methods [Bro87, Rus37]. Metrology [Bro98, Gor98]. Metz [Per03, Per04]. Meurdrac [Tos01]. Mexico [Rap18]. Meyer [Roc19]. Meyerson [Chi17]. Michael [Ano04a, Ano06d, Bur02, Hey69, LRT14, Mos15, Por01, Pri16d, PZ18, Sum07, Szy93, Szy96, Wil12]. Michitake [CS19]. Microcosmos [vF65]. Mid [GW17, Leh09, Nye14, Orl16, Jam95, Leh10]. mid-20th [Orl16]. Mid-Eighteenth [Leh10]. Mid-Eighteenth-century [Leh09]. Mid-Nineteenth-Century [Jam95]. Mid-Twentieth [GW17, Nye14]. Middle [KF70, New23, Rob66, Wil75]. Miguel [Num07]. Military [DLS86]. Milk [SW14, Ver22]. Millions [Bro21]. Milton [Sad77]. Minds [Per19]. Mine [Nye14]. Mineral [App87, Bor15, Col84, Hen18, Lef18a, Mah77, Nor06, Nor07, Pag14].

Mineralogy [OO74, Old74]. Mines [Ben70]. Mining [CE17, Dym08, Kon13, Oak23, RGG23, Sán16a, Vrt23]. Minor [FFS75]. Mirror [Lau19, Gru09]. Miscellanies [Kib60]. Misdated [Tre79]. Mishaps [Fil92]. Misnamed [SD87]. Missing [And12, Bro20b]. Mitte [Orl16]. Mittelstand} [Mun90]. Mixed [Col78, War76a]. Mixis [Via15]. Mixture [dS22]. Mobility [Per19]. Mocked [Ano07e, Day05]. Model [Han76, Mau70, Roc03, War08]. Models [dCH04, Ano06a]. Modern [AF16, All19a, And16, Ano06c, Bro11, CGT18, Cha07, Clu07, Dri16, Dup23, Ebe07, Edd08, Fer22, For13, Fri21a, Gia18, Hal11, Hed13a, HBB^{+}03, HS23, Iri20, Kah10, Kah11, Kim20, KR14, Kni04, LMN01, Maw16, New20b, Nor15, Nor06, Nye02, PT17, Pri21b, Roc22, Roo23, Sch19, SC90, Sha05, Sim22, Sim16c, Sta98, Tom17a, Wal16, Wat17, Wes61, Zha23, Zub18]. Modernity [Duf20]. Modes [Dun81]. Modifications [Tay14a, Tay14b]. Molecular [Ram18, Uss03]. Molecules [Ano97c, Kim96, Mar93, Mor80]. Molella [Tra05]. Moment [Gro21]. Monad [For17, Nor05]. Monas [Clu17b, Clu05b, For05, Jos64, Luh86, Mer17, Van17, Wal76]. Monopolies [Sim06]. Mons [Per84]. Monsieur [Gue76, Sme89c]. Montpellier [Run20, Hun79, Le 82]. Moore [Mor16b]. Moran [Roo23]. Morienus [Sta70]. Moritz [Mor85]. Morley [Rag22]. Morosow [Lau19]. Morphine [Lev21]. Morris [Ano20a, HS23, Ano17a, Ano21a]. Mortars [Sim06]. Morveau [Sme57, Sme61, Sme63, Sme66, Sme68, Sme87c]. Moseley [Jew19]. Moses [App87]. Mossman [Ano97c]. Most [Pas20, WW07]. Mother [Pat82]. Mount [Gue15]. Moureau [New23]. move [Hag16]. Moyen [Zin84]. MS [Dun65, Jam19b, Tim06, Fer09, Fri59]. Ms. [Tho59b]. MSS [Kib60, Ple59]. Müller [Bro20a]. Multidisciplinary [Pul19]. multiple [Chi17]. multiples [Chi17]. München [Ano05a, Mor06]. mundo [GF23]. Münster [Lau19]. Murdoch [Sha05]. Museum [Ano97c, Bro20a, VMT22]. Museums [Ano82]. Museumsinsel [Ano12a]. Music [Wel51]. Musings [Mor12]. Muspratts [Mor16a]. Mutability [Run18]. Mystery [Fer16, Kau83, Per84]. Mystical [Bra85]. Mysticism [Hei95, Mer90]. Mythos [Zim84].

Nâgârjuna [Wuj84]. Nadim [Füc51]. Nahrung [Orl16]. Nakayama [Ano75c]. Names [HS15, HS20, O’N21]. Nancy [Per03, Per04]. Naples [Gia18]. Napoleonic [Sel07]. Narrative [Zub21]. Nathan [Ano75c]. National [Joh18]. Nationalistic [Sch89]. natura [Gan20b]. Natural [Bax13, Dob71, Dob73, Dob74, Duc08, Fer22, Gib60b, Gor21, Gre38, Gue15, Hei73, Kir03, Lef18a, Lla85, Rat66, Ree77b, Roo18, Sch67, Sie08, SK87, Kim20]. Naturalism [PW74]. Nature [Cha17, Cob20, Day05, Deb68, Gou83, GL17, Kno06, Rag22, Web66, Wee07, Ano07e, Hei73]. Naturforschers [SK87, Don88]. Naturphilosophie [Löw80]. Navelbine(R) [Tom18]. Nazi [Kea92, Rei10]. Nearly [Kle15]. Nearly- [Kle15]. Neglected [Sut76]. Nehemiah [Roo07]. Nemours [Sme89b]. Neo [Old74]. Neo-Platonic [Old74]. Neoplatonic [Pag60]. Netherlands [Bro02, Hen18, RK06, Van04]. Nettesheim [MJ75, New82]. network [Mal08]. Networks [Mor21a, Pui04]. Never [Wes80, Dob82a]. Newlands [Gor18]. Newly [LRT14, Pri13]. Newman [Sha05]. News [Ano98a, GC96, Hud98, Lea97]. Newton [Roo18, Sch19, All19b, DP03, Dob79, Dob82b, Geo57b, Hei73, McG68, Mcg67, Tay56, Tha68, Wal80, Wes75, Wes80, Wes84, Dob82a, Ano06b]. Newtonian [Sme71]. Newtons [Gor98]. Nicholas [Lau19, Pow98]. Nick [Ano06a]. Nicolas [Pri16a]. Nicolaus [KLPG06]. Niederer [Mun90]. Nieto [CG22]. Nieto-Galan [CG22]. Nikolai [Bro95]. Nikolaus [Lau19]. Nine [LRT14]. Nineteenth [And10, Ave86, BSGBBV02, BG00, Cha23, DM97, Dol98, Fis18, GB14, Goo69, Jam95, Kra89, Sal13, Sch06, Tra14, Vag19, Wey76, Whi18, BS14, Bro22, CC80, SG23, Gor77, PG21, Hij22, Lac17, Le 15, O’N21, Olt05, MR20, Sim14a]. Nineteenth-Century [BG00, Dol98, Fis18, Sal13, Tra14, Whi18, Sch06, BS14, SG23, Gor77, Hij22, Lac17, Olt05, PG21]. NiO [Bro16b, Bro16b]. NiO-Kristalle [Bro16b]. Nitration [War76a]. Nitre [Por01, Szy93, Szy96]. Nitrogen [Far71, HJ02, Kau72, Mår19, Tra17]. no [Hom19]. Nobel [Van22, BR00, DFB22]. Nobelpreisträger [Ano04a, BR00]. Noblewomen [Wal16]. Nocturnall [Hay77]. Noel [Bro20c]. Nomenclatural [O’N21]. Nomenclature [HS15, HS20, NW18, Sim02, Lef18b]. Nomination [DFB22]. Nominations [Ano05d, Ano07c, Ano07d, Ano17a, Ano20a]. non [Del20]. North [CS19, McV75]. Norton [Rei57]. Norwegian [Lyk20, Lyk22, Nie16]. Notation [O’N21]. Note [Bai00, Le 73, Sme77, Sme78, Wil09]. Notebooks [Ano07b, HRR03]. Notes [Ano04b, Ano05e, Ano05f, Ano05g, Cop92, Leh09, Poo09, Wes61, Vic94]. Notices [Ano70e, Ano70f, Ano72d, Ano73e, Ano73f, Ano74c, Ano75f, Ano75g, Ano77c, Ano89e, Ano98b, Ano01c, Ano04c, Ano06f, Ano06g, Ano07h, Ano07i, Ano10d, Bro97b, Hom01, Pri99, Ano11d]. Notions [Pap90]. Novel [Kea92, Gor19]. Novella [Cri73]. November [LP05]. noviembre [LP05]. Noyes [Ser76]. NS [Joh18]. NS-Herrschaftsapparat [Joh18]. NTNU [Nie16]. Nuclear [Her14]. Nucleic [Hun99]. Numbers [Wil88]. Numerical [Kar93]. Nut [Tha68]. Nut-Shell [Tha68].

Nutrition [dMC15]. Nutritional [PV17]. NY [Tho22]. Nye [Ano06c].

O [Bro02, Ano06d, Dmi04]. Oak [Tra95]. Oakleigh [Vic94]. Obituary [Bro20b, Duv96, Hol56, McK60]. Object [Amb22]. Observations [Per69]. Observed [Kle95]. occidentale [Hed17]. Occult [Dav87, FH05, Num07, Mac67]. Ocherka [Dmi04]. Ocherki [Ano06d]. Odling [Fis96]. Oeconomics [Rob06]. oeuvre [Duc08]. Offensive [Ric04]. Officer [Fis96]. Oganesson [Pul19]. Ohio [Her18, Her18]. Oil [Low18, Mus17]. Oils [Rae06]. Old [PT17, MT12, Rya90, VMT22]. Oldest [Kar90b]. Olfactory [Whi18]. Olschki [Bar19a]. Omodeo [Kim20]. One [Roq17, Ano97c, Moy89, Ree15, GL18]. Open [Ano05c, Pio18]. operation [Bro22]. Operationum [GS89]. Opponent [Rei10]. Opportunity [Pui04, Sut76]. Opposites [Pay89]. Opposition [Ric98]. Optical [HJ02]. Opticks [Tha68]. Orchil [WK16]. Order [BSGBBV02, Lef18a]. ordered [Ano06d, Gor04]. Orders [CGT18, Gia18]. Ordinall [Rei57]. Orfila [GL21, BS14, BG00]. Organic [Ano97c, Bro87, Fis73a, Fis73b, Fis74, HS15, Kap69, Las03, Lev70, Löw80, Mar93, Mun91, Rae89, Ram00, Roc87, Ros02, Rus87, Sch94, Sho03, Wei18]. Organisation [Sla04]. Organised [Lev21]. Organisms [Qui16a]. Organizations [Rae95]. Orgin [Mah77]. Origin [BJJK81, Col78, DeK80, Dub61, Gro21, New90a, Roc83, She70b, Sie82]. Original [Fri59, Sta70]. Originator [MC06]. Origins [Ben70, Del21, Gol80, Mas91, Pol15, Roc03, She62, She70b, Sum07, Tay37b]. Orna [Kra16a]. Ornament [Jam95]. O’Shaughnessy [Gor83]. Oster [Ano05c]. Ostwald [Kra18, DFB22]. Oswald [Sha03]. Other [Bee94, Tho60b, Van17]. Otto [BR00, Bro16c, Ano04a]. Our [Bro21, Sch22]. Ouroboros [She62]. Ours [Nye14]. ouverts [Pio18]. Own [MC06, Van04]. Oxford [Bar19b, Hal11, Kni04, Kra16a, Pul19, GS61, HBB^{+}03, Kni17, Lev81, MT12, Tho59b, VMT22]. Oximes [Kau72]. Oxygen [All92, Cro95]. Ozonised [Rae06].

P [Ano06e, Bro02, Vic94]. P. [Rob06, Sme89b]. PA [Tra05]. Pagan [Hal11, Van09]. Page [Sim22]. Paint [DeK09]. Painted [New20b]. Painters [Ano05b, Cla01]. Painting [Amb22]. Paintings [Dav87, Low18]. Pair [Tay49b]. Paiva [Fil92]. Palaeontology [Kni74]. Palate [Ree22]. Palgrave [Ano05c, Mor06]. Palgrave/The [Ano05c]. Palingenesis [Sec79]. Palissy [Kir61]. Palladium [DU79, KCU81]. Panaceia [Wal16]. Panopolis [Hal09]. Pantheism [Szu88]. Pantheus [Nor05]. Paper [Gor18, HS18, NW18, RT97]. paperback [Ano97c, CS19, Mos22, Rag22]. Papers [Clu07, FFS75, Tay49a, Wil68, Wil70a]. Paracelse [Fra17]. Paracelsian [Con22, Deb60, Deb62a, Gan20b, Har00, Maw16, Mor21b, Mur20, PW69, Ree75a, Ree75b, Tos01]. Paracelsianism [Hed21]. Paracelsians [Deb81]. Paracelsus [Cis08, Dan06, Eam21, Fra17, Gan20a, KH20, Mor96, New20a, Pag60, Pag61, Pag62, PW74, Rat63, Roc85, She48, Spa20, Tit38, Fer22]. Paradigm [Sha20]. Parallel [Min18]. Parent [Mer17]. Pariente [Num07]. Paris [Hom19, Kah21, Pri13, CS79]. Parisian [Per10b]. Parity [Mas91]. Parkes [Kni86]. Parkinson [Kni74]. Part [And12, Ano68a, Cra10, Dob73, Dob74, FFS73, FFS74a, FFS74b, FFS75, FFS76, FFS77, Fis73b, Hey37b, Kah10, Kah11, Kin82, KCU81, McE78a, McE78b, McE78c, Mce79, Per03, Per04, Sin88a, Sin88b, Sme87a, Sme87b, Tay14a, Tay14b, Wil68, Wil70a]. Participation [Kik11]. Partington [And12, Ano84b, Ano85c, Ano88b, Ano07f, Ano07g, Ano16a, Ano16b, Ano16c, Ano19b, Ano20b, Ano22a, Ano23a, Bro19c, New90b]. Parvenir [Zin84]. pasado [GF23]. Passion [Hor18]. Past [Oak23, GF23]. Pasteur [Mas91]. Patents [Gab16]. Path [Big17, Nic12]. Pathology [Arm12]. Pathway [Ant19, Nic12, Roc19]. Pathways [Ano97c]. Patrimoine [Hom19]. Patriotic [Ser22]. Patriotism [Car16a]. Patronage [Mer20, PZ18]. Patterns [Hom05]. Paul [Mor06, Far89, Sán16c]. Paulze [Ant22]. Paulze-Lavoisier [Ant22]. Päuser [Ano06e]. Payen [Mer03]. pbk [Ano05a, Ano06a, Ano06b, Ano06e, Ano07e]. Peculiarities [Gol17]. Pedagogical [Tay14a, Tay14b]. Pedagogy [DFB22]. pensamiento [LP05, Num07]. Pensar [GF23]. People [Ano07e, Day05]. Perchlorate [Tra14]. Peremen [Dmi04, Ano06d]. Pérez [Num07]. Perfect [Low18]. Perfecto [Del21]. Performance [Mor22b]. Perfume [Lon07]. Period [Löw80, Ree20, Sie08, Sel07]. Periodic [Ano06d, Bro16a, Gor04, Gor18, Kaj13, Kra16a, Pul19, Pul20, Res23, FCO15]. Periodicity [Roc19]. Peripheral [Fil92, Pal10]. Peripheries [Hom05]. Perpetual [Hei73]. Perry [Shm94]. Personal [Sme87c]. Personality [Gol17, GLCK76]. Perspective [Car16a, GL15, Kra16b, Pul19, Sim98, WW07]. Perspectives [Ano97c, Bro20d, Mos97, Sim16b, Tim06]. Peru [CE17, New20c, Min18]. Pest [Gay12]. Pesticides [Gay12, Pér16]. Peter [Ano07e, Cha23, HS23, FFS74a, Gol83, Pra91, Rei10, Sum07]. Petersburg [Ano06d, Wer13]. Petroleum [For38, Sta98]. Petrus [Cri73]. Pewterer [Rue22]. Pharmaceutical [Ano06e, BCS03, Pap90, RST03, Sim06, Ano06e]. pharmaciens [CS79]. Pharmacist [For08]. Pharmacists [Sel07]. Pharmacopoeia [Qui16b]. Pharmacopoeias [Sim22]. Pharmacy [Hen18, MB21, Sim98, Sim16c, Bar19a]. Phials [SvOM93]. Philadelphia [Ano06e]. Philalethes [New90a, Wil64, Wil72]. Philanthropy [CA15, GW17]. Philip [Vau04]. Philipot [Roo01]. Philipp [Mor96, Tho88]. Philosopher [McE78a, McE78b, McE78c, Mce79, Por01, Sch67, Szu86, The86, Rus19]. Philosophers [Pol20]. Philosophiae [PLB23]. Philosophica [Ano84a]. Philosophical [Bro02, Cha21, Kle15, Tay14a, Tay14b, Deb80, Old73]. Philosophicall [Jos63]. Philosophies [Vag19]. Philosophorum [Hed22, Hal09, Ple59]. Philosophy [Bax13, Bro02, Cle21, Deb80, Dob71, Dob73, Dob74, Duc08, Gay97, Gib60b, Gor21, Gre38, Hei73, Kaj13, Kim91, Kir03, Kni16c, Ree77b, Roo17a, Sán16b, Szy96, Web67b, Fer22, Roo18]. phlogistic [Sie89]. Phlogiston [All92, Con96, Jen81, Lei75, Mau02, Sud78]. Phoebus [Hun99]. Phosphor [Kel14]. Phosphorus [Sou00]. Photographic [Mer22]. Photography [Amb22]. Photosynthesis [Nic12, Sco70]. Physical [Ano06c, Mcg67, Nye02, Wei18]. Physician [Fig65, Fil92, Rag12, Wil63b]. Physicians [Iac22, Sel07]. Physics [Run18]. Physiology [Col96]. Physitians [Web67a]. Piccardi [Bar19a]. Pickering [Bur02]. Pictures [Pet15]. Pietro [Kim20]. Pigments [Kir23]. Pile [Sud80]. Pillars [Gay00]. Pinner [Ano06b]. Pioneer [Gor83, Sán16c]. Pioneering [Bro20e]. Pioneers [Jam22]. Pittsburgh [Tra05, Tar03, Tar05]. Place [Bre76, Gou83]. Places [Ano07e, Day05, RK06]. Planetary [Car16a]. Plant [Kle05]. Plantis [Tho63b]. Plants [Mor22b, Roo07, Str79]. Plastic [Ano97c, Mei95]. Plastics [Ano97c, Mos97]. Platinum [And15]. Plato [Wil88]. Platonic [Old74]. Plattes [Deb61, Geo62]. Playhouse [Her18]. Please [MC06]. Pliny [Lon07]. Plot [Tay49a]. Plutarch [vL48]. Pneuma [Rin17]. Pneumatic [ML08].

Pocket [Sal19]. Poem [Sal17]. Poetics [Ano08a, Sad77, Söd07]. Poetry [Kah10, Kah11, Lin84, War76b, Tho22]. Poison [BS22, Dav87, Mur22, PC23, Wat17, Orl16]. Poisoned [SMGL18]. Poisoning [Sán21]. Poisonous [NG20]. Poisons [Rya90]. Policies [PV17]. policy [Orl16]. Political [Cha07, RGG23]. Politics [Lei16, Lev81, Sán16a]. Politik [LE03, Mor06]. Pollution [NG20, Ree12, Wat17]. Pollutions [CB23]. Polonium [Her18]. Polymers [Ras18]. Polytechnique [Lan81]. Pont [Sme89b]. Pope [Jos49a]. Popp [PZ18]. Porcelain [Kle13]. Porta [Hed17]. Portable [DeK09, Sme66]. Portrait [Jen04]. Portugal [De 85]. Portuguese [Abr79]. Positions [mKC21]. Post [Hor11]. Post-war [Hor11]. Postage [RT97]. Pot [dS22]. Potassium [Sou00]. Potsdam [Ano04a, BR00, BR00]. Powder [Jon93]. Power [Web67b]. Powers [Bow75, Szu88]. Pp [Ano75c, Ano97c, Ano04a, Ano05b, Ano05a, Ano05c, Ano06c, Ano06a, Ano06b, Ano06d, Ano06e, Ano07a, Ano07b, Ano07e, Ano08a, Bar19a, Bar19b, BS22, Bro20a, Bro97c, Bur02, Cha23, CS19, Cob22, CG22, Dob82a, Don88, Duf23, Fer22, Hal11, Hed22, HS23, Her18, Hom19, Kim20, Kni04, Kno22, Kra16a, Las96, Lau19, Mor06, Mos22, New23, Num07, Pul19, Rag22, Rat92a, Ree22, Roo23, Run20, Sha05, Tay23, Tho22, Tra05, Vag19, Van22, Vau04, Vic94, Iac22]. Practical [Dup23, Hen19, Jam83a, Mor69, Vin09, Wil98]. Practice [Kel12, Mei83, Nor20, PHW22, Rae16, Taa14, War76b, Wil98]. Practices [Eli11, Maw16, PV17]. Practise [Geo57a]. Prague [Kar90a]. Pratique [Nev62]. Pre [Cam91, Mck61b, Nor06]. Pre-History [Cam91]. Pre-Modern [Nor06]. Pre-Publication [Mck61b]. Prediction [GL17]. Preliminary [Ber92, Duc08, Kah10, Kah11, Ple59]. Prelude [Per69]. Preparation [Col84, Por01]. Preparatory [Low18]. Present [Las03, Oak23, GF23, Jam15, Bar19b]. Presentation [dS22]. presente [GF23]. Press [Ano75c, Ano97c, Ano06c, Ano06a, Ano06b, Ano06e, Ano07e, Bar19b, Bro97c, CS19, Cob22, Dob82a, Hal11, Her18, Kni04, Kra16a, Las96, Mor06, Mos22, Num07, Pul19, Rag22, Ree22, Tra05, Van22, Vic94]. Press/Science [Ano97c]. Presses [Run20]. Pressure [Ree12, Ree20]. Pretiosa [Cri73]. Preußen [Hag16]. Previously [Big17, Jen04]. Priestley [Abr64, Con96, McE78a, McE78b, McE78c, Mce79, Mck61a, Sch67, Sme68]. Primam [Big17]. Primary [Ano05c, Ost02a]. Prime [Pag61, PW74]. Primi [Big17]. Principe [Bur02]. principes [Per84]. Principle [Web66]. Principles [Hir81, Kle14, Lev21, McG68, Old73, Riv18]. Print [Clu17a]. Printer [Fri22]. Printing [RT97]. Priority [Ful75, Rae97, Rus19]. Private [Eli11, GW17, Kah21]. Privilege [Mor85]. Prize [Ano07f, Ano07g, Ano16a, Ano16b, Ano16c, Ano19b, Ano20b, Ano22a, Ano23a, Ano84b, Ano85b, Ano85c, Ano88b, DFB22, New90b]. Prizewinners [Ano00b]. Probably [Bri66]. Problem [Cec04, Lev70, LE03, Mor06, Tar02, Tat19, Tra14, Wil64]. Problematic [Kir03]. Problems [Goo69, GL17, Kin81, Kin82]. Proceedings [Rat92a, vM90, Roq17]. Process [vR07]. Processes [Sco70]. Production [Cha81, Gol18, Maw18, Wil75]. Productions [Van17]. Products [Mer03, Sie08]. Profession [Hom99]. Professional [App77, SL22]. Professions [Roc94]. Professor [Kir03, Lyk20, Rea61]. Profiles [Ano97c]. Progeniture [Tay14a, Tay14b]. Progress [Kin81, Kin82, Löw80, Mar93, Ano97c].

Project [Fig73, Her18]. Prolongation [Hed13b]. Prominent [Ano07e, Chu14, Chu13]. Promise [Zub21]. Proofs [Sán21]. Propaganda [CA15]. Propagandist [Har00]. Property [Hem15, Roq16]. Prophecy [Lei16, New90a]. Prophet [Hal11, Van09]. Proponents [Abr64]. Prosecutor [Las94, GL21]. Prospect [Ful19]. Prospecting [Dym08]. Prospectors [Mac88]. Prospectus [Nev60]. Protecting [dV22]. Protein [TR99]. Protestant [Kim20]. Prout [Bro86a]. Providence [Bax13]. Provincial [Le 82, VP18, Per03, Per04]. Prussia [Hag16, Hij21]. Prussian [Kle13]. pseudo [New20a, Gru06, Gru09, Man61, Mar09, Mou09, New14, New15, Szu88, Tho63b, Fer22]. Pseudo-Albertan [Gru06]. Pseudo-Albertus [Gru09]. Pseudo-Aristotelian [Man61]. Pseudo-Avicenna [Mou09]. Pseudo-Democritus [Mar09]. Pseudo-Galen [Tho63b]. Pseudo-Lullian [Szu88]. pseudo-Paracelsus [New20a, Fer22]. Pseudo-Roger [New14, New15]. Psychological [Mar75]. Psychologists [RW88]. Psychology [vF65]. Public [BV17, Bro97d, Eli11, Kah21, Las94, Mil68, Ric98]. Publication [Bai00, Mck61b, Sme89b]. Publications [Ano58a, Ano60a, Ano64a, Ano67b, Ano68c, Ano70a, Ano70b, Ano05b, Bro02, Zot93]. Published [Gib60a]. Publisher [Fri22, Kim20]. Publishers [Ano07b, Bro02]. Publishing [Ano07a, BS22]. Pure [Cob20, Cob22, Uss15, Kni16a]. Puritan [Bil13, Rat63, Web67a]. Purling [Rue22]. Purple [Kir18, WK16]. Puzzles [Wil73]. Pythagorean [Wil88, Wil98]. pythagoricien [Hed22].

q [Car17a]. quête [Pio18]. Qualitative [Jam83a]. Qualities [Clu71, Mcg67]. Quality [PG21, Las10]. Quantifiable [Ban07]. Quantification [Ede96]. Quantified [Tay08]. quantum [Sim14b]. Queen [Pri79, RT97]. Quellen [Ano04a]. Query [Bra85]. Quest [Kle15, Roo17a, Sch22, Pio18]. Question [Cha07]. Quicksilver [RGG23]. Quiet [Mor16b]. Quotation [Ple69]. Quoted [SLT49].

R [Bro02, Bur02, Iac22, Rat92a, Sha05]. Rachel [Tra12b]. Radical [Hed13b, Lev81]. Radio [Tre79]. Radio-Activity} [Tre79]. Radioactivity [Sin85, Sin88a, Sin88b]. Radium [Her16]. Ragland [Iac22]. Rain [NG20, Sán16a]. Rainbow [Ree22]. Raleigh [Rat66, Shi49]. Ramberg [Cha23]. Ramsay [Tre74, Wat95]. Ranelagh [Bou22]. Rapport [Kle95]. Rare [Kra21]. Raritan [Mor06, Tra04]. Rasaratnâkara [Wuj84]. Rationality [HS20]. Ratione [GS89]. Ray [DeK09, Tre79]. Raymund [Pat88]. Rayon [Mor18b]. Razi [Hey38a, New14, New15, Par38a]. Re [For21]. Re-enactment [For21]. Reaction [Cre17, Wei18]. Reactions [Sin59]. Reader [Ost02a, Ost02b, Ano05c, Ano05c]. Reading [Nor05, Sch19]. Real [Gan20b, Zac56]. Realm [McG68]. Reappraisal [Per73]. Reasoning [Lan22]. Receipt [Bay05]. Received [Ano58a, Ano60a, Ano64a, Ano67b, Ano68c, Ano70a, Ano70b, Ano71a, Ano72a, Ano73a, Ano73b, Ano74a, Ano75a, Ano75b, Ano89c]. Reception [For05, GC80, Per82]. Recipe [Ano05b, Cla01]. Recipes [All19a, Mor21a, Car17a]. Recollections [GLCK76]. Recombinant [GS16]. Reconsidered [Con96]. Reconstructing [Dup23]. Reconstruction [For21, Hor11]. Record [DN37]. Records [RST03, Ano06e]. Recta [Gru09]. Red [CT88]. Redemption [She59]. Rediscovered [Fri59]. Rediscovery [Koh75]. Rees [Cob22]. References [Bur92, Gou82, vL48]. Refined [Rue22]. Refining [Rap18]. Reformation [Mon63, dV22]. Reformers [Web67a]. Reforms [Sel07]. Refuge [Naw14]. Regency [LR16]. Regimens [Dob79]. Region [Tar03, Tar05, Tra05]. Regionalgeschichte [Ano04a]. Regius [Lev81]. Regulation [GL15, Ree12, SMGL18]. Rehabilitation [Fin91]. Reich [Joh18, Ano12a, Joh18]. Reign [Kra22a]. Reinterpretation [Kle95, Koh75]. Related [PW69]. Relation [BN80, Col96, Kir03]. Relations [Sad77]. Relationship [Bra85, Bro21, Ree22]. Relationships [BJ21]. Relatives [Mor80]. Relazioni [Car17b]. Reliable [Taa14]. Religion [Mor96, Num13, Son77, dV22]. Religious [Bro48, Gia18]. Reluctant [Per82]. Remane [Ano04a]. Remarks [Ano94c, Cia01, vL38]. Remedies [MB21, Taa14]. Remedy [Riv22]. Remsen [Han76, War08]. Renaissance [Num07, Roo18, Bra85, Deb67, Deb68, Deb72, Deb81, FH05, Gel94, Gui80, Mar04, Mur22, PW74, Sad77, Sec79]. Renewal [Tar03, Tar05, Tra05, Zin84]. Renn [Ano07b]. Renovation [Roc94]. Reorganisation [Fau11, Kik11]. Replication [For21]. Reply [Bro86a]. Report [Hud99, Par37b, Ple59, Tra93]. Reports [Kni94]. Representations [NW18]. Representing [Rag22, Cob19]. Repression [ZH05]. Reproducing [AF16, Hed13b]. Reproduction [Num01]. Rerum [Num07, Gan20b]. Research [Ano07b, Arm12, BCS03, Car16a, Gel78, GW17, GL18, HRR03, Kra21, Lev58, Mor72, Rae89, Rus37, Sla04, Tre79, vR07, Nie16, Ano06e]. Researchers [Chu14, Chu13, Ano07e]. Residues [Tra12b]. Resilience [Sán16a]. Resistance [Cec04]. Resources [Kni19]. Respect [Luh86]. Respiration [McK38]. Response [All19b, Ano09a]. Responses [Bro16a]. Rest [Wes80]. Restless [BV16a].

Restoration [Rat64]. Restrepo [Pul19]. Résumé [Tit38]. Resurrected [Mat13]. Rethinking [Cis08, Dan06]. Retrospect [KP78]. Return [Ben72]. Rev. [Edd01]. Reveal’d [Dob79]. Revelations [Bil13]. Reversal [mKC21]. Review [Ano48, Ano51, Ano75c, Ano05b, Ano05a, Ano06c, Ano06a, Ano06b, Ano06d, Ano06e, Ano07a, Ano07b, Ano07e, Ano08a, Bág16, Bar19a, Bar19b, BS22, Bro02, Bro17b, Bro19a, Bro20a, Bro97c, Bur02, Car17b, Cha23, CS19, Cob22, CG22, Cra93, Dob82a, Don88, Duf23, Fer22, Gol18, Gor19, GL18, Hed22, Hen19, HS23, Her18, Hom19, Hom95, Hor18, Iac22, Jew19, Kah19, Kim20, Kir18, Kle04, Kni16a, Kno22, Kra16a, Kra18, Las96, Lau19, Low18, Min18, Mor06, Mor18a, Mos22, New23, Pat92, Pul19, Rag22, Ram18, Rap18, Rat69, Rat92a, Rat96, Ree22, Roo18, Roo23, Roq16, Run18, Run20, Rus98, Sch19, Sha05, Sme89a, Tak17, Tay19, Tay23, Tho22, Tom18, Tra05, Vag19, Van22, Vau04, Vic94, Whi18]. Reviews [Ano37a, Ano37b, Ano38b, Ano38c, Ano49a, Ano49b, Ano53, Ano56, Ano57a, Ano57b, Ano58b, Ano59a, Ano59b, Ano60b, Ano60c, Ano60d, Ano61a, Ano61b, Ano61c, Ano62b, Ano62c, Ano62d, Ano63a, Ano63b, Ano63c, Ano64b, Ano65b, Ano66a, Ano66b, Ano67c, Ano67d, Ano67e, Ano68d, Ano68e, Ano68f, Ano69, Ano70c, Ano70d, Ano71b, Ano71c, Ano72c, Ano73c, Ano73d, Ano74b, Ano75d, Ano75e, Ano77b, Ano80, Ano89d, Ano90b, Ano94d, Ano97c, Ano02, Ano04a, Ano05c, Ano07a, Ano08b, Ano08c, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano11a, Ano11b, Ano11c, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, BKS^{+}15, Bro80, Bro81, Bro84, Bro86b, Bro90, Bro00b, BZC^{+}15, Buè03, Cle93, Cle98, Col76a, Col83, Cou77, Deb76, Deb84, Dej73, Duc14, Far79, Gav99, Gre90, Hal96]. Reviews [Hal11, HFL^{+}14, Hen03, HS88, Hud02, Hut92, Kar78, Kau93, Kau94, Kel78, Kel89, Kel97, Kni00, Kni04, Kwa03, MTC^{+}14, Mau72, MJD^{+}15, McK59, Mea01, Mei01, Mun97, Pag72, Pag83, Per79, Pri93, Pri94, Pum95, Rae96, Rat78, Rat92b, Ree84, Ree95, Ree97, RHF^{+}14, RR05, Sch01, Sch87, She71, She75, She77, She79, She81a, She81b, She85b, She90, She91, Sho86, Smi80, Sno98, Sta87a, Sta87b, Sut98, Sut99, Tei95, Tor00, Tre82, War85, War89, Web82, Wil70b, Wil76, Wil91, Win83, Win91, Win92, Num07]. Revisited [Rat92a, vM90, Sim14b]. Revisiting [All19b, NG15, Ras18]. Revolution [Cro09, De 85, Gou83, Gue76, GL17, Lla85, Per10b, Rat63, Sim98, Web67a, Per03, Per04, Sme89c]. Revolutionary [Mor16b, Tom17b]. Revolutions [Cro95, Gou82]. Reworking [HRR03, Rob16, Ano07b]. Rey [McK58]. Reynolds [Amb22]. Rheinberger [Ano07b]. Rhetorical [Bro48]. Rhodium [KCU81]. ricettari [Car17a]. Richard [Dob82a, All19b, DU79, HB84, Mau02, Tay08]. Richards [Gay97, Kop76]. Richmond [Ano06e]. Ripley [Bra79, Ram08, Ram10]. Rise [BV16b, Hom99, Mor11, Mun90]. Rival [Han67]. River [Tra04, Mor06]. Road [Ano07e, Chu14, Chu13]. Roads [Pag69]. Robert [Bro16b, Bur02, Sie08, Tay49a, Bou22, Bro16b, Bro22, Bur99, Cle21, Deb67, Eam80, Fis18, Fri22, HS78, HD00, HCP01, BJ21, Jos49b, Jos63, Kim91, Pri92, Pri16b, Sim20, Wes61, Zub21]. Roberti [Cam01]. Roberts [Ano07a, Tra91]. Robinson [Sie08, Sla01]. Rocha [CA15]. Roche [Ano06e]. Rochester [Tho22]. Roger [Bre76, Mat13, New14, New15]. Role [Ede93, Jac00, Jan20, Kop76, Pri87, Rus87, SB92, Ser13, Tra14, Ber11]. Roles [Zub18]. Rolfinck [Pri21a]. Roma [Hed17]. Roman [Jan20, Lon07, Tar18]. Romantic [Kir03, Löw80]. Rome [Hed17]. Root [Wal80]. Roots [Mar21, Sch06]. Rosicrucian [Zub18]. Rotger [GL21]. Rouelle [Leh09]. Rounds [Bee94]. Route [Uss03]. Roy [Ano06e, VG04]. Royal [Ben70, Jam15, Jam19b, Rag22, Ben72, CC80, Clu17b, Gay00, Kle13, Oga08]. Royale [Byc21, Le 82, Str79]. RSC [Ano07a]. Rubber [Mor18a, CS19]. Rudolf [PZ18]. Rule [SD87]. Rummel [Van22]. Russell [Ano07a, Bro97c]. Russia [App79, Bro97d, Bro05, Eli11, Fig65, Gel78]. Russian [Dmi04, Rya90, Sta98, VP18]. Rutgers [Ano97c]. Rutherford [Tre74, Tre79]. Ruysch [Kno06]. Rychard [Har00].

S [All19b, Bar19a, Dob82a, Hay77, Mor06, Sme89b, Tho22, Szu88]. S’-Circle [Szu88]. Sabine [Ano06e]. Saccharin [KP78, War08]. Sage [Hal11, Van09]. Sages [Kel71]. Sainte [Duv48]. Sake [Van04]. Sala [Gel94]. Saline [Roo07]. Salis [Lon07]. Salk [Bág16, DeC15]. Salons [BS14]. Saltpetre [Rob16, Tra14, Wil75]. Salts [Edd01]. Salve [Cam01]. Samuel [Kni86]. Samurai [Kik09]. Santorio [Big17]. Satire [Jam21]. Saturn [Kra22a]. Sayings [SLT49]. Scale [Gol57]. Sceptical [Web66]. Scepticism [Pow14]. Scerri [Bro02, Pul19]. Scheele [For08, Jam22, Tay19]. Scheme [Ano16d, Ano17b, Ano19a, Ano19c, Ano21b, Ano22c, Ano22b, Ano23b]. Schemnitz [Kon13]. Schenzinger [Kea92]. Scholarship [Cha07]. School [Ben70, Bro17a]. Schools [Hor18, Mor72]. Schummer [BS22, Bro02]. Science [All19a, And16, Ano75c, Ano97c, Ano06c, Ano07e, Ano09a, BV16b, BV17, BGH78, Bur16, Cra10, Day05, Deb71, Fil92, For08, FPS16, Fre10, SG23, GL15, HH00, HH05, HBB^{+}03, Hei95, HRR03, Jam95, Jam22, Jew19, Joh11, Kar93, Kir23, Kni16b, Kni17, Las96, Le 15, Lef18a, LK12, Mer04, Min16, Mos22, Moy89, Mur22, NG16, Nor06, Num07, Nye02, Ost02a, Ost02b, Pas20, Pér16, Rag22, Roc94, Roc03, Sán16b, SL22, Sha94, Sim87, Sim14a, Sim16a, Son77, Vag19, VG04, Van09, Van04, Wel82, Wes61, Zot93, Zub21, dCH04, LP05, Ano05c, Ano06e, Ano07e, Hag16, Hal11, Lyk18, ZH05, Ano06a, Ano07b, Kni04, Ano05c]. Sciences [For21, Kim96, Nye02, Byc21, Le 82, Str79, Ano06c, Tom18]. Scientific [Ano07e, Ano08a, BV16b, BS22, Cec04, Cha21, Chu14, Chu13, GL17, Kel71, Lac17, Mal08, Mei83, Mor09, MT16, NG16, Per19, RC16, Ser13, Shi49, Sim14b, Söd07, Zot93]. Scientist [Gol17, Sch06, SL22, SK87]. Scientists [HH05, ZH05, Ano07e]. Scot [Hey69]. Scotland [And10, Rea38]. Scottish [Edd01]. Scribe [Poo09]. Script [Sea56]. Se [Abr79, Per69]. Sea [Kir18, Mac88]. Seabra [De 85]. Search [Pat92, Rae16, Str79]. Sébastien [New23, Nev62]. sec [Car17a]. Second [Mon38, Vag19, Van04]. Secondary [Ano05c, Ost02b]. Secrecy [Mer04, PLB23, Pri92]. Secret [Ebe07, Hal11, Her16, Her18, Rei10, Sea56]. Secretorum [Man61, Rya90]. secrets [Pio18, Dob79, Dob74]. Secretum [Man61, Rya90]. See [ML08]. Seeds [Cis08]. Seen [Leh09]. seinen [Bro20a]. Self [KM90, Kni16b, Sch22]. Self-Mastery [Sch22]. Self-Taught [KM90]. Sell [PC23]. Semi [Ree75a, Ree75b]. Semi-Paracelsian [Ree75a, Ree75b]. Seminal [Riv18]. Semita [Gru09]. Sendivogius [Lin77, Por01, Pri16d, PZ18, Szy93, Szy96]. Sennert [Hir21, Kle15]. Senses [BP04, Ano06e]. Sensors [Ano06e, BP04]. Sergeevich [Ano06d]. series [Ano97c, Roo18]. serve [Bax13]. set [Bur02, Cha23, Duf23, HS23, New23, Roo23, Tay23]. Setting [PV17]. Settles [Pra92]. Seventeenth [Cam01, Las96, Pri21a, Tho22, Bal93, Big17, HVD20, Hed13b, Lin84, Old74, Poo09, Pri16c, Rag12, Ree77a, Rob16, Roo01, Sad77, Sha94, Szy96, VMT22, WW07, Wil68, Wil70a]. Seventeenth-Century [Las96, Pri21a, Tho22, Hed13b, Lin84, Poo09, Rob16, Roo01, Sha94, VMT22, WW07, Wil68, Wil70a]. Seventh [Fin91, Tho38a]. Sevilla [Num07]. Sex [Pay89]. SHAC [Ano22b]. Shadow [Ano06d, FCO15, Gor04, Kra16a]. Shapin [Las96]. Shaping [Kra16b, Sim16c]. Shaw [Gol83, Sum07]. Shell [Tha68]. Shellfish [Kir18]. Sherley [Deb80]. Sherwood [Hol56, Sim87]. Shifting [Hom05]. Shigeru [Ano75c]. Short [Ano70e, Ano70f, Ano72d, Ano73e, Ano73f, Ano74c, Ano75f, Ano75g, Ano77c, Ano89e, Ano98b, Ano01c, Ano04c, Ano06f, Ano06g, Ano07h, Ano07i, Ano10d, Ano11d, Bro97b, Hom01, Pri99, Tom16a]. should [Bax13]. Siberia [Alm98]. Sibiu [Fer17]. Side [Edd08, FCO15, Kra16a]. siècle [Pio18]. siglo [GL21]. Signed [Pra10]. Significance [Deb85, Dob82b]. Silberrad [Sha03]. Silent [Gay12, Tra12a]. Silica [New86]. Silicon [Mor16b]. Silk [Kir18, Mor18b]. Silver [Rap18, Rap18]. Silvius [Clu17b, Mer17, Van17]. simbolo [Hed17]. Simple [Sie82]. Since [Ber11]. Sine [Pow98]. Singapore [Ano07e, BS22]. Singular [Riv22]. Sir [Amb22, App79, Clu93, Deb62b, Dob71, Dob73, Dob74, Gor83, Gre38, Kel12, Pri13, Shi49, Sie08, Tay56, Wat95]. Sister [Bou22]. Site [CA15]. Sites [GB14, Kon13, Le 15, Per13, Rei15]. Sivin [Ano75c]. Sixteenth [Lan66, Nor07, Old74, Ple69, Tho38b]. Sixteenth-Century [Ple69]. Skies [NG20]. Sloane [Maw18]. Small [Ano97c, Mar93, Tay08]. Smeaton [Col02, Rus01]. Smell [Whi18]. Smithells [DeK80]. Smithson [Iri18]. So-Called [Gol80]. Social [Ave86, For21, Kim91, Las96, Mun90, Sha94]. Socialist [Joh18]. Société [CS79, Le 82]. Societies [Ave86, Ser13]. Society [Ano97c, MT92, Rag22, Van17, VP18, Ano84c, Ano16d, Ano17c, Ano17b, Ano19a, Ano19c, Ano21b, Ano22c, Ano23b, Bal93, Bro67, Bro11, Bro20b, Joh18, Oga08, Sim06, Sme87a, Sme87b, Web67a]. Socio [Sim16b]. Socio-Cultural [Sim16b]. Söderqvist [Ano08a]. Sofie [Van22]. Soldier [Zub18]. Solomon [MT12]. Solution [New86]. Solutions [TCPYN59]. Solved [Per84]. Solving [Tra14]. Some [App79, Cia01, Deb72, Far71, Geo57b, GLCK76, Gou82, Lin77, Lov72, MT11, McK38, Mck61b, Mou09, Old74, Per69, Poo09, Sch95, Sco70, Sin59, Tho56, Tho59a, Tit38, War76b, War97, Wil73, vL38, Far75]. Sonnets [Wel82]. Sons [FFS74a]. Sophisticated [Vag19]. Soraya [Ano06a]. Soul [Szu88]. Souls [Tho59b, FH05, Num07]. Source [Hal09]. Sources [Abr94, Ano05c, Kni19, Lev58, Lov72, Mor22a, Ost02a, Ost02b, Ram08]. Soviet [Bro05]. Spa [Edd01]. Spaces [Kim96]. Spain [Rap18, GC80, MB21]. Spanish [GF23, GL21, LP05, Pui04, Ser22, SMGL18]. Spark [Jam83b]. Spas [SM17]. Speaking [Fig73, Vag19]. Special [Ano22b, Hom19, Sad77]. Spectra [DeK80, Jam83b, Jam85]. Spectro [Jam83a]. Spectro-Chemical [Jam83a]. Spectrometry [DeK09]. Spectroscopy [CT88, Sut76]. Spectrum [Rae97, Shm94]. Spéculation [Run20]. Speculations [KLPG06, Roc83, Tre74, War97]. Sphere [BV17]. Spheres [Pol20]. Spiegel [Lau19]. Spirit [Kra16b, Pag62]. Spiritual [Mer90]. Spiritualism [Ric98]. Spiritus [Pri86]. Spread [Roc03]. Sprenger [Fer09]. Spring [Gay12, Tra12a]. Square [Bee94]. Squares [Kar93]. St [Ano06d, Wer13]. Staatsbibliothek [Fer09]. Stability [Zim84]. Stages [Ham82]. Stahl [mKC21, Old73, Roo17b]. Stahlism [Fic71]. Stalin [ZH05]. Stamps [RT97]. Standardization [Kra16b]. Standards [PV17, Qui16b]. Stanford [Ano06a, GS16]. Staple [dMC15]. Stapleton [McK63]. Stark [Don88]. Starkey [Wil63b, Wil73].

State [Bro00a, Her18, Qui16b, RS04]. States [Mil68, Mil70, Tar02]. Statique [Sme77, Sme78]. Status [Mei83, Sim14a]. Stefano [Vag19]. Stephanos [Tay37a, Tay38]. Stephanus [Pap90]. Stephen [Las96, Bax13, Bro08, Kau83]. Stephenson [Štr16, Bro17b]. Stepping [For08]. Steps [Kni67, Tay08]. Stereochemistry [Kau72]. Stevenson [Ano06e]. Stills [BN80, Sim14a]. Stocks [Sim14a]. Stoic [Old74]. Stoicism [Rin17]. Stokes [DeK80, Rae97]. Stone [Col71, Iri17, Por01, Pri16c, Szu86, The86]. Stones [Rya90, Tho63b]. Stories [Ano07e, Chu14, Chu13]. Storm [MC06]. Story [Bro20b, Hor18, PT17]. Strange [Bal93]. Straus [Vau04]. Stringer [App87]. Strong [Wol72]. Structural [HS15]. Structure [HS15, Hun99, Kau72, Lar71, Lev70, Roc83, Sla01, Zim84]. Structures [Hen19]. Strychnine [Sla01]. Student [Leh09]. Students [Bro13]. Studie [LE03, Mor06]. Studien [Ano04a]. Studies [AGJ08, Bro02, Col96, DM97, Dob71, Dob73, Dob74, Fic71, JR19, Pri16b]. Study [Bro11, Deb62b, Duc08, Fis18, Fri21b, Fri22, GC80, Jam83b, KF73, MT12, Maw18, Mer90, Mon37, She62, Sie08, Sme57, Sme87c]. Styles [Dun81]. Subatomic [Roc83]. Subject [Ano38d]. Sublunary [Via15]. Substances [Kle95, Sie82]. Success [Ano07e, Chu14, Chu13, Mau02]. Successful [Kau83]. Suffering [Sch22]. Sulfur [HJ02]. Sullivan [Ano06e]. Sulphur [New14, New15]. Supplementary [Sme78]. Supply [Cha81]. Support [Ano22b, Eli11]. Supporting [New86]. Supposed [Kar80]. Surgeon [Deb62a]. Surrey [Har00]. Survey [Ber92, Gor77, Kah10, Kah11, She60, Tho38a]. Survival [Jon18]. Susan [Ano97c]. Sverige [Lyk18]. Swallowfield [Jos49c]. Sweden [Lyk18]. Swedish [AF16, Lyk18, Sal19]. Swindler [Kau83]. Sydney [Mor06]. Sylvius [Rag12]. symbol [Hed17]. Symbolic [Gol57, Kah19, Kim96]. Symbolism [Dav87, Par37b, She58, She60]. Symbols [Bro86a, O’N21]. Symposium [Res23]. Synthesis [Ant19, Kah10, Kah11, KR14, Ram00, Rus87, Sie08, Sne79]. Synthetic [Bud18, DU79, Ras18, Tra02, Tra17]. System [Ban07, Bro16a, Fis96, Nev60, Sch94]. Systems [Ano06e, BP04, Kar03].

T. [Kop76]. Table

[Gor04, Kaj13, Kle95, Kra16a, Per73, Res23, Sie82, Ano06d, FCO15, Pul19]. Tables [Dun70, Gor18, Pul20]. Tactical [Tay14a, Tay14b]. Takes [Ful19]. Tales [NG16]. talets [Lyk18]. Tamoxifen [Qui20]. Tandridge [Har00]. Tanning [Lev57]. Tantalum [Hij22]. Tardy [Wei18]. Tarr [Tra05]. Tarrega [Pat88]. Tarrega-Marrano [Pat88]. Tasmanian [Hom17a]. Taste [Rag12]. Taught [KM90]. Taxotère(R) [Tom18]. Taylor [Hol56, Sim87]. Teacher [DLS86]. Teaching [And06, And15, Hed21, Iac22]. Tech [Ser76]. Technical [Fer09]. Technik [Hag16]. Technique [Hom19, Zha23]. Technological [Kel71, vL48]. Technologiekonzerns [Die04a, Mor06]. Technologist [Pag14]. Technology [Ano09a, Cra10, Die04a, Die04b, Hag16, Hen20, Joh11, Kea92, Lev57, MTKT^{+}15, Mor06, Tya04, Lyk20, Lyk22]. Technoscience [Hij21]. Telle [KN14]. Telling [PT17]. Tells [Cha17]. Temperature [Gol57]. Temporally [Amb22]. Tenebris [Kah19]. Teniers [Bri66, Dav87]. Tenth [Ric09]. Tenth-Century [Ric09]. Term [BJJK81]. Terms [Fer09, Mah76]. Tetranucleotide [Hun99]. Text [Bro72, Jos49a, KLPG06, Sta70, TCPYN59, Vin09, Tho63b]. Textbook [FFS76]. Textbooks [BSGBBV02, FM07]. Texts [Bur92, Deb68, HVD20, Hal09, MT16, Rob66, Tho59a, Tho59b, Tho60b]. Textual [Gru09, PLB23, Suc16]. Thallium [Jam81]. Theatre [Bee94]. Their [Bro20e, Dav87, Gol19, Sce22, Szu86, App77, Bra85, Gou83, Kle05]. Theme [She59]. Thenard [BBS10]. Theodore [App79, Bar19b, Gay97]. Theodoricus [Poo09]. Theological [Hir21]. Theology [Gan20a]. Theoretical [Sho03]. Theoretische [FM07]. Theories [LMN01, Nor07, Pri86, Roo01, Sim16a, Sha05]. Theory [Col78, Deb61, Don78, Duf15, Fis18, Gro21, Jam19b, BJ21, Kau73, Lar71, Lys37, Mau02, Mei83, New86, Nor06, Per69, Per19, Pra10, Rae16, Ree77b, Roc83, Sud80, Szy93, Szy96, Tay08, Wil98, PV17]. Theresa [Roc21]. Thermochemistry [Ant19]. Thermometry [Sum07]. Thesis [Lar67]. Thine [Nye14]. Thing [Ano06d, Gor04]. Things [BV16a, BSGBBV02, Bro19a, Mor21a, Num07, FH05]. Things-in-the-Making [Mor21a]. Thinking [GF23]. Third [Ano06a, dCH04, Ano12a, Joh18]. Thirteenth [Szu88, Abr79, Suc16]. Thirteenth-Century [Abr79, Suc16]. Thomas [Ano08a, Deb62b, Her18, Lev81, Cam77, Deb80, FFS73, FFS74a, GS61, Gui80, ML08, Mor72, New82, New20b, Pri79, Rei57, Roo01, Sea56, Sha02]. Thomsen [Ant17]. Thomson [Mor72, Sin87, Sin88a, Sin88b]. Thomsons [Sim14b]. those [Tho63b]. Thought [Bra85, Dun81, McE78a, McE78b, McE78c, Mce79, Pri16b, Son77, Tar18, War76b, LP05, Num07]. Thoughts [Wil72]. Thread [Clu05b]. Threat [Ano09a]. Three [Ham82, Ple59, VG04, Shi49]. Thrifty [Pas20]. Thurn [Mor22b]. Thurneisser [Mor22b]. Tick [BV16a]. Tides [Roo01]. Tighter [Ree12]. Timaeus [Wil88]. Time [Kin81, Kin82]. Times [Ano06d, Dmi04, Ebe07, GL17, Kin84, ZH05, Bil13, Hal11]. Tin [Rue22]. Titrimetry [Tra14]. Toh [BS22]. Toleration [Cha07]. Tollens [Gel78]. Tom [BS22]. Tomb [DN37]. Tome [Hal09]. Tool [Ede93, FPS16]. Tools [Gor18, NW18]. Torbern [Jam22]. Toronto [Van22]. Tower [Shi49]. Toxic [Wat17, GF23, GF23]. Toxicants [GL15]. tóxico [GF23]. toxicología [GL21]. Toxicology [BS14, Pér16, GL21]. Tóxicos [GF23, CG22]. Trace [Tra14]. Tracing [RGG23, Tay08]. Tracy [Ano06e]. Trade [Iri20, Ree23, Sim06, SMGL18]. Tradition [Deb81, Gru06, Gui80, HS20, Iac22, Kle14, Mar04, Pag60, Suc16, Ano75c]. Traditions [Maw16]. Traité [Run20, Bai00, Mck61b]. Trajectory [Kra21]. Transatlantic [Min16]. Transcendental [Roc87]. Transference [Szu86]. Transformation [Ede96, Fer16, Gru09, Tei82, Vrt23]. Transformations [Pom62]. Transformed [Kra22a]. Transit [AGCF^{+}15]. Transition [Lef18b, Ree20]. Translated [LK12, Van22]. Translating [Mar17]. Translation [Bai00, Bur99, Cop92, Del21, Jos64, Num07, Pri87, Sme59, Tay14a, Tay14b, Vic94]. Translations [Gor77, Sme77]. Transmutation [Cha07, Duf15, KH20, Kar92, Mcg67, Sin85, Tre74, Wal80]. Transmutational [Spa20]. Transmutations [Byc21]. Travel [Gol19]. través [LP05, Num07]. Travis [Mor06]. Treasure [AGJ08]. Treasures [Kir18]. Treatise [Ano06d, Dmi04, Jos49b]. Treatises [SAHL62, Wel51]. Tree [Hed13b, Szu86, Wal80]. Très [Duv48]. Trials [Iac22, Mur22]. Trinité [Duv48]. Trinity [Dun65, Tim06]. Trismegistus [Hal11, Ebe07]. Trithemius [Bra79, Par38b]. Triumph [Pri87]. Triumphant [Far89]. Trivalent [HJ02, Kau72]. Troubled [Sch22]. True [MC06, Sch20, Rus19]. Truth [Jos49b, Kno01, Las96, Sha94]. Tübingen [Fri59]. Tuck [BS22]. Turba [Hal09, Hed22, Ple59]. Turn [O’N21]. Turner [Bro86a]. Turpentine [Cam91, Kra16b]. Turton [Ano06e]. Twentieth [Bur16, DeK09, GW17, Hom05, Joh11, Le 15, Nye14, Qui16a, Rae16, Rei15, MR20, Sie08, Szu86, Wey76]. Twentieth-Century [Sie08, Hom05, Joh11, Rae16]. Twenty [Bro11]. Twenty-Five [Bro11]. Two [App79, Con22, Cro95, DP03, Jam22, Kar80, Kib60, Kim05, Min18, Nev60, Pri21a, Sme87c, SAHL62, WK16, Zot93, Ano06b]. Typographer [Clu17b]. Typographical [Mer17]. Tyrian [WK16].

u [Car17a]. u.q [Car17a]. UK [Tho22]. Ultimate [Web66]. Ultracentrifuge [Ede96]. Ultramarine [Mer04]. Umail [SLT49, SLT49]. Uncatalogued [Tho59b]. Uncertain [Uss03]. Uncertainties [Sán21]. Understanding [MT16, Ede93]. Uneasy [All03, Tra05]. Unifying [Ree77b]. Union [Bro05, Gol19, Pay89, Joh18]. United [Mil68, Mil70, Ree15, Tar02]. Unity [She62]. Universal [Kle15, Le 73, Old73, Roo17a]. universitaires [Run20]. Universities [Kim20]. University [Ano97c, Ano05c, Ano06c, Ano06a, Ano06e, Bar19b, Bro02, Bro97c, mKC21, CS19, Cob22, DLS86, Dob82a, Fre10, Fri21a, Fri21b, GS16, Hal11, Her18, Iac22, Kni04, Kra16a, Las96, Mor21a, Num07, Pul19, Rag22, Rat92a, Ree22, Roc03, Tra05, Van22, Vic94, vM90, Bro95, Fri21b, Gol16, Kik09, Kir03, Mac89, McV75, Mor69, Tya04]. Unknown [Big17, Bri66]. Unpublished [Jos49b, Wel82]. Unrecognised [Jen04]. Unrecognized [Mah76]. Unrecorded [McK58, Nev60]. Unsound [BJ21]. Unwise [BJ21]. Upon [Hay77, War76b, Par37b]. Urban [Whi18]. Urea [Bro68, Ram00]. Urinary [Col71]. Uroscopy [Dri16]. Ursula [Tay23]. Uruguay [Bro17d]. Use [Cha81, DeK09, Tar18, War97]. Usefulness [Sim20]. Uses [Col84, FM07]. Using [DeK09].

V [Ano07e, HH05]. Vagelos [Ano06e, VG04]. Valdecilla [LP05]. Valence [BJ21]. Valentine [Pri87]. Valerius [Pri86]. Valley [Mor16b]. Value [Hom17a]. Values [Pul20, RS04]. Vatican [Kib60]. Vaughan [Gui80, New82]. Vauquelin [Sou92]. ved [Nie16]. Vegetable [Lef18a]. Venable [McV75]. Venel [Leh09]. Ventures [Far71, Mor22b, Mor16a]. Verbraucherpolitik [Orl16]. Verdict [Bro68, DU79]. verdugo [GL21]. verdwenen [Kno22]. Verein [Joh18]. Verlag [Ano04a, Ano05a, Don88, Kim20, Mor06]. Verloren [Kno22]. Vermande [Kno22]. Vernon [Kin84]. Verse [Rob66]. Versus [Roc87, Rae16]. Verticity [Ree79]. Verville [Zin84]. Very [Pra10, Tom16a]. Vesuvius [Gue15]. VI [Geo57a]. Vicente [De 85]. Victoria [RT97]. Victorian [Bro97c, Kee13, Cob19, DLS86, NG16, Roc22, Rus96, SW14]. Vie [Run20]. Vietnam [CS19]. View [Ful19]. Views [DP03, Jen81, Lei75, Obr09, Ano06b]. vii [BS22]. viii [Ano05b, Sha05]. Villains [Tra12b]. Vilnius [Ano22b]. Vindicated [Lin77]. Violation [Mas91]. Virginia [Kra16a]. Virtues [Rya90]. Viscose [Mor18b]. Visible [Ant22]. Vision [The99]. Visions [Pri21a, Tay37c, Wer13]. Visiting [App79]. Visual [Fri22, Kah19, Sce22]. Vital [Bro68, Kni76, Sou00]. Vitalism [Mer79, Ram00]. Voarchadumia [Nor05]. Vocations [Riv18]. Voigts [Maw18]. Vol [Ano68a, Ano04a, Cha23, Duf23, HS23, New23, Roo23, Tay23]. volatil [Fra17]. volatile [Fra17]. Volcano [Gue15]. Volkhard [Kim20]. vols [Bro02, Bur02]. Voltaic [Sud80]. Volume [Ano07e, Mon38, Bur02, Cha23, Duf23, HS23, New23, Roo23, Tay23, Cra10]. volumes [Bur02]. Vortex [Sin87]. Vremeni [Dmi04, Ano06d]. vs [HS20, Kno06, Mor21a, Sha20].

W [Arm12, Ben70, For08, Gay97, Jac00, Kop76, Rat92a]. W. [KCU81]. Wächersbach [Bro16b]. Wagen [Pri87]. Walker [Edd01]. Wallach [Ano04a, BR00, BR00]. Walter [Shi49, Gel71]. War [Eli11, Fau11, Hor11, Kik11, Min16, Moy89, Ree15, Sla04, Bud20, Van04]. Waraqi [SLT49]. Wares [Cha81]. Warfare [GL18]. Warington [Bro22, Sim20]. Wars [Joh11]. was [Sme71]. Washington [Ano97c]. Wasser [Ano05a, BB04]. Waste [RA02, Tar02]. Water [BB04, Gou83, GDNN72, Ham17, Mar09, RA02, Sme68, Sne79, SM17, Tra02, Tra12a, Wal80, Web66]. Waters [Col84]. Watt [ML08, Tom16b]. Ways [PT17]. Weapon [Cam01, Sin59]. Weapon-Salve [Cam01]. website [Kim20]. Wedel [Pri21a]. Wedgwood [Cha81]. Well [Ano06d, Gor04]. Well-ordered [Ano06d, Gor04]. Wels [Kim20]. Werke [LE03, Mor06]. Werner [Kau72, Kau73]. Wernerian [New86]. West [Ree22]. Western [Kah19, Hed17, Hen17, She85a]. Westfall [Dob82a, All19b]. Wheat [Jos63]. Whispers [AGJ08]. White [Ver22, WW07]. Who [Bax13]. Wider [Ree12]. Wiesbaden [Kim20]. Wijck [New20b]. Wild [Pag62]. Wiley [Cob22]. Wilhelm [Bro16b, Jam22, Kir03, Tay19, Byc21, DFB22, Str79, Kra18]. Will [MC06, Bro22]. Willem [Clu17b, Mer17, Van17]. William [Gor83, Kik09, Kni16a, Sha05, Wat95, Bro86a, Chr94, Col02, FFS74a, FFS74b, FFS75, FFS76, FFS77, Fis96, Jam81, Jos49c, Kah21, Mil70, Pag14, Rea61, Rus01, Sch73, Tay14a, Tay14b, Uss15, WW07]. Williams [Jam81]. Williamson [Kik09, Pri86, Roc22]. Wine [SMGL18]. Winter [Ano84c]. Winthrop [Wil63a, Wil66]. Wisdom [Pol20]. Wissenschaft [Hag16]. Wissenschafts [Ano04a]. Wissenschafts- [Ano04a]. Witalij [Lau19]. Wizard [Shi49]. Wöhler [Bro68, Ram00]. Wohlwill [Sal13].

Wolfenbüttel [Mor22a]. Wolfgang

[Col76b]. Wollaston

[KCU81, Uss15, Kni16a]. Woman

[Mer22, Roc21]. Wombs [Dan06]. Women [Bro20e, Hor11, LR16, Lyk22, New20a, Ros96, SL22, Ser22]. Wonder [Mur22]. Wonderland [NG16]. Woodall [Deb62a]. Woodbridge [Tho22]. Woodward [Sla01]. Word [Dup23, Gue76, Ano97c]. Words [Con22]. Work [Abr79, Bro17b, Deb67, Dob79, Fil92, GC80, Gib60a, GLCK76, Her18, Kni74, McK38, Mer22, New20b, Pap90, Ree15, Ros96, Sch95, Sha03, She48, Štr16, Tay56]. Worker [Hei73]. Working [Poo09]. Workplaces [Sán21]. Works [Bur02, Kni86, Tay37a, Tay38, Tay19, HD00, VP18]. World [Ano07e, BS22, Iri20, Mor18a, Per10b, Rat66, Riv22, Ser22, Sim87, Sim22, Tei82, Tom16b, Van04, Via15, WW07, GF23, Dup23, Fau11, Hor11, Kik11, Moy89, Ree15, Sla04]. Wragge [Rag22]. Wragge-Morley [Rag22]. Write [HVD20]. Writing [Zin84]. Writings [Gib58, Gru06, Mar09, Obr09, Pri16d, Ram10, Wes61]. Written [Dup23]. Wurtz [Car93, Roc94].

X [Ano97c, Ano06b, Bur02, Ano07e, Bar19b, Don88, Num07, Ano06d, DeK09]. X-Ray [DeK09]. xi [Ano06e, Ano07e]. xii [Ano97c, Ano05c, Ano07a, Ano07e, Hal11, Iac22, Rat92a]. xiii [Cha23, Hal11, Tra05, Car17a]. Ximo [CG22]. xiv [Ano97c, Ano05c, Ano08a, HS23, Mor06, Riv18]. xix [Tra05, GL21]. XV [Ano68a, Ano07b, Bur02]. xvi [Ano06a, Ano07e]. XVIe [Pio18]. xvii [CS19, Tra05, Clu93]. XVII-Century [Clu93]. xviii [Dob82a]. XVIth [Pio18]. xx [Ano06d, Bro97c]. XXe [Pio18]. XXII [Jos49a]. XXIX.d.24 [Tho63a]. XXth [Pio18]. xxvi [Ano97c]. xxvii [Ano06c]. xxviii [Kni04]. xxxvii [Ano75c, Kra16a].

Yazid [Sal17]. Years [Ano85a, Bro11, Gol16, GL18, Res23, Roq17, Sme87a, Sme87b, Nie16]. Yego [Dmi04]. yes [Mor06]. Ymage [Run20]. York [Ano05c, Ano06d, Ano06e, Bro97c, Dob82a, Mor06, Pul19, Rat92a, Roo23, Vau04, Vic94]. Younger [Dav87]. Yugoslav [SC90].

Z [Rat92a]. Zealand [Rae95]. Zieglerin [Num01]. Zimmermann [Mor96]. Zinin [Bro95]. Ziolkowski [Bar19b]. Zosimos [Hal09, Tay37c]. Zosimus [Rin17]. zum [Bro20a, LE03, Mor22b, Mor06]. Zunbul [KF73]. Zur [Orl16, Ano04a].

References

Abrahams:1964:PAP

[Abr64] Harold J. Abrahams. Priestley answers the proponents of abiogenesis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 12(1):44–71, 1964. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1964.12.1.44.

Abrahams:1979:TCP

[Abr79]  Harold J. Abrahams.  A

Thirteenth-Century Portuguese work on manuscript illumination (“Libro De Como Se Facem as Cores”). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(2):93–99, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/

abs/10.1179/amb.1979.26.2.
93.

Abrahams:1984:AJF

[Abr84]  Harold J. Abrahams.  Al-

Jawbari on false alchemists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(2):84–88, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.2.
84.

Abraham:1994:SAD

[Abr94] Lyndy Abraham. The sources of Arthur Dee’s Fasciculus Chemicus (1631). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(3):135–141, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1994.41.3.
135.

Ahnfelt:2016:MEM

[AF16] Nils-Otto Ahnfelt and Hjalmar Fors. Making early modern medicine: Reproducing Swedish bitters. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):162–183, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1212886.

Alfonso-Goldfarb:2015:CKT

[AGCF^{+}15] Ana Maria Alfonso-Goldfarb, Hasok Chang, Marcia H. M. Ferraz, Jennifer M. Rampling, and Silvia Waisse. Chemical knowledge in transit. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(4):305–311, 2015. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2015.1129852.

Alfonso-Goldfarb:2008:LWM

[AGJ08] Ana Maria Alfonso-Goldfarb and Safa Abou Chahla Jubran. Listening to the whispers of matter through Arabic hermeticism: New studies on the Book of the Treasure of Alexander. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(2):99–121, 2008. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582308X255426.

Allchin:1992:PAO

[All92] Douglas Allchin. Phlogiston after oxygen. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):110–116, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.3.
110.

Allen:2003:UAC

[All03] Barbara L. Allen. Uneasy Alchemy: Citizens and Experts in Louisiana’s Chemical Corridor Disputes. Urban and industrial environments. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-01203-0 (hardcover), 0-262-51134-7 (paperback). xiii + 211 pp. LCCN GE235.L8 A38 2003.

Allen:2019:REK

[All19a] Katherine Allen. Recipes and everyday knowledge: Medicine, science, and the household in Early Modern England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):356–358, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1589035.

Allen:2019:RIN

[All19b] Meagan S. Allen. Revisiting Isaac Newton’s Index Chemicus: a response to Richard S. Westfall. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):72–81, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1573523.

See [Wes75].

Almgren:1998:DMS

[Alm98] Beverly S. Almgren. D. I.

Mendeleev and Siberia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):50–66, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.2.
50.

Ambrosio:2022:PFS

[Amb22] Chiara Ambrosio. Painting with fire: Sir Joshua Reynolds, photography, and the temporally evolving chemical object. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):85–87, 2022. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2021.1959738.

Anderson:2006:BBE

[And06] Robert G. W. Anderson. Boerhaave to Black: the evolution of chemistry teaching. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3):237– 254, November 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X148092.

Anderson:2010:CBA

[And10] Robert G. W. Anderson. Chemistry beyond the Academy: Diversity in Scotland in the Early Nineteenth Century. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 57(1):84–103, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12629173850041.

Anderson:2011:HHC

[And11] Robert G. W. Anderson. The history of the history of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(3):187–189, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X13129418298893.

Anderson:2012:PMP

[And12] Robert G. W. Anderson. Partington: The missing part. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(1):49–64, 2012. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1179/174582312X13296104891391.

Anderson:2015:TCP

[And15] Robert G. W. Anderson. Teaching the chemistry of platinum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(4):333–344, 2015. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2015.1113068.

Anderson:2016:FHC

[And16] Robert G. W. Anderson.

Fakes!? Hoaxes, counterfeits and deception in early modern science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):273–274, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1246806.

Anderson:2017:MFH

[And17] Robert G. W. Anderson. The matter factory: a history of the chemistry laboratory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):105–106, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1310434.

Anonymous:1937:Ra

[Ano37a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):84–87, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1937.1.1.

Anonymous:1937:Rb

[Ano37b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(2):140–141, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1937.1.2.
                                        140.

Anonymous:1938:AI

[Ano38a] Anonymous. Author index. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):201, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1938.1.3.
                                        201.

Anonymous:1938:Ra

[Ano38b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):50–51, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1938.2.1.
                                        50.

Anonymous:1938:Rb

[Ano38c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(2):112–113, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1938.2.2.
112.

Anonymous:1938:SI

[Ano38d] Anonymous. Subject index.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):202, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1938.1.3.
202.

Anonymous:1948:R

[Ano48] Anonymous. Review. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(1–2):64–67, 1948. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1948.3.1-
2.64.

Anonymous:1949:Ra

[Ano49a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(3–4):151–152, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.3.3-
4.151.

Anonymous:1949:Rb

[Ano49b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(1–2):79–80, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.4.1-
2.79.

Anonymous:1951:R

[Ano51] Anonymous. Review. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(3–4):159, 1951. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1951.4.3-
4.159.

Anonymous:1953:R

[Ano53] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(1–2):53–57, 1953. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1953.5.1-
2.53.

Anonymous:1956:R

[Ano56] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):129–130, 1956. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1956.5.3-
4.129.

Anonymous:1957:Ra

[Ano57a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(1):47–58, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.1.
47.

Anonymous:1957:Rb

[Ano57b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(2):109–111, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.2.
109.

Anonymous:1958:PR

[Ano58a] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3): 164, 1958. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1958.6.3.
164.

Anonymous:1958:R

[Ano58b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3):157–163, 1958. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1958.6.3.
                                        157.

Anonymous:1959:Ra

[Ano59a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(1):51–53, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1959.7.1.
                                        51.

Anonymous:1959:Rb

[Ano59b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(2):116–117, June 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1959.7.2.
                                        116.

Anonymous:1960:PR

[Ano60a] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(3): 182, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1960.8.3.
                                        182.

Anonymous:1960:Ra

[Ano60b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(1):46–51, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.1.
42.

Anonymous:1960:Rb

[Ano60c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(2):118–??, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.2.
118.

Anonymous:1960:Rc

[Ano60d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(3):177–181, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.3.
177.

Anonymous:1961:Ra

[Ano61a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(1):50–52, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.1.
50.

Anonymous:1961:Rb

[Ano61b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(2):115–116, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1961.9.2.
                                        115.

Anonymous:1961:Rc

[Ano61c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(3):168–170, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1961.9.3.
                                        168.

Anonymous:1962:E

[Ano62a] Anonymous. Editorial. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(3):??, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1962.10.3.
                                        i.

Anonymous:1962:Ra

[Ano62b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(1):37–39, 1962. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269862790222915.

Anonymous:1962:Rb

[Ano62c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(2):103–105, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1962.10.2.
103.

Anonymous:1962:Rc

[Ano62d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(3):139–149, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1962.10.3.
139.

Anonymous:1963:Ra

[Ano63a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(1):52–53, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1963.11.1.
52.

Anonymous:1963:Rb

[Ano63b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(2):97–100, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1080/00026980.1963.
11923898.

Anonymous:1963:Rc

[Ano63c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(3):158–160, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1963.11.3.
158.

Anonymous:1964:PR

[Ano64a] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 12(1): 82, 1964. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1964.12.1.
82.

Anonymous:1964:R

[Ano64b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 12(1):72–82, 1964. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1964.12.1.

Anonymous:1965:LE

[Ano65a] Anonymous. Letters to the Editor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(1):52–54, 1965. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1965.13.1.
52.

Anonymous:1965:R

[Ano65b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(1):55–60, 1965. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1965.13.1.
55.

Anonymous:1966:Ra

[Ano66a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):118–120, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2013.13.2.
007.

Anonymous:1966:Rb

[Ano66b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(3):189–192, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1966.13.3.
                                        189.

Anonymous:1967:LE

[Ano67a] Anonymous. Letters to the Editor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(1):60–62, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1967.14.1.
                                        60.

Anonymous:1967:PR

[Ano67b] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(1): 68, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1967.14.1.
                                        68.

Anonymous:1967:Ra

[Ano67c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(1):63–68, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1967.14.1.
                                        63.

Anonymous:1967:Rb

[Ano67d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(2):140–148, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.2.
140.

Anonymous:1967:Rc

[Ano67e] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(3):203–211, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.3.
203.

Anonymous:1968:CAV

[Ano68a] Anonymous. Correction to Ambix Vol. XV, Part 1. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(3):211, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1968.15.3.
211.

Anonymous:1968:LE

[Ano68b] Anonymous. Letters to the Editor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(2):115–124, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1968.15.2.
                                        115.

Anonymous:1968:PR

[Ano68c] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(1): 72, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.1.
                                        72.

Anonymous:1968:Ra

[Ano68d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(1):70–72, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.1.
                                        70.

Anonymous:1968:Rb

[Ano68e] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(2):125–128, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.2.
                                        125.

Anonymous:1968:Rc

[Ano68f] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(3):209–210, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1968.15.3.
209.

Anonymous:1969:R

[Ano69] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(1–2):109– 111, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.1-
2.109.

Anonymous:1970:PRa

[Ano70a] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1): 68, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.1.
68.

Anonymous:1970:PRb

[Ano70b] Anonymous. Publications received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(2): 136, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.2.
136.

Anonymous:1970:Ra

[Ano70c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1):58–66, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.1.
                                        58.

Anonymous:1970:Rb

[Ano70d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(2):127–135, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.2.
                                        127.

Anonymous:1970:SNBa

[Ano70e] Anonymous. Short notices of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1): 66, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.1.
                                        66.

Anonymous:1970:SNBb

[Ano70f]  Anonymous.  Short notices

of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(2):135–136, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.2.
135.

Anonymous:1971:BR

[Ano71a] Anonymous. Books received.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(2):147–148, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1971.18.2.
147.

Anonymous:1971:Ra

[Ano71b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(1):56–68, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1971.18.1.
56.

Anonymous:1971:Rb

[Ano71c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(2):139–146, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1971.18.2.
139.

Anonymous:1972:BR

[Ano72a] Anonymous. Books received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):228, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1972.19.3.
                                        228.

Anonymous:1972:GH

[Ano72b] Anonymous. Gerard Heym, 1888–1972. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):216–217, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1972.19.3.
                                        216.

Anonymous:1972:R

[Ano72c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):218–227, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1972.19.3.
                                        218.

Anonymous:1972:SNB

[Ano72d] Anonymous. Short notices of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3): 227, 1972. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.3.
227.

Anonymous:1973:BRa

[Ano73a] Anonymous. Books received.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1):75, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.1.
75.

Anonymous:1973:BRb

[Ano73b] Anonymous. Books received.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):142, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.2.
142.

Anonymous:1973:Ra

[Ano73c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1):67–74, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.1.
67.

Anonymous:1973:Rb

[Ano73d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):132–140, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.2.
                                        132.

Anonymous:1973:SNBa

[Ano73e] Anonymous. Short notices of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1): 74, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.1.
                                        74.

Anonymous:1973:SNBb

[Ano73f]  Anonymous.  Short notices

of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):140–141, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.2.
                                        140.

Anonymous:1974:BR

[Ano74a] Anonymous. Books received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):91, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1974.21.1.
91.

Anonymous:1974:R

[Ano74b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):86–89, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1974.21.1.
86.

Anonymous:1974:SNB

[Ano74c]  Anonymous.  Short notices

of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):89–90, 1974. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1974.21.1.89.

Anonymous:1975:BRa

[Ano75a] Anonymous. Books received.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):80, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.1.

Anonymous:1975:BRb

[Ano75b] Anonymous. Books received. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):160, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
160.

Anonymous:1975:ERC

[Ano75c] Anonymous. Essay review: Chinese Science. Explorations of an Ancient Tradition. Ed. by Shigeru Nakayama and Nathan Sivin. Pp. xxxvii + 334. Cambridge, Mass & London: M.I.T. Press. 1973. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):71–73, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.1.
71.

Anonymous:1975:Ra

[Ano75d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):74–79, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.1.
74.

Anonymous:1975:Rb

[Ano75e] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):156–158, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
156.

Anonymous:1975:SN

[Ano75f] Anonymous. Short notices.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):159, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
159.

Anonymous:1975:SNB

[Ano75g]  Anonymous.  Short notices

of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):79–80, 1975. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1975.22.1.79.

Anonymous:1977:I

[Ano77a] Anonymous. Index. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):186–188, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.3.
                                        186.

Anonymous:1977:R

[Ano77b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):175–184, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.3.
                                        175.

Anonymous:1977:SNB

[Ano77c]  Anonymous.  Short notices

of books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):184–186, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.3.
                                        184.

Anonymous:1980:R

[Ano80] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(3):210–214, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.3.
                                        004.

Anonymous:1982:EMH

[Ano82] Anonymous. European museums of the history of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(3): 207, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.3.
207.

Anonymous:1983:DA

[Ano83]  Anonymous.  The Dexter

awards. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(3): 162, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.3.
162.

Anonymous:1984:BPH

[Ano84a] Anonymous. Bibliotheca Philosophica Hermetica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(3):137, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.3.
137.

Anonymous:1984:PP

[Ano84b] Anonymous. The Partington Prize. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(1):

48, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.1.
48.

Anonymous:1984:SHA

[Ano84c] Anonymous. The Society for The History of Alchemy and Chemistry: Winter Meeting. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(2):88–89, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.2.
88.

Anonymous:1985:YCL

[Ano85a]  Anonymous.  300 years of

chemistry at Louvain — 1685– 1985. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(1): 45, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.1.
45.

Anonymous:1985:DP

[Ano85b] Anonymous. The Dexter Prize. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(3):148, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/

abs/10.1179/amb.1985.32.3.
148.

Anonymous:1985:PP

[Ano85c] Anonymous. The Partington Prize. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(2): 91, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.2.
91.

Anonymous:1986:Aa

[Ano86a] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(1):42, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.1.
42.

Anonymous:1986:Ab

[Ano86b] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):152, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.2-
3.152.

Anonymous:1987:A

[Ano87a] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):180, 1987.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.
180.

Anonymous:1987:DA

[Ano87b] Anonymous. The Dexter award. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(2):116, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.2.
116.

Anonymous:1988:A

[Ano88a] Anonymous. Announcements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(3):168, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.3.
168.

Anonymous:1988:PP

[Ano88b] Anonymous. The Partington Prize. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(2): 105, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.2.
105.

Anonymous:1989:Aa

[Ano89a] Anonymous. Announcement.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):40, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.1.
40.

Anonymous:1989:Ab

[Ano89b] Anonymous. Announcement.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(2):65, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.2.
65.

Anonymous:1989:JR

[Ano89c]  Anonymous.  Journals re-

ceived. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3): 160, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.3.
160.

Anonymous:1989:R

[Ano89d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3):153–159, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1989.36.3.
                                        153.

Anonymous:1989:SN

[Ano89e] Anonymous. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3):159–160, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1989.36.3.
                                        159.

Anonymous:1990:A

[Ano90a] Anonymous. Announcements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(3):147, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1990.37.3.
                                        147.

Anonymous:1990:ABR

[Ano90b] Anonymous. Announcements book reviews for Ambix. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(2):84, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1990.37.2.
                                        84.

Anonymous:1992:A

[Ano92] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):143, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.3.
143.

Anonymous:1993:A

[Ano93] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(1):26, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.1.
26.

Anonymous:1994:Ab

[Ano94a] Anonymous. Announcement.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(3):168, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1994.41.3.
168.

Anonymous:1994:Aa

[Ano94b] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(2):107, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.2.
                                        107.

Anonymous:1994:CRE

[Ano94c] Anonymous. Chairman’s remarks and Editor’s remarks. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):1–3, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.1.
                                        3.

Anonymous:1994:R

[Ano94d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):48–61, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.1.
                                        48.

Anonymous:1996:A

[Ano96] Anonymous. Announcements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(3):163, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1996.43.3.
                                        163.

Anonymous:1997:Ab

[Ano97a] Anonymous. Announcement.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(3):144, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.3.
144.

Anonymous:1997:Aa

[Ano97b] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(2):95, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.2.
95.

Anonymous:1997:ERJ

[Ano97c] Anonymous. Essay reviews:

“Just one word …, the historiography of plastics”. Early Plastics. Perspectives 1850– 1950. Edited by Susan Mossman. Pp. xii + 292, illus. Leicester University Press/Science Museum: London and Washington. 1997. £65. ISBN 0-7185-0020-2. American Plastic: a Cultural History. By Jeffrey I. Meikle. Pp. xiv + 403, illus. Rutgers University Press: New Brunswick, New Jersey. 1995. $49.95. ISBN 0-8135-2234-X (hardback). $24.00. ISBN 0-8135-2235-8 (paperback). From Small Organic Molecules to Large. A Century of Progress. By Herman F. Mark. (Profiles, Pathways and Dreams series). Pp. xxvi + 148. American Chemical Society: Washington, DC. 1993. $24.95. ISBN 0-8412-1776-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(3):145–148, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.3.
                                        145.

Anonymous:1998:NA

[Ano98a] Anonymous. News and announcements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(3): 190, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.3.
                                        190.

Anonymous:1998:SN

[Ano98b] Anonymous. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(1):44–48, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.1.
                                        44.

Anonymous:2000:Aa

[Ano00a] Anonymous. Announcement. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(2):117, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.2.
117.

Anonymous:2000:AP

[Ano00b] Anonymous. Announcement:

Prizewinners. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(1): 48, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.1.
48.

Anonymous:2000:Ab

[Ano00c] Anonymous. Announcements.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(3):119, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.3.
119.

Anonymous:2001:Aa

[Ano01a] Anonymous. Announcement.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(1):52, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.1.
52.

Anonymous:2001:Ab

[Ano01b] Anonymous. Announcements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):124, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.2.
                                        124.

Anonymous:2001:SN

[Ano01c] Anonymous. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):122–123, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.2.
                                        122.

Anonymous:2002:R

[Ano02] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(3):251–270, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2002.49.3.
                                        251.

Anonymous:2004:BRO

[Ano04a] Anonymous. Book reviews: Otto Wallach 1847–1931: Chemiker und Nobelpreisträger: Lebenserinnerungen: Potsdam, Berlin, Bonn, Göttingen = (Studien und Quellen zur Geschichte der Chemie, Vol. 12). Edited by Günther Beer and Horst Remane. Pp. 270, illus., index. Verlag für Wissenschafts- und Regionalgeschichte Dr. Michael Engel: Berlin–Charlottenburg. 2000. EUR 35. ISBN 3-929134-34-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):187–191, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.2.
187.

Anonymous:2004:NC

[Ano04b] Anonymous. Notes on contributors. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):193–194, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.2.
193.

Anonymous:2004:SN

[Ano04c] Anonymous. Short notices.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):192, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1080/00026980.2004.
11917467.

Anonymous:2005:BRF

[Ano05a]  Anonymous.  Book review:

Feuer, Wasser, Erde, Luft: eine Kulturgeschichte der Elemente. By Gernot Böhme and Hartmut Böhme. Pp. 344, illus., index. Verlag C. H. Beck: München. 2004. EUR 16.90 (pbk). ISBN 3-406-51067-1 (pbk). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):285–295, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X77808.

Anonymous:2005:BRA

[Ano05b] Anonymous. Book review: The Art of All Colours: Medieval Recipe Books for Painters and Illuminators. By Mark Clarke. Pp. viii + 152. Archetype Publications: London. 2001. £24.50; $37.50. ISBN 1-873132-72-7. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):89–101, 2005. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/000269805X52931.

Anonymous:2005:BRS

[Ano05c] Anonymous. Book reviews: Science in Europe, 1500–1800: a Primary Sources Reader. Edited by Malcolm Oster. Pp. xiv + 282, illus., index. Palgrave/The Open University: Basingstoke, New York. 2002. £17.50. ISBN 0-333-97001-2. Science in Europe, 1500–1800: a Secondary Sources Reader.

Edited by Malcolm Oster. Pp. xii + 307, index. Palgrave/The Open University: Basingstoke, New York. 2002. £17.50. ISBN 0-333-97005-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52 (2):179–187, July 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X70139.

Anonymous:2005:CNE

[Ano05d] Anonymous. Call for nominations for the 2006 Edelstein award. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(2):190, July 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X70300.

Anonymous:2005:NCa

[Ano05e] Anonymous. Notes on contributors. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):102–104, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X52940.

Anonymous:2005:NCb

[Ano05f] Anonymous. Notes on contributors. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(2):189, July 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269805X70148.

Anonymous:2005:NCc

[Ano05g] Anonymous. Notes on contributors. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):296, 2005. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/000269805X77817.

Anonymous:2006:BRM

[Ano06a]  Anonymous.  Book review:

Models: the Third Dimension of Science. Edited by Soraya de Chadarevian and Nick Hopwood. Pp. xvi + 464, illus., index. Stanford University Press: Stanford, California. 2004. £17.50. ISBN 0-8047-3972-2 [pbk]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3): 277–298, November 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582306X148858.

Anonymous:2006:BRN

[Ano06b]  Anonymous.  Book review:

Newton’s darkness: two dramatic views. By Carl Djerassi and David Pinner. Pp. 184. Imperial College Press: London. 2003. £18, $24 (hbk); £11, $15 (pbk). ISBN 1-86094-389-6 (hbk), 1-86094-390-X (pbk). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):87–96, 2006. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582306X93228.

Anonymous:2006:BRC

[Ano06c]  Anonymous.  Book review:

The Cambridge History of Science: The Modern Physical and Mathematical Sciences. Edited by Mary Jo Nye. Pp. xxvii + 678, illus., index. Cambridge University Press: Cambridge. 2003. £100. ISBN 0-521-57199-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53 (2):173–189, July 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X117924.

Anonymous:2006:ERC

[Ano06d] Anonymous. Essay review: Chelovek epokh peremen: Ocherki o D. I. Mendeleeve i ego vremeni [A Man in a Changing Epoch: a Treatise on D. I. Mendeleev and his Times]. By Igor Sergeevich Dmitriev. Pp. 576, illus. Khimizdat: St. Petersburg, 2004. ISBN 5-93808-082-7. A Well-ordered Thing: Dmitrii Mendeleev and the Shadow of the Periodic Table. By Michael D. Gordin. Pp. xx + 364, illus. Basic Books: New York. 2004. £11.50. ISBN 0-465-02775-X. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3): 273–276, November 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X148849.

Anonymous:2006:ERP

[Ano06e] Anonymous. Essay review: The Pharmaceutical Industry: a Guide to Historical Records. Edited by Lesley Richmond, Julie Stevenson, and Alison Turton. Pp. ix + 561, index. Ashgate: Aldershot. 2003. £55. ISBN 0-7546-3352-7. Science, Medicine, and Merck. By P. Roy Vagelos and Louis Galambos. Pp. xi + 301, illus., index. Cambridge University Press: Cambridge and New York. 2004. £25; $30. ISBN 0-521-66295-8 (hbk). Pharmaceutical Achievers: the Human Face of Pharmaceutical Research. By Mary Ellen Bowden, Amy Beth Crow, and Tracy Sullivan. Pp. xi + 220, illus., index. Chemical Heritage Press: Philadelphia. 2003. $30. ISBN 0-941901-30-0 (pbk.). Senses, Sensors and Systems: a Journey Through the History of Laboratory Diagnosis. Edited by Sabine Päuser. Pp. 326, illus., index. Editiones Roche: Basel. 2004. ISBN 3-907770-89-7. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53 (2):167–171, July 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X117915.

Anonymous:2006:SNa

[Ano06f] Anonymous. Short notices.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(2):189–192, July 2006. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2006.53.2.
189.

Anonymous:2006:SNb

[Ano06g] Anonymous. Short notices.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3):299, November 2006. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2006.53.3.
299.

Anonymous:2007:BRC

[Ano07a]  Anonymous.  Book review:

Chemical History: Reviews of the Recent Literature. Edited by Colin Archibald Russell and Gerrylynn K. Roberts. Pp. xii + 247, index. RSC Publishing: Cambridge. 2005. £59.95. ISBN 0-85404-464-7. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):217–236, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X239478.

Anonymous:2007:BRR

[Ano07b]  Anonymous.  Book review:

Reworking the Bench: Research Notebooks in the History of Science. Edited by Frederic Lawrence Holmes, Jürgen Renn, and Hans-Jörg Rheinberger. Pp. xv + 325, illus. Kluwer Academic Publishers: Dordrecht, Boston, London. 2003. EUR 127; £81; $119. ISBN 1-4020-1039-7. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):89–112, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X169908.

Anonymous:2007:CNEa

[Ano07c] Anonymous. Call for nominations for the 2008 Edelstein award. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):212, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X219308.

Anonymous:2007:CNEb

[Ano07d] Anonymous. Call for nominations for the 2008 Edelstein award. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):304, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582313X13789813148904.

Anonymous:2007:ERR

[Ano07e] Anonymous. Essay review:

The Road to Scientific Success: Inspiring Life Stories of Prominent Researchers. Volume 1. Edited by Deborah D. L. Chung. Pp. xi + 230, illus. index. World Scientific: Singapore and London. 2006. £33 (hbk); £17 (pbk). ISBN 981-256-600-7 (hbk); 981-256-466-7 (pbk). Nature Not Mocked: Places, People and Science. By Peter Day. Pp. x + 262, illus., index. Imperial College Press: London. 2005. $48. ISBN 1-86094-576-7. Candid Science: Conversations with Famous Chemists. By István Hargittai and edited by Magdolna Hargittai. Pp. xii + 516, illus., index. Imperial College Press: London. 2000. £48 (hbk); £21 (pbk). ISBN 1-86094-151-6 (hbk), 1-86094-228-8 (pbk). Candid Science V: Conversations with Famous Scientists. By Balazs Hargittai and István Hargittai. Pp. xvi + 695, illus., index. Imperial College Press: London. 2005. £34. ISBN 1-86094-506-6 (pbk). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):213–216, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X227183.

Anonymous:2007:PPa

[Ano07f] Anonymous. The Partington Prize 2008. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):211, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X219290.

Anonymous:2007:PPb

[Ano07g] Anonymous. The Partington Prize 2008. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):303, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582313X13789813148986.

Anonymous:2007:SNa

[Ano07h] Anonymous. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):109–112, 2007. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2007.54.1.

Anonymous:2007:SNb

[Ano07i] Anonymous. Short notices.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):236–239, 2007. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2007.54.2.
220.

Anonymous:2008:BRH

[Ano08a] Anonymous. Book review: The History and Poetics of Scientific Biography. Edited by Thomas Söderqvist. Pp. xiv + 270, illus., index. Ashgate: Aldershot. 2007. £55. ISBN 978-0-7546-5181-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):293–308, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174963208X371083.

Anonymous:2008:Ra

[Ano08b] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 55(1):83–96, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X255433.

Anonymous:2008:Rb

[Ano08c] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 55(2):169–187, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X255497.

Anonymous:2009:JUT

[Ano09a] Anonymous. Journals under threat: a joint response from history of science, technology and medicine Editors. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):1–3, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X405183.

Anonymous:2009:Ra

[Ano09b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):89–92, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X411681.

Anonymous:2009:Rb

[Ano09c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(2):193–199, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X441426.

Anonymous:2009:Rc

[Ano09d] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 56(3):273–276, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269809X12529013496455.

Anonymous:2010:Ra

[Ano10a] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 57(1):104–122, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582310X12629173850087.

Anonymous:2010:Rb

[Ano10b] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 57(2):233–247, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582310X12758843013935.

Anonymous:2010:Rc

[Ano10c] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 57(3):311–323, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582313X13789813148788.

Anonymous:2010:SN

[Ano10d] Anonymous. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(1):120–122, 2010. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2010.57.1.
                                        120.

Anonymous:2011:Ra

[Ano11a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(1):82–95, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12849808295823.

Anonymous:2011:Rb

[Ano11b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(2):171–184, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X13008456751189.

Anonymous:2011:Rc

[Ano11c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(3):277–291, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582311X13129418299108.

Anonymous:2011:SN

[Ano11d] Anonymous. Short notices.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(2):184–185, 2011. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2011.58.2.
184.

Anonymous:2012:MGN

[Ano12a] Anonymous. From Museumsinsel to the Grüneburg: New approaches in the history of the Third Reich. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(2):170–173, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13345259996092.

Anonymous:2012:Ra

[Ano12b] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 59(1):65–79, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13296104891517.

Anonymous:2012:Rb

[Ano12c] Anonymous. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 59(2):174–185, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13352956406841.

Anonymous:2012:Rc

[Ano12d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(3):294–307, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13457672281902.

Anonymous:2013:Rb

[Ano13a] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(1):78–93, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698012Z.
00000000022.

Anonymous:2013:Rc

[Ano13b] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):182–201, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000030.

Anonymous:2013:Rd

[Ano13c] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(3):289–308, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000035.

Anonymous:2013:Re

[Ano13d] Anonymous. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(4):415–428, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000041.

Anonymous:2016:PPa

[Ano16a] Anonymous. The Partington Prize 2017. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):84, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1204085.

Anonymous:2016:PPb

[Ano16b] Anonymous. The Partington Prize 2017. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):184, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1218605.

Anonymous:2016:PPc

[Ano16c] Anonymous. The Partington Prize 2017. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):283, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1257274.

Anonymous:2016:SHA

[Ano16d] Anonymous. Society for the History of Alchemy and Chemistry Award Scheme 2017. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):362, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1283131.

Anonymous:2017:MAC

[Ano17a]  Anonymous.  2018 Morris

award: Call for nominations. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):201, 2017. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1346918.

Anonymous:2017:SHAb

[Ano17b]  Anonymous.  Society for

the History of Alchemy and Chemistry Award Scheme 2018. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):390–391, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1429138.

Anonymous:2017:SHAa

[Ano17c] Anonymous. Society for the History of Alchemy and Chemistry: E-mail communications with members. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):202, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1338519.

Anonymous:2019:ASH

[Ano19a] Anonymous. Announcement:

Society for the History of Alchemy and Chemistry Award Scheme 2019. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):93–94, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1550037.

Anonymous:2019:PP

[Ano19b] Anonymous. The Partington Prize 2020. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):264, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1620482.

Anonymous:2019:SHA

[Ano19c]  Anonymous.  Society for

the History of Alchemy and Chemistry Award Scheme 2020. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):370–371, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1682820.

Anonymous:2020:MAC

[Ano20a]  Anonymous.  2021 Morris

Award: Call for nominations. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):421–422, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1822667.

Anonymous:2020:PP

[Ano20b] Anonymous. Partington Prize 2020. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):322, 2020. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2020.1763694.

Anonymous:2021:MA

[Ano21a] Anonymous. Morris Award 2021. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68 (2–3):329–330, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1936751.

Anonymous:2021:SHA

[Ano21b]  Anonymous.  Society for

the History of Alchemy and Chemistry Award Scheme 2021. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(1):116–117, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1890397.

Anonymous:2022:PP

[Ano22a] Anonymous. The Partington Prize 2023. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):201, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2054608.

Anonymous:2022:SSI

[Ano22b]  Anonymous.  SHAC spe-

cial ICHC13 award scheme — grants to support attendance at ICHC13 in Vilnius, May 2023. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):434–435, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2134615.

Anonymous:2022:SHA

[Ano22c] Anonymous. Society for the History of Alchemy and Chemistry Award Scheme 2022. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):93–94, 2022. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2020506.

Anonymous:2023:PP

[Ano23a] Anonymous. The Partington Prize 2023. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):205, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2023.2197328.

Anonymous:2023:SHA

[Ano23b]  Anonymous.  Society for

the History of Alchemy and Chemistry Award Scheme 2023. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):106–107, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2180223.

Antzoulatos:2017:JTL

[Ant17] Vangelis Antzoulatos. Julius Thomsen. A life in chemistry and beyond. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):195–196, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1346955.

Antzoulatos:2019:BPS

[Ant19] Vangelis Antzoulatos. Berthelot’s pathway from synthesis to thermochemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):51–71, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1537326.

Antonelli:2022:BVM

[Ant22] Francesca Antonelli. Becoming visible. Marie-Anne Paulze-Lavoisier and the campaign for the “New Chemistry” (1770s– 1790s). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):221–242, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2091352.

Appleby:1977:ADJ

[App77] John H. Appleby. Arthur Dee and Johannes Bánfi Hunyades: Further information on their alchemical and professional activities. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(2):96–109, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.2.
                                        96.

Appleby:1979:SAD

[App79] John H. Appleby. Some of Arthur Dee’s associations before visiting Russia clarified, including two letters from Sir Theodore Mayerne. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(1):1–15, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1979.26.1.
                                        1.

Appleby:1987:MSF

[App87]  John H. Appleby.  Moses

Stringer (fl. 1695–1713): Iatrochemist and mineral master general. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(1):31–45, 1987. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1987.34.1.31.

Armon:2012:PCB

[Arm12] Rony Armon. From pathology to chemistry and back: James W. Cook and early chemical carcinogenesis research. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(2):152–169, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13345259996056.

Averley:1986:SCE

[Ave86] Gwen Averley. The “social chemists”: English chemical societies in the Eighteenth and Early Nineteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):99–128, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.2-
3.99.

Baguena:2016:BRJ

[Bág16] María-José Báguena. Book review: Jonas Salk. A Life. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):196–197, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1080/00026980.2016.
                                        1227179.

Baird:2000:NDP

[Bai00] John D. Baird. Note on the date of publication of the English translation of Lavoisier’s Traité Élémentaire De Chymie. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(1):47–48, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.1.
                                        47.

Baldwin:1993:ASJ

[Bal93] Martha Baldwin. Alchemy and the Society of Jesus in the Seventeenth Century: Strange bedfellows? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(2):41–64, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.2.
                                        41.

Ball:2002:BEA

[Bal02] Philip Ball. Bright Earth: Art and the Invention of Color. Farrar, Straus and Giroux, New York, NY, USA, 2002. ISBN 0-374-11679-2 (hardcover). ix + 382 + 24 pp. LCCN N7432.7 .B35 2002. URL http:

//www.loc.gov/catdir/bios/
hol051/2001041820.html;
http://www.loc.gov/catdir/
description/hol021/2001041820.
html.

Bandinelli:2007:ISQ

[Ban07] Angela Bandinelli. The isolated system of quantifiable experiences in the 1783 “Mémoire sur la chaleur” of Lavoisier and Laplace. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):274–284, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582307X237038.

Barker:2019:BRF

[Bar19a]  Sheila Barker.  Book re-

view: La farmacia granducale di Firenze, By Giovanni Piccardi. Pp. 131, illus., index. Leo S. Olschki: Florence. 2018. £17.00. ISBN 978-8-82-226566-

1. (Italian) [The grand-ducal pharmacy of Florence]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):368–369, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1611018.

Barth:2019:BRA

[Bar19b] Howard G. Barth. Book review:

The Alchemist in Literature: From Dante to the Present, by Theodore Ziolkowski. Pp. 237 + x, illus., index. Oxford University Press: Oxford. 2015. £74.00. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):364–365, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1592805.

Baxfield:2013:WAW

[Bax13] Christopher Baxfield. “Who is the Almighty that we should serve Him?” Chaos, Providence and natural philosophy in Stephen Hales. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(1):31–53, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698012Z.
00000000018.

Bayer:2005:LMC

[Bay05]  Penny Bayer.  Lady Mar-

garet Clifford’s alchemical receipt book and the John Dee Circle. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):271–284, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X77754.

Bohme:2004:FWE

[BB04] Gernot Böhme and Hartmut Böhme. Feuer, Wasser, Erde, Luft: eine Kulturgeschichte der Elemente. (German) [Fire, Water, Earth, Air: a Cultural History of the Elements], volume 1565 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 2004. ISBN 3-406-51067-1. 344 pp. LCCN BD851 .B575 2004. URL http://d-

nb.info/969614977/04.

Belmar:2010:LJT

[BBS10] Antonio García Belmar and José Ramón Bertomeu-Sánchez. Louis Jacques Thenard’s chemistry courses at the Collège de France, 1804–1835. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(1):48–63, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582310X12629173849962.

Bowden:2003:PAH

[BCS03] Mary Ellen Bowden, Amy Beth Crow, and Tracy Sullivan. Pharmaceutical Achievers: the Human Face of Pharmaceutical Research. Chemical Heritage Press, Philadelphia, PA, USA, 2003. ISBN 0-941901-30-0 (paperback). xi + 220 pp. LCCN RM301.25 .B69

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0630/2001055313-d.html.

Beer:1994:SRO

[Bee94] Gillian Beer. Square Rounds and other awkward fits: Chemistry as theatre. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):33–41, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1994.41.1.
33.

Bentley:1970:CDR

[Ben70] Jonathan Bentley. The Chemical Department of The Royal School of Mines. Its origins and development under A. W. Hofmann. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(3):153–181, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.3.
153.

Bentley:1972:HRG

[Ben72] Jonathan Bentley. Hofmann’s return to Germany from the Royal College of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):197–203, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.3.
197.

Beretta:1992:HCE

[Ber92] Marco Beretta. The historiography of chemistry in the Eighteenth Century: a preliminary survey and bibliography. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(1):1–10, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.1.
1.

Beretta:2011:CRH

[Ber11] Marco Beretta. The changing role of the historiography of chemistry in Continental Europe since 1800. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(3):257–276, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582311X13129418299027.

BertomeuSanchez:2000:MOE

[BG00] José Ramón Bertomeu Sánchez and Antonio García Belmar. Mateu Orfila’s Elémens de Chimie Médicale and the debate about the medical applications of chemistry in Early Nineteenth-Century France. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(1):1–28, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.1.
                                        1.

Buchanan:1978:EHS

[BGH78] Peta Dewar Buchanan, J. F. Gibson, and Marie Boas Hall. Experimental history of science: Boyle’s colour changes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(3):208–210, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1978.25.3.
                                        208.

Bigotti:2017:PUP

[Big17] Fabrizio Bigotti. A previously unknown path to corpuscularism in the Seventeenth Century: Santorio’s Marginalia to the Commentaria in Primam Fen Primi Libri Canonis Avicennae (1625). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):29–42, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1287550.

Bilak:2013:AET

[Bil13] Donna Bilak. Alchemy and the End Times: Revelations from the laboratory and library of John Allin, Puritan alchemist (1623–1683). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(4):390–414, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000040.

Bilak:2023:LTN

[Bil23] Donna Bilak. Living then and now with gold and mercury. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):1–6, 2023. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2023.2192590.

Jewess:2021:URU

[BJ21] William H. Brock and Michael Jewess. Unwise relationships and an unsound valence theory: The chemical career of Robert Fergus Hunter (1904– 1963). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):407–430, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1984623.

Brock:1981:ODT

[BJJK81] William H. Brock, K. A.

Jensen, Christian Klixbüll Jørgensen, and George B. Kauffman. The origin and dissemination of the term “ligand” in chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):171–183, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1981.28.3.
                                        171.

Becker:2015:R

[BKS^{+}15] Daniel Becker, Yoshiyuki Kikuchi, Anna Simmons, Helge Kragh, and Elena Serrano. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(2):193–201, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698015Z.
                                        00000000080.

Butler:1980:ECE

[BN80] Anthony R. Butler and Joseph Needham. An experimental comparison of the East Asian, Hellenistic, and Indian (Gandharan) stills in relation to the distillation of ethanol and acetic acid. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(2):69–76, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.2.
                                        69.

Bohm:1963:JMH

[Böh63] Walter Böhm. John Mayow and his contemporaries. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(3):105–120, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1963.11.3.
105.

Bolzan:1976:CCA

[Bol76] J. E. Bolzan. Chemical combination according to Aristotle. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):134–144, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.3.
134.

Bortolotto:2015:JAC

[Bor15] Andréa Bortolotto. Johann Andreas Cramer and chemical mineral assay in the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(4):312–332, 2015. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2015.1129854.

Bourke:2022:LRI

[Bou22] Evan Bourke. Lady Ranelagh:

The incomparable life of Robert Boyle’s sister. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):190–191, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2026605.

Bowles:1975:JHP

[Bow75] Geoffrey Bowles. John Harris and the powers of matter. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):21–38, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1975.22.1.
                                        21.

Buttner:2004:SSS

[BP04] Johannes Büttner and Sabine Päuser, editors. Senses, Sensors and Systems: a Journey Through the History of Laboratory Diagnosis. Editiones Roche, Verlag der F. Hoffmann-La Roche AG, Basel, Switzerland, 2004. ISBN 3-907770-89-7 (paperback). 326 pp. LCCN ???? URL http:

                                        //www.gbv.de/dms/hbz/toc/
                                        ht014127439.pdf.

Beer:2000:OWC

[BR00] Günther Beer and Horst Remane. Otto Wallach 1847– 1931: Chemiker und Nobelpreisträger: Lebenserinnerungen: Potsdam, Berlin, Bonn, Göttingen. (German) [Otto Wallach 1847–1931: Chemist and Nobel Laureate: Memoirs: Potsdam, Berlin, Bonn, Göttingen], volume 12 of Studien und Quellen zur Geschichte der Chemie. Verlag für Wissenschafts- und Regionalgeschichte Dr. Michael Engel, Berlin, Germany, 2000. ISBN 3-929134-34-9 (paperback). ISSN 0941-8059. 270 pp. LCCN QD22.W35 O88 2000.

Brann:1979:GRA

[Bra79] Noel L. Brann. George Ripley and the Abbot Trithemius: an inquiry into contrasting medical attitudes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):212–220, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.3.
212.

Brann:1985:AMM

[Bra85]  Noel L. Brann.  Alchemy

and melancholy in Medieval and Renaissance thought: a query into the mystical basis of their relationship. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(3):127–148, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.3.
127.

Brehm:1976:RBP

[Bre76] Edmund Brehm. Roger Bacon’s place in the history of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(1):53–58, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/000269876790223568.

Brinkman:1966:UAD

[Bri66] A. A. A. M. Brinkman. An unknown alchemical drawing probably by David Teniers II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(3):187–188, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1966.13.3.
                                        187.

Browne:1948:RRA

[Bro48] C. A. Browne. Rhetorical and religious aspects of Greek alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(1–2):15–25, 1948. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1948.3.1-2.15.

Brock:1967:LCS

[Bro67] W. H. Brock. The London Chemical Society 1824. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(2):133–139, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.1967.14.2.
133.

Brooke:1968:WUV

[Bro68] John H. Brooke. Wöhler’s urea, and its vital force? — a verdict from the chemists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(2):84–114, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1968.15.2.
84.

Brock:1969:LCF

[Bro69] W. H. Brock. Lockyer and the chemists: The first dissociation hypothesis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(1–2):81–99, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.1-
2.81.

Brock:1972:TLL

[Bro72] W. H. Brock. Text: Liebig’s laboratory accounts. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(1):47–58, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.1.
47.

Brock:1980:R

[Bro80] W. H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(1):59–68, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.1.
                                        59.

Brock:1981:R

[Bro81] W. H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(2):108–120, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1981.28.2.
                                        108.

Brock:1984:R

[Bro84] W. H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(2):90–96, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1984.31.2.
                                        90.

Brock:1986:BAC

[Bro86a] W. H. Brock. The British Association Committee on Chemical Symbols 1834: Edward Turner’s letter to British chemists and a reply by William Prout. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(1):33–42, 1986. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1986.33.1.33.

Brock:1986:R

[Bro86b] W. H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(1):43–52, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.1.
43.

Brooke:1987:MMD

[Bro87] John Hedley Brooke. Methods and methodology in the development of organic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):147–155, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.
147.

Brock:1990:R

[Bro90] W. H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(3):148–152, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1990.37.3.
                                        148.

Brooks:1995:NZK

[Bro95] Nathan M. Brooks. Nikolai Zinin at Kazan University. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(3):129–142, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1995.42.3.
                                        129.

Brock:1997:A

[Bro97a] W. H. Brock. In appreciation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(1): 39, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.1.
                                        39.

Brock:1997:SN

[Bro97b]  W. H. Brock.  Short no-

tices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(3): 164, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.3.
                                        164.

Brooke:1997:ERE

[Bro97c]  John Hedley Brooke.  Es-

say review: Edward Frankland. Chemistry, Controversy and Conspiracy in Victorian England. By Colin A. Russell. Pp. xx + 535, illus. Cambridge University Press:Cambridge and New York. 1996. £65. ISBN 0-521-49636-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(2):96–112, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.2.
96.

Brooks:1997:PLC

[Bro97d]  Nathan M. Brooks.  Pub-

lic lectures in chemistry in Russia: 1750–1870. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(1):1–10, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.1.
1.

Brooks:1998:MM

[Bro98] Nathan M. Brooks. Mendeleev and metrology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):116–128, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.2.

Brock:2000:EDC

[Bro00a] W. H. Brock. Ernst Dieffenbach’s comments on the state of chemistry in Britain in 1846. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(3):121–134, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.3.
                                        121.

Brock:2000:R

[Bro00b] W. H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(1):49–66, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.1.
                                        49.

Brock:2002:ERP

[Bro02] William H. Brock. Essay review: The philosophy of chemistry. Foundations of Chemistry. Philosophical, Historical and Interdisciplinary Studies of Chemistry. Edited by Eric R. Scerri, vols. 1 (1991) and 2 (2000). Kluwer Academic Publishers, P. O. Box 322, 3300 AA Dordrecht, The Netherlands. ISSN 1386-4328. Hyle. International Journal of Philosophy of Chemistry. Edited by Joachim Schummer, vols 3 (1997), 4 (1998), 5 (1999), and 6 (2000). Hyle Publications Inc., Institute of Philosophy, University of Karlsruhe, D-76128, Karlsruhe, Germany. ISSN 1433-5158. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(1):67–71, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.1.
67.

Brock:2003:BCG

[Bro03a] William H. Brock. Breeding chemists in Giessen. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(1):25–70, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.1.
25.

Brock:2003:E

[Bro03b] William H. Brock. Editorial.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(1):1–2, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.1.
1.

Brooks:2005:GLB

[Bro05] Nathan M. Brooks. Growing links between chemistry and industry in Russia and the Soviet union, 1900–1953. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):27–43, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269805X52896.

Brock:2008:SFM

[Bro08] William H. Brock. Stephen Finney Mason (1923–2008). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(2):97–98, 2008. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/174582313X13789813148823.

Brock:2011:EEM

[Bro11] W. H. Brock. Exploring early modern chymistry: The first twenty-five years of the Society for the Study of Alchemy & Early Modern Chemistry 1935–1960. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(3):191–214, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X13129418298947.

Brock:2013:BBS

[Bro13] W. H. Brock. Bunsen’s British students. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(3):203–233, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000032.

Brock:2016:ERP

[Bro16a] William H. Brock. Early responses to the periodic system. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):351–352, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279410.

Brock:2016:FAG

[Bro16b] William H. Brock. Friedrich August Genth aus Wächersbach und die Entdeckung der ersten NiO-Kristalle am Marburger chemischen Institut unter Robert Wilhelm Bunsen. (German) [Friedrich August Genth from wächersbach and the discovery of the first NiO crystals at the Marburg Chemical Institute under Robert Wilhelm Bunsen]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):359, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279419.

Brock:2016:OLE

[Bro16c] William H. Brock. Otto Linné Erdmann an Justus von Liebig — kommentierte Briefe von 1853 bis 1867. (German) [Otto Linné Erdmann to Justus von Liebig — annotated letters from 1853 to 1867]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):193–194, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1227175.

Brock:2017:BSC

[Bro17a] W. H. Brock. British school chemistry laboratories, 1830– 1920. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):43–65, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1285853.

Brock:2017:BRH

[Bro17b] William H. Brock. Book review: Holding Hands with Bacteria. The Life and Work of Marjory Stephenson. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):291–292, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1385902.

Brock:2017:CDI

[Bro17c] William H. Brock. The chemistry department at Imperial College London. a history, 1845–2000. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):101–102, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1310427.

Brock:2017:LCI

[Bro17d] William H. Brock. LEMCO. un coloso de la industria cárnica en Fray Bentos, Uruguay. (the meat industry’s colossus in Fray Bentos, Uruguay). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):293, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385905.

Brock:2019:BRM

[Bro19a] William H. Brock. Book review:

The Measure of All Things. A History of Analytical Chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):82, 2019. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2018.1534645.

Brock:2019:DMK

[Bro19b] William H. Brock. David Marcus Knight (1936–2018): an appreciation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):246–263, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1621499.

Brock:2019:GP

[Bro19c] William H. Brock. A German Partington. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):346–352, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1596725.

Brock:2020:BRL

[Bro20a] William H. Brock. Book review: “Liebig lebt!” 100 Jahre Liebig Museum im Laboratorium, Edited by Eduard Alter. Pp. 33, illus. Liebig Museum: Giessen. 2020. EUR 6.Das Liebig-Laboratorium von seinen Anfängen bis in die Gegenwart. Festschrift zum 100Jährigen Jubiläum des Liebigs-Museum. (Berichte der Justus Liebig-Gesellschaft zu Giessen. Band 10). By Franziska Müller. Pp. 173, illus. Justus Liebig-Gesellschaft: Giessen. 2020. EUR 14. ISSN: 0940-3426. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):419–420, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2020.1806517.

Brock:2020:CSC

[Bro20b] William H. Brock. The curious story of the Chemical Society’s missing obituary of John Lloyd Bullock. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):389–399, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1826822.

Brock:2020:NGC

[Bro20c] William H. Brock. Noel George Coley (1927–2020). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):313–314, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1779472.

Brock:2020:PCB

[Bro20d] William H. Brock. Perspectives on chemical biography in the 21st century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):199–200, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1687122.

Brock:2020:PBW

[Bro20e] William H. Brock. Pioneering British women chemists. their lives and contributions. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):315–316, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1728997.

Brock:2021:CAH

[Bro21] William H. Brock. The Chemical Age. How chemists fought famine and disease, killed millions, and changed our relationship with the Earth. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):327–328, 2021. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2021.1919456.

Brock:2022:RWH

[Bro22] Anna Simmons William H. Brock. Robert Warington and Heinrich Will: Friendship and co-operation in chemistry in Nineteenth Century Britain and Germany. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):374–398, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2133807.

Brock:1990:LGB

[BS90] W. H. Brock and Susanne Stark. Liebig, Gregory and the British association, 1837–1842. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(3):134–147, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.3.
134.

Bertomeu-Sanchez:2014:CSA

[BS14] José Ramón Bertomeu-Sánchez.

Classrooms, salons, academies, and courts: Mateu Orfila (1787–1853) and Nineteenth-Century French toxicology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):162–186, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000051.

Bertomeu-Sanchez:2022:BRE

[BS22] José Ramón Bertomeu-Sánchez.

Book review: Ethics of Chemistry. From Poison Gas to Climate Engineering. Edited by Joachim Schummer and Tom Børsen. Pp. vii + 559, index. World Scientific Publishing: Toh Tuck Link (Singapore). 2021. £175.00. ISBN: 978-9-81-123353-1 (hardcover); 978-9-81-123353-8 (ebook). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):337–338, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2071818.

Bertomeu-Sanchez:2002:LOT

[BSGBBV02] José Ramon Bertomeu-Sánchez, Antonio Garcia-Belmar, and Bernadette Bensaude-Vincent. Looking for an order of things: Textbooks and chemical classifications in nineteenth century France. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(3):227–250, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2002.49.3.
                                        227.

Bud:2018:SHL

[Bud18] Robert Bud. Synthetic: How life got made. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2): 194–196, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Bud:2020:FCC

[Bud20] Robert Bud. Fluid careers and Cold War boundaries. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):308–312, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2020.1794701.

Bues:2003:R

[Buè03]  Christiane Buès.  Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(2):231–259, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.2.
231.

Burnett:1992:AAA

[Bur92] Charles Burnett. The Astrologer’s assay of the alchemist: Early references to alchemy in Arabic and Latin texts. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):103–109, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.3.
103.

Burnett:1999:CER

[Bur99] D. Graham Burnett. The cosmogonic experiments of Robert Fludd: A Translation with Introduction and Commentary. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(3):113–170, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1999.46.3.
113.

Burns:2002:ERW

[Bur02] D. Thorburn Burns. Essay review: The Works of Robert Boyle. 14 volumes. Edited by Michael Hunter and Edward R. Davis. Pp. dccxvii + 7388, index. Pickering & Chatto: London. 1999–2000. £1190/ $1950. ISBN 1-85196-522-X (vols. 1–7), ISBN 1-85196-109-7 (14-volume set). The Correspondence of Robert Boyle. 6 volumes. Edited by Michael Hunger, Antonio Clericuzio, and Lawrence M. Principe. Pp. xv + 3249, index. Pickering & Chatto: London. 2001. £495/$740. ISBN 1-85196-123-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(2):158–160, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.2.
158.

Burney:2016:HFS

[Bur16] Ian Burney. A history of forensic science: British beginnings in the Twentieth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):353–355, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279412.

Buttner:2000:JLH

[Büt00]  Johannes Büttner.  Justus

Von Liebig and his influence on clinical chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(2):96–117, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.2.
96.

Bensaude-Vincent:2016:RCH

[BV16a] Bernadette Bensaude-Vincent.

The restless clock: a history of the centuries-long argument over what makes living things tick. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):348–350, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279408.

Bensaude-Vincent:2016:SBL

[BV16b] Bernadette Bensaude-Vincent.

Scientific Babel: The language of science from the fall of Latin to the rise of English. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):80–82, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1201309.

Bensaude-Vincent:2017:SPS

[BV17] Bernadette Bensaude-Vincent. Science in the public sphere: a history of lay knowledge and expertise. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):199–200, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1346958.

Bycroft:2021:TCW

[Byc21] Michael Bycroft. The transmutations of chymistry: Wilhelm Homberg and the Académie Royale des Sciences. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):323–325, 2021. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1893519.

Brock:2015:R

[BZC^{+}15] William H. Brock, Mike A. Zuber, Amy Eisen Cislo, Sarah Lowengard, Nuno Figueiredo, Ana Simões, Jeffrey Allan Johnson, and Jonathan Simon. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1):98–106, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/

full/10.1179/0002698014Z.
00000000073.

Carneiro:2015:PPI

[CA15] Ana Carneiro and Isabel Amaral. Propaganda and philanthropy: The Institute Bento da Rocha Cabral, the Lisbon site of biochemistry (1925–1953). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(2):138–166, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823415Y.
0000000002.

Campbell:1977:CLT

[Cam77] W. A. Campbell. The chemical library of Thomas Britton (1654–1714). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):143–148, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1977.24.3.
143.

Cammidge:1991:BTP

[Cam91] E. M. Cammidge. Benzene and turpentine: The prehistory of drycleaning. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(2):79–84, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1991.38.2.
                                        79.

Camenietzki:2001:JAE

[Cam01] Carlos Ziller Camenietzki. Jesuits and alchemy in the Early Seventeenth Century: Father Johannes Roberti and the weapon-salve controversy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):83–101, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.2.
                                        83.

Carneiro:1993:AWA

[Car93] Ana Carneiro. Adolphe Wurtz and the atomism controversy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(2):75–95, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.2.
                                        75.

Carneiro:2016:LPP

[Car16a] Ana Carneiro. From local patriotism to a planetary perspective: Impact crater research in Germany, 1930s–1970s. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):77–78, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2016.1201306.

Carrier:2016:EIK

[Car16b] Marcus B. Carrier. Experimentalisierung und internationale Kommunikationen: Der Fall Curare. (German) [Experimentalization and international communications: The curare case]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):73–74, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1201301.

Carlotta:2017:RDC

[Car17a] Vincenzo Carlotta. I ricettari del codice 52 della Historical Medical Library di New Haven (XIII sec. u.q.). (Italian) [Recipes of code 52 of the Historical Medical Library of New Haven (XIII Century. u. q.)]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):379–380, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1405894.

Carneiro:2017:RAA

[Car17b]  Ana Carneiro.  Review:

Amadeo Avogadro. Relazioni accademiche. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):103, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1310429.

Corral-Broto:2023:CEH

[CB23] Pablo Corral-Broto. The contamination of the Earth. A history of pollutions in the Industrial Age. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):104–105, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2156109.

Chilton:1980:LRI

[CC80] Donovan Chilton and Noel G. Coley. The laboratories of the Royal Institution in the Nineteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(3):173–203, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.3.
                                        002.

Canizares-Esguerra:2017:MEM

[CE17] Jorge Cañizares-Esguerra. The matter of empire: Metaphysics and mining in Colonial Peru. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):286–287, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385907.

Ceccatti:2004:BCI

[Cec04] John S. Ceccatti. Biology in the chemical industry: Scientific approaches to the problem of insecticide resistance, 1920s–1960s. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):135–147, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.2.
135.

Correa-Gomez:2022:TIC

[CG22] María José Correa-Gómez.

Book review: Tóxicos Invisibles. La construcción de la ignorancia ambiental. Edited by Ximo Guillem-Llobat and Agustí Nieto-Galan. Pp. 336, illus., index. Icaria: Barcelona. 2020. £19.50. ISBN: 978-84-9888-976-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):332–334, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2076875.

Campbell:2018:AMO

[CGT18] Andrew Campbell, Lorenza Gianfrancesco, and Neil Tarrant.

Alchemy and the Mendicant orders of Late Medieval and Early Modern Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):201–209, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1512778.

Chaldecott:1981:WCW

[Cha81]  J. A. Chaldecott.  Wedg-

wood’s ceramic wares for chemical use production and supply from 1779 to 1794. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):184–205, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1981.28.3.
                                        184.

Chang:2007:TAP

[Cha07] Ku-Ming (Kevin) Chang. Toleration of alchemists as a political question: Transmutation, disputation, and early modern scholarship on alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):245–273, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X237047.

Chang:2017:WHT

[Cha17] Hasok Chang. What history tells us about the distinct nature of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):360–374, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1412135.

Chang:2021:PCG

[Cha21] Hasok Chang. Philosophical chemistry: Genealogy of a scientific field. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):455–456, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1907089.

Chang:2023:BRC

[Cha23] Hasok Chang. Book review: A Cultural History of Chemistry in the Nineteenth Century (vol. 5). Edited By Peter Ramberg. Pp. xiii + 272, illus., index. Bloomsbury: London. 2022. £395.00 (6-volume set). ISBN: 978-1-4742-9480-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):194–196, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2186584.

Chimisso:2017:IME

[Chi17] Cristina Chimisso. Les identités multiples d’Émile Meyerson. (French) [The multiple identities of Émile Meyerson]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):197–199, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1346957.

Christie:1994:HCE

[Chr94] J. R. R. Christie. Historiography of chemistry in the Eighteenth Century: Hermann Boerhaave and William Cullen. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):4–19, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.1.
                                        4.

Chung:2013:RSS

[Chu13] Deborah D. L. Chung. Road to Scientific Success: Inspiring Life Stories of Prominent Researchers, volume 2. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2013. ISBN 981-4541-91-5 (paperback), 981-4541-92-3 (e-book). viii + 206 pp. LCCN ????

Chung:2006:RSS

[Chu14] Deborah D. L. Chung, editor. The Road to Scientific Success: Inspiring Life Stories of Prominent Researchers, volume 1.

World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2006–2014. ISBN 981-256-600-7 (hardcover), 981-256-466-7 (paperback). xi + 230 pp. LCCN Q141 .R518 2006.

Ciardi:2001:AAC

[Cia01] Marco Ciardi. Amedeo Avogadro’s concept of the atom: Some new remarks. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(1):17–24, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.1.
17.

Cislo:2008:PCS

[Cis08] Amy Eisen Cislo. Paracelsus’s conception of seeds: Rethinking Paracelsus’s ideas of body and matter. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):274–282, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X255488.

Clarke:2001:AAC

[Cla01] Mark Clarke. The Art of All Colours: Medieval Recipe Books for Painters and Illuminators. Archetype Publications, London, UK, 2001. ISBN 1-873132-72-7. viii + 152 pp. LCCN ND141 .C58 2001.

Clericuzio:1993:R

[Cle93] Antonio Clericuzio. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(3):171–176, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.3.
                                        171.

Clericuzio:1998:R

[Cle98] Antonio Clericuzio. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(3):193–200, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.3.
                                        193.

Clericuzio:2021:CPR

[Cle21] Antonio Clericuzio. The chemical philosophy of Robert Boyle: Mechanicism, chymical atoms, and emergence. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):450–451, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1930366.

Clulee:1971:JDM

[Clu71] N. H. Clulee. John Dee’s mathematics and the grading of compound qualities. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(3):178–211, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1971.18.3.
178.

Clucas:1993:CXC

[Clu93] Stephen Clucas. The correspondence of a XVII-Century ‘chymicall gentleman’: Sir Cheney Culpeper and the chemical interests of the Hartlib Circle. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(3):147–170, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.3.
147.

Clucas:2005:E

[Clu05a]  Stephen Clucas.  Editorial.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):195, 2005. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/000269805X77745.

Clulee:2005:MHA

[Clu05b] Nicholas H. Clulee. The Monas Hieroglyphica and the alchemical thread of John Dee’s career. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):197–215, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X77763.

Clucas:2007:AEM

[Clu07] Stephen Clucas. Alchemy and Early Modern Chemistry: Papers from Ambix. edited by Allen G. Debus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):87–88, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582307X169890.

Clucas:2011:MCM

[Clu11]  Stephen Clucas.  Margaret

Cavendish’s materialist critique of Van Helmontian chymistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(1):1–12, 2011. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582311X12947034675596.

Clucas:2017:JDA

[Clu17a] Stephen Clucas. John Dee, alchemy, and print culture. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):107–114, 2017. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

full/10.1080/00026980.2017.
1356632.

Clucas:2017:RTA

[Clu17b] Stephen Clucas. The Royal Typographer and the alchemist: John Dee, Willem Silvius, and the diagrammatic alchemy of the Monas Hieroglyphica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):140–156, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1356972.

Cobbold:2019:AAR

[Cob19] Carolyn Cobbold. Adulation or adulteration? Representing chemical dyes in the Victorian media. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):23–50, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1574385.

Cobbold:2020:PAC

[Cob20] Carolyn Cobbold. Pure adulteration: Cheating on nature in the age of manufactured food. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):318–319, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1770442.

Cobbold:2022:CFH

[Cob22] Carolyn Cobbold. Book review: The Chemistry of Fear: Harvey Wiley’s Fight for Pure Food. By Jonathan Rees. Pp. 320, illus., index. Johns Hopkins University Press: Baltimore, MA. 2021. £28.00 (hardcover). ISBN: 978-1-42-143995-2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):428–429, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2085752.

Coley:1971:ACU

[Col71] N. G. Coley. Animal chemists and urinary stone. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(2):69–93, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.2.
                                        69.

Collins:1975:HDH

[Col75] Peter Collins. Humphry Davy and heterogeneous catalysis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(3):205–217, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.3.
205.

Coley:1976:R

[Col76a] N. G. Coley. Reviews. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 23(1):59–64, 1976. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269876790223577.

Collins:1976:JWD

[Col76b] Peter Collins. Johann Wolfgang Döbereiner and heterogeneous catalysis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(2):96–115, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.2.
96.

Cole:1978:DMG

[Col78] Theron Cole. Dalton, mixed gases, and the origin of the chemical atomic theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(2):117–130, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.2.

Coley:1983:R

[Col83] N. G. Coley. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(2):117–120, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1983.30.2.
                                        117.

Coley:1984:PUA

[Col84] Noel G. Coley. The preparation and uses of artificial mineral Waters (ca. 1680–1825). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(1):32–48, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1984.31.1.
                                        32.

Coley:1988:MCG

[Col88]  Noel G. Coley.  Medical

chemistry at Guy’s Hospital (1770–1850). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(3):155–168, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1988.35.3.
                                        155.

Coley:1996:SHA

[Col96] Noel G. Coley. Studies in the history of animal chemistry and its relation to physiology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(3):164–187, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.3.
164.

Coley:2002:WAB

[Col02] Noel Coley. William Arthur (“Bill”) Smeaton (1925–2001). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(3):177–184, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.3.
177.

Conlin:1996:JPA

[Con96] Michael F. Conlin. Joseph Priestley’s American defense of phlogiston reconsidered. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(3):129–145, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.3.
129.

Considine:2022:BEP

[Con22]  John Considine.  The be-

ginnings of English Paracelsian lexicography: Two collections of words from Elizabethan Cambridge. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 69(2):163–189, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.2012315.

Copenhaver:1992:HGC

[Cop92] Brian P. Copenhaver. Hermetica: the Greek Corpus Hermeticum and the Latin Asclepius in a New English Translation, with Notes and Introduction. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36144-3 (hardcover), 0-521-42543-3 (paperback), 1-4619-3820-1 (e-book), 1-107-05007-3 (ebook). lxxxiii + 320 pp. LCCN BF1600 .H475 1992.

                                        URL http://www.loc.gov/
                                        catdir/description/cam024/
                                        91025703.html;   http://
                                        www.loc.gov/catdir/toc/cam028/
                                        91025703.html.

Court:1972:AC

[Cou72] Susan Court. The Annales de chimie, 1789–1815. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(2):113–128, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1972.19.2.
                                        113.

Coulter:1977:LRE

[Cou77] Harris L. Coulter. Letters to the reviews Editor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(1):63–68, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1977.24.1.
63.

Craven:1993:ER

[Cra93] Kenneth Craven. Essay review.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(2):96–97, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.2.
96.

Craddock:2010:SCC

[Cra10] Paul Craddock. Science and civilisation in China. Volume

5. Chemistry and chemical technology. Part 11: Ferrous metallurgy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(2):231–232, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582310X12719003720368.

Creager:2017:CRH

[Cre17] Angela N. H. Creager. A chemical reaction to the historiography of biology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):343–359, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1412136.

Crisciani:1973:CAE

[Cri73] Chiara Crisciani. The conception of alchemy as expressed in the Pretiosa Margarita Novella of Petrus Bonus of Ferrara. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):165–181, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.3.
                                        165.

Crosland:1981:LGL

[Cro81] Maurice Crosland. The library of Gay-Lussac. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):158–170, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1981.28.3.
                                        158.

Crosland:1995:LTF

[Cro95] Maurice Crosland. Lavoisier, the two French revolutions and ‘the Imperial despotism of oxygen’. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(2):101–118, 1995. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1995.42.2.
101.

Crosland:2009:LAM

[Cro09] Maurice Crosland. Lavoisier’s achievement; more than a chemical revolution. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(2):93–114, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X441417.

Court:1979:FJS

[CS79] Susan Court and W. A.

Smeaton. Fourcroy and the Journal de la Société des pharmaciens de Paris. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(1):39–55, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.1.
39.

Clarence-Smith:2019:BRR

[CS19] William G. Clarence-Smith.

Book review: Rubber and the Making of Vietnam: an Ecological History, 1897–1975, by Michitake Aso. Pp. 405 + xvii, illus., index. The University of North Carolina Press: Chapel Hill. 2018. £70 (hardcover), £25.30 (paperback), £18.40 (ebook). ISBN: 978-1-4696-3714-3 (hardcover), 978-1-4696-3715-0 (paperback), 978-1-4696-3716-7 (ebook). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):367–368, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1611017.

Cesaro:1988:EAI

[CT88] S. Nunziante Cesaro and E. Torracca. Early applications of infra-red spectroscopy to chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(1):39–47, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1988.35.1.
                                        39.

Damon:1977:BA

[Dam77] S. Foster Damon. De Brahm: Alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(2):77–88, 1977. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1977.24.2.77.

Daniel:2006:IWR

[Dan06]  Dane T. Daniel.  Invisible

wombs: Rethinking Paracelsus’s concept of body and matter. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53 (2):129–142, July 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X117870.

Davidson:1987:PDD

[Dav87] Jane P. Davidson. “I am the poison dripping dragon”: Iguanas and their symbolism in the alchemical and occult paintings of David Teniers the Younger. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(2):62–80, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.2.
62.

Day:2005:NMP

[Day05] Peter Day. Nature Not Mocked:

Places, People and Science. Imperial College Press, London, UK, 2005. ISBN 1-86094-576-7. x + 262 pp. LCCN Q181 .D363 2005.

deChadarevian:2004:MTD

[dCH04] Soraya de Chadarevian and Nick Hopwood, editors. Models: the Third Dimension of Science. Writing science. Stanford University Press, Stanford, CA, USA, 2004. ISBN 0-8047-3971-4 (hardcover), 0-8047-3972-2 (paperback). xvi + 464 pp. LCCN Q223 .M58 2004. URL

                                        http://www.e-streams.com/
                                        es0803/es0803_4012.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0710/2003025168-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0710/2003025168-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0411/2003025168.
                                        html.

DeGouveia:1985:VSC

[De 85] A. J. A. De Gouveia. Vicente De Seabra and the chemical revolution in Portugal. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(3):97–109, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1985.32.3.
                                        97.

Debus:1960:PCE

[Deb60] Allen G. Debus. The Paracelsian compromise in Elizabethan England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(2):71–97, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1960.8.2.
                                        71.

Debus:1961:GPH

[Deb61] Allen G. Debus. Gabriel Plattes and his chemical theory of the formation of the Earth’s crust. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(3):162–165, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.3.
162.

Debus:1962:JWP

[Deb62a] Allen G. Debus. John Woodall, Paracelsian surgeon. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(3):108–118, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1962.10.3.
108.

Debus:1962:STB

[Deb62b] Allen G. Debus. Sir Thomas Browne and the study of colour indicators. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(1):29–36, 1962. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269862790222933.

Debus:1963:FCI

[Deb63] Allen G. Debus. A forgotten chapter in the introduction of the new chemistry in Italy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(3):153–157, 1963. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1963.11.3.
153.

Debus:1967:RCW

[Deb67]  Allen G. Debus.  Renais-

sance chemistry and the work of Robert Fludd. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(1):42–59, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.1.
42.

Debus:1968:MNC

[Deb68] Allen G. Debus. Mathematics and nature in the chemical texts of the Renaissance. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(1):1–28, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1968.15.1.
1.

Debus:1971:HCH

[Deb71] Allen G. Debus. The history of chemistry and the history of science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(3):169–177, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1971.18.3.
169.

Debus:1972:SCC

[Deb72] Allen G. Debus. Some comments on the contemporary Helmontian Renaissance. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(2):145–150, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.2.
145.

Debus:1976:R

[Deb76] Allen G. Debus. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(2):124–127, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.2.
124.

Debus:1980:TSP

[Deb80]  Allen G. Debus.  Thomas

Sherley’s Philosophical Essay (1672): Helmontian mechanism as the basis of a new philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(2):124–135, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.2.
124.

Debus:1981:PEC

[Deb81] Allen G. Debus. The Paracelsians in Eighteenth Century France: A Renaissance tradition in the age of the enlightenment. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(1):36–54, 1981. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1981.28.1.36.

Debus:1984:R

[Deb84] Allen G. Debus. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(1):49–56, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1984.31.1.
                                        49.

Debus:1985:SCH

[Deb85] Allen G. Debus. The significance of chemical history. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(1):1–14, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1985.32.1.
                                        1.

DeCroesJacobs:2015:JSL

[DeC15] Charlotte DeCroes Jacobs. Jonas Salk: a Life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-933441-2 (hardcover). x + 559 + 16 pp. LCCN QR31.S25 J33 2015.

Dejong:1973:R

[Dej73] H. M. E. Dejong. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):246–250, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.3.
246.

DeKosky:1980:GGS

[DeK80] Robert K. DeKosky. George Gabriel Stokes, Arthur Smithells and the origin of spectra in flames. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(2):103–123, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.2.
103.

DeKosky:2009:DCI

[DeK09] Robert DeKosky. Developing chemical instrumentation for environmental use in the late Twentieth Century: Detecting lead in paint using portable X-ray fluorescence spectrometry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(2):138–162, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582309X441408.

Delwiche:2020:FIN

[Del20] Theodore R. Delwiche. Fuit Ille non Empiricus Mercenarius: apprehensions to alchemy in Colonial New England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):346–365, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1826820.

Delmeulle:2021:OPM

[Del21]  Guillaume Delmeulle.  At

the origins of the De Perfecto Magisterio: a translation from Arabic or a Latin composition? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):431–441, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1983693.

DosSantosPereira:2022:WOP

[DFB22] Letícia Dos Santos Pereira, Olival Freire Júnior, and Gisela Boeck. Wilhelm Ostwald’s pedagogy: an analysis of the Nobel Prize nomination letters. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):139–162, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2052529.

DiGennaroSplendore:2022:MMC

[Di 22] Barbara Di Gennaro Splendore. Merchants of medicines. The commerce and coercion of health in Britain’s long Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):84–85, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1943143.

Dickinson:1972:LEJ

[Dic72] Robert Dickinson. Letter to Editor Josef Hawliczek. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(1):59, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.1.
59.

Dienel:2004:LAG

[Die04a] Hans-Liudger Dienel. Die Linde AG: Geschichte eines Technologiekonzerns 1879–2004. (German) [The Linde AG: History of a Technology Corporation 1879–2004]. C. H. Beck, München, Germany, 2004. ISBN 3-406-51484-7 (hardcover). 512 + 16 pp. LCCN HD9971.5.R444 L563 2004.

Dienel:2004:LHT

[Die04b] Hans-Liudger Dienel. Linde: History of a Technology Corporation, 1879–2004. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2004. ISBN 1-4039-2033-8. xiv + 369 pp. LCCN TA217 .D48 2004.

                                        URL http://www.h-net.org/
                                        review/hrev-a0e5s9-aa;
                                        http://www.loc.gov/catdir/
                                        bios/hol059/2004042838.html;
                                         http://www.loc.gov/catdir/
                                        description/hol053/2004042838.
                                        html;   http://www.loc.gov/
                                        catdir/toc/hol052/2004042838.
                                        html.

Davies:1986:CLB

[DLS86] D. I. Davies, D. C. Lyon, and R. J. Spring. Charles Loudon Bloxam — a Victorian university and military academy chemistry teacher. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(1):11–32, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1986.33.1.
                                        11.

Decet:1997:SCA

[DM97] Fabio Decet and Rosario Mosello. Studies on the chemistry of atmospheric deposition in the eighteenth and nineteenth centuries. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(2):57–84, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.2.
57.

Couto:2015:CDM

[dMC15] Cristiana Loureiro de Mendonça Couto. The chemistry of diet: Medicine, nutrition, and staple foods in Imperial Brazil. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(4):345–362, 2015. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2015.1129855.

Dmitriev:2004:CEP

[Dmi04] I. S. (Igor Sergeevich) Dmitriev.

Chelovek Epokhi Peremen: Ocherka o D.I. Mendeleyeve i Yego Vremeni. (Russian) [A Man in a Changing Epoch: a Treatise on D. I. Mendeleev and his Times]. Khimizdat, St. Petersburg, Russia, 2004. ISBN 5-93808-082-7. 574 pp. LCCN QD22.M43 D65 2004.

Davis:1937:TJJ

[DN37] Tenney L. Davis and Rokuro Nakaseko. The tomb of Jofuku or Joshi, the earliest alchemist of historical record. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(2):109–115, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1937.1.2.
                                        109.

Dobbs:1971:SNP

[Dob71]  Betty Jo Dobbs.  Studies

in the natural philosophy of Sir Kenelm Digby. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(1):1–25, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.1.
                                        1.

Dobbs:1973:SNP

[Dob73]  Betty Jo Dobbs.  Studies

in the natural philosophy of Sir Kenelm Digby. Part II. Digby and alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):143–163, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.3.
                                        143.

Dobbs:1974:SNP

[Dob74]  Betty Jo Dobbs.  Studies

in the natural philosophy of Sir Kenelm Digby. Part III. Digby’s experimental alchemy — The Book of Secrets. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):1–28, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1974.21.1.
1.

Dobbs:1979:NCS

[Dob79] B. L. T. Dobbs. Newton’s copy of “Secrets Reveal’d” and the regimens of the work. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):145–169, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.3.
145.

Dobbs:1982:ERN

[Dob82a] B. J. T. Dobbs. Essay review:

Never at Best. A Biography of Isaac Newton. By Richard S. Westfall. Pp. xviii + 908, with many illustrations. Cambridge, London, New York, etc.: Cambridge University Press. 1981. £25. ISBN 0-521-23143-4. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(1):62–68, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.1.
62.

Dobbs:1982:NCN

[Dob82b] B. J. T. Dobbs. Newton’s “Clavis”: New evidence on its dating and significance. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(3):198–202, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1982.29.3.
                                        198.

Dolan:1998:BBH

[Dol98] Brian P. Dolan. Blowpipes and batteries: Humphry Davy, Edward Daniel Clarke, and experimental chemistry in Early Nineteenth-Century Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(3):137–162, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.3.
                                        137.

Donovan:1978:JHJ

[Don78] Arthur Donovan. James Hutton, Joseph Black and the chemical theory of heat. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(3):176–190, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1978.25.3.
                                        176.

Donner:1988:ERE

[Don88]  Henriette Donner.  Essay

review: Erinnerungen eines deutschen Naturforschers. By Johannes Stark, with an introduction by Andreas Kleinert. Pp. x + 153. Bionomica-Verlag: Mannheim, 1997. DM 20. ISBN 3-88208-060-0. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(3):169–176, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.3.
169.

Djerassi:2003:NDT

[DP03] Carl Djerassi and David Pinner.

Newton’s Darkness: Two Dramatic Views. Imperial College Press, London, UK, 2003. ISBN 1-86094-389-6 (hardcover), 1-86094-390-X (paperback). 184 pp. LCCN PS3554.J47 N49 2003.

Driggers:2016:UEM

[Dri16] Edward Allen Driggers. Uroscopy in Early Modern Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):188–189, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/00026980.2016.
1227167.

daSilveira:2022:MPC

[dS22] Fabiana Lopes da Silveira. In the melting pot: Cultural mixture and the presentation of alchemical knowledge in the Letter from Isis to Horus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):49–64, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.2023961.

Dean:1979:SPR

[DU79] Philip A. W. Dean and Melvyn C. Usselman. The ‘synthetic’ palladium of Richard Chenevix: A verdict on the chemist and the chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(2):100–115, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1979.26.2.
                                        100.

Dubs:1961:OA

[Dub61] Homer H. Dubs. The origin of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(1):23–36, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1961.9.1.
                                        23.

Ducheyne:2008:PSA

[Duc08] Steffen Ducheyne. A preliminary study of the appropriation of Van Helmont’s oeuvre in Britain in chymistry, medicine and natural philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(2):122–135, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X255479.

Duchacek:2014:R

[Duc14] Suzanne Sutherland Duchacek.

Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(1):100–108, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000046.

Dufault:2015:TTG

[Duf15] Olivier Dufault. Transmutation theory in the Greek alchemical corpus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(3):215–244, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/1745823415Y.
0000000003.

Dufault:2020:GAL

[Duf20] Olivier Dufault. Greek alchemy from Late Antiquity to Early Modernity. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):104–106, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1687121.

Dufault:2023:BRC

[Duf23] Olivier Dufault. Book review: A Cultural History of Chemistry in Antiquity (vol. 1), Edited by Marco Beretta. Pp. 294, illus., index. Bloomsbury: London. 2022. £395.00 (6-volume set). ISBN: 978-1-4742-9453-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):186–188, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2023.2186577.

Dunleavy:1965:CAT

[Dun65] Gareth W. Dunleavy. The Chaucer Ascription in Trinity College, Dublin MS. D.2.8. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(1):2–21, 1965. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1965.13.1.
                                        2.

Duncan:1970:FAT

[Dun70] A. M. Duncan. The functions of affinity tables and Lavoisier’s list of elements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1):28–42, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.1.
28.

Duncan:1981:SLM

[Dun81] A. M. Duncan. Styles of language and modes of chemical thought. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(2):83–107, 1981. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1981.28.2.83.

Dupre:2023:LEW

[Dup23] Sven Dupré. From lived experience to the written word: Reconstructing practical knowledge in the Early Modern World. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):201–202, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2170541.

Duveen:1948:LTS

[Duv48] Denis Duveen. Le Livre de la Très Sainte Trinité. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(1–2):26–32, 1948. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1948.3.1-
                                        2.26.

Duveen:1996:O

[Duv96]  Denis Ian Duveen.  Obit-

uary. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(1): 31, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1996.43.1.
                                        31.

deVries:2022:PAR

[dV22] Lyke de Vries. Protecting academia and religion: Andreas Libavius’s criticism of a general reformation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):34–48, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.2023962.

Heathcote:1958:LE

[dVH58] N. H. de V. Heathcote. Letters to the Editor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3):155–156, 1958. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1958.6.3.
                                        155.

Dym:2008:AMM

[Dym08] Warren Alexander Dym. Alchemy and mining: Metallogenesis and prospecting in early mining books. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):232–254, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X358105.

Eamon:1980:NLR

[Eam80] William Eamon. New light on Robert Boyle and the discovery of colour indicators. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(3):204–209, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.3.
003.

Eamon:2021:PAL

[Eam21] William Eamon. Paracelsus: an alchemical life. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):321–323, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1874196.

Ebeling:2007:SHH

[Ebe07] Florian Ebeling. The Secret History of Hermes Trismegistus: Hermeticism from Ancient to Modern Times. Cornell University Press, Ithaca, NY, USA, 2007. ISBN 0-8014-4546-9 (hardcover). xiii + 158 pp. LCCN BF1589 .E2513 2007. URL http:

//www.loc.gov/catdir/toc/
ecip076/2006101315.html.

Eddy:2001:DSR

[Edd01]  M. D. Eddy.  The ‘Doc-

trine of Salts’ and Rev. John Walker’s analysis of a Scottish spa (1749–1761). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(3):137–160, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.3.
137.

Eddy:2008:DSC

[Edd08] M. D. Eddy. The dark side of collecting — early modern chemistry, humanism and classification. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):283–292, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X358132.

Ede:1993:WTT

[Ede93] Andrew Ede. When is a tool not a tool? Understanding the role of laboratory equipment in the early colloidal chemistry laboratory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(1):11–24, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.1.
11.

Ede:1996:CQU

[Ede96] Andrew Ede. Colloids and quantification: The ultracentrifuge and its transformation of colloid chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(1):32–45, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.1.
32.

Edmondson:2019:CC

[Edm19] Hattie Lloyd Edmondson.

Chivalrous chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):103–120, 2019. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2019.1616931.

Elina:2011:PIP

[Eli11] Olga Y. Elina. Private initiatives, public support, and war practices: Development of fertilisers in Russia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(1):29–61, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X12947034675479.

Elliott:2022:CCN

[Ell22] George D. Elliott. Chemistry in 17th-Century New England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):91–92, 2022. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1969763.

Farrar:1971:SEV

[Far71] W. V. Farrar. Some early ventures in the fixation of atmospheric nitrogen. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(2):123–138, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.2.
                                        123.

Farrar:1975:ESC

[Far75] W. V. Farrar. Examination of some compounds isolated from madder, 1845–1855. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):151–153, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
151.

Farrar:1979:R

[Far79] Kathleen R. Farrar. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(1):74–80, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.1.
74.

Farkas:1989:PHT

[Far89] Adalbert Farkas, Dr. Phil. Nat., Dr. Ing. Paul Harteck: the triumphant decade 1925–1934. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(2):91–102, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.2.
91.

Fara:2020:ALD

[Far20] Patricia Fara. American lucifers: The dark history of artificial light. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):200–202, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1689615.

Fauque:2011:FCI

[Fau11] Danielle M. E. Fauque. French chemists and the international reorganisation of chemistry after World War I. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(2):116–135, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X13008456751071.

Fontani:2015:LEP

[FCO15] Marco Fontani, Mariagrazia Costa, and Mary Virginia Orna. The Lost Elements: the Periodic Table’s Shadow Side. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-938334-0. xxxvii + 531 pp. LCCN QD467 .F66 2015.

Ferrario:2009:ADT

[Fer09] Gabriele Ferrario. An Arabic dictionary of technical alchemical terms: MS Sprenger 1908 of the Staatsbibliothek zu Berlin (fols. 3r–6r). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):36–48, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582309X405219.

Ferrario:2016:AAM

[Fer16] Gabriele Ferrario. Art and alchemy: The mystery of transformation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):71–72, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1201297.

Ferrario:2017:MAA

[Fer17] Gabriele Ferrario. La messe alchimique attribuée à Melchior de Sibiu. (French) [The alchemical mass attributed to Melchior de Sibiu]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):99–100, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1310425.

Fernandez:2022:BRP

[Fer22] María Belén Boned Fernández.

Book review: Pseudo-Paracelsus: Forgery and Early Modern Alchemy, Medicine, and Natural Philosophy. Edited by Didier Kahn and Hiro Hirai. Pp. 490, illus., index. Brill: Leiden. 2021. £122.00. ISBN: 978-90-04-50337-3. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):426–427, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2105491.

Farrar:1973:HMP

[FFS73] W. V. Farrar, Kathleen R. Farrar, and E. L. Scottt. The Henrys of Manchester. Part I: Thomas Henry (1734–1816). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):183–208, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.3.
                                        183.

Farrar:1974:HMPa

[FFS74a] W. V. Farrar, Kathleen R. Farrar, and E. L. Scott. The Henrys of Manchester: Part 2. Thomas Henry’s sons: Thomas, Peter and William. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(2–3):179–207, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/000269874790223605.

Farrar:1974:HMPb

[FFS74b] W. V. Farrar, Kathleen R. Farrar, and E. L. Scott. The Henrys of Manchester: Part 3. William Henry and John Dalton. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21 (2–3):208–246, 1974. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/000269874790223588.

Farrar:1975:HMP

[FFS75] W. V. Farrar, Kathleen R. Farrar, and E. L. Scott. The Henrys of Manchester. Part 4: William Henry: Hydrocarbons and the gas industry: Minor chemical papers. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(3):186–204, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.3.
186.

Farrar:1976:HMP

[FFS76] W. V. Farrar, Kathleen R. Farrar, and E. L. Scott. The Henrys of Manchester: Part 5: William Henry: Contagion and cholera; the textbook. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(1):27–52, 1976. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269876790223559.

Farrar:1977:HMP

[FFS77] W. V. Farrar, Kathleen R. Farrar, and E. L. Scott. The Henrys of Manchester. Part 6. William Charles Henry: The magnesia factory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(1):1–26, 1977.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.1.
                                        1.

Forrester:2005:JFH

[FH05] J. M. (John M.) Forrester and John Henry. Jean Fernel’s “On the hidden causes of things”: forms, souls, and occult diseases in Renaissance medicine, volume 6 of Medieval and early modern science. Brill, Leiden, The Netherlands, 2005. ISBN 90-04-14128-6 (hardcover). x + 779 pp. LCCN R128.6 .F38 2005.

Fichman:1971:FSC

[Fic71]  Martin Fichman.  French

Stahlism and chemical studies of air, 1750–1770. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(2):94–122, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.2.
                                        94.

Figurovski:1965:APA

[Fig65] N. A. Figurovski. The alchemist and physician Arthur Dee (Artemii Ivanovich Dii): an episode in the history of chemistry and medicine in Russia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(1):35–51, 1965. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1965.13.1.
35.

Figala:1973:LEP

[Fig73] Karin Figala. Letter to the Editor project for cataloguing alchemical manuscripts in German-speaking areas. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):245–246, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.3.
245.

Filgueiras:1992:MPS

[Fil92] Carlos A. L. Filgueiras. The mishaps of peripheral science: The life and work of Manoel Joaquim Henriques De Paiva, Luso–Brazilian chemist and physician of the late Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(2):75–90, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.2.
75.

Finlay:1991:RAC

[Fin91] Mark R. Finlay. The rehabilitation of an agricultural chemist: Justus Von Liebig and the Seventh Edition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(3):155–167, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1991.38.3.
                                        155.

Fisher:1973:OCBa

[Fis73a] N. W. Fisher. Organic classification before Kekulé. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):106–131, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.2.
                                        106.

Fisher:1973:OCBb

[Fis73b] N. W. Fisher. Organic classification before Kekulé: Part II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):209–233, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.3.
                                        209.

Fisher:1974:KOC

[Fis74] N. W. Fisher. Kekulé and organic classification. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):29–52, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1974.21.1.
29.

Fisher:1996:WOI

[Fis96] Susie Fisher. William Odling:

‘interpreter and liaison-officer’ advocate of a new system of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(3):146–163, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.3.
146.

Fisher:2018:RHT

[Fis18] Amy Fisher. Robert Hare’s theory of Galvanism: a study of heat and electricity in early nineteenth-century American chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2): 169–189, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Frercks:2007:ITC

[FM07] Jan Frercks and Michael Markert. The invention of Theoretische Chemie: Forms and uses of German chemistry textbooks, 1775–1820. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):146–171, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X212378.

Forbes:1938:PBA

[For38] R. J. Forbes. Petroleum and bitumen in antiquity. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(2):68–92, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1938.2.2.
                                        68.

Forshaw:2005:EAR

[For05]  Peter J. Forshaw.  The

early alchemical reception of John Dee’s Monas Hieroglyphica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):247–269, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269805X77772.

Fors:2008:STS

[For08] Hjalmar Fors. Stepping through science’s door: C. W. Scheele, from pharmacist’s apprentice to man of science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(1):29–49, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582308X255398.

Forshaw:2013:CCC

[For13]  Peter J. Forshaw.  Cabala

Chymica or Chemia Cabalistica — early modern alchemists and Cabala. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(4):361–389, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000039.

Forshaw:2017:HFC

[For17] Peter J. Forshaw. The hermetic frontispiece: Contextualising John Dee’s hieroglyphic monad. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):115–139, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1353247.

Fors:2021:RRR

[For21] Hjalmar Fors. Reconstruction, replication and re-enactment in the humanities and social sciences. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):453–454, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1927365.

Fors:2016:LLB

[FPS16] Hjalmar Fors, Lawrence M. Principe, and H. Otto Sibum. From the library to the laboratory and back again: Experiment as a tool for historians of science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):85–97, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1213009.

Franckowiak:2011:MCE

[Fra11] Rémi Franckowiak. Mechanical and chemical explanations in du clos’ chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(1):13–28, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X12947034675550.

Franckowiak:2017:FVC

[Fra17] Rémi Franckowiak. Le fixe et le volatil. Chimie et alchimie, de Paracelse à Lavoisier. (French) [The fixed and the volatile. Chemistry and alchemy, from Paracelsus to Lavoisier]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):191–192, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2017.1346952.

Frercks:2010:DFF

[Fre10] Jan Frercks. Demonstrating the facticity of facts: University lectures and chemistry as a science in Germany around 1800. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(1):64–83, 2010. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582310X12629173850005.

Frick:1959:ROM

[Fri59]  Karl Frick.  The rediscov-

ered original MS. “Hehrenrettung Der Alchymie” of the Tübingen alchemist Johann Conrad Creiling (1673–1752). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(3):164–167, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1959.7.3.
164.

Frietsch:2021:AEM

[Fri21a] Ute Frietsch. Alchemy and the early modern university: an introduction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):119–134, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2021.1936800.

Frietsch:2021:MUF

[Fri21b] Ute Frietsch. Making university fields for chymistry: a case study of Helmstedt University. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68 (2–3):273–301, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1922018.

Frietsch:2022:RFV

[Fri22] Ute Frietsch. Robert Fludd’s visual and artisanal episteme: A case study of Fludd’s interaction with his engraver, his printer–publisher, and his Amanuenses. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):341–373, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2133809.

Fuck:1951:ALA

[Füc51] J. W. Fück. The Arabic literature on alchemy according to an-Nadim (A.D. 987). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(3–4):81–144, 1951. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1951.4.3-
4.81.

Fullmer:1975:DPI

[Ful75] Juke Z. Fullmer. Davy’s priority in the iodine dispute: Further documentary evidence. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):39–51, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.1.
39.

Fulford:2019:DTH

[Ful19] Tim Fulford. Davy takes to the hills: Dialogic enquiry and the aesthetics of the prospect view. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66 (2–3):158–180, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1616946.

Furukawa:2021:EHC

[Fur21]  Yasu Furukawa.  Exploring

the history of chemistry in Japan. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68 (2–3):302–317, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1930675.

Gabriel:2016:BBD

[Gab16] Joseph M. Gabriel. Bioproperty, biomedicine and deliberative governance: Patents as discourse of life. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):79–80, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1201308.

Gantenbein:2020:CCA

[Gan20a] Urs Leo Gantenbein. Cross and crucible: Alchemy in the theology of Paracelsus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):88–99, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1723944.

Gantenbein:2020:RFN

[Gan20b] Urs Leo Gantenbein. Real or fake? New light on the Paracelsian De natura rerum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):4–29, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1720339.

Gavroglu:1999:R

[Gav99] Kostas Gavroğlu. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(2):106–112, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1999.46.2.
106.

Gay:1997:CPT

[Gay97]  Hannah Gay.  The chemi-

cal philosophy of Theodore W. Richards. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(1):19–38, 1997. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1997.44.1.19.

Gay:2000:PCA

[Gay00] Hannah Gay. ‘pillars of the college’: Assistants at the Royal College of chemistry, 1846– 1871. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(3):135–169, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.3.
135.

Gay:2012:BAS

[Gay12] Hannah Gay. Before and after Silent Spring: From chemical pesticides to biological control and integrated pest management — Britain, 1945–1980. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(2):88–108, 2012. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1179/174582312X13345259995930.

Gee:1991:CJJ

[GB91] Brian Gee and William H. Brock. The case of John Joseph Griffin. from artisan– chemist and author–instructor to business-leader. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(1):29–62, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1991.38.1.
                                        29.

Garcia-Belmar:2014:SCN

[GB14] Antonio García-Belmar. Sites of chemistry in the nineteenth century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):109–114, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/0002698014Z.
                                        00000000048.

Gago:1980:BSR

[GC80] Ramón Gago and Juan L. Carrillo. A bibliographical study of the reception of Lavoisier’s work in Spain. Addenda to A Bibliography by Duveen and Klickstein. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(1):19–25, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.1.
19.

Goodrick-Clarke:1996:N

[GC96] Nicholas Goodrick-Clarke. News.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(2):97–98, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.2.
97.

Gwei-Djen:1972:CAW

[GDNN72] Lu Gwei-Djen, Joseph Needham, and Dorothy Needham. The coming of ardent water. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(2):69–112, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.2.
69.

Gelbart:1971:IDW

[Gel71] Nina Rattner Gelbart. The intellectual development of Walter Charleton. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(3):149–168, 1971.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.3.
                                        149.

Gelman:1978:BTH

[Gel78]  Z. E. Gel’man.  Bernhard

Tollens and his influence on research into carbohydrates in Russia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(1):56–62, 1978. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1978.25.1.56.

Gelman:1994:ASI

[Gel94]  Zahkare E. Gelman.  An-

gelo Sala, an iatrochemist of the Late Renaissance. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(3):142–160, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.3.
                                        142.

Geoghegan:1957:LHV

[Geo57a] D. Geoghegan. A licence of Henry VI to practise alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(1):10–17, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.1.
10.

Geoghegan:1957:SIN

[Geo57b]  D. Geoghegan.  Some in-

dications of Newton’s attitude towards alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(2):102–106, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.2.
102.

Geoghegan:1962:GPC

[Geo62] D. Geoghegan. Gabriel Plattes’ caveat for alchymists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(2):97–102, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1962.10.2.
97.

Gil-Farrero:2023:TPP

[GF23] Judit Gil-Farrero. Tóxicos:

pasado y presente. pensar históricamente un mundo tóxico. (Spanish) [Toxic: past and present. Thinking historically about a toxic world]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):100–102, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2140979.

Grapi:1997:BCC

[GI97] Père Grapí and Mercè Izquierdo. Berthollet’s conception of chemical change in context. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(3):113–130, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.3.
                                        113.

Gianfrancesco:2018:BGE

[Gia18] Lorenza Gianfrancesco. Books, gold, and elixir: Alchemy and religious orders in early modern Naples. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):250–274, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1512777.

Gibbs:1958:BCW

[Gib58] F. W. Gibbs. Boerhaave’s chemical writings. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3):117–135, 1958. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1958.6.3.
                                        117.

Gibbs:1960:DJF

[Gib60a] F. W. Gibbs. Dr. Johnson’s first published work? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(1):24–34, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.1.
24.

Gibbs:1960:ILN

[Gib60b] F. W. Gibbs. Itinerant lecturers in natural philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(2):111–117, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.2.
111.

Gibbs:1966:LE

[Gib66] F. W. Gibbs. Letters to the Editor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):108–117, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2013.13.2.
006.

Guillem-Llobat:2015:SRT

[GL15] Ximo Guillem-Llobat. Science and the regulation of toxicants in historical perspective. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(3):287–291, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698015Z.
00000000084.

Guillem-Llobat:2017:ANP

[GL17] Ximo Guillem-Llobat. Autonomous nature: Problems of prediction and control from ancient times to the scientific revolution. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):194–195, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1346954.

Guillem-Llobat:2018:BRO

[GL18] Ximo Guillem-Llobat. Book review: One Hundred Years of Chemical Warfare: Research, Deployment, Consequences. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):411–412, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1520677.

Guillem-Llobat:2021:FVM

[GL21] Ximo Guillem-Llobat. ‘?Entre el fiscal y el verdugo? Mateu Orfila i Rotger (1787–1853) y la toxicología del siglo XIX.

(Spanish) [Between the prosecutor and the executioner? Mateu Orfila i Rotger (1787– 1853) and the toxicology of the XIX century]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):325–326, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1919457.

Golovnya:1976:IIC

[GLCK76] V. A. Golovnya, T. N.

Leonova, Wayne Craig, and George B. Kauffman. Il’ya Il’ich Chernyaev (1893–1966): Some recollections of his work and personality. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):187–198, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.3.
187.

Goldschmidt:1957:SMF

[Gol57] Leontine Goldschmidt. The symbolic meaning of Fahrenheit’s temperature scale. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(2):107–108, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.2.

Goldsmith:1980:OCD

[Gol80] Robert H. Goldsmith. Origins of the so-called Dewar benzene formula. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(2):136–141, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.2.
                                        136.

Golinski:1983:PSC

[Gol83] Jan V. Golinski. Peter Shaw: Chemistry and communication in Augustan England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(1):19–29, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1983.30.1.
                                        19.

Golinski:2016:CCE

[Gol16] Jan Golinski. The cradle of chemistry: The early years of chemistry at the University of Edinburgh. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):189–190, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1227168.

Golinski:2017:PHC

[Gol17] Jan Golinski. The personality of Henry Cavendish: a great scientist with extraordinary peculiarities. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):287–288, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385899.

Golinski:2018:BRC

[Gol18] Jan Golinski. Book review:

Compound Histories: Materials, Governance and Production, 1760–1840. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):402–404, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1488125.

Golinski:2019:FTU

[Gol19] Jan Golinski. “The fitness of their union”: Travel and health in the letters of Humphry and Jane Davy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):181–194, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1616945.

Gomori:1977:NIJ

[Gom77] George Gomori. New information on Janos Banfihunyadi’s life. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):170–174, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.3.
                                        170.

Goodman:1969:PCE

[Goo69] D. C. Goodman. Problems in crystallography in the early nineteenth century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(3):152–166, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1969.16.3.
                                        152.

Gorman:1977:SCT

[Gor77] Mel Gorman. A survey of the chemical translations of John Fryer in Nineteenth-Century China. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(2):89–95, 1977. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1977.24.2.89.

Gorman:1983:SWB

[Gor83] Mel Gorman. Sir William B.

O’Shaughnessy, pioneer chemical educator in India. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(2):107–116, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.2.
107.

Gordin:1998:MNM

[Gor98] Michael D. Gordin. Making Newtons: Mendeleev, metrology, and the chemical ether. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):96–115, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.2.
96.

Gordin:2004:WOT

[Gor04]  Michael D. Gordin.  A

Well-ordered Thing: Dmitrii Mendeleev and the Shadow of the Periodic Table. Basic Books, New York, NY, USA, 2004. ISBN 0-465-02775-X. xx + 364 pp. LCCN QD22.M43 G67 2004. URL

http://www.loc.gov/catdir/
enhancements/fy0832/2003025533-
b.html;   http://www.loc.
gov/catdir/enhancements/
fy0832/2003025533-d.html;
 http://www.loc.gov/catdir/
toc/ecip0411/2003025533.
html.

Gordin:2018:PTP

[Gor18] Michael D. Gordin. Paper tools and periodic tables: Newlands and Mendeleev draw grids. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(1):30–51, 2018. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1418251.

Gorman:2019:BRF

[Gor19] Cassandra Gorman. Book review: Fictional Matter: Empiricism, Corpuscles, and the Novel. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):87–88, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1554616.

Gorman:2021:GNP

[Gor21] Cassandra Gorman. Grounds of natural philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):451–453, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1921450.

Gough:1982:SER

[Gou82] Jerry B. Gough. Some early references to revolutions in chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(2):106–109, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.2.
106.

Gough:1983:LMN

[Gou83] J. B. Gough. Lavoisier’s memoirs on the nature of water and their place in the chemical revolution. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(2):89–106, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.2.
89.

Gregory:1938:CAN

[Gre38] Joshua C. Gregory. Chemistry and alchemy in the natural philosophy of Sir Francis Bacon, 1561–1626. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(2):93–111, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1938.2.2.
93.

Greenaway:1990:R

[Gre90] Frank Greenaway. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(1):54–60, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1990.37.1.
                                        54.

Grossman:2021:JDA

[Gro21] Mark I. Grossman. John Dalton’s “Aha” moment: the origin of the chemical atomic theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(1):49–71, 2021. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1868861.

Grund:2006:FMA

[Gru06] Peter Grund. “ffor to make Azure as Albert biddes”: Medieval English alchemical writings in the Pseudo-Albertan tradition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):21–42, 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582306X93174.

Grund:2009:TAT

[Gru09] Peter Grund. Textual alchemy: The transformation of Pseudo-Albertus Magnus’s Semita Recta into the Mirror of Lights. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(3):202–225, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269809X12529013496257.

Gibbs:1961:TBO

[GS61] F. W. Gibbs and W. A.

Smeaton. Thomas Beddoes at Oxford. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(1):47–49, 1961. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1961.9.1.47.

Guerrini:1989:JKO

[GS89] Anita Guerrini and Jole R.

Shackelford. John Keill’s de Operationum Chymicarum Ratione Mechanica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3):138–152, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.3.
138.

Garcia-Sancho:2016:RUG

[GS16] Miguel García-Sancho. The recombinant university: Genetic engineering and the emergence of Stanford biotechnology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):275–276, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1246809.

Guerlac:1976:CRW

[Gue76] Henry Guerlac. The chemical revolution: a word from Monsieur Fourcroy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(1): 1–4, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.1.
1.

Guerra:2015:IYD

[Gue15] Corinna Guerra. If you don’t have a good laboratory, find a good volcano: Mount Vesuvius as a natural chemical laboratory in Eighteenth-Century Italy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(3):245–265, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823415Y.
0000000005.

Guinsburg:1980:HMT

[Gui80] Arlene Miller Guinsburg. Henry more, Thomas Vaughan and the Late Renaissance magical tradition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(1):36–58, 1980. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.1.
36.

Gortler:2017:PPB

[GW17] Leon Gortler and Stephen J.

Weininger. Private philanthropy and basic research in mid-Twentieth Century America: The Hickrill Chemical Research Foundation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):66–94, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1305150.

Hagendijk:2016:HPW

[Hag16] Thijs Hagendijk. Humboldts Preußen. Wissenschaft und Technik im Aufbruch. (German) [Humboldt’s Prussia. Science and technology on the move]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):190–191, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1227170.

Hall:1996:R

[Hal96] Marie Boas Hall. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(2):101–128, 1996. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.2.
101.

Hallum:2009:TIA

[Hal09] B. C. Hallum. The Tome of Images: an Arabic compilation of texts by Zosimos of Panopolis and a source of the Turba Philosophorum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):76–88, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X405255.

Hallum:2011:ERS

[Hal11] Bink Hallum. Essay reviews: The Secret History of Hermes Trismegistus: Hermeticism from Ancient to Modern Times. By Florian Ebeling, with a foreword by Jan Assmann. Pp. xiii + 158, illus., index. Cornell University Press: Ithaca and London. 2007. $29.95; £15.95. ISBN 978-0-8014-4546-0 (hbk). The Arabic Hermes: from Pagan Sage to Prophet of Science. By Kevin Van Bladel. Pp. xii + 278, index. Oxford University Press: Oxford. 2009. £45. ISBN 978-0-19-537613-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(1):78–81, 2011. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582311X12947034675631.

Hamarneh:1982:AIA

[Ham82] Sami K. Hamarneh. Arabic– Islamic alchemy — three intertwined stages. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(2):74–87, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.2.
74.

Hamlin:2017:HLW

[Ham17] Christopher Hamlin. The history of the London water industry, 1580–1820. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):381–382, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1405895.

Hannaway:1967:JCB

[Han67] O. Hannaway. Johann Conrad Barchusen (1666—1723)

— contemporary and rival of Boerhaave. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(2):96–111, 1967. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1967.14.2.96.

Hannaway:1976:GMC

[Han76] Owen Hannaway. The German model of chemical education in America: Ira Remsen at Johns Hopkins (1876–1913). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):145–164, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1976.23.3.
                                        145.

Harley:2000:RBT

[Har00] David Harley. Rychard Bostok of Tandridge, Surrey (c. 1530– 1605), M.P., Paracelsian propagandist and friend of John Dee. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(1):29–36, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.1.
                                        29.

Haschmi:1961:BAA

[Has61] Mohamed Yahia Haschmi. The beginning of Arab alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(3):155–161, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1961.9.3.
                                        155.

Hayes:1977:AIJ

[Hay77] Thomas W. Hayes. Alchemical imagery in John Donne’s “A Nocturnall Upon S. Lucies Day”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(1):55–62, 1977. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1977.24.1.55.

Hunt:1984:RKF

[HB84] Leslie B. Hunt and Peta D.

Buchanan. Richard Knight (1768–1844): a forgotten chemist and apparatus designer. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(2):57–67, 1984. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1984.31.2.57.

Heilbron:2003:OCH

[HBB^{+}03] J. L. Heilbron, James Bartholomew, Jim Bennett, Frederic L. Holmes, Rachel Laudan, and Giuliano Pancaldi, editors. The Oxford Companion to the History of Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-511229-6. xxviii + 941 pp. LCCN Q125 .O86

2003. URL http://www.loc.
gov/catdir/enhancements/

                                        fy0613/2002153783-d.html;
                                         http://www.oxfordreference.
                                        com/views/BOOK_SEARCH.html?
                                        book=t124.

Hunter:2001:CRB

[HCP01] Michael Cyril William Hunter, Antonio Clericuzio, and Lawrence Principe, editors. The Correspondence of Robert Boyle, 1636–1691. Pickering and Chatto, London, UK, 2001. ISBN 0-19-263241-8 (paperback), 0-19-829329-1, 0-19-850786-0 (hardcover), 0-19-850787-9 (paperback), 0-19-263242-6 (hardcover), 1-85196-125-9, 0-19-850608-2, 1-85196-125-9 (set). li + 521 (vol. 1); viii + 674 (vol. 2); vii + 445 (vol. 3); vii + 528 (vol. 4); viii + 474 (vol. 5); viii + 616 (vol. 6) pp. LCCN Q143.B77 A4 2001.

Hunter:1999:WRB

[HD00] Michael Cyril William Hunter and Edward Bradford Davis, editors. The works of Robert Boyle. Pickering and Chatto, London, UK, 1999–2000. ISBN 1-85196-109-7 (set), 1-85196-522-X (vol. 1–7 of set), 1-85196-523-8 (vol. 8–14 of set). ???? pp. LCCN Q155 .B667 1999.

Hedesan:2013:AEM

[Hed13a] Georgiana Hedesan. Alchemy in Early Modern England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(3):285–288, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000034.

Hedesan:2013:RTL

[Hed13b] Georgiana D. Hedesan. Reproducing the tree of life: Radical prolongation of life and biblical interpretation in Seventeenth-Century medical alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(4):341–360, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000038.

Hedesan:2017:PMR

[Hed17] Georgiana Hedesan. La Porta Magica di Roma, simbolo dell’alchimia occidentale. (Italian) [The Magic Door of Rome, symbol of Western alchemy]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):284–285, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385898.

Hedesan:2021:ILT

[Hed21] Georgiana D. Hedesan. The influence of Louvain teaching on Jan Baptist Van Helmont’s adoption of Paracelsianism and alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):231–246, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1922020.

Hedesan:2022:BRT

[Hed22] Georgiana Hedesan. Book review: Turba Philosophorum: Congrès pythagoricien sur l’art d’Hermès. By Grégoire Lacaze. Pp. 663, index. Brill: Leiden. 2018. £243.00. ISBN: 978-90-04-36032-7. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):326–327, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2066258.

Heimann:1973:NPW

[Hei73] P. M. Heimann. “Nature is a perpetual worker”: Newton’s aether and eighteenthcentury natural philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1):1–25, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.1.
                                        1.

Heinecke:1995:MSJ

[Hei95] Berthold Heinecke. The mysticism and science of Johann Baptista Van Helmont (1579–1644). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(2):65–78, 1995. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1995.42.2.65.

Heitsch:2016:DCH

[Hei16] Dorothea Heitsch. Descartes, cardiac heat, and alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):285–303, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/00026980.2016.
1279401.

HemmungsWirten:2015:MMC

[Hem15] Eva Hemmungs Wirtén. Making Marie Curie: Intellectual Property and Celebrity Culture in an Age of Information. Science.culture. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-23584-X (hardcover), 0-226-23598-X (e-book). 223 (est.) pp. LCCN QD22.C8 H46 2015.

Hendry:2003:R

[Hen03] Robin Findlay Hendry. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(1):116–143, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.1.
116.

Henry:2017:MWC

[Hen17] John Henry. Magic in Western culture: From antiquity to the Enlightenment. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):190–191, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1346951.

Hendriksen:2018:BMC

[Hen18] Marieke M. A. Hendriksen. Boerhaave’s mineral chemistry and its influence on Eighteenth-Century pharmacy in The Netherlands and England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):303–323, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1488099.

Hendriksen:2019:BRS

[Hen19] Marieke Hendriksen. Book review: The Structures of Practical Knowledge. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):88–90, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2018.1557790.

Hendriksen:2020:TCH

[Hen20] Marieke Hendriksen. Technology: Critical history of a concept. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):101–103, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1662578.

Hershbell:1987:DBG

[Her87] Jackson P. Hershbell. Democritus and the beginnings of Greek alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(1):5–20, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.1.
5.

Herran:2014:NMN

[Her14] Néstor Herran. Nuclear markets, nuclear bodies. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):407–410, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000066.

Herran:2016:RSL

[Her16] Nestor Herran. Radium and the secret of life. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):355–356, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279413.

Herran:2018:BRP

[Her18] Nestor Herran. Book review: Polonium in the Playhouse: The Manhattan Project’s Secret Chemistry Work in Dayton, Ohio. Linda Carrick Thomas. Pp. 247, illus., index. The Ohio State University Press: Columbus. 2017. £23 (cloth). ISBN: 978-0-8142-1338-4. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):407–408, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1518681.

Heym:1937:ACH

[Hey37a]  Gerard Heym.  The Au-

rea Catena Homeri. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):78–83, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1937.1.1.
78.

Heym:1937:IBA

[Hey37b] Gerard Heym. An introduction to the bibliography of alchemy — Part I. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):48–60, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1937.1.1.
48.

Heym:1938:ARA

[Hey38a] Gerard Heym. Al-Razi and alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):184–191, 1938. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1938.1.3.184.

Heym:1938:AJE

[Hey38b] Gerard Heym. An alchemical journal of the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):197–199, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1938.1.3.
197.

Heym:1969:MS

[Hey69] Gerard Heym. Michael Scot.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(3):167–172, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.3.
167.

Hedesan:2014:R

[HFL^{+}14] Georgiana D. Hedesan, Gabriele Ferrario, Christine Lehman, Luc Peterschmitt, Barbara Orland, Ana Carneiro, Charlotte Bigg, Ximo Guillem-Llobat, David Knight, and Anna Simmons. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):199–209, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000054.

Hargittai:2000:CSC

[HH00] István Hargittai and Magdolna Hargittai, editors. Candid Science: Conversations with Famous Chemists. Imperial College Press, London, UK, 2000. ISBN 1-86094-151-6 (hardcover), 1-86094-228-8 (paperback). xii + 516 pp. LCCN QD21 .H294 2000.

Hargittai:2005:CSV

[HH05] Balazs Hargittai and István Hargittai, editors. Candid Science V: Conversations with Famous Scientists. Imperial College Press, London, UK, 2005. ISBN 1-86094-505-8 , 1-86094-506-6 (paperback). xvi + 695 pp. LCCN Q141 .H264 2005. URL http://www.

loc.gov/catdir/toc/fy0602/
2004062538.html.

Hijmans:2021:THP

[Hij21] Sarah Hijmans. Technoscience in history: Prussia, 1750-1850. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68 (2–3):326–327, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1893516.

Hijmans:2022:TMN

[Hij22]  Sarah N. Hijmans.  The

tantalum metals (1801–1866): Nineteenth-Century analytical chemistry and the identification of chemical elements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):399–419, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2133806.

Hill:1975:IL

[Hil75] C. R. Hill. The iconography of the laboratory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):102–110, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1975.22.2.
                                        102.

Hirsh:1981:CPD

[Hir81] Richard F. Hirsh. A conflict of principles: The discovery of argon and the debate over its existence. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):121–130, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1981.28.3.
                                        121.

Hirai:2021:DSC

[Hir21] Hiro Hirai. Daniel Sennert, chymistry, and theological debates. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68 (2–3):198–213, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1936354.

Harle:2002:HEA

[HJ02] Anna E. J. Harle and Claus Jacob. Historical and epistemological aspects of optical isomerism on trivalent sulfur and nitrogen atoms. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(2):127–147, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.2002.49.2.
127.

Ho:1959:LEE

[HN59] Ho Ping-Yü and Joseph Needham, F. R. S. The laboratory equipment of the Early Mediæval Chinese alchemists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(2):58–112, June 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1959.7.2.
58.

Holmyard:1956:OFS

[Hol56]  E. J. Holmyard.  Obitu-

ary: Frank Sherwood Taylor (1897–1956). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):57–58, 1956. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1956.5.3-
4.57.

Holloway:1989:ECM

[Hol89] S. W. F. Holloway. Eighteenth century medics. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3):152–153, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.3.
152.

Home:1995:ERC

[Hom95] R. W. Home. Essay review: The chemistry of light. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(1):50–51, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1995.42.1.
                                        50.

Homburg:1999:RAC

[Hom99] Ernst Homburg. The rise of analytical chemistry and its consequences for the development of the German chemical profession (1780–1860). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(1):1–32, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1999.46.1.
                                        1.

Homburg:2001:SN

[Hom01] Ernst Homburg. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(3):209–212, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.3.
                                        209.

Homburg:2005:SCE

[Hom05] Ernst Homburg. Shifting centres and emerging peripheries: Global patterns in Twentiethcentury chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):3–6, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X52878.

Homburg:2017:AVT

[Hom17a]  Ernst Homburg.  Adding

value: Tasmanian alkaloids 1975–2015. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):385–386, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1405898.

Homburg:2017:LGE

[Hom17b] Ernst Homburg. Aus der Luft gewonnen. Die Entwicklung der globalen Gaseindustrie, 1880 bis 2012/Building on air. The international industrial gases industry, 1886–2006. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):289–291, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385889.

Homburg:2019:BRP

[Hom19] Ernst Homburg. Book review: Le Patrimoine Industriel de la Chimie. A Special Issue of Patrimoine Industriel

— Archéologie — Technique

— Mémoire, no. 69, by Florence Hachez-Leroy. Pp. 152, illus. Cilac: Paris. 2016. £22.00. ISSN: 0220-5521. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):362–364, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1592804.

Horne:1960:AAG

[Hor60] R. A. Horne. Atomism in Ancient Greece and India. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(2):98–110, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1960.8.2.
                                        98.

Horrocks:2011:WWI

[Hor11] Sally Horrocks. World War II, post-war reconstruction and British women chemists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(2):150–170, 2011. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/174582311X13008456751026.

Horrocks:2018:BRC

[Hor18] Sally M. Horrocks. Book review: A Chemical Passion: The Forgotten Story of Chemistry at British Independent Girls’ Schools, 1820s–1930s. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):398–399, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1499428.

Holmes:2003:RBR

[HRR03] Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors. Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1039-7 (hardcover), 0-306-48152-9 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xv + 325 pp. LCCN Q225.5 .R475

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0814/2002041330-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0814/2002041330-
t.html.

Huffman:1978:RFD

[HS78] William H. Huffman and Robert A. Seelinger. Robert Fludd’s “Declaratio Brevis” to James I. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(2):69–92, 1978. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1978.25.2.69.

Hollister-Short:1988:R

[HS88] Graham Hollister-Short. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(1):47–52, 1988. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1988.35.1.47.

Hepler-Smith:2015:JSF

[HS15] Evan Hepler-Smith. “Just as the structural formula does”: Names, diagrams, and the structure of organic chemistry at the 1892 Geneva Nomenclature Congress. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1):1–28, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/1745823414Y.
                                        0000000006.

Hepler-Smith:2018:PCF

[HS18]  Evan Hepler-Smith.  Paper

chemistry: François Dagognet and the chemical graph. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(1):76–98, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1418232.

Hepler-Smith:2020:ECN

[HS20] Evan Hepler-Smith. The etymology of chemical names: Tradition and convenience vs. rationality in chemical nomenclature. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):202–203, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1718303.

Hepler-Smith:2023:BRC

[HS23] Evan Hepler-Smith. Book review: A Cultural History of Chemistry in the Modern Age (vol. 6). Edited by Peter Morris. Pp. 271 + xiv, illus., index. Bloomsbury: London, 2022. £395.00 (6-volume set). ISBN: 978-1-4742-9481-2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):197–199, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2186585.

Hudson:1998:NA

[Hud98] John Hudson. News and announcements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):129–130, 1998.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.2.
                                        129.

Hudson:1999:RMN

[Hud99] John Hudson. Report of meeting: ‘new chemical biography’. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(2):104–105, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1999.46.2.
                                        104.

Hudson:2001:A

[Hud01]  John Hudson.  Announce-

ment. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(3): 190, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.3.
                                        190.

Hudson:2002:R

[Hud02] John Hudson. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(1):72–94, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2002.49.1.

Hunt:1979:FM

[Hun79] Leslie B. Hunt. The Figuiers of Montpellier. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):221–223, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.3.
221.

Hunter:1999:PLT

[Hun99] Graeme K. Hunter. Phoebus Levene and the tetranucleotide structure of nucleic acids. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(2):73–103, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1999.46.2.
73.

Hutchison:1992:R

[Hut92] Keith Hutchison. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):144–157, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.3.
144.

Hagendijk:2020:MFT

[HVD20] Thijs Hagendijk, Márcia Vilarigues, and Sven Dupré. Materials, furnaces, and texts:

How to write about making glass colours in the Seventeenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):323–345, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1826823.

Iacob:2022:MPT

[Iac22] Anisia Alis Iacob. Book review: Making Physicians: Tradition, Teaching, and Trials at Leiden University, 1575–1639. by Evan R. Ragland, Brill: Leiden, 2022. pp. 457+xii, £120.00, illus., index. ISBN: 978-90-04-46511-4. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):431–433, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2107318.

Irish:2013:BCC

[Iri13] Stephen T. Irish. Brodie’s calculus and chemical classification. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(3):234–254, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000031.

Irish:2017:CSC

[Iri17] Stephen T. Irish. The corundum stone and crystallographic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):301–325, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1413163.

Irish:2018:JSC

[Iri18] Stephen T. Irish. James Smithson on the calamines: Chemical combination in crystals. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):373–397, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1520521.

Irish:2020:GEM

[Iri20] Stephen T. Irish. Gems in the early modern world: Materials, knowledge and global trade, 1450–1800. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):100–101, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1619028.

Jacobsen:2000:WHR

[Jac00] Anja Skaar Jacobsen. A. W.

Hauch’s role in the introduction of antiphlogistic chemistry into Denmark. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(2):71–95, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.2.
71.

James:1981:LWC

[Jam81] Frank A. J. L. James. The letters of William Crookes to Charles Hanson Greville Williams 1861–2: The detection and isolation of thallium. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):131–157, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.3.
131.

James:1983:ESC

[Jam83a] Frank A. J. L. James. The establishment of spectro-chemical analysis as a practical method of qualitative analysis, 1854– 1861. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(1):30–53, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.1.
30.

James:1983:SSS

[Jam83b] Frank A. J. L. James. The study of spark spectra 1835– 1859. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(3):137–162, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.3.
137.

James:1985:DLS

[Jam85] Frank A. J. L. James. The discovery of line spectra. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(2):53–70, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.2.
53.

James:1995:SCO

[Jam95] Frank A. J. L. James. Science as a cultural ornament: Bunsen, Kirchhoff and Helmholtz in Mid-Nineteenth-Century Baden. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(1): 1–9, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1995.42.1.
1.

James:2015:ACP

[Jam15] Frank A. J. L. James. “Agricultural chymistry is at present in it’s infancy”: The Board of Agriculture, The Royal Institution and Humphry Davy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(4):363–385, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/00026980.2015.
1107953.

James:2019:CHD

[Jam19a] Frank A. J. L. James. Constructing Humphry Davy’s biographical image. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):214–238, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1620986.

James:2019:HDE

[Jam19b] Frank A. J. L. James. Humphry Davy’s early chemical knowledge, theory and experiments: an edition of his 1798 manuscript, “An Essay on Heat and the Combinations of Light” from The Royal Institution of Cornwall, Courtney Library, MS DVY/2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):303–345, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2019.1684080.

James:2021:CSC

[Jam21] Frank A. J. L. James. A chemical satire on the 1809 change of government in Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):442–446, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1983694.

James:2022:CWS

[Jam22] Frank A. J. L. James. Carl Wilhelm Scheele and Torbern Bergman: The science, lives and friendship of two pioneers in chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):196–197, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2043613.

Janacek:2020:HRA

[Jan20] Bruce Janacek. The historic role of alchemy in the Holy Roman Empire. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):408–415, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1806518.

Jas:2014:CEH

[Jas14] Nathalie Jas. Chemicals and environmental history. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):194–198, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698014Z.
                                        00000000053.

Jennings:1981:LVP

[Jen81] Richard C. Jennings. Lavoisier’s views on phlogiston and the matter of fire before about 1770. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):206–209, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1981.28.3.
                                        206.

Jensen:2004:PUP

[Jen04] William B. Jensen. A previously unrecognised portrait of Joan Baptista Van Helmont (1579–1644). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(3):263–268, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2004.51.3.
                                        263.

Jewess:2019:BRS

[Jew19] Michael Jewess. Book review:

For Science, King and Country: The Life and Legacy of Henry Moseley. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):85–87, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1548411.

Johnson:2011:CCC

[Joh11] Jeffrey Allan Johnson. Crisis, change and creativity in science and technology: Chemistry in the aftermath of Twentiethcentury global wars. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(2):101–115, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582311X13008456750982.

Johnson:2018:CID

[Joh18] Jeffrey A. Johnson. Chemiker im “Dritten Reich”: Die Deutsche Chemische Gesellschaft und der Verein Deutscher Chemiker im NS-Herrschaftsapparat. (German) [Chemists in the “Third Reich”: The German Chemical Society and the German Chemical Union in the National Socialist apparatus of domination]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2):

193–194, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Jones:1993:JLE

[Jon93] Paul R. Jones. Justus Von Liebig, Eben Horsford and the development of the baking powder industry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(2):65–74, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.2.
                                        65.

Jones:2018:SFM

[Jon18] Peter Murray Jones. The survival of the Frater Medicus? English friars and alchemy, ca. 1370–ca. 1425. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):232–249, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1512780.

Jones:2021:EFI

[Jon21] Peter Murray Jones. The experimental fire: Inventing English alchemy, 1300–1700. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):319–321, 2021. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2021.1919458.

Josten:1949:TJD

[Jos49a]  C. H. Josten.  The text

of John Dastin’s ‘Letter to Pope John XXII’. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(1–2):34–51, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1949.4.1-
2.34.

Josten:1949:TGH

[Jos49b] C. H. Josten. Truth’s golden harrow: an unpublished alchemical treatise of Robert Fludd in the Bodleian Library. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(3–4):91–150, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.3.3-
4.91.

Josten:1949:WBS

[Jos49c] C. H. Josten. William Backhouse of Swallowfield. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(1–2):1–33, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.4.1-
2.1.

Josten:1963:RFP

[Jos63] C. H. Josten. Robert Fludd’s “philosophicall key” and his alchemical experiment on wheat. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(1):1–23, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1963.11.1.
                                        1.

Josten:1964:TJD

[Jos64] C. H. Josten. A translation of John Dee’s “Monas Hieroglyphica” (Antwerp, 1564), with an introduction and annotations. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 12(2–3):84–221, 1964. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269864790223101.

James:2019:NSH

[JR19] Frank A. J. L. James and Sharon Ruston. New studies on Humphry Davy: Introduction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):95–102, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1625238.

Kahn:2010:APM

[Kah10] Didier Kahn. Alchemical poetry in Medieval and Early Modern Europe: a preliminary survey and synthesis. Part I — preliminary survey. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(3):249–274, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582310X12849808295788.

Kahn:2011:APM

[Kah11] Didier Kahn. Alchemical poetry in Medieval and Early Modern Europe: a preliminary survey and synthesis. Part II

— synthesis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(1):62–77, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582311X12947034675514.

Kahn:2013:THJ

[Kah13] Didier Kahn. Towards a history of Joseph Du Chesne’s manuscripts. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(1):25–30, 2013. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698012Z.
00000000021.

Kahn:2019:BRL

[Kah19] Didier Kahn. Book review: Lux in Tenebris: The Visual and the Symbolic in Western Esotericism. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):82–83, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1529899.

Kahn:2021:FPP

[Kah21] Didier Kahn. The first private and public courses of chymistry in Paris (and Italy) from Jean Beguin to William Davisson. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):247–272, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1922019.

Kaji:2013:PCP

[Kaj13] Masanori Kaji. Philosophy of chemistry and the periodic table. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):179–181, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000029.

Kapoor:1969:DOC

[Kap69] Satish C. Kapoor. Dumas and organic classification. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(1–2):1–65, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.1-
2.1.

Karpenko:1978:R

[Kar78] Vladimír Karpenko. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(3):213–222, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.3.
213.

Karpenko:1980:DSN

[Kar80] V. Karpenko. The discovery of supposed new elements: Two centuries of errors. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(2):77–102, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.2.
77.

Karpenko:1988:CMM

[Kar88] V. Karpenko. Coins and medals made of alchemical metal. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(2):65–76, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.2.
65.

Karpenko:1990:CBL

[Kar90a] Vladimir Karpenko. Christoph Bergner: The last Prague alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(3):116–120, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.3.
116.

Karpenko:1990:OAM

[Kar90b] VladimÍr Karpenko. The oldest alchemical manuscript in the Czech language. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(2):61–73, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.2.
61.

Karpenko:1992:CMT

[Kar92] Vladimír Karpenko. The chemistry and metallurgy of transmutation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(2):47–62, 1992. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1992.39.2.47.

Karpenko:1993:BMS

[Kar93] Vladimír Karpenko. Between magic and science: Numerical magic squares. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(3):121–128, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.3.
121.

Karpenko:2003:SMA

[Kar03] Vladimír Karpenko. Systems of metals in alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(2):208–230, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.2.
208.

Karpenko:2007:AGG

[Kar07] Vladimír Karpenko. Not all that glitters is gold: Gold imitations in history. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):172–191, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582307X212323.

Kauffman:1972:STN

[Kau72] George B. Kauffman. The stereochemistry of trivalent nitrogen compounds: Alfred Werner and the controversy over the structure of oximes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(2):129–144, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.2.
129.

Kauffman:1973:AWT

[Kau73]  George B. Kauffman.  Al-

fred Werner’s theory of acids, bases, and hydrolysis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1):53–66, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.1.
53.

Kauffman:1983:MSH

[Kau83] George B. Kauffman. The mystery of Stephen H. Emmens: Successful alchemist or ingenious swindler? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(2):65–88, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.2.
65.

Kauffman:1993:R

[Kau93] George B. Kauffman. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(2):99–118, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.2.
99.

Kauffman:1994:R

[Kau94] George B. Kauffman. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(2):108–119, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1994.41.2.
108.

Kronberg:1981:AHD

[KCU81] B. I. Kronberg, L. L. Coatsworth, and M. C. Usselman. The artifact as historical document. Part 2: The palladium and rhodium of W. H. Wollaston. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(1):20–35, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.1.

Keas:1992:KAS

[Kea92] Michael N. Keas. Karl Aloys Schenzinger’s novel, Anilin: Chemistry and Chemical Technology in Nazi Literaturpolitik. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):127–140, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1992.39.3.
                                        127.

Keene:2013:CCF

[Kee13] Melanie Keene. From candles to cabinets: “familiar chemistry” in Early Victorian Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(1):54–77, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698012Z.
                                        00000000020.

Keller:1971:STS

[Kel71] Alex G. Keller. The scientific and technological sages of Ancient China. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(1):49–55, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.1.
                                        49.

Keller:1978:R

[Kel78] A. G. Keller. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(2):148–152, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.2.
148.

Keller:1989:R

[Kel89] A. G. Keller. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(2):102–108, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.2.
102.

Keller:1997:R

[Kel97] A. G. Keller. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(3):148–163, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.3.
148.

Keller:2012:APA

[Kel12]  Vera Keller.  The author-

ity of practice in the alchemy of Sir John Heydon (1588– 1653). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(3):197–217, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13457672281740.

Keller:2014:HAC

[Kel14] Vera Keller. Hermetic atomism: Christian Adolph Balduin (1632–1682), Aurum Aurae, and the 1674 phosphor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):366–384, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823414Y.
0000000003.

Klein-Franke:1970:KAL

[KF70] Felix Klein-Franke. The knowledge of Aristotle’s lapidary during the Latin Middle Ages. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(3):137–142, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.3.
137.

Klein-Franke:1973:GAB

[KF73] Felix Klein-Franke. The geomancy of Ahmad B. ’Ali Zunbul: a study of the Arabic Corpus Hermeticum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1):26–35, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.1.
26.

Kahn:2020:PFT

[KH20] Didier Kahn and Hiro Hirai.

Paracelsus, forgeries and transmutation: Introduction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):1–3, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1721764.

Kibre:1960:TAM

[Kib60] Pearl Kibre. Two alchemical miscellanies: Vatican Latin MSS. 4091, 4092. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(3):167–176, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.3.
167.

Kikuchi:2009:SCC

[Kik09] Yoshiyuki Kikuchi. Samurai chemists, Charles Graham and Alexander William Williamson at University College London, 1863–1872. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(2):115–137, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582309X441390.

Kikuchi:2011:WWI

[Kik11]  Yoshiyuki Kikuchi.  World

War I, international participation and reorganisation of the Japanese chemical community. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(2):136–149, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582311X13008456751107.

Kim:1991:ALR

[Kim91] Yung Sik Kim. Another look at Robert Boyle’s acceptance of the mechanical philosophy: Its limits and its chemical and social contexts. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(1):1–10, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1991.38.1.
                                        1.

Kim:1996:CSS

[Kim96] Mi Gyung Kim. Constructing symbolic spaces: Chemical molecules in the Académie des sciences. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(1):1–31, 1996. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1996.43.1.1.

Kim:2005:TCT

[Kim05] Dong-Won Kim. Two chemists in two Koreas. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):67–84, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X52913.

Kim:2020:BRN

[Kim20] Fred Kim. Book review: Natural knowledge and Aristotelianism at Early Modern Protestant Universities, Edited by Pietro Daniel Omodeo and Volkhard Wels. Pp. 342, illus. Harrassowitz Verlag: Wiesbaden. 2019. £65, free ebook on publisher’s website. ISBN 978-3-447-11265-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):418–419, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1806516.

King:1981:ETP

[Kin81] M. Christine King. Experiments with time: Progress and problems in the development of chemical kinetics.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(2):70–82, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.2.
70.

King:1982:ETP

[Kin82] M. Christine King. Experiments with time: Progress and problems in the development of chemical kinetics. Part 2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(1):49–61, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.1.
49.

King:1984:CCC

[Kin84] M. Christine King. The course of chemical change: the life and times of Augustus G. Vernon Harcourt (1834–1919). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(1):16–31, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.1.
16.

Kirsop:1961:LBP

[Kir61]  Wallace Kirsop.  The leg-

end of Bernard Palissy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(3):136–154, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.3.
136.

Kirschke:2003:LHU

[Kir03] Martin Kirschke. Liebig, his University Professor Karl Wilhelm Gottlob Kastner (1783– 1857), and his problematic relation with romantic natural philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(1):3–24, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.1.
3.

Kirby:2018:BRT

[Kir18] Jo Kirby. Book review: Treasures from the Sea: Sea Silk and Shellfish Purple Dye in Antiquity. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):405–407, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1508948.

Kirby:2023:MPC

[Kir23] Jo Kirby. March of the pigments: Color history, science and impact. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):202–204, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2175959.

Klein:1995:FGT

[Kle95] Ursula Klein. E. F. Geoffroy’s table of different ‘rapports’ observed between different chemical substances — a reinterpretation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(2):79–100, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1995.42.2.
79.

Klein:2004:ER

[Kle04] Ursula Klein. Essay review. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(1):77–78, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.1.
77.

Klein:2005:CLL

[Kle05] Ursula Klein. Contexts and limits of Lavoisier’s analytical plant chemistry: Plant materials and their classification. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52 (2):107–157, July 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X70102.

Klein:2013:CER

[Kle13] Ursula Klein. Chemical experts at the Royal Prussian porcelain manufactory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):99–121, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000025.

Klein:2014:CEP

[Kle14] Joel A. Klein. Corporeal elements and principles in the learned German chymical tradition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):345–365, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823414Y.
0000000002.

Klein:2015:DSP

[Kle15] Joel A. Klein. Daniel Sennert, the philosophical hen, and the epistolary quest for a (nearly-)universal medicine. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1):29–49, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/1745823414Y.
                                        0000000007.

Kiss:2006:AMN

[KLPG06] Farkas Gábor Kiss, Benedek Láng, and Cosmin Popa-Gorjanu. The alchemical mass of Nicolaus Melchior Cibinensis: Text, identity and speculations. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53 (2):143–159, July 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582306X117889.

Kauffman:1990:ACS

[KM90] George B. Kauffman and Ester Molayem. Alfonso Cossa (1833–1902), a self-taught Italian chemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(1):20–34, 1990. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1990.37.1.20.

Kahn:2014:JT

[KN14] Didier Kahn and William R. Newman. Joachim Telle (1939–2013). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):187–193, 2014.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000052.

Knight:1967:STD

[Kni67] David M. Knight. Steps towards a dynamical chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(3):179–197, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.3.
179.

Knight:1974:CPW

[Kni74] David M. Knight. Chemistry in palaeontology: The work of James Parkinson (1755–1824). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):78–85, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1974.21.1.
78.

Knight:1976:VF

[Kni76] D. M. Knight. The vital flame.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(1):5–15, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/

                                        abs/10.1179/amb.1976.23.1.
                                        5.

Knight:1986:ADC

[Kni86] David Knight. Accomplishment or dogma: Chemistry in the introductory works of Jane Marcet and Samuel Parkes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):94–98, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1986.33.2-
                                        3.94.

Knight:1994:R

[Kni94] David Knight. Reports. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):42–44, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.1.
                                        42.

Knight:2000:R

[Kni00] David Knight. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(2):118, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.2.
                                        118.

Knight:2004:BRO

[Kni04] David Knight. Book reviews:

The Oxford companion to the History of Modern Science. Edited by J. L. Heilbron. Pp. xxviii + 941, illus., index. Oxford University Press: Oxford. 2003. £80. ISBN 0-19-511229-6. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(1):79–95, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.1.
79.

Knight:2016:BRP

[Kni16a]  David Knight.  Book re-

view: Pure Intelligence: The Life of William Hyde Wollaston. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):191–192, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1227172.

Knight:2016:ESH

[Kni16b] David Knight. The experimental self: Humphry Davy and the making of a man of science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):350–351, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1279409.

Knight:2016:PCG

[Kni16c] David Knight. Philosophy of chemistry: Growth of a new discipline. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):82–83, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1201318.

Knight:2017:OIH

[Kni17]  David Knight.  The Ox-

ford illustrated history of science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):378–379, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1405892.

Knight:2019:SRD

[Kni19] David Knight. Sources and resources for Davy: 1960 and now. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66 (2–3):239–245, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1622259.

Knoeff:2001:MCC

[Kno01] Rina Knoeff. The making of a Calvinist chemist: Herman Boerhaave, God, fire and truth. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):102–111, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.2.
102.

Knoeff:2006:CMM

[Kno06] Rina Knoeff. Chemistry, mechanics and the making of anatomical knowledge: Boerhaave vs. Ruysch on the nature of the glands. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3): 201–219, November 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X148074.

Knoeff:2022:BRC

[Kno22] Rina Knoeff. Book review: De chemist. De geschiedenis van een verdwenen beroepsgroep, 1600–1800. By Henk Vermande. Pp. 662, illus., index. Verloren: Hilversum. 2021. £41.00. ISBN: 978-9-08-704968-3. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):331–332, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2071815.

Kohler:1975:LRA

[Koh75] Robert E. Kohler. Lavoisier’s rediscovery of the air from Mercury calx: a reinterpretation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):52–57, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1975.22.1.
                                        52.

Konecny:2013:SCS

[Kon13] Peter Konecný. Sites of chemistry in the Schemnitz Mining Academy and the Eighteenth-Century Habsburg Mining Administration. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):160–178, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000028.

Kopperl:1976:WRR

[Kop76] Sheldon J. Kopperl. T. W. Richards’ role in American graduate education in chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):165–174, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.1976.23.3.
165.

Kauffman:1978:DSC

[KP78] George B. Kauffman and Paul M. Priebe. The discovery of saccharin: a centennial retrospect. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(3):191–207, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.3.
191.

Klein:2014:IAS

[KR14] Joel A. Klein and Evan R.

Ragland. Introduction analysis and synthesis in Medieval and Early Modern Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):319–326, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000065.

Kragh:1989:ALN

[Kra89] Helge Kragh. The aether in late nineteenth century chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(2):49–65, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1989.36.2.
49.

Kragh:2016:BRL

[Kra16a] Helge Kragh. Book review: The Lost Elements: The Periodic Table’s Shadow Side. By Marco Fontani, Mariagrazia Costa, and Mary Virginia Orna. Pp. xxxvii + 531. Oxford University Press: Oxford. 2015. ISBN 0-19-938334-0. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):74–75, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1201303.

Krasnodebski:2016:DSF

[Kra16b] Marcin Krasnodebski. From distillation to standardization: a French perspective on the shaping of turpentine spirit (1909–1976). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):326–346, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1271574.

Kragh:2018:BRW

[Kra18] Helge Kragh. Book review: Wilhelm Ostwald. The Autobiography. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):296–297, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1469311.

Krasnodebski:2021:MLR

[Kra21] Marcin Krasnodebski. The meandering life of a research trajectory: Rare Earths in the aubervilliers research centre (1953-2020). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):385–406, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1983692.

Kraft:2022:RST

[Kra22a] Alexander Kraft. The reign of Saturn transformed into an age of gold. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):199–200, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2055902.

Kragh:2022:BHC

[Kra22b] Helge Kragh. Biographical histories of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):420–425, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2112850.

Kwa:2003:R

[Kwa03] Chunglin Kwa. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(3):310–343, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2003.50.3.
                                        310.

Lacey:2017:CCE

[Lac17] Andrew Lacey. The Chemical Club: an early Nineteenth-Century scientific dining club. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):263–282, 2017. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1080/00026980.2017.
                                        1398855.

Lacey:2019:NLJ

[Lac19] Andrew Lacey. New light on John Davy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):195–213, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1620985.

Lange:1966:ASC

[Lan66] Erwin F. Lange. Alchemy and the Sixteenth Century metallurgists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):92–95, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2013.13.2.
004.

Langins:1981:DCE

[Lan81] Janis Langins. The decline of chemistry at the École Polytechnique (1794–1805). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(1):1–19, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.1.
1.

Lang:2022:MRA

[Lan22] Sarah Lang. A machine reasoning algorithm for the digital analysis of alchemical language and its Decknamen. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):65–83, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2038428.

Larder:1967:ACB

[Lar67] David F. Larder. Alexander Crum Brown and his doctoral thesis of 1861. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(2):112–132, 1967. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1967.14.2.
                                        112.

Larder:1971:DCB

[Lar71] David F. Larder. A dialectical consideration of Butlerov’s theory of chemical structure. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(1):26–48, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1971.18.1.
                                        26.

Laszlo:1994:CPP

[Las94] Pierre Laszlo. Le Châtelier, ‘public prosecutor’ of Darzens. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(2):65–74, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.2.
                                        65.

Laszlo:1996:ERS

[Las96]  Pierre Laszlo.  Essay re-

view: A Social History of Truth. Civility and Science in Seventeenth-Century England. By Stephen Shapin. Pp. 483. University of Chicago Press: Chicago. 1994. £23.95, $34.50. ISBN 0-226-75018-3. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(1):59–64, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.1.
59.

Laszlo:2003:FMO

[Las03] Pierre Laszlo. Features of mechanistic organic chemistry already present in 1910. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(3):261–273, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.3.
261.

Laszlo:2010:QIG

[Las10] Pierre Laszlo. Quality information from the Grapevine. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(2):202–215, 2010. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/174582310X12719003720287.

Laube:2019:BRE

[Lau19]  Stefan Laube.  Book re-

view: Das Erbe des Nikolaus von Kues im Spiegel der Alchemie, by Witalij Morosow. Pp. 268, illus., index. Aschendorff: Münster. 2018. EUR 39.00. ISBN: 978-3-402-16005-

3. (German) [The legacy of Nicholas of Cusa in the mirror of alchemy]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):358–360, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1590915.

LeGrand:1973:NFA

[Le 73] H. E. Le Grand. A note on fixed air: The universal acid. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):88–94, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.2.
88.

LeGrand:1975:CCB

[Le 75] H. E. Le Grand. The “conversion” of C.-L. Berthollet to Lavoisier’s chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):58–70, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.1.
58.

LeGrand:1982:CPC

[Le 82] Homer E. Le Grand. Chemistry in a provincial context: The Montpellier Société Royale des Sciences in the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(2):88–105, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.2.
88.

LeRoux:2015:SIS

[Le 15] Muriel Le Roux. From science to industry: The sites of aluminium in France from the Nineteenth to the Twentieth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(2):114–137, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/1745823415Y.
0000000001.

Lorentz:2003:CPG

[LE03] Bernhard Lorentz and Paul Erker. Chemie und Politik: die Geschichte der Chemischen Werke Hüls 1938 bis 1979: eine Studie zum Problem der corporate governance. C. H. Beck, München, Germany, 2003. ISBN 3-406-50962-2 (hardcover). 461 pp. LCCN HD9654.9.C48 C484 2003.

Leaback:1997:N

[Lea97] David H. Leaback. News.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(1):40, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.1.
                                        40.

Lefevre:2018:AVM

[Lef18a] Wolfgang Lefèvre. Animal, vegetable, mineral? How eighteenth-century science disrupted the natural order. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2):191–192, 2018. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Lefevre:2018:MNC

[Lef18b]  Wolfgang Lefèvre.  The

Méthode de nomenclature chimique (1787): a document of transition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(1):9–29, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1418233.

Lehman:2009:MEC

[Leh09]  Christine Lehman.  Mid-

eighteenth-century chemistry in France as seen through student notes from the courses of Gabriel-François Venel and Guillaume–François Rouelle. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(2):163–189, 2009. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/174582309X441381.

Lehman:2010:ICC

[Leh10] Christine Lehman. Innovation in chemistry courses in France in the Mid-Eighteenth Century: Experiments and affinities. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(1):3–26, 2010. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582310X12629173849881.

Leicester:1975:LVC

[Lei75] Henry M. Leicester. Lomonosov’s views on combustion and phlogiston. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1): 1–9, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.1.
1.

Leitao:2016:APP

[Lei16] José Vieira Leitão. Alchemy, prophecy, and politics in Eighteenth-Century Iberia: Anselmo Castelo Branco’s critique of Benito Feijoo. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):304–325, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1283148.

Levey:1957:TTA

[Lev57] Martin Levey. Tanning technology in Ancient Mesopotamia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(1):35–46, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1957.6.1.
                                        35.

Levey:1958:RSA

[Lev58] Martin Levey. Research sources in Ancient Mesopotamian chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3):149–154, 1958. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1958.6.3.149.

Levere:1970:ASE

[Lev70] Trevor H. Levere. Affinity or structure: an early problem in organic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(2):111–126, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.2.
                                        111.

Levere:1981:DTB

[Lev81] Trevor H. Levere. Dr. Thomas Beddoes at Oxford: Radical politics in 1788–1793 and the fate of the Regius Chair in Chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(2):61–69, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.2.
61.

Levitt:2021:MDA

[Lev21]  Theresa Levitt.  Morphine

dreams: Auguste Laurent and the active principles of organised matter. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(1):97–115, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1867786.

Lewicki:1987:A

[Lew87] W. Lewicki. Announcement.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(1):4, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.1.
4.

Linden:1977:JSS

[Lin77] Stanton J. Linden. Jonson and Sendivogius: Some new light on Mercury Vindicated from the Alchemists at Court.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(1):39–54, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.1.
                                        39.

Linden:1984:AES

[Lin84] Stanton J. Linden. Alchemy and eschatology in Seventeenth-Century poetry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(3):102–124, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1984.31.3.
                                        102.

Luna:2012:ECT

[LK12] Fernando J. Luna and Lorelai B. Kury. Enlightenment chemistry translated by a Brazilian man of science in Lisbon. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(3):218–240, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1179/174582312X13457672281786.

Llana:1985:CNH

[Lla85] James W. Llana. A contribution of natural history to the chemical revolution in France.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(2):71–91, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.2.
71.

Luthy:2001:LME

[LMN01] Christoph Herbert Lüthy, John Emery Murdoch, and William R. Newman, editors. Late Medieval and Early Modern Corpuscular Matter Theories, volume 1 of Medieval and early modern science. Brill, Leiden, The Netherlands, 2001. ISBN 90-04-11516-1 (hardcover). ISSN 1567-8393. viii + 610 pp. LCCN QC173.38 .L38 2001.

Longhurst:2007:IPF

[Lon07] Ian Longhurst. The identity of Pliny’s Flos Salis and Roman perfume. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):299–302, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582307X237056.

Love:1972:SSH

[Lov72] Rosaleen Love. Some sources of Herman Boerhaave’s concept of fire. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):157–174, 1972. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.3.
157.

Low:1980:POC

[Löw80] Reinhard Löw. The progress of organic chemistry during the period of German romantic Naturphilosophie (1795–1825). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(1):1–10, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.1.
1.

Lowengard:2018:BRP

[Low18] Sarah Lowengard. Book review: A Perfect Ground: Preparatory Layers for Oil Paintings 1550–1900. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):404–405, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1493866.

LopezPerez:2005:ACP

[LP05] Miguel López Pérez and Joaquín Pérez Pariente. Alquimia: ciencia y pensamiento a través de los libros: [catálogo de la exposición celebrada en la Biblioteca Histórica “Marqués de Valdecilla”, 7 de noviembre de 2005–31 de enero de 2006].

(Spanish) [Alchemy: science and thought through the books: [catalog of the exhibition held in the Historical Library “Marqués de Valdecilla”, November 7, 2005–31 January 2006]]. Servicio de Publicaciones de la Universidad Complutense de Madrid [etc.], Madrid, Spain, 2005. ISBN 84-472-1046-4 (Universidad de Sevilla), 84-7491-792-1 (UCM). 155 + 6 pp. LCCN QD26.S7 P47 2005; QD26 .A54 2005.

Leigh:2016:WCR

[LR16] G. Jeffery Leigh and Alan J.

Rocke. Women and chemistry in Regency England: New light on the Marcet Circle. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):28–45, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1186584.

Lenke:2014:MMN

[LRT14] Nils Lenke, Nicolas Roudet, and Hereward Tilton. Michael Maier — nine newly discovered letters. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(1):1–47, 2014. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000043.

Luhrmann:1986:IFF

[Luh86] T. M. Luhrmann. An interpretation of the Fama Fraternitatis with respect to Dee’s Monas Hieroglyphica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(1):1–10, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1986.33.1.
                                        1.

Lykknes:2018:KOK

[Lyk18] Annette Lykknes. Kunskap och kemisk industri i 1800-talets Sverige. (Swedish) [Science and chemical industry in 19th century Sweden]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2): 196–197, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Lykknes:2020:CPC

[Lyk20] Annette Lykknes. The chemistry professor as consultant at the Norwegian Institute of Technology, 1910–1930. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):271–288, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1794697.

Lykknes:2022:ECW

[Lyk22] Annette Lykknes. Enabling circumstances: Women chemical engineers at the Norwegian Institute of Technology, 1910–1943. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):262–290, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2098897.

Lysaght:1937:HTC

[Lys37] D. J. Lysaght. Hooke’s theory of combustion. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(2):93–108, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1937.1.2.
93.

MacPhail:1967:MCA

[Mac67] Ian MacPhail. The Mellon Collection of Alchemy and the Occult. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(3):198–202, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.3.
198.

MacLeod:1988:GSA

[Mac88] Roy M. MacLeod. Gold from the sea: Archibald Liversidge, F.R.S., and the “chemical prospectors”; 1870–1970.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(2):53–64, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1988.35.2.
                                        53.

Maccoll:1989:ACU

[Mac89]  Allan Maccoll.  Australian

chemists at University College London 1899–1988. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(2):82–90, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1989.36.2.
                                        82.

Mahdihassan:1976:ETE

[Mah76] S. Mahdihassan. Early terms for elixir hitherto unrecognized in Greek alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):129–133, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1976.23.3.
                                        129.

Mahdihassan:1977:EMO

[Mah77] S. Mahdihassan. Elixirs of mineral orgin in Greek alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):133–142, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1977.24.3.
133.

Malaquias:2008:ASN

[Mal08] Isabel Malaquias. Aspects of the scientific network and communication of John Hyacinth de Magellan in Britain, Flanders and France. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):255–273, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X358097.

Manzalaoui:1961:JDP

[Man61] Mahmoud Manzalaoui. John Dastin and the Pseudo-Aristotelian Secretum, Secretorum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(3):166–167, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.3.
166.

Martin:1975:HPI

[Mar75] Luther H. Martin, Jr. A history of the psychological interpretation of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(1):10–20, 1975. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1975.22.1.
                                        10.

Mark:1993:SOM

[Mar93] H. F. (Herman Francis) Mark. From Small Organic Molecules to Large: a Century of Progress. Profiles, pathways, and dreams. American Chemical Society, Washington, DC, USA, 1993. ISBN 0-8412-1776-9 (hardcover), 0-8412-1802-1 (paperback). ISSN 1047-8329. xxvi + 148 pp. LCCN QD22.M36 A3

                                        1993. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0604/92030099-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0604/92030099-
                                        t.html.

Martin:2004:ARC

[Mar04] Craig Martin. Alchemy and the Renaissance commentary tradition on Meteorologica IV. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(3):245–262, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2004.51.3.
                                        245.

Martelli:2009:DWA

[Mar09] Matteo Martelli. “Divine water” in the alchemical writings of Pseudo-Democritus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):5–22, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X405192.

Martelli:2017:TAA

[Mar17] Matteo Martelli. Translating ancient alchemy: Fragments of Graeco–Egyptian alchemy in Arabic compendia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):326–342, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1412137.

Maarald:2019:NCG

[Mår19] Erland Mårald. Nitrogen capture: The growth of an international industry (1900– 1940). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):353–354, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1573793.

Martelli:2021:HAA

[Mar21]  Matteo Martelli.  Hathor’s

alchemy: The ancient roots of the hermetic art. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):447–448, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1927539.

Martelli:2023:MLB

[Mar23] Vincenzo Carlotta Matteo Martelli. Metals as living bodies. founts of mercury, amalgams, and chrysocolla. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):7–30, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2023.2173833.

Mason:1991:PPV

[Mas91] Stephen F. Mason. From Pasteur to parity violation: Cosmic dissymmetry and the origins of biomolecular handedness. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(2):85–108, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1991.38.2.
                                        85.

Matus:2013:RBR

[Mat13] Zachary Matus. Resurrected bodies and Roger Bacon’s elixir. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(4):323–340, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

full/10.1179/0002698013Z.
00000000037.

Mauskopf:1970:HMC

[Mau70] Seymour H. Mauskopf. Haüy’s model of chemical equivalence: Daltonian doubts exhumed. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(3):182–191, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.3.
182.

Mauskopf:1972:R

[Mau72] Seymour H. Mauskopf. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(1):60–67, 1972. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1972.19.1.60.

Mauskop:2002:RKP

[Mau02] Seymour Mauskop. Richard Kirwan’s phlogiston theory: Its success and fate. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(3):185–205, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.3.

Mawer:2016:BTA

[Maw16] Judith Mawer. Bridging traditions: Alchemy, chemistry, and Paracelsian practices in the Early Modern Era. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):185–186, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1227164.

Mawer:2018:AMC

[Maw18]  Judith Mawer.  Alchemy,

medicine, and commercial book production: A codicological and linguistic study of the Voigts–Sloane manuscript group. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2): 197–199, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Molella:2003:IE

[MB03] Arthur P. Molella and Joyce Bedi, editors. Inventing for the Environment. Lemelson Center studies in invention and innovation. MIT Press in association with the Lemelson Center, Smithsonian Institution, Cambridge, MA, USA, 2003. ISBN 0-262-13427-6 (hardcover). xix + 398 pp. LCCN T173.8 .I6185 2003. URL http://www.

                                        loc.gov/catdir/toc/fy043/
                                        2002035775.html.

Moctezuma-Burns:2021:CRG

[MB21] Olin Moctezuma-Burns. Compound remedies: Galenic pharmacy from the Ancient Mediterranean to New Spain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):318– 319, 2021. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/00026980.2021.
1919459.

McConnell:2006:WTO

[MC06] Anita McConnell and Philip Collins. Will the true originator of the storm glass please own up. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):67–75, 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X93200.

McDonald:1966:CBL

[McD66] E. McDonald. The collaboration of Bucquet and Lavoisier. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):74–83, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2013.13.2.
002.

McEvoy:1978:JPAa

[McE78a] John G. McEvoy. Joseph Priestley, “aerial philosopher”: Metaphysics and methodology in Priestley’s chemical thought, from 1762 to 1781. Part 1. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(1):1–55, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.1.
1.

McEvoy:1978:JPAb

[McE78b] John G. McEvoy. Joseph Priestley, “aerial philosopher”: Metaphysics and methodology in Priestley’s chemical thought, from 1772 to 1781. Part II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(2):93–116, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.2.
93.

McEvoy:1978:JPAc

[McE78c] John G. McEvoy. Joseph Priestley, “aerial philosopher”: Metaphysics and methodology in Priestley’s chemical thought, from 1772 to 1781. Part III. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(3):153–175, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.3.
153.

Mcevoy:1979:JPA

[Mce79] John G. Mcevoy. Joseph Priestley, “aerial philosopher”: Metaphysics and methodology in Priestley’s chemical thought, from 1772 to 1781. Part IV. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(1):16–38, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.1.
16.

Mcguire:1967:TIN

[Mcg67] J. E. Mcguire. Transmutation and immutability: Newton’s doctrine of physical qualities. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(2):69–95, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.2.
69.

McGuire:1968:FAP

[McG68] J. E. McGuire. Force, active principles, and Newton’s invisible realm. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(3):154–208, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.3.
                                        154.

McKie:1938:SEW

[McK38] Douglas McKie. Some early work on combustion, respiration and calcination. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):143–165, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1938.1.3.
                                        143.

McKie:1958:FHU

[McK58]  Douglas McKie.  On five

hitherto unrecorded copies of Jean Rey’s Essays. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3):136–139, 1958. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1958.6.3.
                                        136.

McK:1959:R

[McK59] D. McK. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(3):168–171, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1959.7.3.
                                        168.

McKie:1960:OEJ

[McK60] D. McKie. Obituary: Eric John Holmyard (1891–1959). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(1):1–5, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.1.
1.

Mckie:1961:JPC

[Mck61a] Douglas Mckie. Joseph Priestley and The Copley Medal. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(1):1–22, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.1.
1.

Mckie:1961:SPP

[Mck61b] Douglas Mckie. On some prepublication copies of Lavoisier’s Traité (1789). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(1):37–46, 1961. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1961.9.1.37.

McKie:1963:HES

[McK63]  D. McKie.  Henry Ernest

Stapleton (1878–1962). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(3):101–104, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1963.11.3.
101.

McVaugh:1975:VCH

[McV75] Michael McVaugh. The “venable collection” in the history of chemistry at the University of North Carolina. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):154–155, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
154.

Meadows:2001:R

[Mea01] Jack Meadows. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(1):41–52, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.1.
41.

Meinel:1983:TPE

[Mei83] Christoph Meinel. Theory or practice? The Eighteenth-Century debate on the scientific status of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(3):121–132, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.3.
121.

Meikle:1995:APC

[Mei95] Jeffrey L. Meikle. American Plastic: a Cultural History. Rutgers University Press, New Brunswick, NJ, USA, 1995. ISBN 0-8135-2234-X (hardcover), 0-8135-2235-8 (paperback), 0-585-20693-7 (e-book). xiv + 403 pp. LCCN TP1117 .M45 1995.

Meinel:2001:R

[Mei01] Christoph Meinel. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(3):191–208, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.3.
191.

Merchant:1979:VFM

[Mer79] Carolyn Merchant. The vitalism of Francis Mercury Van Helmont: Its influence on Leibniz. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):170–183, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.3.

Merkur:1990:SSA

[Mer90] Daniel Merkur. The study of spiritual alchemy: Mysticism, gold-making, and esoteric hermeneutics. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(1):35–45, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.1.
35.

Mertens:2003:APL

[Mer03] Joost Mertens. Anselme Payen (1795–1871), learned manufacturer of chemical products. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(2):182–207, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.2.
182.

Mertens:2004:HAU

[Mer04]  Joost Mertens.  The his-

tory of artificial ultramarine (1787–1844): Science, industry and secrecy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(3):219–244, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.3.
219.

Mertens:2017:WST

[Mer17] Manuel Mertens. Willem Silvius: “typographical parent” of John Dee’s Monas Hieroglyphica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):175–189, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1355594.

Merianos:2020:EGA

[Mer20] Gerasimos Merianos. Early Greek alchemy, patronage and innovation in late antiquity. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):319–321, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1775971.

Mercelis:2022:MDL

[Mer22] Joris Mercelis. “Men don’t like to work under a woman”: Female chemists in the photographic manufacturing industry, ca. 1918–1950. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):291–319, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2097980.

Miles:1968:PLC

[Mil68] Wyndham D. Miles. Public lectures on chemistry in the United States. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(3):131–153, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.3.
                                        131.

Miles:1970:WJM

[Mil70] Wyndham D. Miles. William James Macneven and early laboratory instruction in the United States. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(3):143–152, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.3.
                                        143.

Minor:2016:CWS

[Min16] Adriana Minor. Cold war science and the transatlantic circulation of knowledge. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):357–358, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1279415.

Minor:2018:BRP

[Min18] Adriana Minor. Book review:

Parallel Lives: Two Hoosier Chemists from Peru. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):297–298, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1469312.

Muller-Jahncke:1975:AAN

[MJ75] Wolf Dieter Müller-Jahncke.

The attitude of Agrippa von Nettesheim (1486–1535) towards alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):134–150, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
134.

Mawer:2015:R

[MJD^{+}15] Judith Mawer, Sean F. Johnston, Mariachiara Di Matteo, Catarina Madruga, Ana Simões, Josep Simon, Joel Vargas-Domínguez, Robert Bud, Simon Werrett, Sacha Tomic, and David Knight. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(3):292–303, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/0002698015Z.
                                        00000000085.

Chang:2021:UCR

[mKC21] Ku ming (Kevin) Chang. From university to court: The reversal of Stahl’s positions on gold-making. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):214–230, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1930677.

Miller:2008:IIS

[ML08] David Philip Miller and Trevor H. Levere. “Inhale it and see?” The collaboration between Thomas Beddoes and James Watt in pneumatic medicine. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(1):5–28, 2008. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/174582313X13789813148869.

Mond:1937:SA

[Mon37] Robert Mond. The study of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1): 1–2, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1937.1.1.
                                        1.

Mond:1938:ISV

[Mon38] Robert Mond. Introduction to the second volume. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):1–2, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1938.2.1.
1.

Montgomery:1963:CCC

[Mon63] John Warwick Montgomery.

Cross, constellation, and crucible: Lutheran astrology and alchemy in the age of the reformation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(2):65–86, 1963. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1080/00026980.1963.
11923895.

Morrell:1969:PCU

[Mor69] J. B. Morrell. Practical chemistry in the University of Edinburgh, 1799–1843. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(1–2):66–80, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.1-
2.66.

Morrell:1972:CBR

[Mor72] J. B. Morrell. The chemist breeders: The research schools of Liebig and Thomas Thomson. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(1):1–46, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1972.19.1.
                                        1.

Morselli:1980:MAE

[Mor80]  Mario A. Morselli.  The

manuscript of Avogadro’s “Essai d’une Manière de Déterminer Les Masses Relatives des Molécules Élémentaires”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(3):147–172, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.3.
                                        001.

Moran:1985:PCC

[Mor85] Bruce T. Moran. Privilege, communication, and chemiatry: the hermetic–alchemical circle of Moritz of Hessen-Kassel. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(3):110–126, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.1985.32.3.
110.

Moran:1996:PRD

[Mor96] Bruce T. Moran. Paracelsus, religion, and dissent: The case of Philipp Homagius and Georg Zimmermann. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(2):65–79, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.2.
65.

Morris:2002:Eb

[Mor02] Peter Morris. Erratum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(3):271, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.3.
271.

Morris:2004:E

[Mor04]  Peter J. T. Morris.  Edi-

torial. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(1): 3–4, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.1.
3.

Morris:2006:ERY

[Mor06] Peter Morris. Essay review: yes made in America: 1915– 1980; the Calco Chemical Company, American Cyanamid, and the Raritan River. By Anthony S. Travis. Pp. xiv + 582, illus., index. Sydney M. Edelstein Center, and Hexagon Press. Jerusalem. 2004. $115. ISBN 965-555-149-0. Chemie und Politik: die Geschichte der Chemischen Werke Hüls 1938 bis 1979: eine Studie zum Problem der corporate governance. By Bernhard Lorentz and Paul Erker. Pp. 461, illus., index. Verlag C. H. Beck. München. 2003. EUR 34.90 (hbk.). ISBN 3-406-50962-2. Die Linde AG: Geschichte eines Technologiekonzerns 1879–2004. By Hans-Liudger Dienel. Pp. 512, illus., index. Verlag C. H. Beck. München. 2004. EUR 34.90. ISBN 3-406-51484-7. Linde: History of a Technology Corporation, 1879–2004. By Hans-Liudger Dienel. Pp. xiv + 352, illus., index. Palgrave Macmillan. Basingstoke and New York. 2004. £45. ISBN 1-4039-2033-8. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):83–86, 2006. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582306X93219.

Morris:2009:NDS

[Mor09] Peter Morris. New dictionary of scientific biography. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(2):190–192, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X441372.

Morris:2011:FRH

[Mor11] Peter J. T. Morris. The fall and rise of the history of recent chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(3):238–256, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582311X13129418298983.

Morris:2012:MDE

[Mor12] Peter J. T. Morris. Musings from a departing Editor on the 75th anniversary of Ambix. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(3):189–196, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13484746071949.

Morris:2016:EVC

[Mor16a] Peter Morris. Entrepreneurial ventures in chemistry: The Muspratts of Liverpool, 1793– 1934. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):352–353, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1279411.

Morris:2016:MLL

[Mor16b] Peter Morris. Moore’s law: The life of Gordon Moore, silicon Valley’s quiet revolutionary. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):277, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1246810.

Morris:2018:BRW

[Mor18a] Peter Morris. Book review: A World History of Rubber: Empire, Industry, and the Everyday. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):399, 2018. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1469315.

Morris:2018:FSL

[Mor18b] Peter Morris. Fake silk: The lethal history of viscose rayon. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2):199–200, 2018. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Moran:2021:CAU

[Mor21a] Bruce T. Moran. Court authority and the university: Networks, recipes, and things-in-the-making vs. the abstractions of made things. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):135–153, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1922008.

Moreau:2021:LCC

[Mor21b] Elisabeth Moreau. Learning the chymical compromise: Paracelsian and Galenic medicine in Marburg disputations on Chymiatria. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):154–179, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1930676.

Moran:2022:ASD

[Mor22a] Bruce T. Moran. Alchemy, sources, and digital exploring at the Herzog August Bibliothek Wolfenbüttel. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):320–325, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2085764.

Moran:2022:MPA

[Mor22b] Bruce T. Moran. Medical performance and the alchemy of plants in the ventures of Leonhard Thurneisser zum Thurn. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):95–117, 2022. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2042058.

Mossman:1997:EPP

[Mos97] S. T. I. (Susan T. I.) Mossman, editor. Early Plastics: Perspectives, 1850–1950. Leicester University Press, London, UK, 1997. ISBN 0-7185-0020-2 (hardcover). x + 292 pp. LCCN TP1116 .E27 1997.

Moshenska:2015:MFC

[Mos15] Gabriel Moshenska. Michael Faraday’s contributions to archaeological chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(3):266–286, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/1745823415Y.
                                        0000000004.

Mossman:2022:BRB

[Mos22] Susan Mossman. Book review: Beyond Bakelite: Leo Baekeland and the Business of Science and Invention. By Joris Mercellis. Pp. 378, illus., index.

The MIT Press: Cambridge, Massachusetts. 2020. £45.00. ISBN: 978-0-26-253869-5 (paperback). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):429–431, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2095717.

Moureau:2009:SCC

[Mou09]  Sébastien Moureau.  Some

considerations concerning the alchemy of the De anima in arte alchemiae of Pseudo-Avicenna. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):49–56, 2009. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582309X405228.

Moy:1989:EFC

[Moy89] Timothy D. Moy. Emil Fischer as “chemical mediator”: Science, industry, and government in World War One. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3):109–120, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.3.
109.

MaiaNeto:2002:BC

[MP02] José R. Maia Neto and Elene C. Pereira Maia. Boyle’s Carneades. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(2):97–111, 2002. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.2002.49.2.97.

Roberts:2020:CCL

[MR20] Robin Mackie and Gerrylynn Roberts. Consultancy as a career in Late Nineteenth and Twentieth Century Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):214–233, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1794699.

Morris:1992:CSL

[MT92] Peter. J. T. Morris and Anthony S. Travis. The Chemical Society of London and the dye industry in the 1860s. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):117–126, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1992.39.3.
                                        117.

Morris:2002:Ea

[MT02] Peter J. T. Morris and Anthony S. Travis. Editorial.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(1):1–3, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.1.
1.

Martinon-Torres:2011:SRD

[MT11] Marcos Martinón-Torres. Some recent developments in the historiography of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 58(3):215–237, 2011. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582311X13129418299063.

Martinon-Torres:2012:ISH

[MT12] Marcos Martinón-Torres. Inside Solomon’s house: an archaeological study of the Old Ashmolean Chymical Laboratory in Oxford. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(1):22–48, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13296104891436.

Moureau:2016:UTH

[MT16] Sébastien Moureau and Nicolas Thomas. Understanding texts with the help of experimentation: The example of cupellation in Arabic scientific literature. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):98–117, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1216691.

Martelli:2014:R

[MTC^{+}14] Matteo Martelli, Anke Timmermann, Amy Eisen Cislo, Jonathan Simon, Willem Vijvers, Pierre Laszlo, Seymour Mauskopf, John K. Smith, Peter Morris, and David A. Kirby. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):411–421, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000067.

Martinon-Torres:2015:LAA

[MTKT^{+}15] Marcos Martinón-Torres, David Knight, Sacha Tomic, Jeffrey Allan Johnson, William H. Brock, Néstor Herran, Marina Maestrutti, and Hjalmar Fors. Laboratories of art. Alchemy and art technology from antiquity to the 18th Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(4):386–396, 2015. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2015.1129853.

Multhauf:1975:AAM

[Mul75] Robert P. Multhauf. Aurifiction, aurifaction and macrobiotics. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(3):218–220, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.3.
218.

Munday:1990:SCT

[Mun90] Pat Munday. Social climbing through chemistry: Justus Liebig’s rise from the Niederer Mittelstand to the Bildungsbürgertum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(1):3–19, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.1.
3.

Munday:1991:LMO

[Mun91] Pat Munday. Liebig’s metamorphosis: From organic chemistry to the chemistry of agriculture. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(3):135–154, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1991.38.3.
                                        135.

Munday:1997:R

[Mun97] Pat Munday. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(1):41–54, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.1.
                                        41.

Murase:2020:HPL

[Mur20] Amadeo Murase. The Homunculus and the Paracelsian Liber de imaginibus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):47–61, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1720359.

Murphy:2022:PTW

[Mur22] Hannah Murphy. The poison trials: Wonder drugs, experiment, and the battle for authority in Renaissance science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):191–193, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2030123.

Musso:2017:COE

[Mus17]  Marta Musso.  Coal, oil,

and empire. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):95–98, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1310424.

Nawa:2014:RIC

[Naw14] Christine Nawa. A refuge for inorganic chemistry: Bunsen’s Heidelberg Laboratory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):115–140, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000049.

Neville:1960:UDT

[Nev60] Roy G. Neville. Unrecorded Daltoniana: Two letters to John Bostock and a prospectus to the “new system”, 1808. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(1):42–45, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/000269860790225004.

Neville:1962:PCS

[Nev62] Roy G. Neville. The “Pratique de Chymie” of Sébastien Matte La Faveur. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(1):14–28, 1962. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269862790222924.

Newman:1982:TVI

[New82] William Newman. Thomas Vaughan as an interpreter of Agrippa Von Nettesheim. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(3):125–140, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.3.
125.

Newcomb:1986:LES

[New86] Sally Newcomb. Laboratory evidence of silica solution supporting Wernerian theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):88–93, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.2-
3.88.

Newman:1990:PAO

[New90a] William Newman. Prophecy and alchemy: The origin of Eirenaeus Philalethes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(3):97–115, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.3.
97.

Newmark:1990:PP

[New90b] Ann Newmark. The Partington Prize. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(2): 96, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.2.
96.

Newman:2014:MSA

[New14] William R. Newman. Mercury and sulphur among the high Medieval alchemists: From Razi and Avicenna to Albertus Magnus and Pseudo-Roger Bacon. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):327–344, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823414Y.

0000000004. See erratum [New15].

Newman:2015:EMS

[New15] William R. Newman. Erratum:

Mercury and sulphur among the high Medieval alchemists: From Razi and Avicenna to Albertus Magnus and Pseudo-Roger Bacon. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1): 107, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698014Z.

00000000074. See [New14].

Newman:2020:BCB

[New20a] William R. Newman. Bad chemistry: Basilisks and women in Paracelsus and pseudo-Paracelsus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):30–46, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1720341.

Newman:2020:PAE

[New20b] William R. Newman. Painted alchemists: Early modern artistry and experiment in the work of Thomas Wijck. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):316–317, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1769929.

Newson:2020:ACM

[New20c] Linda A. Newson. Alchemy and chemical medicines in Early Colonial Lima, Peru. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):107–134, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1747306.

Newman:2023:BRC

[New23] William Royall Newman. Book review: A Cultural History of Chemistry in the Middle Ages (vol. 2). Edited by Charles Burnett and Sébastien Moureau. Pp. 214, illus., index. Bloomsbury: London. 2022. £395.00 (6-volume set). ISBN: 978-1-4742-9454-6. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):188–190, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2186580.

Nieto-Galan:2015:RCH

[NG15] Agustí Nieto-Galan. Revisiting colour history. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1):94–97, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000072.

Nieto-Galan:2016:SWS

[NG16] Agustí Nieto-Galan. Science in Wonderland: The scientific fairy tales of Victorian Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):274–275, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1246808.

Nieto-Galan:2020:PSA

[NG20] Agustí Nieto-Galan. Poisonous skies: Acid rain and the globalization of pollution. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):103–104, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1671523.

Nickelsen:2012:PCP

[Nic12] Kärin Nickelsen. The path of carbon in photosynthesis: How to discover a biochemical pathway. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(3):266–293, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1179/174582312X13457672281867.

Nielsen:2016:AOI

[Nie16] Anita Kildebæk Nielsen. Akademi og industri. Kjemiutdanning og -forskning ved NTNU gjennom 100 år. (Norwegian) [Academy and industry. Chemical education and chemical research at NTNU over 100 years]. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 63(3):277–278, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1246823.

Norrgren:2005:IHM

[Nor05] Hilde Norrgrén. Interpretation and the hieroglyphic monad: John Dee’s reading of Pantheus’s Voarchadumia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(3):217–245, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X77781.

Norris:2006:MET

[Nor06] John A. Norris. The mineral exhalation theory of metallogenesis in pre-modern mineral science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):43–65, 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X93183.

Norris:2007:ETA

[Nor07] John A. Norris. Early theories of aqueous mineral genesis in the Sixteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):69–86, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X165434.

Normark:2015:FIL

[Nor15] Daniel Normark. Flexibility or inexactitude? The “Lab 60” at Karolinska Institutet: From medical disciplines towards the modern biomedical complex. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(2):167–188, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698015Z.
                                        00000000078.

Norrgren:2020:AGH

[Nor20] Hilde Norrgrén. An alchemist in Greenland: Hans Egede (1686– 1758) and alchemical practice in the Colony of Hope. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):153–173, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1747305.

Nummedal:2001:ARC

[Num01] Tara E. Nummedal. Alchemical reproduction and the career of Anna Maria Zieglerin. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):56–68, 2001. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.2.
56.

Nummedal:2007:BRA

[Num07] Tara E. Nummedal. Book reviews: Alquimia: ciencia y pensamiento a través de los libros (Alchemy: Science and thought through the books). Edited by Joaquín Pérez Pariente and Miguel López Pérez. Pp. 166, illus. index. Complutense University Press and Sevilla University Press: Madrid and Sevilla. 2006. EUR 15. ISBN 84-7491-792-1 and 84-472-1046-4 Jean Fernel’s “On the hidden causes of things”: forms, souls, and occult diseases in Renaissance medicine, with an Edition and Translation of Fernel’s “De Abditis Rerum Causis”. Edited by John Henry and John M. Forrester. Pp. x + 779, index. Brill: Leiden and Boston. 2005. EUR 172; $232. ISBN 90-04-14128-6. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):305–320, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582313X13789813148940.

Nummedal:2013:ARC

[Num13] Tara Nummedal. Alchemy and religion in Christian Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(4):311–322, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000036.

Nye:2018:PTN

[NW18] Mary Jo Nye and Stephen J. Weininger. Paper tools from the 1780s to the 1960s: Nomenclature, classification, and representations. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(1):1–8, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1419651.

Nye:2002:CHS

[Nye02] Mary Jo Nye, editor. The Cambridge History of Science: The Modern Physical and Mathematical Sciences, volume 5. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-57244-4 (vol. 3: hardback), 0-521-57243-6 (vol. 4), 0-521-57199-5 (vol. 5), 0-521-59442-1 (vol. 7). xxvii + 678 pp. LCCN Q125 .C32 2003. URL http://

                                        www.h-net.org/review/hrev-
                                        a0d3l3-aa;   http://www.h-
                                        net.org/review/hrev-a0d7k4-
                                        aa;   http://www.h-net.org/
                                        review/hrev-a0f4e9-aa;
                                        http://www.loc.gov/catdir/
                                        description/cam021/2001025311.
                                        html;   http://www.loc.gov/
                                        catdir/enhancements/fy0731/

2001025311-b.html;   http:
//www.loc.gov/catdir/samples/
cam033/2001025311.html;
http://www.loc.gov/catdir/
toc/cam026/2001025311.html.

Nye:2014:MTO

[Nye14] Mary Jo Nye. Mine, thine, and ours: Collaboration and co-authorship in the material culture of the mid-Twentieth Century chemical laboratory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(3):211–235, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000055.

Oakley:2023:MMH

[Oak23] Peter Oakley. Making mercury’s histories: Mercury in gold mining’s past and present. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):77–98, 2023. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2023.2192125.

Obrist:2009:VHM

[Obr09] Barbara Obrist. Views on history in Medieval alchemical writings. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(3):226–238, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269809X12529013496293.

Ogawa:2008:LRA

[Oga08] Mariko Ogawa. Liebig and the Royal Agricultural Society meeting at Bristol, 1842. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(2):136–152, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582308X255460.

Oldroyd:1973:EGS

[Old73] David Oldroyd. An examination of G. E. Stahl’s Philosophical Principles of Universal Chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(1):36–52, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1973.20.1.
                                        36.

Oldroyd:1974:SNP

[Old74] D. R. Oldroyd. Some neoplatonic and stoic influences on mineralogy in the Sixteenth and Seventeenth Centuries. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(2–3):128–156, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/000269874790223614.

Oltra:2005:FJH

[Olt05] Lluís Garrigós Oltra. François Joseph Houtou de Labillardière (1796–1867): a case of desertion in Nineteenth-Century chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52 (2):159–173, July 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X70111.

ONeil:2021:GNB

[O’N21] Sean O’Neil. Good names but better symbols: The establishment of chemical notation as a nomenclatural corrective at the turn of the Nineteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):331–364, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1983695.

Oldroyd:1974:MM

[OO74] D. R. Oldroyd and D. R.

Oldroyd. Mechanical mineralogy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21 (2–3):157–178, 1974. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269874790223597.

Orland:2016:GNG

[Orl16] Barbara Orland. Gift in der Nahrung. Zur Genese der Verbraucherpolitik Mitte des 20. Jahrhunderts. (German) [Poison in the food. The genesis of consumer policy in the mid-20th century]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):78–79, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1201307.

Oster:2002:SEP

[Ost02a] Malcolm Oster, editor. Science in Europe, 1500–1800: a Primary Sources Reader. Palgrave, Houndmills, Basingstoke, Hampshire, UK, 2002. ISBN 0-333-97001-2, 0-333-97002-0 (paperback). xiv + 282 pp. LCCN Q127.E8 S352 2002. URL http:/

                                        /www.loc.gov/catdir/bios/
                                        hol058/2001036833.html;
                                        http://www.loc.gov/catdir/
                                        description/hol052/2001036833.
                                        html;   http://www.loc.gov/
                                        catdir/toc/hol051/2001036833.
                                        html.

Oster:2002:SES

[Ost02b] Malcolm Oster, editor. Science in Europe, 1500–1800: a Secondary Sources Reader.

Palgrave, Houndmills, Basingstoke, Hampshire, UK, 2002. ISBN 0-333-97005-5 (hardcover), 0-333-97006-3 (paperback). xii + 307 pp. LCCN Q127.E8 S354 2002. URL

http://www.loc.gov/catdir/
bios/hol058/2001045149.html;
http://www.loc.gov/catdir/
description/hol052/2001045149.
html;   http://www.loc.gov/
catdir/toc/fy02/2001045149.
html.

Owen:2001:CBD

[Owe01]  Dawn Owen.  The con-

stant battery and the Daniell– Becquerel–Grove controversy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(1):25–40, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.1.
25.

Pagel:1960:PNG

[Pag60] Walter Pagel. Paracelsus and the Neoplatonic and Gnostic tradition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(3):125–166, 1960. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1960.8.3.125.

Pagel:1961:PMP

[Pag61]  Walter Pagel.  The prime

matter of Paracelsus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(3):117–135, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.3.
117.

Pagel:1962:WSG

[Pag62]  Walter Pagel.  The “wild

spirit” (gas) of John Baptist Van Helmont (1579–1644) and Paracelsus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(1):1–13, 1962. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269862790222906.

Pagel:1969:CCR

[Pag69] Walter Pagel. Chemistry at the cross-roads: The ideas of Joachim Jungius. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(1–2):100–108, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.1-
2.100.

Pagel:1972:R

[Pag72] Walter Pagel. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(2):151–155, 1972. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.2.
151.

Pagel:1983:R

[Pag83] Walter Pagel. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(3):163–168, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.3.
163.

Page:2014:LKM

[Pag14] Frederick G. Page. The little known Mineral and Chemical History of Iron and A History of Brass by an eighteenthcentury technologist, William Lewis (1708–1781). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(3):299–304, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000057.

Pallo:2010:ADP

[Pal10] Gábor Palló. The advantage and disadvantage of peripheral ignorance: The gas adsorption controversy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(2):216–230, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12719003720241.

Papathanassiou:1990:SAP

[Pap90] Maria Papathanassiou. Stephanus of Alexandria: Pharmaceutical notions and cosmology in his alchemical work. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(3):121–133, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1990.37.3.
                                        121.

Partington:1937:AMA

[Par37a] J. R. Partington. Albertus Magnus on alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):3–20, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1937.1.1.
                                        3.

Partington:1937:RDU

[Par37b] J. R. Partington. Report of discussion upon chemical and alchemical symbolism. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):61–77, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/

abs/10.1179/amb.1937.1.1.
61.

Partington:1938:CR

[Par38a] J. R. Partington. The chemistry of Razi. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):192–196, 1938. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1938.1.3.192.

Partington:1938:TA

[Par38b] J. R. Partington. Trithemius and alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(2):53–59, 1938. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1938.2.2.53.

Paskins:2020:TSM

[Pas20]  Mat Paskins.  Thrifty sci-

ence: Making the most of materials in the history of experiment. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):203–205, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1727647.

Patai:1982:MJF

[Pat82] Raphael Patai. Maria the Jewess — founding mother of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(3):177–197, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.3.
177.

Patai:1988:RTM

[Pat88] Raphael Patai. Raymund De Tarrega-Marrano, heretic, alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(1):14–30, 1988. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1988.35.1.14.

Paton:1992:ERS

[Pat92]  William Paton.  Essay re-

view: In search of a cure. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(3):141–142, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.3.
141.

Payne:1989:SGU

[Pay89] Richard K. Payne. Sex and gestation, the union of opposites in European and Chinese alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(2):66–81, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.2.
66.

Perez-Criado:2023:HSP

[PC23] Silvia Pérez-Criado. How to sell a poison. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):200–201, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2157097.

Perrin:1969:PLT

[Per69] C. E. Perrin. Prelude to Lavoisier’s theory of calcination: Some observations on Mercurius Calcinatus Per Se. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(3):140–151, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.3.
140.

Perrin:1973:LTE

[Per73] C. E. Perrin. Lavoisier’s table of the elements: a reappraisal. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):95–105, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1973.20.2.
                                        95.

Perrin:1979:R

[Per79] C. E. Perrin. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(2):137–144, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1979.26.2.
                                        137.

Perrin:1982:RCJ

[Per82] C. E. Perrin. A reluctant catalyst: Joseph Black and the Edinburgh reception of Lavoisier’s chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(3):141–176, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1982.29.3.
                                        141.

Perrin:1984:JBV

[Per84] C. E. Perrin. J. B. Van Mons’ Essai sur les principes de la chimie antiphlogistique: a mystery solved. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(1): 1–5, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1984.31.1.

Perrin:1989:LBC

[Per89] C. E. Perrin. The Lavoisier– Bucquet collaboration: a conjecture. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):5–13, 1989. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1989.36.1.5.

Perkins:2003:CCP

[Per03] John Perkins. Creating chemistry in Provincial France before the Revolution: The examples of Nancy and Metz. Part 1. Nancy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(2):145–181, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.2.
145.

Perkins:2004:CCP

[Per04] John Perkins. Creating chemistry in Provincial France before the Revolution: The examples of Nancy and Metz. Part 2. Metz. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(1):43–75, 2004. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.2004.51.1.43.

Perkins:2010:CCC

[Per10a]  John Perkins.  Chemistry

courses and the construction of chemistry, 1750–1830. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(1):1, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12629173849845.

Perkins:2010:CCP

[Per10b]  John Perkins.  Chemistry

courses, the Parisian chemical world and the chemical revolution, 1770–1790. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(1):27–47, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12629173849926.

Perkins:2013:SCE

[Per13] John Perkins. Sites of chemistry in the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):95–98, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000024.

Perez:2016:BHP

[Pér16] Silvia Pérez. Banned: a history of pesticides and the science of toxicology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):194–195, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1227176.

Perucchetti:2019:MMM

[Per19] Dr Laura Perucchetti. Metals, minds and mobility: Integrating scientific data with archaeological theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):354–356, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1586307.

Peterschmitt:2015:NBP

[Pet15] Luc Peterschmitt. New big pictures of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(2):189–192, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698015Z.
00000000079.

Guesnon:2021:GCF

[PG21] Arnaud Page and Maxime Guesnon. Glutenophilia: Chemistry and flour quality in Nineteenth-century France and Great Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):365–384, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1983696.

Piorko:2022:APL

[PHW22] Megan Piorko, Marieke Hendriksen, and Simon Werrett. Alchemical practice: Looking towards the chemical humanities. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):1–18, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2035572.

Pilcher:1938:BL

[Pil38] Richard B. Pilcher. ‘Boyle’s laboratory’. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):17–20, 1938. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1938.2.1.17.

Piorko:2018:LLO

[Pio18] Megan Piorko. L’alchimie à livres ouverts: En quête des secrets de la matière: livres et manuscrits du XVIe au XXe siècle. (French) [Alchemy in open books: a quest for the secrets of matter: books and manuscripts of the XVIth to XXth century]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2): 192–193, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Plass:1982:GAF

[Pla82]  Paul Plass.  A Greek al-

chemical formula. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(2):69–73, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.2.
69.

Piorko:2023:DHP

[PLB23] Megan Piorko, Sarah Lang, and Richard Bean. Deciphering the Hermeticae Philosophiae Medulla: Textual cultures of alchemical secrecy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):150–183, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2201744.

Plessner:1959:TPP

[Ple59]  M. Plessner.  The Turba

Philosophorum: a preliminary report on three Cambridge Mss. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(3):159–163, 1959. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1959.7.3.
159.

Plessner:1969:GJI

[Ple69]  M. Plessner.  Geber and

Jabir ibn Hayyan: an authentic Sixteenth-Century quotation from Jabir. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(3):113–118, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.3.
113.

Plessner:1972:HAL

[Ple72] Martin Plessner. The history of Arabic literature. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):209–215, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.3.
209.

Pollard:2015:LCH

[Pol15] A. M. Pollard. Letters from China: a history of the origins of the chemical analysis of ceramics. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1):50–71, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/1745823414Y.
0000000008.

Polloni:2020:MPS

[Pol20] Nicola Polloni. A matter of philosophers and spheres: Medieval glosses on Artephius’s Key of Wisdom. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):135–152, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1747307.

Pomper:1962:LDL

[Pom62] Philip Pomper. Lomonosov and the discovery of the Law of the Conservation of Matter in chemical transformations. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(3):119–127, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1962.10.3.
119.

Poole:2009:TGF

[Poo09] William Poole. Theodoricus Gravius (fl. 1600–1661): Some biographical notes on a German chymist and scribe working in Seventeenth-Century England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(3):239–252, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269809X12529013496338.

Porto:2001:MSN

[Por01] Paulo Alves Porto. Michael Sendivogius on nitre and the preparation of the philosophers’ stone. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(1):1–16, 2001. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.2001.48.1.1.

Powers:1998:ASA

[Pow98] John C. Powers. ‘Ars Sine Arte:’ Nicholas Lemery and the end of alchemy in Eighteenth-Century France. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(3):163–189, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.3.
                                        163.

Powers:2014:FAE

[Pow14] John C. Powers. Fire analysis in the Eighteenth Century: Herman Boerhaave and scepticism about the elements. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(4):385–406, 2014. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823414Y.
0000000005.

Pratt:1991:PCJ

[Pra91]  Herbert T. Pratt.  Peter

Crosthwaite: John Dalton’s “friend and colleague”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(1):11–28, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1991.38.1.
11.

Pratt:1992:JDS

[Pra92] Herbert T. Pratt. John Dalton settles in Manchester. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(1):17–20, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.1.
17.

Pratt:2010:LSV

[Pra10] Herbert T. Pratt. A letter signed: The very beginnings of Dalton’s atomic theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(3):301–310, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12849808295742.

Pritchard:1979:TCB

[Pri79]  Allan Pritchard.  Thomas

Charnock’s book dedicated to Queen Elizabeth. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(1):56–73, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1979.26.1.
                                        56.

Priesner:1986:SAF

[Pri86]  Claus Priesner.  Spiritus

Aethereus — formation of ether and theories on etherification from Valerius Cordus to Alexander Williamson. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):129– 152, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1986.33.2-
                                        3.129.

Principe:1987:CTR

[Pri87] Lawrence Principe. “Chemical translation” and the role of impurities in alchemy: Examples from Basil Valentine’s Triumph–Wagen. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(1):21–30, 1987. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.1.
21.

Principe:1992:RBA

[Pri92] L. M. Principe. Robert Boyle’s alchemical secrecy: Codes, ciphers and concealments. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(2):63–74, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1992.39.2.
63.

Principe:1993:R

[Pri93] Lawrence M. Principe. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(1):27–38, 1993. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1993.40.1.27.

Principe:1994:R

[Pri94] Lawrence M. Principe. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(3):161–168, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1994.41.3.
161.

Principe:1999:SN

[Pri99] Lawrence M. Principe. Short notices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(1):63–64, 1999. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1999.46.1.63.

Principe:2006:NHC

[Pri06]  Lawrence M. Principe.  A

new History of Chemistry Library at the Chemical Heritage Foundation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):77–82, 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269805X77790.

Principe:2009:AGD

[Pri09] Lawrence M. Principe. Allen G. Debus (1926–2009): an appreciation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(3):201, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269809X12529013496419.

Principe:2013:SKD

[Pri13] Lawrence M. Principe. Sir Kenelm Digby and his alchemical circle in 1650s Paris: Newly discovered manuscripts. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(1):3–24, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698012Z.
00000000019.

Priesner:2016:LAL

[Pri16a] Claus Priesner. Legends about legends: Abraham Eleazar’s adaptation of Nicolas Flamel. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):1–27, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2016.1163631.

Principe:2016:BSA

[Pri16b] Lawrence M. Principe. Boyle studies: Aspects of the life and thought of Robert Boyle (1627-91). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):72–73, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1201299.

Principe:2016:CES

[Pri16c] Lawrence M. Principe. Chymical exotica in the Seventeenth Century, or, how to make the Bologna stone. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):118–144, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1213011.

Prinke:2016:NLA

[Pri16d] Rafał T. Prinke. New light on the alchemical writings of Michael Sendivogius (1566– 1636). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):217–243, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1246845.

Principe:2021:CVC

[Pri21a] Lawrence M. Principe. The changing visions of chymistry at seventeenth-century Jena: The two Brendels, Rolfinck, Wedel, and others. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(2–3):180–197, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2021.1922021.

Prinke:2021:CBB

[Pri21b] Rafał T. Prinke. The correspondence of Boldizsár Batthyány. The everyday life of an early modern alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(4):448–450, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1932965.

Pardo-Tomas:2017:ADE

[PT17] José Pardo-Tomás. American drugs and Early Modern Europe: New ways of telling an old story. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):375–377, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385897.

Puig:2004:NIN

[Pui04] Nuria Puig. Networks of innovation or networks of opportunity? The making of the Spanish antibiotics industry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):167–185, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.2.
167.

Pulkkinen:2019:BRM

[Pul19] Karoliina Pulkkinen. Book review: Mendeleev to Oganesson: A Multidisciplinary Perspective on the Periodic Table, edited by Eric Scerri and Guillermo Restrepo. Pp. 328, illus., index.

Oxford University Press: New York. 2018. £80.00. ISBN: 978-0-19-066853-2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):365–367, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1610598.

Pulkkinen:2020:VDE

[Pul20] Karoliina Pulkkinen. Values in the development of early periodic tables. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(2):174–198, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1747325.

Pumfrey:1995:R

[Pum95] Stephen Pumfrey. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(2):119–127, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1995.42.2.
                                        119.

Pohl-Valero:2017:SNS

[PV17] Stefan Pohl-Valero. Setting nutritional standards. Theory, policies, practices. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):386–388, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1405901.

Pascual-Valderrama:2017:AMD

[PVPP17] Ignacio-Miguel Pascual-Valderrama and Joaquín Pérez-Pariente. The alchemical manuscripts of David Lindsay (1587– 1641), Lord Lindsay of Balcarres. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):234–262, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1370875.

Pagel:1969:EAR

[PW69] Walter Pagel and Marianne Winder. The eightness of Adam and related “gnostic” ideas in the Paracelsian corpus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(3):119–139, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.3.
119.

Pagel:1974:HEP

[PW74] Walter Pagel and Marianne Winder. The higher elements and prime matter in Renaissance naturalism and in Paracelsus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(2–3):93–127, 1974. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269874790223623.

Prinke:2018:APM

[PZ18] Rafal T. Prinke and Mike A. Zuber. Alchemical patronage and the making of an adept: Letters of Michael Sendivogius to Emperor Rudolf II and his Chamberlain Hans Popp. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):324–355, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1512776.

Reed:2020:CCC

[QR20] Viviane Quirke and Peter Reed. Chemistry, consultants, and companies, c. 1850–2000: Introduction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):207–213, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1794700.

Quirke:2016:BHA

[Qui16a] Viviane Quirke. Biologics. A history of agents made from living organisms in the Twentieth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):279–280, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1246824.

Quirke:2016:BPM

[Qui16b] Viviane Quirke. The British pharmacopoeia, 1864 to 2014: Medicines, international standards and the state. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):197, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1227180.

Quirke:2020:ICI

[Qui20]  Viviane Quirke.  Imperial

Chemical Industries and Craig Jordan, “the first Tamoxifen consultant,” 1960s–1990s. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):289–307, 2020. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/00026980.2020.
1794675.

Ross:2002:DCW

[RA02] Benjamin Ross and Steven Amter. Deregulation, chemical waste, and ground water: a 1949 debate. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(1):51–66, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2002.49.1.
                                        51.

Rae:1989:ROC

[Rae89] Ian D. Rae. The research in organic chemistry of Aleksandr Borodin (1833–1887). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(3):121–137, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1989.36.3.
                                        121.

Rae:1995:COA

[Rae95] Ian D. Rae. Chemical organizations in Australia and New Zealand. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(1):28–49, 1995. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1995.42.1.28.

Rae:1996:R

[Rae96] Ian D. Rae. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(3):189–191, 1996. CODEN AMBXAO.

ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.3.
189.

Rae:1997:SAP

[Rae97] Ian D. Rae. Spectrum analysis: The priority claims of Stokes and Kirchhoff. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(3):131–144, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.3.
131.

Rae:2005:CAG

[Rae05]  Ian D. Rae.  Chemistry

in Australia: Growing up, down under. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):7–25, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X52887.

Rae:2006:OOD

[Rae06] Ian D. Rae. Ozonised oils as disinfectants. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(1):3–20, 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X93192.

Rae:2016:TVP

[Rae16] Ian D. Rae. Theory versus practice in the Twentiethcentury search for the ideal anaesthetic gas. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):46–65, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1163638.

Ragland:2012:CTS

[Rag12]  Evan Ragland.  Chymistry

and taste in the Seventeenth Century: Franciscus Dele Boë Sylvius as a chymical physician between Galenism and Cartesianism. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(1):1–21, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1179/174582312X13296104891472.

Ragland:2022:BRA

[Rag22] Evan Ragland. Book review: Aesthetic Science: Representing Nature in the Royal Society of London, 1650–1720. By Alexander Wragge-Morley. Pp. 272, illus., index. University of Chicago Press: Chicago. 2020. £30.00 (paperback). ISBN: 978-0-22-668086-6. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):329–331, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2049499.

Ramberg:2000:DVB

[Ram00] Peter J. Ramberg. The death of vitalism and the birth of organic chemistry: Wöhler’s urea synthesis and the disciplinary identity of organic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(3):170–195, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2000.47.3.
170.

Rampling:2008:ECG

[Ram08] Jennifer M. Rampling. Establishing the canon: George Ripley and his alchemical sources. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):189–208, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X358114.

Rampling:2010:CRC

[Ram10] Jennifer M. Rampling. The catalogue of the Ripley Corpus: Alchemical writings attributed to George Ripley (d. ca. 1490). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(2):125–201, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12719003720322.

Rampling:2013:NDA

[Ram13] Jennifer M. Rampling. New developments for Ambix. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(1):1–2, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/0002698012Z.
                                        00000000023.

Rampling:2016:EMA

[Ram16] Jennifer M. Rampling. The Englishing of Medieval alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):268–272, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1080/00026980.2016.
                                        1256108.

Rampling:2017:FHC

[Ram17] Jennifer M. Rampling. The future of the history of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):295–300, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2017.1434970.

Ramberg:2018:BRC

[Ram18]  Peter J. Ramberg.  Book

review: The Chemical Century: Molecular Manipulation and Its Impact on the 20th Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):300–301, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1469314.

Raphael:2018:BRS

[Rap18] Renée J. Raphael. Book review: Silver by Fire, Silver by Mercury: a Chemical History of Silver Refining in New Spain and Mexico, 16th to 19th Centuries. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):408–410, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1518683.

Rasmussen:2018:REH

[Ras18] Seth C. Rasmussen. Revisiting the early history of synthetic polymers: Critiques and new insights. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):356–372, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1512775.

Ratiansi:1963:PPR

[Rat63] P. M. Ratiansi. Paracelsus and the Puritan revolution. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(1):24–32, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1963.11.1.
                                        24.

Rattansi:1964:HGC

[Rat64] P. M. Rattansi. The Helmontian– Galenist controversy in Restoration England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 12(1):1–23, 1964. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1964.12.1.1.

Rattansi:1966:ANM

[Rat66]  P. M. Rattansi.  Alchemy

and natural magic in Raleigh’s “history of the world”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(3):122–138, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1966.13.3.

Rattansi:1969:R

[Rat69] P. M. Rattansi. Review. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 16(3):173–175, 1969. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1969.16.3.
173.

Rattansi:1978:R

[Rat78] P. M. Rattansi. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(1):63–68, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1978.25.1.
63.

Rattansi:1992:ERA

[Rat92a] P. M. Rattansi. Essay review:

Alchemy revisited: Alchemy Revisited. Proceedings of the International Conference in the History of Alchemy at the University of Groningen 17–19 April 1989. Edited by Z. R. W. M. Von Martels. Pp. xii + 284. E. J. Brill: Leiden, New York, etc. 1990. Gld. 140, ca $80. ISBN 90-04-09287-0. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(2):91–95, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

                                        abs/10.1179/amb.1992.39.2.
                                        91.

Rattansi:1992:R

[Rat92b] P. M. Rattansi. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(2):95–102, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1992.39.2.
                                        95.

Rattansi:1996:ER

[Rat96] P. M. Rattansi. Essay review. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(2):99–100, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1996.43.2.
                                        99.

Ruiz-Castell:2016:SID

[RC16] Pedro Ruiz-Castell. Scientific instruments on display. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):75–76, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1201304.

Read:1938:AUJ

[Rea38] John Read. Alchemy under James IV of Scotland. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(2):60–67, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1938.2.2.
60.

Read:1961:WDA

[Rea61] J. Read. William Davidson of Aberdeen the first British Professor of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(2):70–101, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1961.9.2.
70.

Rees:1975:FBSa

[Ree75a] Graham Rees. Francis Bacon’s semi-Paracelsian cosmology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):81–101, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
81.

Rees:1975:FBSb

[Ree75b] Graham Rees. Francis Bacon’s semi-Paracelsian cosmology and the Great Instauration. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(3):161–173, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1975.22.3.
                                        161.

Rees:1977:FBC

[Ree77a]  Graham Rees.  The fate

of Bacon’s cosmology in the Seventeenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(1):27–38, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.1.
                                        27.

Rees:1977:MTU

[Ree77b] Graham Rees. Matter theory: a unifying factor in Bacon’s natural philosophy? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(2):110–125, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.2.
                                        110.

Rees:1979:FBV

[Ree79] Graham Rees. Francis Bacon on verticity and the bowels of the Earth. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):202–211, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/

abs/10.1179/amb.1979.26.3.
202.

Rees:1984:R

[Ree84] G. C. Rees. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(3):138–144, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.3.
138.

Rees:1995:R

[Ree95] Graham Rees. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(1):52–62, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1995.42.1.
52.

Rees:1997:R

[Ree97] Graham Rees. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(2):98–112, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.2.
98.

Reed:2012:AIP

[Ree12] Peter Reed. The alkali inspectorate 1874–1906: Pressure for wider and tighter pollution regulation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(2):131–151, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13345259996001.

Reed:2015:MWW

[Ree15] Peter Reed. Making war work for industry: The United Alkali Company’s Central Laboratory during World War One. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(1):72–93, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/1745823414Y.
0000000009.

Reed:2020:GDT

[Ree20] Peter Reed. George E. Davis (1850–1907): Transition from consultant chemist to consultant chemical engineer in a period of economic pressure. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):252–270, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1794674.

Rees:2022:BRR

[Ree22] Jonathan Rees. Book review: A Rainbow Palate: How Chemical Dyes Changed the West’s Relationship With Food. By Carolyn Cobbold. Pp. 288, illus., index. University of Chicago Press: Chicago. 2020. £30. ISBN: 978-0-22-672705-9. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):334–335, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2066257.

Reed:2023:GDE

[Ree23] Peter Reed. George E. Davis:

Editing the Chemical Trade Journal, 1887–1906. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):131–149, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2201747.

Reidy:1957:TNO

[Rei57] J. Reidy. Thomas Norton and the Ordinall of Alchimy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(2):59–85, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.2.
59.

Reiding:2010:PDN

[Rei10] Jurrie Reiding. Peter Debye: Nazi collaborator or secret opponent? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 57(3):275–300, 2010. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582310X12849808295706.

Reinhardt:2015:SCT

[Rei15] Carsten Reinhardt. Sites of chemistry in the Twentieth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(2):109–113, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/0002698015Z.
                                        00000000077.

Restrepo:2023:YPT

[Res23]  Guillermo Restrepo.  150

years of the periodic table: a commemorative symposium. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):102–104, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2137978.

Rubiano-Galvis:2023:AHT

[RGG23] Sebastián Rubiano-Galvis and Jimena Diaz Leiva Ruth Goldstein. Amalgamated histories: Tracing quicksilver’s legacy through environmental and political bodies in Andean and Amazonian gold mining. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):54–76, 2023. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2023.2189387.

Roos:2014:R

[RHF^{+}14] Anna Marie Roos, Georgiana D. Hedesan, Robert Fox, Bernadette Bensaude-Vincent, Jeffrey Allan Johnson, Duncan Wilson, Robert G. W. Kirk, Alfredo Menéndez-Navarro, and William H. Brock. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(3):305–318, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000059.

Rice:1998:MPO

[Ric98] Richard E. Rice. Mendeleev’s public opposition to spiritualism. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):85–95, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.2.

Rice:2004:HAO

[Ric04]  Richard E. Rice.  Henry

Armstrong on the offensive: Association as an alternative to dissociation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(1):5–21, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2004.51.1.
                                        5.

Richter:2009:WKA

[Ric09] Tonio Sebastian Richter. What kind of alchemy is attested by Tenth-Century Coptic manuscripts? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):23–35, 2009. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/174582309X405200.

Rinotas:2017:SAL

[Rin17] Athanasios Rinotas. Stoicism and alchemy in Late Antiquity: Zosimus and the concept of Pneuma. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):203–219, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1412134.

Rivest:2018:CCL

[Riv18] Justin Rivest. The chymical Capuchins of the Louvre: Seminal principles and charitable vocations in France under Louis XIV. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):275–295, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1512781.

Rivest:2022:SRC

[Riv22] Justin Rivest. A singular remedy: Cinchona across the Atlantic world, 1751–1820. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):89–91, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1964797.

Roberts:2006:PCE

[RK06] Lissa Roberts and Rina Knoeff. The places of chemistry in eighteenth-century Great Britain and The Netherlands. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3):197– 200, November 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X148056.

Robbins:1966:ATM

[Rob66] Rossell Hope Robbins. Alchemical texts in Middle English verse: Corrigenda and addenda. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):62–73, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2013.13.2.
                                        001.

Roberts:2006:PJK

[Rob06] Lissa Roberts. P. J. Kasteleyn and the “oeconomics” of Dutch chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3): 255–272, November 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582306X148065.

Robertson:2016:RSC

[Rob16] Haileigh Robertson. Reworking Seventeenth-Century saltpetre. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):145–161, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1218129.

Rocke:1983:SSO

[Roc83] Alan J. Rocke. Subatomic speculations and the origin of structure theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(1):1–18, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.1.
1.

Rocke:1985:APC

[Roc85] A. J. Rocke. Agricola, Paracelsus, and “chymia”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(1):38–45, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.1.
38.

Rocke:1987:KVT

[Roc87] A. J. Rocke. Kolbe versus the “transcendental chemists”: the emergence of classical organic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):156–168, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.
156.

Rocke:1994:HSH

[Roc94] A. J. Rocke. History and science, history of science: Adolphe Wurtz and the renovation of the academic professions in France. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):20–32, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.1.
                                        20.

Rocke:2003:OSG

[Roc03] Alan J. Rocke. Origins and spread of the “Giessen model” in university science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(1):90–115, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2003.50.1.
                                        90.

Rocke:2019:LMP

[Roc19] Alan J. Rocke. Lothar Meyer’s pathway to periodicity. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):265–302, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1677976.

Rocke:2021:WLA

[Roc21] Alan J. Rocke. A woman’s life alongside chemistry: The memoirs of Theresa Kopp Baumann. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(1):1–27, 2021.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2020.1867785.

Rocke:2022:AWV

[Roc22]  Alan J. Rocke.  Alexan-

der Williamson: a Victorian chemist and the making of modern Japan. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):195–196, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.1999630.

Roos:2001:TPC

[Roo01] Anna Marie Roos. Thomas philipot and chemical theories of the tides in Seventeenth-Century England. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(3):125–136, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.3.
125.

Roos:2004:MLF

[Roo04] Anna Marie Roos. Martin Lister (1639–1712) and fools’ gold. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(1):23–41, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2004.51.1.
                                        23.

Roos:2007:NGS

[Roo07] Anna Marie Roos. Nehemiah Grew (1641–1712) and the saline chymistry of plants. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):51–68, 2007. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X165443.

Roos:2017:AQU

[Roo17a] Anna Marie Roos. An alchemical quest for universal knowledge: The ‘Christian philosophy’ of Jan Baptist Van Helmont (1579–1644). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):100–101, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1310426.

Roos:2017:LSC

[Roo17b] Anna Marie Roos. The Leibniz-Stahl controversy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):192–193, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1346953.

Roos:2018:BRI

[Roo18] Anna Marie Roos. Book review: Isaac Newton and Natural Philosophy. Renaissance Lives Series. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):399–400, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1469316.

Roos:2023:BRC

[Roo23] Anna Marie Roos. Book review: A Cultural History of Chemistry in the Early Modern Age (vol. 3). Edited by Bruce Moran. Pp. 265, illus., index. Bloomsbury: London and New York. 2022. £395.00 (6-volume set). ISBN 978-1-4742-9459-1. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):190–192, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2186581.

Roque:2016:BRM

[Roq16] Xavier Roqué. Book review:

Making Marie Curie: Intellectual Property and Celebrity Culture in an Age of Information. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):356–357, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1279414.

Roque:2017:OHY

[Roq17]  Xavier Roqué.  One hun-

dred years of the Bohr atom. Proceedings from a conference. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):384–385, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1405897.

Rossiter:1996:CLK

[Ros96] Margaret W. Rossiter. Chemical librarianship: a kind of ‘women’s work’ in America. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(1):46–58, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1996.43.1.
                                        46.

Rosner:2002:OCH

[Ros02] Robert Rosner. Organic chemistry in the Habsburg Empire between 1845–1865. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(2):112–126, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.2002.49.2.
112.

Russell:2005:CHR

[RR05] Colin Archibald Russell and Gerrylynn K. Roberts, editors. Chemical History: Reviews of the Recent Literature. Royal Society of Chemistry, Cambridge, UK, 2005. ISBN 0-85404-464-7 (hardcover). xii + 247 pp. LCCN QD11 .C478 2005.

Reinhardt:2004:AIC

[RS04] Carsten Reinhardt and Harm G.

Schröter. Academia and industry in chemistry: The impact of state intervention, and the effects of cultural values. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):99–106, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.2.
99.

Richmond:2003:PIG

[RST03] Lesley Richmond, Julie Stevenson, and Alison Turton, editors. The Pharmaceutical Industry: a Guide to Historical Records. Studies in British business archives. Ashgate, Aldershot, Hampshire, UK, 2003. ISBN 0-7546-3352-7. ix + 561 pp. LCCN HD9667.5 .P465 2003.

Reinhardt:1997:IAD

[RT97] Carsten Reinhardt and Anthony S. Travis. The introduction of aniline dyes to paper printing and Queen Victoria’s postage stamps. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(1):11–18, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1997.44.1.
11.

Ruellet:2022:PCM

[Rue22] Aurélien Ruellet. The pewterer and the chymist: Major Erasmus Purling and his refined tin. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):118–138, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2050102.

Runstedler:2018:BRC

[Run18] Curtis Runstedler. Book review: Chaucer the Alchemist: Physics, Mutability, and the Medieval Imagination. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):410–411, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1520676.

Runstedler:2020:BRY

[Run20] Curtis Runstedler. Book review: Ymage de Vie: Spéculation et Expérimentation dans un Traité d’Alchimie Médiévale, by Geneviève Dumas. Pp. 293, illus., index. Presses universitaires de la Méditerranée: Montpellier. 2019. £19. ISBN: 978-2-36781-323-3. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):416–417, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1806514.

Ruska:1937:MRH

[Rus37] Julius Ruska. Methods of research in the history of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):21–29, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1937.1.1.
21.

Russell:1987:CRS

[Rus87] C. A. Russell. The changing role of synthesis in organic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):169–180, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.

Russell:1996:EFC

[Rus96] Colin Archibald Russell. Edward Frankland: Chemistry, Controversy, and Conspiracy in Victorian England. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49636-5 (hardcover). xx + 535 pp. LCCN QD22.F65 R86 1996.

                                        URL http://www.loc.gov/
                                        catdir/description/cam027/
                                        95040319.html;   http://
                                        www.loc.gov/catdir/toc/cam021/
                                        95040319.html.

Russell:1998:ERC

[Rus98] Colin A. Russell. Essay review: The Chemical Gatekeeper. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(3):191–192, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.3.
                                        191.

Russell:2001:WAS

[Rus01] Colin A. Russell. William Arthur Smeaton: an appreciation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):53–55, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.2.
                                        53.

Ruston:2019:HDA

[Rus19]  Sharon Ruston.  Humphry

Davy: Analogy, priority, and the “true philosopher”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(2–3):121–139, 2019. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2019.1616947.

Rudofsky:1988:PDA

[RW88] Susanna F. Rudofsky and John H. Wotiz. Psychologists and the dream accounts of August Kekulé. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(1):31–38, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.1.
31.

Ryan:1990:AMP

[Rya90] W. F. Ryan. Alchemy, magic, poisons and the virtues of stones in the Old Russian Secretum Secretorum. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(1):46–54, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.1.
46.

Ryding:1994:IAA

[Ryd94] Karin C. Ryding. Islamic alchemy according to Al-Khwarizmi. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(3):121–134, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.3.
                                        121.

Sadler:1975:AGF

[Sad75] Lynn Veach Sadler. Alchemy and Greene’s Friar Bacon and Friar Bungay. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):111–124, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1975.22.2.
                                        111.

Sadler:1977:RBA

[Sad77] Lynn Veach Sadler. Relations between alchemy and poetics in the Renaissance and Seventeenth Century, with special glances at Donne and Milton. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(2):69–76, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.2.
                                        69.

Stapleton:1962:TAT

[SAHL62] H. E. Stapleton, R. F. Azo, M. Hidayat Husain, and G. L.

Lewis. Two alchemical treatises attributed to Avicenna. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(2):41–82, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1962.10.2.
41.

Salvia:2013:EWE

[Sal13] Stefano Salvia. Emil Wohlwill’s “Entdeckung des Isomorphismus”: A nineteenth-century “material biography” of crystallography. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(3):255–284, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698013Z.
00000000033.

Saliba:2017:NAP

[Sal17] George Saliba. A new alchemical poem attributed to Khalid b. Yazid (d. ca. 705). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):220–233, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1412133.

Salomon:2019:PLB

[Sal19] Charlotte A. Abney Salomon.

The pocket laboratory: The blowpipe in Eighteenth-Century Swedish chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):1–22, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1573497.

Sanchez:2016:ARF

[Sán16a] José Ramón Bertomeu Sánchez. Alchemy in the rain forest: Politics, ecology, and resilience in a New Guinea mining area. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):360, 2016. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1279421.

Sanchez:2016:CEC

[Sán16b] José Ramón Bertomeu Sánchez. Chemistry education and contributions from history and philosophy of science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):360–361, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1279422.

Sanchez:2016:PJC

[Sán16c] José Ramón Bertomeu Sánchez. Paul J. Crutzen: a pioneer on atmospheric chemistry and climate change in the Anthropocene. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):359–360, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279420.

Sanchez:2017:AAL

[Sán17] José Ramón Bertomeu Sánchez.

Amedeo Avogadro. Lettere. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):289, 2017. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2017.1385906.

Sanchez:2021:LPF

[Sán21] José Ramón Bertomeu Sánchez.

Lead poisoning in France around 1840: Managing proofs and uncertainties in laboratories, courtrooms, and workplaces. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(1):72–96, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1868862.

Schling-Brodersen:1992:LRE

[SB92] Uschi Schling-Brodersen. Liebig’s role in the establishment of agricultural chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(1):21–31, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1992.39.1.
                                        21.

Sencar-Cupovic:1990:FFM

[SC90] Ilinka Sencar-Cupović. The foundation of the first modern chemical laboratories in Yugoslav countries. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(2):74–84, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1990.37.2.
                                        74.

Scerri:2022:EVH

[Sce22]  Eric R. Scerri.  The ele-

ments: A visual history of their discovery. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(4):427–428, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2091819.

Schofield:1967:JPN

[Sch67] Robert E. Schofield. Joseph Priestley, natural philosopher. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(1):1–15, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1967.14.1.
1.

Schuler:1973:WBE

[Sch73] Robert M. Schuler. William Blomfild, Elizabethan alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(2):75–87, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.2.
75.

Schneider:1987:R

[Sch87] Hans-Georg Schneider. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):181–184, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.
181.

Schneider:1989:FCE

[Sch89] Hans-Georg Schneider. The “fatherland of chemistry”: Early nationalistic currents in late Eighteenth Century German chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):14–21, 1989. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1989.36.1.14.

Schofield:1994:DIS

[Sch94]  K. Schofield.  The devel-

opment of Ingold’s system of organic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(2):87–107, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1994.41.2.
                                        87.

Schofield:1995:SAW

[Sch95] K. Schofield. Some aspects of the work of Arthur Lapworth. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(3):160–186, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1995.42.3.
                                        160.

Schiemann:2001:R

[Sch01] Gregor Schiemann. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):112–121, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.2.

Schummer:2006:HRM

[Sch06] Joachim Schummer. Historical roots of the “mad scientist”: Chemists in nineteenth-century literature. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53 (2):99–127, July 2006. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582306X117898.

Schilt:2019:BRR

[Sch19] Cornelis J. Schilt. Book review: Reading Newton in Early Modern Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):83–85, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1534334.

Schilt:2020:TA

[Sch20] Cornelis J. Schilt. A true adept. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):400–407, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1798618.

Schmidt:2022:CIE

[Sch22] Susanne Schmidt. Chemically imbalanced: Everyday suffering, medication, and our troubled quest for self-mastery. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):87–89, 2022. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2021.1959739.

Scott:1970:MDA

[Sco70] E. L. Scott. The “Macbridean Doctrine” of air: An Eighteenth-Century explanation of some biochemical processes, including photosynthesis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1):43–57, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.1.
43.

Sliwinska:1987:GRE

[SD87] Ludmila Sliwinska and Burtron H. Davis. The Gurvitsch rule: an example of a rule misnamed? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(2):81–88, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.2.
81.

Seaton:1956:THS

[Sea56] E. Seaton. Thomas Hariot’s secret script. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):111–114, 1956. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1956.5.3-
4.111.

Secret:1979:PAM

[Sec79] François Secret. Palingenesis, alchemy and Metempsychosis in Renaissance medicine. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(2):81–92, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.2.
81.

Seligardi:2007:BGP

[Sel07] Raffaella Seligardi. Between grocers and physicians: Pharmacists, the new chemistry, and institutional reforms in Bologna during the Napoleonic Period. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):31–50, 2007. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582307X165425.

Servos:1976:KCN

[Ser76] John W. Servos. The knowledge corporation: A. A. Noyes and chemistry at Cal-Tech, 1915– 1930. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):175–186, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1976.23.3.
                                        175.

Serrano:2013:CCS

[Ser13]  Elena Serrano.  Chemistry

in the city: The scientific role of female societies in late Eighteenth-Century Madrid. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):139–159, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000026.

Serrano:2022:PWC

[Ser22]  Elena Serrano.  Patriotic

women: Chemistry and gender in the Eighteenth-Century Spanish world. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):243–261, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2022.2097492.

Garrioch:2023:ASI

[SG23] Barry Sturman and David Garrioch. Amateur science and innovation in fireworks in Nineteenth-Century Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):109–130, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2201743.

Shapin:1994:SHT

[Sha94] Steven Shapin. A Social History of Truth: Civility and Science in Seventeenth-Century England. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-75018-3. xxxi + 483 pp. LCCN Q175.52.G7 S48 1994. URL

http://www.loc.gov/catdir/
description/uchi052/93041950.
html;   http://www.loc.gov/
catdir/toc/uchi051/93041950.
html.

Sharp:2002:CTA

[Sha02] Robert Sharp. The correspondents of Thomas Andrews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(2):148–157, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.2.
148.

Sharp:2003:OSW

[Sha03] Robert Sharp. Oswald Silberrad: the work of a forgotten chemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(3):302–309, 2003.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.2003.50.3.
                                        302.

Shackelford:2005:ERL

[Sha05] Jole Shackelford. Essay review: Late Medieval and Early Modern Corpuscular Matter Theories, Edited by Christoph Lüthy, John E. Murdoch, and William R. Newman. Pp. viii + 611, index. Brill: Leiden, Boston, Köln. 2001. EUR 160; $186. ISBN 90-04-11516-1. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52 (2):175–178, July 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269805X70120.

Shackelford:2020:CPV

[Sha20]  Jole Shackelford.  Chemi-

cal paradigm vs. biological paradigm in the biological clock controversy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(4):366–388, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1826821.

Sherlock:1948:CWP

[She48] T. P. Sherlock. The chemical work of Paracelsus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(1–2):33–63, 1948. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1948.3.1-
2.33.

Sheppard:1957:GA

[She57]  H. J. Sheppard.  Gnosti-

cism and alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(2):86–101, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.2.
86.

Sheppard:1958:ESA

[She58] H. J. Sheppard. Egg symbolism in alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(3):140–148, 1958. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1958.6.3.
140.

Sheppard:1959:RTH

[She59]  H. J. Sheppard.  The re-

demption theme and Hellenistic Alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(1):42–46, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1959.7.1.
                                        42.

Sheppard:1960:SAH

[She60] H. J. Sheppard. A survey of alchemical and hermetic symbolism. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(1):35–41, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1960.8.1.
                                        35.

Sheppard:1962:OUM

[She62] H. J. Sheppard. The Ouroboros and the unity of matter in alchemy: A study in origins. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(2):83–96, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1962.10.2.
                                        83.

Shea:1970:GAH

[She70a] William R. Shea. Galileo’s atomic hypothesis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1):13–27, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/

abs/10.1179/amb.1970.17.1.
13.

Sheppard:1970:AOO

[She70b] H. J. Sheppard. Alchemy:

Origin or origins? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(2):69–84, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.2.
69.

Sheppard:1971:R

[She71] H. J. Sheppard. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(3):215–223, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1971.18.3.
215.

Sheppard:1975:R

[She75] H. J. Sheppard. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(3):221–222, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.3.
221.

Sheppard:1977:R

[She77] H. J. Sheppard. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(2):126–132, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1977.24.2.
126.

Sheppard:1979:R

[She79] H. J. Sheppard. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):224–230, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.3.
224.

Sheppard:1981:Ra

[She81a] H. J. Sheppard. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(1):55–60, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.1.
55.

Sheppard:1981:Rb

[She81b] H. J. Sheppard. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 28(3):210–217, 1981. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1981.28.3.
210.

Sheppard:1985:CWA

[She85a]  H. J. Sheppard.  Chinese

and Western alchemy: The link through definition. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(1):32–37, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.1.
32.

Sheppard:1985:R

[She85b] H. J. Sheppard. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(2):92–96, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.2.
92.

Sheppard:1990:R

[She90] H. J. Sheppard. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 37(2):85–95, 1990. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1990.37.2.
85.

Sheppard:1991:R

[She91] H. J. Sheppard. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(1):63, 1991.

CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1991.38.1.
63.

Shirley:1949:SES

[Shi49] John William Shirley. The scientific experiments of Sir Walter Raleigh, the wizard Earl, and the Three Magi in the Tower 1603–1617. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(1–2):52–66, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.4.1-
2.52.

Shmurak:1994:EPC

[Shm94] Carole B. Shmurak. Emma Perry Carr: The spectrum of a life. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(2):75–86, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1994.41.2.
75.

Shorter:1986:R

[Sho86] John Shorter. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):153– 160, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.2-
3.153.

Shorter:2003:BJF

[Sho03]  John Shorter.  Bernhard

Jacques Flürscheim (1874– 1955), organic and theoretical chemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(3):274–301, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2003.50.3.
274.

Siegfried:1982:LTS

[Sie82] Robert Siegfried. Lavoisier’s table of simple substances: Its origin and interpretation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(1):29–48, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.1.
29.

Siegfried:1989:LPC

[Sie89] Robert Siegfried. Lavoisier and the phlogistic connection. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):31–40, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1989.36.1.
31.

Siegel:2008:SRR

[Sie08] Andre Siegel. Sir Robert Robinson’s “anthocyanin period”: 1922–1934 — a case study of an Early Twentieth-Century natural products synthesis. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(1):62–82, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X255406.

Simcock:1987:AWS

[Sim87]  A. V. Simcock.  Alchemy

and the world of science: an intellectual biography of Frank Sherwood Taylor. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):121–139, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.
121.

Simon:1998:CRP

[Sim98] Jonathan Simon. The chemical revolution and pharmacy: a disciplinary perspective. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(1):1–13, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/

abs/10.1179/amb.1998.45.1.
1.

Simon:2002:AAM

[Sim02] Jonathan Simon. Authority and authorship in the method of chemical nomenclature. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(3):206–226, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.3.
206.

Simmons:2006:MMM

[Sim06] Anna Simmons. Medicines, monopolies and mortars: the chemical laboratory and pharmaceutical trade at the Society of Apothecaries in the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(3):221–236, November 2006. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582306X148083.

Simmons:2014:SSS

[Sim14a] Anna Simmons. Stills, status, Stocks and science: The laboratories at Apothecaries’ Hall in the Nineteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(2):141–161, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698014Z.
                                        00000000050.

Simoes:2014:SBR

[Sim14b] Ana Simões. Scientific biographies revisited: Thomsons’ electrons and Bohr’s quantum atoms. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(1):95–99, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1179/0002698014Z.
                                        00000000047.

Simoes:2016:MCS

[Sim16a] Ana Simões. Making 20th century science. How theories became knowledge. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):281–282, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1246826.

Simon:2016:GDM

[Sim16b] Jonathan Simon. Gendered drugs and medicine. Historical and socio-cultural perspectives. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):280–281, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2016.1246825.

Simon:2016:SMP

[Sim16c] Jonathan Simon. The shaping of modern pharmacy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(1):66–70, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1201296.

Simmons:2020:LCH

[Sim20] Anna Simmons. A life of “continuous and honourable usefulness:” chemical consulting and the career of Robert Warington (1807–1867). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(3):234–251, 2020. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2020.1794698.

Simmons:2022:DPP

[Sim22] Anna Simmons. Drugs on the page: Pharmacopoeias and healing knowledge in the Early Modern Atlantic world. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):193–195, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.2019419.

Singer:1959:CCC

[Sin59] Dorothea Waley Singer. On a 16th century cartoon concerning the devilish weapon of gunpowder: Some Medieval reactions to guns and gunpowder: Some Medieval reactions to guns and gunpowder. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(1):25–33, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1959.7.1.
25.

Sinclair:1985:CRI

[Sin85] S. B. Sinclair. Crookes and radioactivity: from inorganic evolution to atomic transmutation. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(1):15–31, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.1.
15.

Sinclair:1987:JJT

[Sin87] S. B. Sinclair. J. J. Thomson and the chemical atom: From ether vortex to atomic decay. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(2):89–116, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1987.34.2.
89.

Sinclair:1988:JJTa

[Sin88a] S. B. Sinclair. J. J. Thomson and radioactivity: Part I. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(2):91–104, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.2.
91.

Sinclair:1988:JJTb

[Sin88b] S. B. Sinclair. J. J. Thomson and radioactivity: Part II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(3):113–126, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.3.
113.

Stark:1987:EDN

[SK87] Johannes Stark and Andreas Kleinert. Erinnerungen eines deutschen Naturforschers. (German) [Memories of a German Natural Scientist]. Bionomica-Verlag, Mannheim, Germany, 1987. ISBN 3-88208-060-0, 3-88208-011-6. x + 153 pp. LCCN QC16.S764 A3 1987.

Serrano:2022:LSC

[SL22] Elena Serrano and Joris Mercelis Annette Lykknes. “I am not a Lady, I am a scientist.” chemistry, women, and gender in the enlightenment and the era of professional science. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):203–220, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2100954.

Slater:2001:WRS

[Sla01]  Leo B. Slater.  Wood-

ward, Robinson, and strychnine: Chemical structure and chemists’ challenge. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(3):161–189, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2001.48.3.
161.

Slater:2004:MCK

[Sla04]  Leo B. Slater.  Malaria

chemotherapy and the “kaleidoscopic” organisation of biomedical research during World War II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):107–134, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.2.
107.

Stapleton:1949:SHQ

[SLT49] H. E. Stapleton, G. L. Lewis, and F. Sherwood Taylor. The sayings of Hermes quoted in the Ma Al-Waraqi of Ibn Umail of ibn Umail. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(3–4):69–90, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.3.3-
4.69.

Suay-Matallana:2017:HWE

[SM17] Ignacio Suay-Matallana. Healing with water: English spas and the water cure, 1840– 1960. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):196–197, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1346956.

Smeaton:1957:BGM

[Sme57] W. A. Smeaton. L. B. Guyton De Morveau (1737–1816): A bibliographical study. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(1):18–34, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.1.

Smeaton:1959:FJB

[Sme59] W. A. Smeaton. F.-J. Bonjour and his translation of Bergman’s “Disquisitio De Attractionibus Electivis”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(1):47–50, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1959.7.1.
                                        47.

Smeaton:1961:GMC

[Sme61] W. A. Smeaton. Guyton De Morveau’s course of chemistry in the Dijon Academy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(2):53–69, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1961.9.2.
                                        53.

Smeaton:1963:GMC

[Sme63] W. A. Smeaton. Guyton De Morveau and chemical affinity. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(2):55–64, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1080/00026980.1963.
                                        11923894.

Smeaton:1966:PCL

[Sme66] W. A. Smeaton. The portable chemical laboratories of Guyton De Morveau, Cronstedt and Göttling. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):84–91, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2013.13.2.
003.

Smeaton:1968:WCA

[Sme68] W. A. Smeaton. Is water converted into air? Guyton de Morveau acts as arbiter between Priestley and Kirwan. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(2):73–83, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1968.15.2.
73.

Smeaton:1971:FGW

[Sme71] W. A. Smeaton. E. F. Geoffroy was not a Newtonian chemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 18(3):212–214, 1971. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1971.18.3.
212.

Smeaton:1977:BES

[Sme77] W. A. Smeaton. Berthollet’s Essai de Statique Chimique and its translations: a bibliographical note and a Daltonian doubt. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):149–158, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.3.
                                        149.

Smeaton:1978:BES

[Sme78]  W. A. Smeaton.  Berthol-

let’s Essai de Statique Chimique: a supplementary note. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(3):211–212, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1978.25.3.
                                        211.

Smeaton:1987:SFFa

[Sme87a] W. A. Smeaton. The Society’s first fifty years: Part I — Ambix. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(1): 1–4, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1987.34.1.

Smeaton:1987:SFFb

[Sme87b] W. A. Smeaton. The Society’s first fifty years: Part II — members and meetings. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(2):57–61, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.2.
57.

Smeaton:1987:TBA

[Sme87c] W. A. Smeaton. Two books are added to Guyton De Morveau’s library: a study of personal and academic communications in 1785. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(3):140–146, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.3.
140.

Smeaton:1989:ER

[Sme89a] W. A. Smeaton. Essay review.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):41–42, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.1.
41.

Smeaton:1989:MLP

[Sme89b] W. A. Smeaton. Madame Lavoisier, P. S. and E. I. Du Pont De Nemours and the publication of Lavoisier’s “Mémoires De Chimie”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):22–30, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.1.
22.

Smeaton:1989:MML

[Sme89c] W. A. Smeaton. Monsieur and Madame Lavoisier in 1789: The Chemical Revolution and the French Revolution. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):1–4, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1989.36.1.
1.

Suay-Matallana:2018:PWR

[SMGL18] Ignacio Suay-Matallana and Ximo Guillem-Llobat. Poisoned wine: Regulation, chemical analyses, and Spanish–French trade in the 1930s. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2):99–121, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Smith:1980:R

[Smi80] Crosbie Smith. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(2):142–146, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1980.27.2.
142.

Snelders:1979:AEA

[Sne79] H. A. M. Snelders. The Amsterdam experiment on the analysis and synthesis of water (1789). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(2):116–133, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.2.
116.

Snobelen:1998:R

[Sno98] Stephen D. Snobelen. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):130–135, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.2.
130.

Soderqvist:2007:HPS

[Söd07] Thomas Söderqvist, editor. The History and Poetics of Scientific Biography. Science, technology and culture, 1700–1945. Ashgate, Aldershot, Hampshire, UK, 2007. ISBN 0-7546-5181-9. xiv + 270 pp. LCCN Q141 .H575 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0620/2006029937.html.

Sonntag:1977:RST

[Son77] Otto Sonntag. Religion and science in the thought of Liebig. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 24(3):159–169, 1977. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1977.24.3.
                                        159.

Sourkes:1992:VBC

[Sou92] Theodore L. Sourkes. Vauquelin, Balzac and the chemical analysis of hair. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(1):11–16, 1992. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1992.39.1.11.

Sourkes:2000:DEJ

[Sou00] Theodore L. Sourkes. Devitalising the elements: Johann Friedrich John (1782– 1847) and the liberation of phosphorus and potassium from a vital force. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(1):37–46, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.2000.47.1.
37.

Sparling:2020:PTA

[Spa20] Andrew Sparling. Paracelsus, a transmutational alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 67(1):62–87, 2020. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2020.1720358.

Stapleton:1953:AA

[Sta53] H. E. Stapleton. The antiquity of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(1–2):1–43, 1953. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1953.5.1-
2.1.

Stapleton:1957:G

[Sta57] H. E. Stapleton. The gnomon.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 6(1):1–9, 1957. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1957.6.1.
1.

Stavenhagen:1970:OTL

[Sta70] Lee Stavenhagen. The original text of the Latin Morienus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(1):1–12, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1970.17.1.
1.

Stanley:1987:Ra

[Sta87a] Michael Stanley. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(1):45–56, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.1.
45.

Stanley:1987:Rb

[Sta87b] Michael Stanley. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 34(2):117–120, 1987. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1987.34.2.
117.

Stanley:1996:ED

[Sta96] Michael Stanley. An error in dating. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(3): 188, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1996.43.3.

Stackenwalt:1998:DIM

[Sta98] Francis Michael Stackenwalt.

Dmitrii Ivanovich Mendeleev and the emergence of the modern Russian petroleum industry, 1863–1877. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(2):67–84, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.2.
67.

Stroup:1979:WHS

[Str79] Alice Stroup. Wilhelm Homberg and the search for the constituents of plants at the 17th-Century Academie Royale des Sciences. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(3):184–201, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.3.
184.

Strbanova:2016:HHB

[Štr16] Soňa Štrbáňová. Holding Hands with Bacteria: the Life and Work of Marjory Stephenson. SpringerBriefs in molecular science. History of chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-49734-4, 3-662-49736-0 (e-book). ISSN 2212-991X. xx + 145 pp. LCCN QP511.8.S74 S77 2016.

Succurro:2016:LCA

[Suc16] Maria Chiara Succurro. The Liber Compostelle attributed to Friar Bonaventura of Iseo: The textual tradition of a Thirteenth-Century alchemical encyclopedia. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):199–216, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1246299.

Sudduth:1978:ECI

[Sud78] William M. Sudduth. Eighteenthcentury identifications of electricity with phlogiston. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 25(2):131–147, 1978. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1978.25.2.
                                        131.

Sudduth:1980:VPE

[Sud80] William M. Sudduth. The Voltaic pile and Electrochemical theory in 1800. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(1):26–35, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.1980.27.1.
26.

Sumner:2007:MCP

[Sum07] James Sumner. Michael Combrune, Peter Shaw and commercial chemistry: the Boerhaavian chemical origins of brewing thermometry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(1):5–29, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582307X165452.

Sutton:1976:SCN

[Sut76] M. A. Sutton. Spectroscopy and the chemists: a neglected opportunity? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(1):16–26, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.1.
16.

Sutton:1983:EA

[Sut83] Michael A. Sutton. Editorial announcement. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(2): 116, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.2.

Sutton:1998:R

[Sut98] Michael A. Sutton. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(1):34–44, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.1.
                                        34.

Sutton:1999:R

[Sut99] Michael A. Sutton. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(1):52–62, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1999.46.1.
                                        52.

Soukup:1993:ACP

[SvOM93] R. W. Soukup, S. von Osten, and H. Mayer. Alembics, cucurbits, phials, crucibles: a 16th-century docimastic laboratory excavated in Austria. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(1): 25, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.1.
                                        25.

Steere-Williams:2014:CAA

[SW14] Jacob Steere-Williams. A conflict of analysis: Analytical chemistry and milk adulteration in Victorian Britain. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(3):279–298, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1179/0002698014Z.
00000000058.

Szulakowska:1986:TAI

[Szu86] Urszula Szulakowska. The tree of Aristotle: Images of the philosophers’ stone and their transference in alchemy from the Fifteenth to the Twentieth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):53–77, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1986.33.2-
3.53.

Szulakowska:1988:TCM

[Szu88] Urszula Szulakowska. Thirteenth century material pantheism in the Pseudo-Lullian ‘S’- circle of the powers of the soul. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(3):127–154, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.3.

Szydlo:1993:AMS

[Szy93] Zbigniew Szydlo. The alchemy of Michael Sendivogius: His central nitre theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(3):129–146, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.3.
                                        129.

Szydlo:1996:ICN

[Szy96] Zbigniew Szydlo. The influence of the central nitre theory of Michael Sendivogius on the chemical philosophy of the Seventeenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 43(2):80–96, 1996. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1996.43.2.
                                        80.

Taape:2014:DRR

[Taa14] Tillmann Taape. Distilling reliable remedies: Hieronymus Brunschwig’s Liber de arte distillandi (1500) between alchemical learning and craft practice. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(3):236–256, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

full/10.1179/0002698014Z.
00000000060.

Taape:2023:CHC

[Taa23] Tillmann Taape. A cultural history of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):184–186, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2186576.

Takabayashi:2017:BRK

[Tak17] Akinobu Takabayashi. Book review: Kagakushi Jiten [Encyclopedic Dictionary of the History of Chemistry]. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):388–389, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1405905.

Tarr:2002:IWD

[Tar02] Joel A. Tarr. Industrial waste disposal in the United States as a historical problem. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(1):4–20, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2002.49.1.

Tarr:2003:DRE

[Tar03] Joel A. (Joel Arthur) Tarr, editor. Devastation and Renewal: an Environmental History of Pittsburgh and Its Region. University of Pittsburgh Press, Pittsburgh, PA, USA, 2003. ISBN 0-8229-4156-2. viii + 281 pp. LCCN GE155.P4 D48 2003.

URL http://www.h-net.org/
review/hrev-a0d8d3-aa;
http://www.loc.gov/catdir/
enhancements/fy0809/2003015126-
d.html;   http://www.loc.
gov/catdir/toc/ecip046/2003015126.
html.

Tarr:2005:DRE

[Tar05] Joel A. (Joel Arthur) Tarr, editor. Devastation and Renewal: an Environmental History of Pittsburgh and Its Region. University of Pittsburgh Press, Pittsburgh, PA, USA, 2005. ISBN 0-8229-5892-9. viii + 281 pp. LCCN GE155.P4 D48 2005. URL http:

//catdir.loc.gov/catdir/
enhancements/fy0809/2003015126-
d.html;   http://catdir.
loc.gov/catdir/toc/ecip046/
2003015126.html.

Tarrant:2018:BAE

[Tar18] Neil Tarrant. Between Aquinas and Eymerich: The Roman Inquisition’s use of Dominican thought in the censorship of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):210–231, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2018.1512779.

Tate:2019:HDP

[Tat19] Gregory Tate. Humphry Davy and the problem of analogy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66 (2–3):140–157, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1616944.

Taylor:1937:AWS

[Tay37a] F. Sherwood Taylor. The alchemical works of Stephanos of Alexandria. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(2):116–139, 1937. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1937.1.2.116.

Taylor:1937:OGA

[Tay37b] F. Sherwood Taylor. The origins of Greek alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):30–47, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1937.1.1.

Taylor:1937:VZ

[Tay37c]  F. Sherwood Taylor.  The

visions of Zosimos. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(1):88–92, 1937. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1937.1.1.
                                        88.

Taylor:1938:AWS

[Tay38] F. Sherwood Taylor. The alchemical works of Stephanos of Alexandria. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):39–49, 1938. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1938.2.1.39.

Taylor:1949:APD

[Tay49a] F. Sherwood Taylor. Alchemical papers of Dr. Robert Plot. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(1–2):67–76, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1949.4.1-
                                        2.67.

Taylor:1949:PAI

[Tay49b] F. Sherwood Taylor. A pair of alchemical ivory figures. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(1–2):77–78, 1949. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1949.4.1-
2.77.

Taylor:1956:AWS

[Tay56] F. Sherwood Taylor. An alchemical work of Sir Isaac Newton. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):59–84, 1956. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1956.5.3-
4.59.

Taylor:2008:TIS

[Tay08] Georgette Taylor. Tracing influence in small steps: Richard Kirwan’s quantified affinity theory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(3):209–231, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X358088.

Taylor:2014:PPTa

[Tay14a] Georgette Taylor. Pedagogical progeniture or tactical translation? George Fordyce’s additions and modifications to William Cullen’s philosophical chemistry — Part I. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(1):48–66, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000045.

Taylor:2014:PPTb

[Tay14b] Georgette Taylor. Pedagogical progeniture or tactical translation? George Fordyce’s additions and modifications to William Cullen’s philosophical chemistry — Part II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(3):257–278, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698014Z.
                                        00000000056.

Taylor:2019:BRC

[Tay19] Georgette Taylor. Book review: The Chemical Works of Carl Wilhelm Scheele. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(1):90–92, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2019.1579849.

Taylor:2023:BRC

[Tay23] Georgette Taylor. Book review: A Cultural History of Chemistry in the Eighteenth Century (vol. 4). Edited by Matthew Daniel Eddy and Ursula Klein. Pp. 238, illus. index. Bloomsbury: London, 2022. £395.00 (6-volume set). ISBN: 978-1-4742-9465-2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(2):192–194, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2186583.

Tien-Chin:1959:EMC

[TCPYN59] Ts’ao T’ien-Ch’in, Ho Ping-Yü, and Joseph Needham. An early mediaeval Chinese alchemical text on aqueous solutions. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(3):121–155, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1959.7.3.
121.

Teich:1982:CTC

[Tei82] Mikulás Teich. Circulation, transformation, conservation of matter and the balancing of the biological world in the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(1):17–28, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.1.

Teich:1995:R

[Tei95] Mikulás Teich. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(3):187–192, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1995.42.3.
                                        187.

Thackray:1968:MNS

[Tha68] Arnold Thackray. “Matter in a nut-shell”: Newton’s Opticks and Eighteenth-Century chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(1):29–53, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.1.
                                        29.

Theissen:1986:JDL

[The86] Wilfred R. Theissen. John Dastin’s letter on the philosophers’ stone. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 33(2–3):78–87, 1986. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1986.33.2-
                                        3.78.

Theisen:1991:JDA

[The91] Wilfred Theisen. John Dastin: The alchemist as co-creator.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(2):73–78, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1991.38.2.
73.

Theisen:1999:JDA

[The99] Wilfred Theisen. John Dastin’s alchemical vision. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(2):65–72, 1999. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1999.46.2.
65.

Thiesen:2008:LJD

[Thi08] Wilfred Thiesen. The letters of John Dastin. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(2):153–168, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582308X255389.

Thompson:1938:SCA

[Tho38a] R. Campbell Thompson. A survey of the chemistry of Assyria in the Seventh Century B.C. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):3–16, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1938.2.1.
                                        3.

Thorndike:1938:ADF

[Tho38b]  Lynn Thorndike.  Alchemy

during the first half of the Sixteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):26–38, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1938.2.1.
                                        26.

Thorndike:1956:SAM

[Tho56] Lynn Thorndike. Some alchemical manuscripts at Bologna and Florence. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):85–110, 1956. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1956.5.3-
                                        4.85.

Thorndike:1959:SMT

[Tho59a] Lynn Thorndike. Some Medieval texts on colours. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(1):1–24, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/

abs/10.1179/amb.1959.7.1.
1.

Thorndike:1959:UTM

[Tho59b]  Lynn Thorndike.  Uncata-

logued texts in Ms. All Souls College 81, Oxford. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 7(1):34–41, 1959. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1959.7.1.
34.

Thorndike:1960:L

[Tho60a] Lynn Thorndike. De lapidibus.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(1):6–23, 1960. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1960.8.1.
6.

Thorndike:1960:OTC

[Tho60b] Lynn Thorndike. Other texts on colours. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 8(2):53–70, 1960. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1960.8.2.53.

Thorndike:1963:AMK

[Tho63a] Lynn Thorndike. An alchemical manuscript: Klagenfurt, Bischoöfl. bibl. XXIX.d.24. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(2):95–96, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1080/00026980.1963.
11923897.

Thorndike:1963:PGP

[Tho63b] Lynn Thorndike. The Pseudo-Galen, De Plantis (with Latin text of chapters on stones and those of chemical interest). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(2):87–94, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1080/00026980.1963.
11923896.

Thomas:1988:PLG

[Tho88] Ulrike Thomas. Philipp Lorenz Geiger and Justus Liebig. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(2):77–90, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.2.
77.

Thompson:2022:BRA

[Tho22] Helen Thompson. Book review: The Atom in Seventeenth-Century Poetry. By Cassandra Gorman. Pp. 251, index. D. S. Brewer: Woodbridge, UK, and Rochester, NY. 2021. £75.00. ISBN: 978-1-84-384593-5. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):327–329, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2082151.

Timmermann:2006:NPC

[Tim06] Anke Timmermann. New perspectives on “The Chaucer Ascription in Trinity College, Dublin MS D.2.8”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 53(2):161–165, July 2006. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582306X117906.

Titley:1938:PRS

[Tit38] A. F. Titley. Paracelsus. a Résumé of some controversies. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 1(3):166–183, 1938. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1938.1.3.
166.

Taylor:1953:JBH

[TJ53] F. Sherwood Taylor and C. H.

Josten. Johannes Banfi Hunyades 1576–1650. Ambix:

Journal of the Society for the History of Alchemy and Chemistry, 5(1–2):44–52, 1953. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1953.5.1-
                                        2.44.

Taylor:1956:JBH

[TJ56] F. Sherwood Taylor and C. H. Josten. Johannes Banfi Hunyades. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):115, 1956. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1956.5.3-
                                        4.115.

Tomory:2009:LIB

[Tom09] Leslie Tomory. Let it burn: Distinguishing inflammable airs 1766–1790. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(3):253–272, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/000269809X12529013496374.

Tomic:2016:HCV

[Tom16a] Sacha Tomic. The history of chemistry: a very short introduction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(4):347–348, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1279407.

Tomory:2016:JWM

[Tom16b] Leslie Tomory. James Watt:

Making the world anew. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):192–193, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1227174.

Tomic:2017:HMC

[Tom17a] Sacha Tomic. A history of modern chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(3):283–284, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1385904.

Tomory:2017:IEB

[Tom17b] Leslie Tomory. The imagined empire: Balloon enlightenments in revolutionary Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):382–384, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1405896.

Tomic:2018:BRN

[Tom18] Sacha Tomic. Book review: La Navelbine(R) et le Taxotère(R): Histoires de sciences. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(3):298–300, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2018.1469313.

Torrens:2000:R

[Tor00] Hugh S. Torrens. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 47(3):196–205, 2000. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2000.47.3.
                                        196.

Tosi:2001:MMP

[Tos01] Lucia Tosi. Marie Meurdrac: Paracelsian chemist and feminist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 48(2):69–82, 2001. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2001.48.2.
                                        69.

Tanford:1999:PCB

[TR99] Charles Tanford and Jacqueline Reynolds. Protein chemists bypass the colloid/macromolecule debate. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 46(1):33–51, 1999. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1999.46.1.33.

Travis:1991:HCR

[Tra91] Anthony S. Travis. Heinrich Caro at Roberts, Dale & Co. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(3):113–134, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1991.38.3.
113.

Travis:1993:R

[Tra93] A. S. Travis. Report. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(2):98–99, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1993.40.2.
98.

Travis:1995:ADJ

[Tra95] Anthony S. Travis. Artificial dyes in John Lightfoot’s Broad Oak Laboratory. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(1):10–27, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1995.42.1.
                                        10.

Travis:2002:CEW

[Tra02]  Anthony S. Travis.  Con-

taminated Earth and water: a legacy of the synthetic dyestuffs industry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 49(1):21–50, 2002. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2002.49.1.
                                        21.

Travis:2004:DMA

[Tra04]  Anthony S. Travis.  Dyes

made in America: 1915–1980; the Calco Chemical Company, American Cyanamid, and the Raritan River. Sidney M. Edelstein Center for the History and Philosophy of Science, Technology and Medicine at the Hebrew Univ. of Jerusalem, Jerusalem, 2004. ISBN 965-555-149-0. xiv + 582 pp. LCCN HD9660.D844 C35 2004; TP914.H3.

Travis:2005:ERD

[Tra05] Anthony S. Travis. Essay review: Devastation and Renewal: an Environmental History of Pittsburgh and Its Region. Edited by Joel E. Tarr. Pp. xvii + 281 illus., index. University of Pittsburgh Press: Pittsburgh, PA, 2004. ISBN 0-8229-4156-2. Uneasy Alchemy: Citizens and Experts in Louisiana’s Chemical Corridor Disputes. By Barbara L. Allen. Pp. xiii + 211, illus., index. MIT Press: Cambridge, Mass. and London. 2003. £14.95. ISBN 0-262-51134-7. Inventing for the environment. Edited by Arthur Molella and Joyce Bedi. Pp. xix + 398, illus., index. MIT Press: Cambridge, Mass. £19.95. ISBN 0-262-13427-6. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):85–87, 2005. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/000269805X52922.

Travis:2012:SSE

[Tra12a] Anthony S. Travis. Silent Spring at 50: Earth, water, and air. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(2):83–87, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13345259995886.

Travis:2012:DCH

[Tra12b] Anthony S. Travis. Detecting chlorinated hydrocarbon residues: Rachel Carson’s villains. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(2):109–130, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1179/174582312X13345259995967.

Travis:2014:ERT

[Tra14] Anthony S. Travis. The emerging role of titrimetry in late nineteenth-century industrial problem solving: The example of trace analysis for perchlorate in Chile saltpetre. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(1):67–94, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000044.

Travis:2017:GSN

[Tra17] Anthony S. Travis. Globalising synthetic nitrogen: The interwar inauguration of a new industry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):1–28, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2017.1325585.

Trenn:1974:JTS

[Tre74] Thaddeus J. Trenn. The justification of transmutation: Speculations of Ramsay and experiments of Rutherford. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 21(1):53–77, 1974. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1974.21.1.
53.

Trenn:1979:RRA

[Tre79] Thaddeus J. Trenn. Rutherford’s Radio-Activity and alpha ray research: The case of a misdated letter. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 26(2):134–136, 1979. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1979.26.2.
134.

Trevisani:1982:R

[Tre82] Francesco Trevisani. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(2):110–124, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1982.29.2.
110.

Tyabji:2004:EAI

[Tya04] Nasir Tyabji. Exemplar of academia–industry interchange: The Department of Chemical Technology at Bombay University. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(2):149–166, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2004.51.2.
                                        149.

Usselman:2003:LAA

[Uss03] Melvyn C. Usselman. Liebig’s alkaloid analyses: the uncertain route from elemental content to molecular formulae. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 50(1):71–89, 2003. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.2003.50.1.
                                        71.

Usselman:2015:PIL

[Uss15] Melvyn C. Usselman. Pure Intelligence: the Life of William Hyde Wollaston. Synthesis. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-24573-X (hardcover), 0-226-24587-X (e-book). xv + 413 pp. LCCN Q143.W795 U87 2015.

Vagelli:2019:BRW

[Vag19] Matteo Vagelli. Book review: When Historiography Met Epistemology: Sophisticated Histories and Philosophies of Science in French-speaking Countries in the Second Half of the Nineteenth Century, by Stefano Bordoni. Pp. 335, index. Brill: Leiden. 2017. £130.00. ISBN 978-90-04-31522-8. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 66(4):360–362, 2019. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2019.1591571.

VanHelVoort:2004:ABN

[Van04] Ton Van HelVoort. Articulating biochemistry in The Netherlands after the Second World War: Science for its own sake. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(3):199–218, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.3.
199.

VanBladel:2009:AHP

[Van09] Kevin Thomas Van Bladel.

The Arabic Hermes: from Pagan Sage to Prophet of Science. Oxford studies in late antiquity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-537613-7. xii + 278 pp. LCCN BF1598.H5 B53

2009. URL http://www.loc.
gov/catdir/enhancements/
fy1007/2008050497-b.html;
http://www.loc.gov/catdir/
enhancements/fy1007/2008050497-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1007/2008050497-t.html.

VandenBroecke:2017:IKS

[Van17] Steven Vanden Broecke. The ideal of a knowledge society in Dee’s Monas Hieroglyphica (1564) and other productions by Willem Silvius. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(2):157–174, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1353230.

VanTiggelen:2022:BRN

[Van22] Brigitte Van Tiggelen. Book review: A Nobel Affair: The Correspondence Between Alfred Nobel and Sofie Hess. Edited and translated by Erika Rummel. Pp. 304, illus., index. University of Toronto Press: Toronto. 2017. £62.00. ISBN: 978-1-48-750177-8. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(3):336–337, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2066253.

Vaupel:2004:BRB

[Vau04] Elisabeth Vaupel. Book review: Bright Earth: Art and the Invention of Color. By Philip Ball. Pp. ix + 384, illus., index. Farrar, Straus and Giroux:New York. 2002. $30. ISBN 0-374-11679-2. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 51(3):269–291, 2004. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.2004.51.3.
269.

Verwaal:2022:WBH

[Ver22] Ruben E. Verwaal. White blood: a history of human milk. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(2):197–198, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2022.2058175.

vonFranz:1965:IMM

[vF65] Marie-Louise von Franz. The idea of the macro- and microcosmos in the light of Jungian psychology. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(1):22–34, 1965. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1965.13.1.
22.

Vagelos:2004:MSM

[VG04] P. Roy Vagelos and Louis Galambos. Medicine, Science, and Merck: the first three careers of Roy Vagelos. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-66295-8 (hardcover), 0-511-16420-3, 0-511-16615-X (e-book), 0-511-16545-5 (e-book), 0-511-51167-1 (ebook), 1-280-43759-6. xi + 301 pp. LCCN R154.V34 A3 2004.

URL http://www.loc.gov/
catdir/description/cam032/
2003051233.html;   http://
www.loc.gov/catdir/samples/
cam041/2003051233.html;
http://www.loc.gov/catdir/
toc/cam032/2003051233.html.

Viano:2015:MDA

[Via15] Cristina Viano. Mixis and Diagnôsis: Aristotle and the “chemistry” of the sublunary world. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 62(3):203–214, 2015. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1179/0002698015Z.
00000000086.

Vickers:1994:ERH

[Vic94] Brian Vickers. Essay review: Hermitica. The Greek Corpus Hermeticum and the Latin Asclepius in a new English translation, with notes and introduction. By Brian P. Copenhaver. Pp. lxxxiii + 320. Cambridge University Press: Cambridge, New York, Oakleigh, 1992. £45, $69.95. ISBN 0-521-36144-3. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 41(1):45–48, 1994. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/

abs/10.1179/amb.1994.41.1.
45.

Vinciguerra:2009:AAF

[Vin09] Antony Vinciguerra. The Ars alchemie: the first Latin text on practical alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):57–67, 2009. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582309X405237.

vonLippmann:1938:SRH

[vL38] Edmund O. von Lippmann.

Some remarks on Hermes and Hermetica. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 2(1):21–25, 1938. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/amb.1938.2.1.21.

vonLippmann:1948:CTR

[vL48] Edmund O. von Lippmann.

Chemical and technological references in Plutarch. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 3(1–2):1–14, 1948. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1948.3.1-
2.1.

vonMartels:1990:ARP

[vM90] Z. R. W. M. von Martels, editor. Alchemy Revisited: Proceedings of the International Conference on the History of Alchemy at the University of Groningen, 17–19 April 1989, volume 33 of Collection de travaux de l’Académie internationale d’histoire des sciences. E. J. Brill, Leiden, The Netherlands, 1990. ISBN 90-04-09287-0. ISSN 0169-7897. LCCN QD23.3 .A43 1990.

Veronesi:2022:OAM

[VMT22] Umberto Veronesi and Marcos Martinón-Torres. The Old Ashmolean Museum and Oxford’s Seventeenth-Century chymical community: a material culture approach to laboratory experiments. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 69(1):19–33, 2022. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2021.2012314.

Vinogradov:2018:CIE

[VP18] Andrei Vinogradov and Stanislav Petriashin. Chemical industry, the environment, and Russian provincial society: The case of the Kokshan Chemical Works (1850–1925). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2): 143–168, 2018. CODEN AM-BXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

vanRooij:2007:IRC

[vR07] Arjan van Rooij. Industrial research as a “corporate counterculture”? The development of the HPO caprolactam process at DSM, 1956–77. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):192–210, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582307X212341.

Vrtis:2023:EAM

[Vrt23]  Donna Bilak  George Vr-

tis. Environmental alchemy: Mercury–gold amalgamation mining and the transformation of the Earth. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):31–53, 2023. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2023.2192131.

Walton:1976:JDM

[Wal76]  Michael T. Walton.  John

Dee’s Monas Hieroglyphica: Geometrical Cabala. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(2):116–123, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1976.23.2.
                                        116.

Walton:1980:BNT

[Wal80] Michael T. Walton. Boyle and Newton on the transmutation of water and air, from the root of Helmont’s tree. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 27(1):11–18, 1980. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1980.27.1.
                                        11.

Wallis:2016:PDN

[Wal16]  Patrick Wallis.  Panaceia’s

daughters: Noblewomen as healers in Early Modern Germany. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(2):186–187, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2016.1227166.

Ward:1976:LEI

[War76a] E. R. Ward. Letter to the Editor industrial mixed acid nitration. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):199–200, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/

abs/10.1179/amb.1976.23.3.
199.

Ward:1976:WFF

[War76b] Robert Ward. What forced by fire: Concerning some influences of chemical thought and practice upon English poetry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(2):80–95, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.2.
80.

Ward:1984:DCB

[War84] Edward Ward. The death of Charles Blachford Mansfield (1819–1855). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(2):68–69, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.2.
68.

Ward:1985:R

[War85] Robert Ward. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 32(1):46–52, 1985. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1985.32.1.
46.

Ward:1989:R

[War89] Robert Ward. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 36(1):42–48, 1989. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1989.36.1.
                                        42.

Ward:1997:BAB

[War97] Robert Ward. Before and after the Bomb — some literary speculations on the use of the atomic Bomb. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 44(2):85–95, 1997. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1997.44.2.
                                        85.

Warner:2008:IRS

[War08] Deborah J. Warner. Ira Remsen, saccharin, and the linear model. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 55(1):50–61, 2008. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582308X255415.

Watson:1995:CEC

[Wat95] Katherine D. Watson. The chemist as expert: The consulting career of Sir William Ramsay. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 42(3):143–159, 1995. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1995.42.3.
143.

Watson:2017:THP

[Wat17] Katherine D. Watson. Toxic histories: Poison and pollution in modern India. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(1):104–105, 2017. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1310430.

Webster:1966:WUP

[Web66] C. Webster. Water as the ultimate principle of nature: the background to Boyle’s Sceptical Chymist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(2):96–107, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.2013.13.2.
005.

Webster:1967:EMR

[Web67a] C. Webster. English medical reformers of the Puritan revolution: a background to the “Society of Chymical Physitians”.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(1):16–41, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1967.14.1.
                                        16.

Webster:1967:HPE

[Web67b] C. Webster. Henry Power’s experimental philosophy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 14(3):150–178, 1967. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1967.14.3.
                                        150.

Webster:1982:R

[Web82] C. Webster. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(3):203–206, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1982.29.3.
                                        203.

Weeks:2007:FBA

[Wee07] Sophie Weeks. Francis Bacon and the art–nature distinction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(2):117–145, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1179/174582307X212332.

Weininger:2018:DRT

[Wei18] Stephen J. Weininger. Delayed reaction: The tardy embrace of physical organic chemistry by the German chemical community. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(1):52–75, 2018. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2017.1418709.

Wellesz:1951:MTG

[Wel51] Egon Wellesz. Music in the treatises of Greek gnostics and alchemists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 4(3–4):145–158, 1951. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1951.4.3-
4.145.

Welles:1982:UAS

[Wel82] Elizabeth B. Welles. The unpublished alchemical sonnets of Felice Feliciano: an episode in science and humanism in 15th Century Italy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 29(1):1–16, 1982. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1982.29.1.
                                        1.

Werrett:2013:GCS

[Wer13] Simon Werrett. Green is the colour: St. Petersburg’s chemical laboratories and competing visions of chemistry in the Eighteenth Century. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 60(2):122–138, 2013. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1179/0002698013Z.
                                        00000000027.

West:1961:NIA

[Wes61] Muriel West. Notes on the importance of alchemy to modern science in the writings of Francis Bacon and Robert Boyle. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 9(2):102–114, 1961. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1961.9.2.
                                        102.

Westfall:1975:INI

[Wes75] Richard S. Westfall. Isaac Newton’s Index Chemicus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(3):174–185, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.3.

174. See response [All19b].

Westfall:1980:NRB

[Wes80] Richard S. Westfall. Never at Rest: a Biography of Isaac Newton. Cambridge University Press, Cambridge, UK, 1980. ISBN 0-521-23143-4 (hardcover), 0-521-27435-4 (paperback), 1-107-34066-7 (e-book). xviii + 908 pp. LCCN QC16.N7 W35.

URL http://www.loc.gov/
catdir/description/cam041/
79026294.html;   http://
www.loc.gov/catdir/toc/cam041/
79026294.html.

Westfall:1984:ANL

[Wes84] Richard S. Westfall. Alchemy in Newton’s library. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(3):97–101, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.3.
97.

Weyer:1976:IAN

[Wey76] Jost Weyer. The image of alchemy in nineteenth and Twentieth Century histories of chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(2):65–79, 1976. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1976.23.2.65.

Whitney:2018:BRS

[Whi18] Kristoffer Whitney. Book review: Smell Detectives: An Olfactory History of Nineteenth-Century Urban America. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(4):400–402, 2018. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1080/00026980.2018.
                                        1478178.

Wilkinson:1962:NEL

[Wil62] Ronald Sterne Wilkinson. New England’s last alchemists. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 10(3):128–138, 1962. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1962.10.3.
                                        128.

Wilkinson:1963:ALJ

[Wil63a] Ronald Sterne Wilkinson. The alchemical library of John Winthrop, Jr. (1606–1676) and his descendants in Colonial America. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(1):33–51, 1963. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/amb.1963.11.1.33.

Wilkinson:1963:GSP

[Wil63b] Ronald Sterne Wilkinson. George Starkey, physician and alchemist. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 11(3):121–152, 1963. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1963.11.3.
121.

Wilkinson:1964:PIE

[Wil64] Ronald Sterne Wilkinson. The problem of the identity of Eirenaeus Philalethes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 12(1):24–43, 1964. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1964.12.1.
24.

Wilkinson:1966:ALJ

[Wil66] Ronald Sterne Wilkinson. The alchemical library of John Winthrop, Jr. (1606–1676) and his descendants in Colonial America. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 13(3):139–186, 1966. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/

                                        abs/10.1179/amb.1966.13.3.
                                        139.

Wilkinson:1968:HPS

[Wil68] Ronald Sterne Wilkinson. The Hartlib papers and Seventeenth-Century chemistry. Part I. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 15(1):54–69, 1968. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1968.15.1.
                                        54.

Wilkinson:1970:HPS

[Wil70a] Ronald Sterne Wilkinson. The Hartlib papers and Seventeenth-Century chemistry. Part II. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(2):85–110, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.2.
                                        85.

Wilkinson:1970:R

[Wil70b] Ronald Sterne Wilkinson. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 17(3):192–200, 1970. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1970.17.3.

Wilkinson:1972:FTI

[Wil72] Ronald S. Wilkinson. Further thoughts on the identity of “Eirenaeus Philalethes”. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):204–208, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.3.
204.

Wilkinson:1973:SBP

[Wil73] Ronald Sterne Wilkinson. Some bibliographical puzzles concerning George Starkey. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 20(3):235–244, 1973. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1973.20.3.
235.

Williams:1975:PSM

[Wil75] A. R. Williams. The production of saltpetre in the Middle Ages. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 22(2):125–133, 1975. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1975.22.2.
125.

Williams:1976:R

[Wil76] Trevor I. Williams. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 23(3):201–206, 1976. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1976.23.3.
201.

Wilson:1988:JNP

[Wil88] C. Anne Wilson. Jabirian numbers, Pythagorean numbers and Plato’s Timaeus. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 35(1):1–13, 1988. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1988.35.1.
1.

Williams:1991:R

[Wil91] Trevor I. Williams. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(3):168–172, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1991.38.3.
168.

Wilson:1998:PTD

[Wil98] C. Anne Wilson. Pythagorean theory and Dionysian practice: The cultic and practical background to chemical experimentation in Hellenistic Egypt. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(1):14–33, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1998.45.1.
14.

Williams:2009:NLI

[Wil09] Alan Williams. A note on liquid iron in Medieval Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 56(1):68–75, 2009. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/174582309X405246.

Wilson:2012:CMF

[Wil12]  Jennifer Wilson.  Celebrat-

ing Michael Faraday’s discovery of benzene. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 59(3):241–265, 2012. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1179/174582312X13457672281821.

Winder:1983:R

[Win83] Marianne Winder. Reviews.

Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(1):54–64, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.1.
54.

Winder:1991:R

[Win91] Marianne Winder. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 38(2):100–111, 1991. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1991.38.2.
                                        100.

Winder:1992:R

[Win92] Marianne Winder. Reviews. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 39(1):32–46, 1992. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1992.39.1.
                                        32.

Winder:1998:C

[Win98] Marianne Winder. Correction. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 45(1): 33, 1998. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1998.45.1.
                                        33.

Whitworth:2016:OTP

[WK16] Isabella Whitworth and Zvi C. Koren. Orchil and Tyrian purple: Two centuries of Bedfords from Leeds. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 63(3):244–267, 2016. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/00026980.2016.1246160.

Wolfenden:1972:ASE

[Wol72]  John H. Wolfenden.  The

anomaly of strong electrolytes. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 19(3):175–196, 1972. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1972.19.3.
175.

Wotiz:1983:KK

[WR83] John H. Wotiz and Susanna Rudofsky. Kekulé or Kekule? Ambix: Journal of the Society for the History of Alchemy and Chemistry, 30(3):133–136, 1983. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1983.30.3.
133.

Wujastyk:1984:AGR

[Wuj84] Dominik Wujastyk. An alchemical ghost: The Rasaratnâkara by Nâgârjuna. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(2):70–83, 1984. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/abs/
                                        10.1179/amb.1984.31.2.70.

White:2007:MEF

[WW07] Bruce D. White and Walter W. Woodward. “A most exquisite fellow” — William White and an Atlantic world perspective on the Seventeenth-Century chymical furnace. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 54(3):285–298, 2007. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1179/174582307X237065.

Zacharias:1956:CRH

[Zac56] Procopios D. Zacharias. Chymeutike: The real Hellenic chemistry. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 5(3–4):116–128, 1956. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.1179/amb.1956.5.3-
                                        4.116.

Zaitseva:2005:CCU

[ZH05] Elena A. Zaitseva and Ernst Homburg. Catalytic chemistry under Stalin. Science and scientists in times of repression. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 52(1):45–65, 2005. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/abs/
10.1179/000269805X52904.

Zhao:2023:IMM

[Zha23] Wenrui Zhao. Ingenuity in the making: Matter and technique in Early Modern Europe. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 70(1):99–100, 2023. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/00026980.2022.2146352.

Zimmermann:1984:SMC

[Zim84] Rainer E. Zimmermann. The structure of mythos: on the cultural stability of alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(3):125–137, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.3.
125.

Zinguer:1984:ALR

[Zin84] Ilana Zinguer. Alchemy, “locus” of renewal for writing in the Moyen de Parvenir of Béroalde de Verville (1610). Ambix: Journal of the Society for the History of Alchemy and Chemistry, 31(1):6–15, 1984. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

/www.tandfonline.com/doi/
abs/10.1179/amb.1984.31.1.

Zott:1993:DSS

[Zot93] Regine Zott. The development of science and scientific communication: Justus Liebig’s two famous publications of 1840. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 40(1):1–10, 1993. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        abs/10.1179/amb.1993.40.1.
                                        1.

Zuber:2018:DSF

[Zub18] Mike A. Zuber. The duke, the soldier of fortune, and a Rosicrucian legacy: Exploring the roles of manuscripts in earlymodern alchemy. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 65(2):122–142, 2018. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic).

Zuber:2021:APF

[Zub21]  Mike A. Zuber.  Alchemi-

cal promise, the fraud narrative, and the history of science from below: a German adept’s encounter with Robert Boyle and Ambrose Godfrey. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 68(1):28–48, 2021. CO-DEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/00026980.2020.1847750.